### v1

# ContextFlow: Context-Aware Flow Matching for Trajectory Inference from Spatial Omics Data

Abstract
Inferring trajectories from longitudinal spatially-resolved omics data is fundamental to understanding the dynamics of structural and functional tissue changes in development, regeneration and repair, disease progression, and response to treatment. We propose ContextFlow, a novel context-aware flow matching framework that incorporates prior knowledge to guide the inference of structural tissue dynamics from spatially resolved omics data. Specifically, ContextFlow integrates local tissue organization and ligand-receptor communication patterns into a transition plausibility matrix that regularizes the optimal transport objective. By embedding these contextual constraints, ContextFlow generates trajectories that are not only statistically consistent but also biologically meaningful, making it a generalizable framework for modeling spatiotemporal dynamics from longitudinal, spatially resolved omics data. Evaluated on three datasets, ContextFlow consistently outperforms state-of-the-art flow matching methods across multiple quantitative and qualitative metrics of inference accuracy and biological coherence. Our code is available at: https://github.com/santanurathod/ContextFlow
1 Introduction
Flow matching (Lipman et al., 2023) is an emerging paradigm that provides an efficient approach for learning the complex latent dynamics, or normalizing flows (Papamakarios et al., 2021), of a system of variables, while enabling parametric flexibility to model data distributions. Inferring the underlying dynamics from sparse and noisy observations is a central challenge in many domains (Gontis et al., 2010; Brunton et al., 2016; Pandarinath et al., 2018; Li et al., 2025), where continuous trajectories are rarely captured; instead, cross-sectional snapshots, collected at discrete time points, are typically available. In single-cell RNA sequencing (scRNA-seq), this challenge becomes especially critical as the destructive nature of profiling technologies yields only unpaired population-level snapshots over time. Uncovering temporal dynamics from such snapshot data is essential for understanding developmental processes, disease progression, treatment and perturbation responses (Wagner & Klein, 2020). Traditional approaches often rely on heuristics or computationally intensive likelihood-based generative models, which struggle with scalability and flexibility in high-dimensional single-cell data. Flow matching overcomes these challenges by directly learning continuous latent dynamics that are constrained to match observed population-level distributions at sampled time points.
The state and function of cells within a tissue are affected by interactions with neighboring cells, extracellular matrix components, and local signaling gradients (Rao et al., 2021). Recent advances in spatial omics technologies, particularly spatial transcriptomics (ST), allow gene expression profiling without tissue dissociation, thereby preserving spatial context and providing a complementary view of cellular organization. The dynamics of complex cellular processes is affected by the tissue microenvironment, where cells engage in reciprocal communication with their neighbors (Dimitrov et al., 2022; Tanevski et al., 2025). A growing body of work highlights the critical role of spatial cell–cell communication patterns in shaping cellular phenotypes (Armingol et al., 2021). In particular, location-specific communication circuits between distinct cell types dynamically interact to reprogram cellular states and influence tissue-level behavior (Mayer et al., 2023; Aguadé-Gorgorió et al., 2024; Zheng et al., 2025). These insights, made possible by the spatiotemporal resolution of transcriptomics data, pave the way for understanding the mechanisms by which cellular interactions drive tissue organization and function in organogenesis (Chen et al., 2022), regeneration (Ben-Moshe et al., 2021; Wei et al., 2022), disease progression (Kukanja et al., 2024), and treatment response (Liu et al., 2024).
Optimal transport (OT) has become a foundational framework to align spatially resolved samples and infer putative developmental or temporal couplings  (Zeira et al., 2022; Liu et al., 2023). As a result, state-of-the-art flow matching frameworks such as minibatch-OT flow matching (MOTFM) (Tong et al., 2024) use OT-derived couplings to define conditional paths to train velocity fields, thus overcoming the lack of generative capabilities in optimal transport. The OT formulation adopted in MOTFM, however, does not account for the contextual richness present in spatial transcriptomics and can result in trajectories that are statistically optimal yet biologically implausible (see Figure 3(a) in Appendix G.1 for an illustration). While recent studies have extended widely-used OT objectives (Halmos et al., 2025; Ceccarelli et al., 2025) for spatial transcriptomics, they primarily focus on pairwise alignment of populations across conditions or modalities and do not explicitly incorporate the cell–cell communication patterns that drive cellular state transitions.
To address the above limitations, we introduce a novel flow matching-based framework, ContextFlow, that incorporates spatial priors for modeling temporal tissue dynamics (Figure 1). By encoding local tissue organization and ligand-receptor-derived spatial communication patterns into prior-regularized optimal transport formulations, ContextFlow fully exploits the contextual richness of spatial omics data and embeds both structural and functional aspects of tissue organization into its objective, thereby generating more biologically informed trajectories. In summary, our contributions are as follows:
  • We leverage local tissue organization and local ligand–receptor communication patterns to extract biologically meaningful features from spatial omics data, and encode them into a biologically-informed transition plausibility matrix to constrain temporal dynamics (Section 3.2).
  • We design two novel integration schemes—cost-based and entropy-based—that incorporate the prior knowledge into an OT-coupled flow matching framework, both amenable to efficient Sinkhorn optimization and scalable on modern hardware (Section 3.3).
  • Comprehensive experiments on regeneration and developmental datasets demonstrate that ContextFlow consistently outperforms baseline methods under both interpolation and extrapolation settings across metrics that capture biological plausibility and statistical fidelity (Section 4).
Figure 1: ContextFlow integrates local tissue organization and ligand-receptor communications to learn biologically meaningful trajectories from spatial omics data. Prior knowledge acts as a soft filter that discourages implausible transitions while preserving flexibility in trajectory inference.
2 Preliminaries
2.1 Flow Matching Basics
Flow matching (Lipman et al., 2023) is a simulation-free and sample-efficient generative framework for training continuous normalizing flows (Chen et al., 2018). Given a pair of source and target data distributions over $\mathbb{R}^{d}$ with densities $q_{0}=q(\bm{x}_{0})$ and $q_{1}=q(\bm{x}_{1})$, the problem task is to learn a time-varying velocity vector field $u_{\theta}:[0,1]\times\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$, whose continuous evolution is captured by a function in the form of a neural-net-based model with weights $\theta$, that can transform $q_{0}$ to $q_{1}$ through integration via an ordinary differential equation (ODE). To be more specific, flow matching (FM) seeks to optimize $\theta$ by minimizing a simple regression loss between $u_{\theta}$ and a target time-varying velocity vector field $u_{t}:[0,1]\times\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$ as follows:
$\min_{\theta}\>\mathbb{E}_{t\sim\mathcal{U}(0,1),\bm{x}\sim p_{t}(\bm{x})}\big\|u_{\theta}(t,\bm{x})-u_{t}(\bm{x})\big\|^{2}.$ (1)
Here, $\mathcal{U}(0,1)$ is the uniform distribution over $[0,1]$, and $p_{t}:[0,1]\times\mathbb{R}^{d}\rightarrow\mathbb{R}_{+}$ denotes a time-varying probability path induced by $u_{t}$ such that (i) $p_{t}$ is a probability density function for any $t\in[0,1]$, (ii) $p_{t}$ satisfies the two boundary conditions: $p_{t=0}=q_{0}$ and $p_{t=1}=q_{1}$, and (iii) the connection between $p_{t}$ and $u_{t}$ can be characterized by the transport equation (Villani et al., 2008): $\frac{\partial p_{t}(\bm{x})}{\partial t}=-\nabla\cdot(u_{t}(\bm{x})p_{t}(\bm{x}))$, where $\nabla$ is the divergence operator. From a dynamical system’s view, $u_{t}$ defines an ODE system $d\bm{x}=u_{t}(\bm{x})dt$. The corresponding solution to the ODE, usually termed as the probability flow, can then transport any $\bm{x}_{0}\sim q_{0}$ to a point $\bm{x}_{1}\sim q_{1}$ along $u_{t}$ from $t=0$ to $t=1$. While the flow matching objective in Equation 1 is simple and intuitive, it is generally intractable in practice: the closed-form velocity vector field $u_{t}$ is unknown for arbitrary source and target distributions ($q_{0}$ and $q_{1}$), and multiple valid probability paths $p_{t}$ may exist between them.
2.2 Conditional Flow Matching
The central idea of conditional flow matching is to express the target probability path $p_{t}$ via a mixture of more manageable conditional probability paths (Lipman et al., 2023). By marginalizing over some conditioning variable $z$, both $p_{t}$ and $u_{t}$ can be constructed using their conditional counterparts:
$p_{t}(\bm{x})=\int p_{t}(\bm{x}|\bm{z})q(\bm{z})d\bm{z},\quad u_{t}(\bm{x})=\int u_{t}(\bm{x}|\bm{z})\frac{p_{t}(\bm{x}|\bm{z})q(\bm{z})}{p_{t}(\bm{x})}d\bm{z},$ (2)
where $q(\bm{z})$ denotes the distribution of the conditioning variable $\bm{z}$, and $p_{t}(\bm{x}|\bm{z})$ is selected such that the boundary conditions are satisfied: $\int p_{t=0}(\bm{x}|\bm{z})q(\bm{z})=q_{0}$ and $\int p_{t=1}(\bm{x}|\bm{z})q(\bm{z})=q_{1}$. Theorem 1 of Lipman et al. (2023) proves that $p_{t}$ and $u_{t}$ defined by Equation 2 satisfy the transport equation, suggesting that $p_{t}$ is a valid probability path generated by $u_{t}$. To avoid the intractable integrals, Lipman et al. (2023) proposed the following optimization of conditional flow matching (CFM), and proved its equivalence to the original flow matching objective in terms of gradient computation:
$\min_{\theta}\>\mathbb{E}_{t\sim\mathcal{U}(0,1),z\sim q(\bm{z}),\bm{x}\sim p_{t}(\bm{x}|\bm{z})}\big\|u_{\theta}(t,\bm{x})-u_{t}(\bm{x}|\bm{z})\big\|^{2}.$ (3)
By choosing an appropriate conditional velocity vector field $u_{t}(\bm{x}|\bm{z})$, we can train the neural network using Equation 3 without requiring a closed-form solution of the conditional probability path $p_{t}(\bm{x}|\bm{z})$, thus avoiding the intractable integration operation. Therefore, the remaining task is to define the conditional probability path and velocity vector field properly such that we can sample from $p_{t}(\bm{x}|\bm{z})$ and compute $u_{t}(\bm{x}|\bm{z})$ efficiently for solving the optimization problem in Equation 3.
Gaussian Conditional Probability Paths. A specific choice proposed in Lipman et al. (2023) is Gaussian conditional probability paths and their corresponding conditional velocity vector fields:
$p_{t}(\bm{x}|\bm{z})=\mathcal{N}(\bm{x}\>|\>\mu_{t}(\bm{z}),\sigma_{t}(\bm{z})^{2}\mathbf{I}),\quad u_{t}(\bm{x}|\bm{z})=\frac{\sigma_{t}^{\prime}(\bm{z})}{\sigma_{t}(\bm{z})}(\bm{z}-\mu_{t}(\bm{z}))+\mu_{t}^{\prime}(\bm{z}),$ (4)
where $\mu_{t}:[0,1]\times\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$ denotes the time-varying mean of the Gaussian distribution, $\mu_{t}^{\prime}$ is its derivative with respect to time, $\sigma_{t}:[0,1]\times\mathbb{R}^{d}\rightarrow\mathbb{R}_{+}$ stands for the time-varying scalar standard deviation, and $\sigma_{t}^{\prime}$ stand for the corresponding derivative. In particular, Lipman et al. (2023) set $q(\bm{z})=q(\bm{x}_{1})$, $\mu_{t}(\bm{z})=t\bm{x}_{1}$, and $\sigma_{t}(\bm{z})=1-(1-\sigma)t$. Then, we can see that $u_{t}(\bm{x}|\bm{z})$ transports the standard Gaussian distribution $p_{t=0}(\bm{x}|\bm{z})=\mathcal{N}(\bm{x};0,I)$ to a Gaussian distribution with mean $x_{1}$ and standard deviation $\sigma$, namely $p_{t=1}(\bm{x}|\bm{z})=\mathcal{N}(\bm{x};\bm{x}_{1},\sigma^{2})$ for any target point $\bm{x}_{1}$. By letting $\sigma\rightarrow 0$, the marginal boundary conditions can easily be verified. Tong et al. (2024) further generalized the application scope to arbitrary source distributions, by setting
$q(\bm{z})=q(\bm{x}_{0})q(\bm{x}_{1}),\quad\mu_{t}(\bm{z})=(1-t)\bm{x}_{0}+t\bm{x}_{1},\quad\sigma_{t}(\bm{z})=\sigma.$ (5)
This choice satisfies the boundary conditions $p_{t=0}(\bm{x})=q_{0}$ and $p_{t=1}(\bm{x})=q_{1}$ when $\sigma\rightarrow 0$. Based on Equation 4, the conditional velocity vector field has a simple analytical form $u_{t}(\bm{x}|\bm{z})=\bm{x}_{1}-\bm{x}_{0}$.
2.3 Flow Matching with Optimal Transport Couplings
The conditionals construction specified by Equation 5 corresponds to the simplest choice of independent coupling, where $\bm{z}=(\bm{x}_{0},\bm{x}_{1})$ with source $\bm{x}_{0}$ and target $\bm{x}_{1}$ are independently sampled from $q(\bm{z})=q(\bm{x}_{0})q(\bm{x}_{1})$. The use of couplings for constructing the sampling paths in the CFM framework naturally connects to the optimal transport theory (Villani et al., 2008). Choosing OT-based couplings has several advantages over independent coupling, including smaller training variance and more efficient sampling (Pooladian et al., 2023; Tong et al., 2024).
Since the classical Kantorovich’s formulation (refer Appendix E) has computational complexity that is cubic with respect to the sample size, a popular alternative is to add an extra regularization term, resulting in entropic optimal transport (EOT), to approximately solve the optimal transport problem while reducing the computational costs from cubic to quadratic:
$\pi^{*}_{\mathrm{eot}}(\epsilon):={\arg\!\min}_{\pi\in\Pi(q_{0},q_{1})}\int_{\mathbb{R}^{d}\times\mathbb{R}^{d}}\|\bm{x}_{0}-\bm{x}_{1}\|_{2}^{2}\>d\pi(\bm{x}_{0},\bm{x}_{1})+\epsilon H(\pi\>|\>q_{0}\otimes q_{1}),$ (6)
where $\epsilon>0$ is the regularization parameter, and $H(\pi\>|\>q_{0}\otimes q_{1})$ denotes the relative entropy (or Kullback-Leibler divergence) with respect to $\pi$ and the product measure $q_{0}\otimes q_{1}$. The optimization problem in Equation 6 can be viewed as a special case of the static Schrödinger bridge problem (Bernton et al., 2022), which can be efficiently solved in a mini-batch fashion via the Sinkhorn algorithm (Cuturi, 2013). Theoretically, one can prove that $\pi^{*}_{\mathrm{eot}}(\epsilon)$ recovers the Kantorovich’s OT coupling $\pi^{*}_{\mathrm{ot}}$ when $\epsilon\rightarrow 0$ (see Equation 16 in Appendix E for its formal definition) and $\pi^{*}_{\mathrm{eot}}(\epsilon)$ corresponds to the independent coupling $q_{0}\otimes q_{1}$ when $\epsilon\rightarrow\infty$.
3 Regularizing the Flow with Spatial Priors
3.1 Problem Formulation
We focus on the task of inferring spatiotemporal trajectories, i.e., inferring the dynamic evolution of the cell states across time from spatially resolved gene expression data. Let $0=t_{1}<t_{2}<\ldots<t_{m+1}=1$ be a sequence of normalized time points. For simplicity, we use $[m]$ to denote the set $\{1,2,\ldots,m\}$. For any $i\in[m+1]$, let $q_{i}$ be the data distribution at time $t_{i}$ over $\mathbb{R}^{d}$. Given $\{\mathbf{X}_{t_{i}}\}_{i\in[m+1]}$, where $\mathbf{X}_{t_{i}}=[\bm{x}_{i}(k)]_{k\in[n_{i}]}$ is the expression matrix at time $t_{i}$ consisting of $n_{i}$ snapshot data sampled from $q_{i}$, the objective is to learn a neural velocity vector field $u_{\theta}:[0,1]\times\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$ to faithfully characterize the temporal evolution of spatially resolved tissues over time, such that the induced probability path $p_{t}$ can describe the state of each cell at time $t\in[0,1]$. This task can be viewed as a continuous temporal generalization of the pairwise generative modeling task described in Section 2.1.
A promising candidate solution is conditional flow matching with entropic OT couplings (EOT-CFM), by targeting linear conditional velocity vector fields for each pair of consecutive time points. Specifically, for any $t\in[0,1]$ satisfying $t\in[t_{i},t_{i+1}]$, define
$p(\bm{x}|\bm{z})=\mathcal{N}\left(\frac{t_{i+1}-t}{t_{i+1}-t_{i}}\bm{x}_{i}+\frac{t-t_{i}}{t_{i+1}-t_{i}}\bm{x}_{i+1},\sigma^{2}\mathbf{I}\right),\quad u_{t}(\bm{x}|\bm{z})=\frac{\bm{x}_{i+1}-\bm{x}_{i}}{t_{i+1}-t_{i}},$ (7)
where the conditioning variable is selected as $\bm{z}=(\bm{x}_{i},\bm{x}_{i+1})$, and $p(\bm{z})$ is the joint probability measure with marginals $q_{i}$ and $q_{i+1}$ corresponding to the EOT coupling $\pi^{*}_{\mathrm{eot}}(\epsilon)$ defined in Equation 6. It can be easily verified that the above construction satisfies the boundary condition at each time point $t_{i}$. To train $u_{\theta}$, we can randomly sample a mini-batch of data at each time, run the Sinkhorn algorithm (Cuturi, 2013) to obtain the entropic OT couplings for each consecutive pair, and iteratively update the model weights $\theta$ using stochastic gradient descent with CFM regression loss (Equation 3).
Despite their enhanced ability to model system dynamics, state-of-the-art OT-CFM frameworks lack provisions to fully exploit the contextual richness and integrate the biological prior knowledge that can be inferred from other associated data modalities. Existing approaches can generate statistically optimal trajectories by targeting probability paths induced by (entropic) OT couplings along the temporal dimension. However, they may overlook important functional or structural prior information, leading to biologically implausible trajectories (see Figure 3(a) in Appendix G.1 for an illustration).
3.2 Introducing Spatial Priors & Transitional Plausibility
To faithfully model the spatial context and cellular organization of spatial omics data, we introduce two types of spatial priors and explain how they relate to the transitional plausibility between locations and cell states at different time points.
Spatial Smoothness. Tissues are well-organized systems. Within a microenvironment, neighboring cells respond to the same set of external mechanical stimuli and intercellular communication, which affects their states in a similar manner and results in local smoothness of cell-type-specific expression. Due to tissue heterogeneity, we cannot assume a common reference coordinate frame across tissue samples or even slices at $t_{i}$ and $t_{j}$ at a larger scale. However, the same heterogeneity allows us to consider the spatial coherence and neighborhood consistency (Greenwald et al., 2024; Ceccarelli et al., 2025) as a proxy for relative cell localization, which cannot change significantly across short time intervals. Therefore the aggregate expression within the microenvironment of each cell can be used to quantify the transitional plausibility in consecutive time points.
Specifically, let $c_{i}=(\bm{x}_{i},\bm{s}_{i})$ and $c_{j}=(\bm{x}_{j},\bm{s}_{j})$ be cells at time points $t_{i}$ and $t_{j}$, respectively, where $\bm{x}_{i},\bm{x}_{j}\in\mathbb{R}^{d}$ denote their gene expression profiles, and $\bm{s}_{i},\bm{s}_{j}\in\mathbb{R}^{2}$ denote their spatial coordinates in the relative tissue reference frame. Let $\mathrm{TP}(c_{i},c_{j})$ denote the transitional plausibility, i.e., the likelihood that $c_{i}$ evolves to $c_{j}$ between $t_{i}$ and $t_{j}$. Spatial smoothness suggests that $\mathrm{TP}(c_{i},c_{j})$ is inversely related to the difference between the average expression profiles of their local neighborhoods:
$\mathrm{SS}(c_{i},c_{j})=\bigg\|\>\frac{1}{|\mathcal{N}_{r}(c_{i})|}\sum_{c\in\mathcal{N}_{r}(c_{i})}\bm{x}(c)-\frac{1}{|\mathcal{N}_{r}(c_{j})|}\sum_{c\in\mathcal{N}_{r}(c_{j})}\bm{x}(c)\>\bigg\|_{2}^{2},$ (8)
where $\mathcal{N}_{r}(c_{i})=\{c:\|\bm{s}(c)-\bm{s}(c_{i})\|_{2}\leq r\}$ denotes the set of neighboring cells of $c_{i}$ in the same tissue slice, $|\mathcal{N}_{r}(c_{i})|$ is the cardinality of $\mathcal{N}_{r}(c_{i})$, and $\bm{x}(c)$ is the gene expression profile of cell $c$.
Cell-Cell Communication Patterns. Cell–cell communication (CCC) has a critical role in the regulation of numerous biological processes, including development, apoptosis, and the maintenance of homeostasis in health and disease (Armingol et al., 2024). A major type of CCC is ligand–receptor (LR) signaling, in which ligands expressed by one cell bind to cognate receptors on another, initiating intracellular cascades that ultimately affect the state of the cell (i.e., its expression profile)  (Armingol et al., 2021). There are numerous databases of prior knowledge of ligand-receptor binding and computational methods that use these databases to systematically link gene expression with the activity of ligand-receptor-mediated communication.
Specifically, we can represent each cell $c_{i}$ by a vector $f_{\mathrm{LR}}\in\mathbb{R}^{p}$, where each entry corresponds to one of $p$ possible ligand–receptor pairs and encodes the extent of $c_{i}$’s participation in communication through that pair. The $\mathrm{TP}(c_{i},c_{j})$ between cells in different tissue slices is higher when they exhibit similar ligand-receptor communication patterns $f_{\mathrm{LR}}$ (see Figure 7 for an illustration). We define $\mathrm{LR}(c_{i},c_{j})$, the dissimilarity between the ligand–receptor communication patterns in the microenvironments of cells $c_{i}$ and $c_{j}$, as:
$\mathrm{LR}(c_{i},c_{j})=\left\|\>f_{\mathrm{LR}}(\mathcal{N}_{r}(c_{i}))-f_{\mathrm{LR}}(\mathcal{N}_{r}(c_{j}))\right\|_{2}^{2},$ (9)
3.3 ContextFlow: CFM with Context-Aware OT Couplings
Our proposed framework, graphically depicted in Figure 1, consists of the following three main steps:
Transitional Plausibility Matrix. First, we create a sequence of transitional plausibility matrices (TPMs) to encode the biological priors for each pair of consecutive time points. Specifically, let $\mathbf{M}_{i,i+1}\in\mathbb{R}^{n_{i}\times n_{i+1}}$ be the TPM with respect to the set of cells measured at time $t_{i}$ and at time $t_{i+1}$, with size $n_{i}$ and $n_{i+1}$ respectively, where the $(k,l)$-th entry of $\mathbf{M}_{i,i+1}$ indicates how plausibly the $k$-th cell measured at $t_{i}$ will evolve to the $l$-th cell measured at $t_{i+1}$, defined as follows:
$\big[\mathbf{M}_{i,i+1}\big]_{kl}=\lambda\cdot\mathrm{SS}\left(c_{i}(k),c_{i+1}(l)\right)+(1-\lambda)\cdot\mathrm{LR}\left(c_{i}(k),c_{i+1}(l)\right),$ (10)
where $\lambda\in[0,1]$ is a trade-off hyperparameter that balances the contribution of the spatial smoothness prior ($\mathrm{SS}$) and the ligand–receptor communication prior ($\mathrm{LR}$).
Prior-Regularized OT Couplings. The transitional plausibility matrices capture our spatially informed prior on cell-cell transitions between consecutive time points, which can naturally be incorporated in the EOT formulation (Equation 6) to promote couplings that maintain the structural and functional properties of the tissue organization. We propose two techniques for prior integration:
Prior-Aware Cost Matrix (PACM). Consider the empirical counterpart of Equation 6 with respect to time $t_{i}$ and time $t_{i+1}$. Our first approach incorporates the transitional plausibility matrix directly into the transport cost:
$\min_{\mathbf{\Pi}\in\mathbb{R}^{n_{i}\times n_{i+1}}}\sum_{k,l}\Pi_{kl}\underbrace{\Big[\alpha\cdot\|\bm{x}_{i}(k)-\bm{x}_{i+1}(l)\|_{2}^{2}+(1-\alpha)\cdot\big[\mathbf{M}_{i,i+1}\big]_{kl}\Big]}_{\text{Prior-Aware Cost Function}}-\epsilon\sum_{k,l}\Pi_{kl}\log\Pi_{kl},$ (11)
where the transport plan $\Pi$ satisfies the boundary conditions: $\sum_{l}\Pi_{kl}=1/n_{i}$ for any $k\in[n_{i}]$, and $\sum_{k}\Pi_{kl}=1/n_{i+1}$ for any $l\in[n_{i+1}]$, and $\alpha\in[0,1]$ controls the trade-off between the original Euclidean cost and the prior-aware cost derived from the transitional plausibility. If $\mathbf{M}_{i,i+1}]_{kl}$ is high, Equation 11 will impose a higher transport cost between the $k$-cell at time $i$ to the $j$-cell at time $i+1$. This aligns with our assumption that such transitions are implausible.
Prior-Aware Entropy Regularization (PAER). While the prior-aware cost matrix approach penalizes couplings in accordance with our spatial priors, it defines a different OT problem characterized by a modified cost function. Consequently, the standard interpretation of OT as minimizing the transport energy between two transcriptomic distributions no longer holds. Since the scales of the pairwise distances often differ, normalization of the cost terms is required to enable meaningful comparison. This normalization, however, may result in couplings that deviate from their original counterparts (Proposition 1 and Corollary 1 in the Appendix C). Besides, selecting an appropriate $\alpha$ in Equation 11 introduces an additional layer of tuning, increasing computational overhead. Therefore, we propose a second approach to integrate the biological priors without introducing additional hyperparameters:
$\min_{\mathbf{\Pi}\in\mathbb{R}^{n_{i}\times n_{i+1}}}$ $\sum_{k,l}\Pi_{kl}\|\bm{x}_{i}(k)-\bm{x}_{i+1}(l)\|_{2}^{2}-\epsilon\sum_{k,l}\underbrace{\Pi_{kl}\log(\Pi_{kl}/[\widehat{\mathbf{M}}_{i,i+1}]_{kl})}_{\text{Prior-Aware Entropy Regularization}},$ (12)
where $\widehat{\mathbf{M}}_{i,i+1}]_{kl}=\exp(-[\mathbf{M}_{i,i+1}]_{kl})/\sum_{l}\exp(-[\mathbf{M}_{i,i+1}]_{kl})$ denotes the prior joint probability matrix induced by $\mathbf{M}_{i,i+1}$. Intuitively, the lower the cost $\mathbf{M}{i,i+1}]_{kl}$, the larger the entry $\widehat{\mathbf{M}}{i,i+1}]_{kl}$, reflecting a higher plausibility of the transition from cell $k$ at $t_{i}$ to cell $l$ at $t_{i+1}$. The entropy regularization term in Equation 12 thus biases the learned transport plan toward the prior $\widehat{\mathbf{M}}_{i,i+1}$ rather than a uniform baseline, providing a soft mechanism for incorporating biological prior knowledge.
ContextFlow. Finally, we apply the Sinkhorn algorithm (Cuturi, 2013) to solve the optimization problem in Equation 11 or Equation 12 to obtain the spatial context-aware EOT couplings, and train the neural velocity vector field $u_{\theta}$ based on stochastic gradient descent by minimizing the multi-time generalization of Equation 3 with respect to conditionals $p_{t}(\bm{x}|\bm{z})$ and $u_{t}(\bm{x}|\bm{z})$ defined according to Equation 7. The pseudocode for the proposed method, named Conditional Flow Matching with Context-Aware OT Couplings (ContextFlow), is detailed in Algorithm 1 in Appendix D.
In particular, to apply the Sinkhorn algorithm to solve our prior-aware entropy regularization problem in Equation 12, we make use of the following theorem, a generalized result of Peyré et al. (2019).
Theorem 1.
Let $\mathbf{C}\in\mathbb{R}^{n_{0}\times n_{1}}$ be a cost matrix and $\mathbf{M}\in\mathbb{R}^{n_{0}\times n_{1}}$ be a prior transition probability matrix. Suppose $\mathbf{\Pi}^{*}_{\mathrm{CTF-H}}$ is the solution to the following prior-aware optimal transport problem:
$\mathbf{\Pi}^{*}_{\mathrm{CTF-H}}={\arg\!\min}_{\mathbf{\Pi}\in\mathbb{R}^{n_{0}\times n_{1}}}$ $\sum_{k,l}\Pi_{kl}C_{kl}+\epsilon\sum_{k,l}\Pi_{kl}\log(\Pi_{kl}/{M}_{kl}),$
where $\epsilon>0$ is the regularization parameter. Then, we can show that $\mathbf{\Pi}^{*}_{\mathrm{CTF-H}}$ can be computed by Sinkhorn and takes the form $\mathrm{diag}(\bm{u})\cdot\mathbf{M}\odot\exp(-\mathbf{C}/\epsilon)\cdot\mathrm{diag}(\bm{v})$, where $\odot$ denotes element-wise multiplication, and $\bm{u}\in\mathbb{R}^{n_{0}},\bm{v}\in\mathbb{R}^{n_{1}}$ are vectors satisfying the marginalization constraints.
Theorem 1, proven in Appendix B, suggests a new Gibbs kernel $\mathbf{K}=\mathbf{M}\odot\exp(-\mathbf{C}/\epsilon)$, which combines both the transport cost and the prior joint probability matrices. When $\epsilon\rightarrow 0$, $\mathbf{\Pi}^{*}_{\mathrm{CTF}}\rightarrow\mathbf{\Pi}_{\mathrm{ot}}^{*}$, thereby recovering the standard OT couplings in Equation 16. When $\epsilon\rightarrow\infty$, the optimal coupling $\mathbf{\Pi}^{*}_{\mathrm{CTF}}\rightarrow\mathrm{diag}(\bm{u})\cdot\mathbf{M}\cdot\mathrm{diag}(\bm{v})$, which corresponds to a plan that aligns with the prior defined by $\mathbf{M}$ rather than the independent couplings obtained with Entropic-OT (Section 2.3). This has the same effect as constraining our transport plan through the proposed prior and, by extension, the flow. By varying the parameter $\epsilon$, we can thus efficiently optimize for a desirable coupling via the Sinkhorn algorithm.
4 Experiments
Datasets. We evaluate ContextFlow on three longitudinal spatial transcriptomics datasets: Axolotl Brain Regeneration (Wei et al., 2022), Mouse Embryo Organogenesis (Chen et al., 2022), and Liver Regeneration (Ben-Moshe et al., 2021). For all the datasets, the gene expression values are log-normalized, and we extract the top $50$ principal components (PCs) as feature vectors. The strength of ligand–receptor interactions in the microenvironment was inferred using spatially informed bivariate statistics implemented in LIANA+ (Dimitrov et al., 2024), where we applied the cosine similarity metric to gene expression profiles. Interaction evidence was aggregated using the consensus of multiple curated ligand–receptor resources, ensuring robustness of the inferred signals.
Baselines & Metrics. We benchmark ContextFlow using its two prior integration strategies—cost-regularized (CTF-C) and entropy-regularized (CTF-H)—against several baselines. As a non-spatial baseline, we include conditional flow matching (CFM), which uses only transcriptomic data with random couplings. We further compare against minibatch-OT flow matching (MOTFM), which leverages OT-derived couplings but does not incorporate spatial priors. For evaluation, we employed 2-Wasserstein distance ($\mathcal{W}_{2}$), a commonly used OT-based metric, and metrics such as MMD and Energy Distance for statistical fidelity. Furthermore, to assess the biological plausibility of our predicted dynamics, we evaluate them using a cell-type-weighted Wasserstein distance (Weighted $\mathcal{W}_{2}$), where the weights correspond to the relative frequency of each cell type in the dataset. Exact metric definitions are present in the Appendix F. All reported metrics are averaged across $10$ runs.
Sampling. A trained velocity field can be evaluated through the samples it generates. We consider two variants. Initial value problem sampling (IVP) integrates the learned gradient starting from the first observed batch of cells and evolves them toward a later time point. IVP provides the most comprehensive evaluation of flow quality, as errors can accumulate across steps. In contrast, next-step sampling (Next Step) integrates the gradient only from the most recently observed batch of cells, thus limiting error propagation but providing a less stringent test of long-term trajectory fidelity.
Table 1: Interpolation at the middle holdout time point for the Brain Regeneration dataset.
Sampling Method $\bm{\lambda}$ $\bm{\alpha}$ Weighted $\bm{\mathcal{W}_{2}}$ $\bm{\mathcal{W}_{2}}$ MMD Energy
Next Step CFM $2.618\pm 0.142$ $2.579\pm 0.197$ $0.043\pm 0.003$ $12.505\pm 1.271$
MOTFM $2.567\pm 0.088$ $2.476\pm 0.161$ $0.040\pm 0.003$ $11.269\pm 1.388$
CTF-C 11 $0.8$ $423\pm 0.164 2.2$$293\pm 0.103$ $0.037\pm 0.001$ $9.874\pm 0.659$
0 $0.2$ $2.396\pm 0.028$ $2.100\pm 0.102$ $0.033\pm 0.003$ $8.577\pm 0.976$
$0.5$ $0.8$ $2.442\pm 0.173$ $2.353\pm 0.241$ $0.035\pm 0.004$ $9.008\pm 2.094$
CTF-H 0 $2.528\pm 0.143$ $2.534\pm 0.180$ $0.040\pm 0.004$ $11.192\pm 1.304$
11 $\bm{2.316\pm 0.141}$ $\bm{1.969\pm 0.221}$ $\bm{0.030\pm 0.004}$ $\bm{6.359\pm 1.336}$
$0.5$ $2.519\pm 0.167$ $2.412\pm 0.158$ $0.039\pm 0.004$ $10.304\pm 1.808$
IVP CFM $4.216\pm 0.463$ $4.266\pm 0.308$ $0.170\pm 0.029$ $32.413\pm 5.122$
MOTFM $198\pm 0.319 4.4$$452\pm 0.243$ $0.173\pm 0.017$ $33.149\pm 3.321$
CTF-C 11 $0.8$ $3.603\pm 0.300$ $3.816\pm 0.310$ $0.127\pm 0.018$ $24.271\pm 3.992$
0 $0.2$ $\bm{3.465\pm 0.232}$ $\bm{3.641\pm 0.320}$ $0.119\pm 0.025$ $23.055\pm 5.939$
$0.5$ $0.8$ $4.015\pm 0.351$ $3.974\pm 0.442$ $0.140\pm 0.038$ $27.592\pm 6.669$
CTF-H 0 $3.925\pm 0.267$ $4.375\pm 0.297$ $0.164\pm 0.013$ $32.034\pm 3.270$
11 $3.905\pm 0.395$ $4.188\pm 0.685$ $\bm{0.074\pm 0.014}$ $\bm{18.728\pm 2.689}$
$0.5$ $3.917\pm 0.343$ $4.159\pm 0.455$ $0.147\pm 0.022$ $29.613\pm 4.822$
4.1 Axolotl Brain Regeneration
We first evaluate ContextFlow on longitudinal Stereo-seq spatial transcriptomic data coming from a post-traumatic brain regeneration study of the Salamander (axolotl telencephalon) species (Wei et al., 2022). The dataset contains samples from five developmental stages, with replicates collected from different individual organisms at each stage. For our CTF-C method, we present the best ablated $\alpha$ in the main text, with full ablation results across different $\alpha$ values provided in Appendix H.
For interpolation, we hold out the middle time point during training and evaluate it using samples generated by the trained velocity field $u_{\theta}$ via both IVP and next-step sampling. Table 1 presents the results. Across multiple evaluation metrics, ContextFlow with entropy regularization (CTF-H) produces trajectories that most closely match the ground truth. CTF-H consistently achieves the best or comparable performance relative to CTF-C, despite the latter being explicitly tuned across multiple $\alpha$ values. This highlights the computational efficiency and superior generalization ability of CTF-H, as it avoids the need for additional hyperparameter tuning while maintaining strong performance.
For extrapolation, we evaluate generation on the last holdout time point, representing the most challenging test of generalizability for the velocity fields $u_{\theta}$, as it lies outside the training time horizon. As shown in Table 2, CTF-H again consistently achieves the best overall performance, particularly under IVP-Sampling, where errors are most likely to accumulate. This result further reinforces the robustness and reliability of CTF-H across the entire sampling horizon. Finally, Figure 3 (Appendix G.1) demonstrates that incorporating spatial priors enables ContextFlow to produce substantially fewer biologically implausible couplings compared to its context-free counterpart.
Table 2: Extrapolation on the last holdout time point for the Brain Regeneration dataset.
Sampling Method $\bm{\lambda}$ $\bm{\alpha}$ Weighted $\bm{\mathcal{W}_{2}}$ $\bm{\mathcal{W}_{2}}$ MMD Energy
Next Step CFM $7.124\pm 0.443$ $7.133\pm 0.533$ $0.275\pm 0.011$ $76.947\pm 5.661$
MOTFM $619\pm 0.611 7.7$$769\pm 0.763$ $0.272\pm 0.007$ $85.352\pm 8.140$
CTF-C 11 $0.5$ $68\pm 0.608 6.969$$69\pm 0.628$ $0.265\pm 0.009$ $77.025\pm 6.056$
0 $0.5$ $7.244\pm 0.804$ $7.146\pm 0.775$ $0.265\pm 0.003$ $80.424\pm 10.376$
$0.5$ $0.5$ $7.188\pm 0.391$ $\bm{6.931\pm 0.260}$ $0.267\pm 0.005$ $78.992\pm 6.195$
CTF-H 0 $\bm{6.914\pm 0.471}$ $7.198\pm 0.726$ $0.266\pm 0.009$ $\bm{76.149\pm 8.436}$
11 $7.505\pm 0.667$ $7.338\pm 0.601$ $\bm{0.263\pm 0.006}$ $83.425\pm 8.793$
$0.5$ $7.243\pm 0.479$ $7.157\pm 0.641$ $0.270\pm 0.007$ $79.826\pm 8.067$
IVP CFM $6.633\pm 1.312$ $7.116\pm 1.084$ $0.143\pm 0.037$ $60.573\pm 21.756$
MOTFM $6.503\pm 0.720$ $6.352\pm 0.592$ $0.162\pm 0.038$ $56.452\pm 15.932$
CTF-C 11 $0.5$ $6.260\pm 0.616$ $7.681\pm 4.003$ $0.157\pm 0.039$ $52.478\pm 12.010$
0 $0.5$ $6.614\pm 0.710$ $6.854\pm 0.740$ $0.201\pm 0.023$ $70.370\pm 9.099$
$0.5$ $0.5$ $6.696\pm 0.427$ $6.481\pm 0.387$ $0.195\pm 0.024$ $66.212\pm 3.542$
CTF-H 0 $6.243\pm 0.760$ $6.220\pm 0.751$ $0.195\pm 0.020$ $61.316\pm 10.288$
11 $\bm{5.277\pm 0.936}$ $6.021\pm 1.192$ $\bm{0.099\pm 0.007}$ $\bm{27.777\pm 8.621}$
$0.5$ $6.254\pm 0.819$ $\bm{5.973\pm 0.757}$ $0.156\pm 0.025$ $54.330\pm 12.089$
Table 3: Interpolation (time 5) and extrapolation (time 8) results on the Organogenesis dataset.
Method $\bm{\lambda}$ $\bm{\alpha}$ Next Step (Interpolation) IVP (Interpolation) Next Step (Extrapolation)
Weighted $\bm{\mathcal{W}_{2}}$ $\bm{\mathcal{W}_{2}}$ Weighted $\bm{\mathcal{W}_{2}}$ $\bm{\mathcal{W}_{2}}$ Weighted $\bm{\mathcal{W}_{2}}$ $\bm{\mathcal{W}_{2}}$
MOTFM $1.892\pm 0.028$ $1.873\pm 0.086$ $3.251\pm 0.676$ $3.418\pm 0.727$ $1.626\pm 0.066$ $1.682\pm 0.096$
CTF-C 11 $0.5$ $\bm{1.865\pm 0.030}$ $1.852\pm 0.093$ $3.137\pm 0.407$ $4.093\pm 1.187$ $1.685\pm 0.096$ $1.714\pm 0.160$
0 $0.8$ $1.882\pm 0.022$ $1.869\pm 0.049$ $2.938\pm 0.476$ $3.904\pm 1.120$ $1.773\pm 0.053$ $1.880\pm 0.180$
$0.5$ $0.8$ $1.888\pm 0.033$ $\bm{1.839\pm 0.134}$ $3.200\pm 0.403$ $3.555\pm 0.637$ $1.768\pm 0.058$ $1.858\pm 0.120$
11 $0.2$ $1.880\pm 0.020$ $1.922\pm 0.078$ $3.260\pm 0.880$ $5.264\pm 3.060$ $1.683\pm 0.058$ $1.803\pm 0.117$
0 $0.2$ $1.900\pm 0.035$ $1.912\pm 0.057$ $2.953\pm 0.425$ $3.816\pm 0.970$ $1.715\pm 0.123$ $1.860\pm 0.267$
CTF-H 0 $1.884\pm 0.027$ $1.862\pm 0.123$ $3.244\pm 0.713$ $3.946\pm 1.671$ $\bm{1.505\pm 0.057}$ $\bm{1.397\pm 0.088}$
11 $1.898\pm 0.029$ $1.866\pm 0.097$ $5.200\pm 0.799$ $6.306\pm 1.037$ $1.890\pm 0.046$ $1.877\pm 0.103$
$0.5$ $1.871\pm 0.030$ $1.919\pm 0.067$ $\bm{2.814\pm 0.414}$ $\bm{3.233\pm 0.567}$ $1.636\pm 0.060$ $1.684\pm 0.099$
(a) Brain Regeneration
(b) Organogenesis
Figure 2: KL-Divergence between predicted and ground-truth cell type distributions.
Table 4: Interpolation results on the middle holdout time point for the Liver Regeneration dataset.
MOTFM CTF-C CTF-H
($\lambda$, $\alpha$) (11, $0.5$) (0, $0.5$) ($0.5$, $0.8$) (0, –) (11, –) ($0.5$, –)
$\bm{\mathcal{W}_{2}}$ $4.303\pm 1.448 33$$3.506\pm 1.148$ $2.741\pm 1.864 33$$3.045\pm 1.644$ $\bm{32.682\pm 1.472}$ $3.481\pm 1.001 33$$3.414\pm 0.995$
4.2 Mouse Embryo Organogenesis
We further evaluated ContextFlow on the larger Mouse Organogenesis Spatiotemporal Atlas (MOSTA) Stereo-seq dataset (Chen et al., 2022) spanning measurements from $8$ developmental time points. For the interpolation study of this dataset, we held out time point 5 during training and evaluated its generation during testing. Table 3 shows the evaluation results. We observe that ContextFlow, with both integration strategies, outperforms MOTFM across all metrics, showcasing the effectiveness of the contextual information. While CTF-C shows stronger performance under next-step sampling—albeit only after fine-tuning the trade-off parameter $\alpha$—CTF-H consistently outperforms it in the more challenging IVP-Sampling setting. On the extrapolation task, integrating to the final time point, CTF-H again achieves the strongest performance, underscoring that the entropy-regularized formulation not only removes the need for additional parameter tuning but also offers more robust generalization to unseen temporal horizons.
Figure 2 reports the KL-Divergence between normalized histograms of predicted and ground-truth cell types from ContextFlow and MOTFM. In both cases, CTF exhibits lower divergence on average across time points, indicating that the trajectories generated by our model better preserve the biological composition of cell types over time. The cell type progression is further visualized in Figure 9 (Appendix H.6). We show that ContextFlow predicts temporal cell type trajectories that evolve smoothly and consistently across consecutive developmental stages. Early progenitor populations, such as neural crest and mesenchyme, progressively diminish as development advances, while terminal fates, including muscle, cartilage primordium, and liver, emerge at later stages. Major lineages such as brain, heart, and connective tissue remain continuous throughout, demonstrating that ContextFlow captures biologically coherent and temporally consistent developmental dynamics.
4.3 Liver Regeneration
Finally, we evaluate ContextFlow on a Visium spatial transcriptomics dataset profiling the temporal dynamics of mouse liver regeneration following acetaminophen-induced injury (Ben-Moshe et al., 2021), collected across three distinct regeneration stages. Unlike the earlier datasets resolved at single-cell resolution, Visium data is captured at the level of $55$ micron diameter spots, capturing the joint expression of multiple cells. Since direct cell-type information is unavailable, we restrict evaluation to the 2-Wasserstein distance. Moreover, since evaluation is performed on the middle of the three time points, IVP and next-step predictions coincide. Table 4 presents the results. Consistent with the previous findings, CTF-H achieved the lowest reconstruction error, indicating that incorporating contextual information improves trajectory estimation even in aggregated spot-level measurements.
5 Conclusion
We introduced ContextFlow, a contextually aware flow matching framework that leverages spatial priors and biologically motivated constraints to learn more plausible trajectories from snapshot spatial transcriptomic data, addressing a central challenge of existing methods. The entropic variant of ContextFlow is theoretically grounded, which always yields OT couplings constrained by prior knowledge, promoting stability and consistency with the imposed contextual constraints. Across three diverse datasets, we showed that ContextFlow consistently improves over state-of-the-art baselines even in challenging Initial Value Problem sampling settings, underscoring the importance of our contextually informed priors. In addition, we demonstrated that our framework reduces the number of biologically implausible couplings and results in coherent and temporally consistent developmental trajectories while maintaining strong quantitative performance across Wasserstein, MMD, and Energy metrics. These results highlight the value of embedding biological context into generative flow models. Future works can adapt our methods to reconstruct tissues and learn spatial latent dynamics by formulating the flow in space (rather than time), or leverage multi-marginal OT formulations for optimizing temporal flows. Looking forward, ContextFlow offers a principled foundation for modeling perturbations and disease progression, bridging generative power with biological interpretability.
Appendix A Related Work
A.1 Flow Matching
Normalizing flows provide a parametric framework for characterizing transformations of a random variable into desired distributions (Papamakarios et al., 2021). These transformations can be realized through either finite (Rezende & Mohamed, 2015) or continuous compositions (Chen et al., 2018). The loss functions used in such formulations typically require computing Jacobians or integrating the flows at each forward pass, making them computationally expensive. Flow matching (FM) (Lipman et al., 2023; Albergo & Vanden-Eijnden, 2022; Liu et al., 2022) addresses this limitation by reducing the training of the velocity field to a regression problem, thereby making normalizing flows substantially more scalable. To ensure valid conditional paths at intermediate time points, samples are coupled either randomly or via optimal transport (Pooladian et al., 2023; Tong et al., 2024). Owing to this scalability, FM has been rapidly adopted across scientific domains, including biology and the life sciences (Li et al., 2025). In transcriptomics, for example, Klein et al. (2024) employed an FM backbone to approximate OT maps for drug response modeling and cross-modal translation tasks. Entropic OT formulations have also been applied to infer cellular trajectories (Tong et al., 2024; Rohbeck et al., 2025), generate imaging-based cell morphology changes (Zhang et al., 2025), and simulate spatial transcriptomics data from histology images (Huang et al., 2025).
Despite these advances, existing work does not address how to meaningfully incorporate biological prior knowledge to constrain the velocity field, limiting the biological plausibility of inferred trajectories.
A.2 Optimal Transport
Omics studies frequently generate uncoupled measurements across conditions, modalities, or time points, which must be integrated into a unified representation to provide a more comprehensive view of the underlying biology. Optimal transport (OT) has recently gained popularity for this task, as it provides a geometry-based approach to couple probability distributions (Bunne et al., 2024; Klein et al., 2025). In spatial transcriptomics (ST), several OT formulations have been introduced depending on context. For instance, Zeira et al. (2022) and Liu et al. (2023) proposed PASTE and PASTE2 to align ST data from adjacent tissue slices, while DeST-OT (Halmos et al., 2025) integrates spatio-temporal slices by modeling cell growth and differentiation. Rahimi et al. (2024) developed DOT, a multi-objective OT framework for mapping features across scRNA-seq and spatially resolved assays, and Ceccarelli et al. (2025) introduced TOAST, a spatially regularized OT framework for slice alignment and annotation transfer.
While these methods are primarily designed to align biological data across space, time, or modality, they do not address the problem of trajectory inference toward biologically plausible solutions, leveraging biological priors to constrain or bias the transport plan.
Appendix B Proofs of Main Theoretical Results
Proposition 1.
Let $\mathbf{C}\in\mathbb{R}^{n_{0}\times n_{1}}$ be a cost matrix and $\mathbf{M}\in\mathbb{R}^{n_{0}\times n_{1}}$ a prior transition matrix with positive entries. Consider the entropy-regularized OT formulation:
$\Pi^{*}$ $={\arg\!\min}_{\Pi\geq 0}\;\sum_{k,l}\Pi_{kl}C_{kl}+\epsilon\sum_{k,l}\Pi_{kl}\log\!\left(\Pi_{kl}\right).$
Let $\tilde{\Pi}^{*}$ be the EOT-coupling where the cost is scaled by a normalization constant $c$ or $\tilde{C}_{ij}=\frac{C_{ij}}{c}$. Let the regularization parameter $\epsilon>0$ be the same in both cases. Then, for indices $(i,j)$ and $(k,l)$,
$\frac{\tilde{\Pi}^{*}_{ij}}{\tilde{\Pi}^{*}_{kl}}\leq\gamma\,\left(\frac{\Pi^{*}_{ij}}{\Pi^{*}_{kl}}\right)^{\frac{1}{c}},$
where $\gamma$ depends on $\Pi^{*}_{ij},c$ and the OT marginal constraints $a,b$.
Proof.
For the original optimal transport (OT) formulation, we note:
$\Pi^{*}_{ij}=u_{i}K_{ij}v_{j},\quad K_{ij}=e^{-C_{ij}/\epsilon},$
with the constraints $\Pi^{*}\mathbf{1}=a$ and $\Pi^{*\top}\mathbf{1}=b$.
Let
$\Pi_{ij}^{*1/c}=u_{i}^{1/c}K_{ij}^{1/c}v_{j}^{1/c},$
where:
$\tilde{K}_{ij}=K_{ij}^{1/c}=\exp\left(-C_{ij}/(c\epsilon)\right)$
is the kernel for the scaled/normalized OT formulation. Let $\tilde{\Pi}^{*}_{ij}$ be the coupling for the scaled version, then:
$\tilde{\Pi}^{*}_{ij}=\tilde{u}_{i}\tilde{K}_{ij}\tilde{v}_{j}.$
Thus, there exist scaling factors $\alpha_{i},\beta_{j}\in\mathbb{R}$ such that:
$\tilde{u}_{i}$ $=\alpha_{i}u^{\frac{1}{c}}_{i},$
$\tilde{v}_{j}$ $=\beta_{j}v^{\frac{1}{c}}_{j}.$
This implies:
$\tilde{\Pi}^{*}_{ij}=(\alpha_{i}u_{i}^{1/c})\tilde{K}_{ij}(\beta_{j}v_{j}^{1/c}),$
$\implies\tilde{\Pi}^{*}=\operatorname{diag}(\alpha u^{1/c})\tilde{K}\operatorname{diag}(\beta v^{1/c}),$ (g1)
$\implies\tilde{\Pi}^{*}=\operatorname{diag}(\alpha)\Pi^{1/c}\operatorname{diag}(\beta).$
Subject to the constraints:
$\sum_{i}\alpha_{i}\beta_{j}\Pi_{ij}^{*1/c}=a_{i},\quad\sum_{i}\alpha_{i}\beta_{j}\Pi_{ij}^{*1/c}=b_{j}.$
For any pair (i,j)&(k,l)(i,j)\&(k,l), we can express:
$\frac{\tilde{\Pi}^{*}_{ij}}{\tilde{\Pi}^{*}_{kl}}=\frac{\alpha_{i}}{\alpha_{k}}\frac{\beta_{j}}{\beta_{l}}\left(\frac{\Pi^{*}_{ij}}{\Pi^{*}_{kl}}\right)^{1/c}.$
Taking logarithms on both sides, we have:
$\log\left(\frac{\tilde{\Pi}^{*}_{ij}}{\tilde{\Pi}^{*}_{kl}}\right)=\log(\alpha_{i})-\log(\alpha_{k})+\log(\beta_{j})-\log(\beta_{l})+\frac{1}{c}\log\left(\frac{\Pi^{*}_{ij}}{\Pi^{*}_{kl}}\right).$
Let $\log(\alpha)=\phi$ and $\log(\beta)=\psi$, then:
$\log\left(\frac{\tilde{\Pi}^{*}_{ij}}{\tilde{\Pi}^{*}_{kl}}\right)=(\phi_{i}-\phi_{k})+(\psi_{j}-\psi_{l})+\frac{1}{c}\log\left(\frac{\Pi^{*}_{ij}}{\Pi^{*}_{kl}}\right).$
This implies:
$\left|\log\left(\frac{\tilde{\Pi}^{*}_{ij}}{\tilde{\Pi}^{*}_{kl}}\right)-\frac{1}{c}\log\left(\frac{\Pi^{*}_{ij}}{\Pi^{*}_{kl}}\right)\right|\leq|\phi_{i}|+|\phi_{k}|+|\psi_{j}|+|\psi_{l}|.$
From Proposition 3 Proposition 2, we have:
$\max_{i}\phi_{i}\leq E,\quad\max_{i}\psi_{i}\leq E.$
Thus:
$\left|\log\left(\frac{\tilde{\Pi}^{*}_{ij}}{\tilde{\Pi}^{*}_{kl}}\right)-\frac{1}{c}\log\left(\frac{\Pi^{*}_{ij}}{\Pi^{*}_{kl}}\right)\right|\leq 4E.$
Therefore:
$4E+\frac{1}{c}\log\left(\frac{\Pi^{*}_{ij}}{\Pi^{*}_{kl}}\right)\leq\log\left(\frac{\tilde{\Pi}^{*}_{ij}}{\tilde{\Pi}^{*}_{kl}}\right)\leq 4E+\frac{1}{c}\log\left(\frac{\Pi^{*}_{ij}}{\Pi^{*}_{kl}}\right).$
This implies:
$\frac{\tilde{\Pi}^{*}_{ij}}{\tilde{\Pi}^{*}_{kl}}\leq\exp(4E)\left(\frac{\Pi^{*}_{ij}}{\Pi^{*}_{kl}}\right)^{1/c}.$
Let $\gamma=\exp(4E)$, then:
$\frac{\tilde{\Pi}^{*}_{ij}}{\tilde{\Pi}^{*}_{kl}}\leq\gamma\left(\frac{\Pi^{*}_{ij}}{\Pi^{*}_{kl}}\right)^{1/c}.$
Corollary 1.
Let $\mathbf{C}\in\mathbb{R}^{n_{0}\times n_{1}}$ be a cost matrix and $\mathbf{M}\in\mathbb{R}^{n_{0}\times n_{1}}$ a prior transition matrix with positive entries. Consider the entropy-regularized OT formulation:
$\Pi^{*}$ $={\arg\!\min}_{\Pi\geq 0}\;\sum_{k,l}\Pi_{kl}C_{kl}+\epsilon\sum_{k,l}\Pi_{kl}\log\!\left(\Pi_{kl}\right).$
Let $\tilde{\Pi}^{*}$ be the EOT-coupling in the case when cost is scaled by a normalization constant $c$ or $\tilde{C}_{ij}=\frac{C_{ij}}{c}$. Let the regularization parameter $\epsilon>0$ be the same in both cases. Then:
$H(\tilde{\Pi}_{ij})\geq mH(\Pi_{ij})-s,$
where $m$ and $s$ are constants that depend on $\Pi^{*}$, the marginalization constants $a,b$ and the normalization constant $c$.
Proof.
From equation (g1) in Proposition 1 above, we know that:
$\tilde{\Pi}^{*}_{ij}=(\Pi^{*}_{ij})^{1/c}\cdot\exp(\phi_{i},\psi_{j})$
and from Proposition 2, we have that,
$\tilde{\Pi}^{*}_{ij}\leq(\Pi^{*}_{ij})^{1/c}\cdot e^{2E}$
$\Rightarrow\log(\tilde{\Pi}^{*}_{ij})\leq\frac{1}{c}\log(\Pi^{*}_{ij})+2E$
$\Rightarrow-\tilde{\Pi}^{*}_{ij}\log(\tilde{\Pi}^{*}_{ij})\geq-\frac{1}{c}(\Pi^{*}_{ij})^{1/c-1}\cdot\Pi^{*}_{ij}\log(\Pi^{*}_{ij})\cdot e^{2E}-2E\cdot e^{2E}\cdot(\Pi^{*}_{ij})^{1/c}$
For $c\gg 1$, $\frac{1}{c}\to 0$:
$\Rightarrow-\tilde{\Pi}^{*}_{ij}\log(\tilde{\Pi}^{*}_{ij})$ $\geq-\frac{1}{c\Pi^{*}_{ij}}\cdot\Pi^{*}_{ij}\log(\Pi^{*}_{ij})\cdot e^{2E}-2E\cdot e^{2E}\cdot(\Pi^{*}_{ij})^{1/c}$
$\Rightarrow-\tilde{\Pi}^{*}_{ij}\log(\tilde{\Pi}^{*}_{ij})$ $\geq-\frac{1}{c\Pi^{*}_{\min}}\cdot\Pi^{*}_{ij}\log(\Pi^{*}_{ij})\cdot e^{2E}-2E\cdot e^{2E}\cdot(\Pi^{*}_{ij})^{1/c}$
Summing for all $(i,j)$ we get,
$H(\tilde{\Pi^{*}})\geq mH(\Pi^{*})-s,$
where $m=\frac{e^{2E}}{c\Pi^{*}_{\min}}$ and $s=2E\cdot e^{2E}$. ∎
Proposition 2.
Let $\mathbf{C}\in\mathbb{R}^{n_{0}\times n_{1}}$ be a cost matrix and $\mathbf{M}\in\mathbb{R}^{n_{0}\times n_{1}}$ a prior transition matrix with positive entries. Consider the entropy-regularized OT formulation:
$\Pi^{*}$ $={\arg\!\min}_{\Pi\geq 0}\;\sum_{k,l}\Pi_{kl}C_{kl}+\epsilon\sum_{k,l}\Pi_{kl}\log\!\left(\Pi_{kl}\right).$
Let $\tilde{\Pi}^{*}$ be the EOT-coupling in the case when cost is scaled by a normalization constant $c$ or $\tilde{C}_{ij}=\frac{C_{ij}}{c}$. Let the regularization parameter $\epsilon>0$ be the same in both cases. Consider the scaling factors $\alpha,\beta$ such that: $\tilde{u}_{i}=\alpha_{i}u_{i}^{1/c}$, $\tilde{v}_{j}=\beta_{j}v_{j}^{1/c}$ where $u,v$ are the Sinkhorn algorithm converged vectors for the original setting and $\tilde{u},\tilde{v}$ are for the cost-scaled version. Then, we have
$\max\{\left\|\phi\right\|_{\infty},\left\|\psi\right\|_{\infty}\}\leq\|M^{-1}\|_{\infty}\cdot\left\|\begin{pmatrix}\Delta_{a}\\ \Delta_{b}\end{pmatrix}\right\|_{\infty}$,
where $\phi=\log(\alpha)$ and $\psi=\log(\beta)$. We also have that,
$\max_{i}|\alpha_{i}-1|,\max_{i}|\beta_{i}-1|\leq\|M^{-1}\|_{\infty}\max(\|\Delta_{a}\|_{\infty},\|\Delta_{b}\|_{\infty}),$
where $M,\Delta_{a},\Delta_{b}$ depend on $\Pi^{*}$, marginalization constants $a,b$ and normalization constant $c$.
Proof.
Let $X_{ij}=\Pi_{ij}^{*1/c}$ and $X=\Pi^{*1/c}$. Consider the exponentiated versions of $\alpha$ and $\beta$:
$\phi=\log(\alpha)\in\mathbb{R}^{n},\quad\psi=\log(\beta)\in\mathbb{R}^{m}.$
From the marginal constraints, we have:
$\sum_{j}X_{ij}e^{\phi_{i}+\psi_{j}}=a_{i},\quad\sum_{i}X_{ij}e^{\phi_{i}+\psi_{j}}=b_{j}.$
Applying a first-order Taylor expansion gives:
$\sum_{j}X_{ij}(1+\phi_{i}+\psi_{j})=a_{i}\quad\implies\quad\sum_{j}X_{ij}(\phi_{i}+\psi_{j})=a_{i}-\sum_{j}X_{ij},$
$\sum_{i}X_{ij}(1+\phi_{i}+\psi_{j})=b_{j}\quad\implies\quad\sum_{i}X_{ij}(\phi_{i}+\psi_{j})=b_{j}-\sum_{i}X_{ij}.$
Define:
$\Delta_{a_{i}}=a_{i}-\sum_{j}X_{ij},\quad\Delta_{b_{j}}=b_{j}-\sum_{i}X_{ij}.$
Thus, we have:
$\sum_{j}X_{ij}(\phi_{i}+\psi_{j})=\Delta a_{i},\quad\sum_{i}X_{ij}(\phi_{i}+\psi_{j})=\Delta b_{j}.$
This implies:
$\phi_{i}\left(\sum_{j}X_{ij}\right)+\sum_{j}X_{ij}\psi_{j}=\Delta a_{i},$
$\sum_{i}X_{ij}\phi_{i}+\psi_{j}\left(\sum_{i}X_{ij}\right)=\Delta b_{j}.$
Let:
$D_{r}=\operatorname{diag}(X\mathbf{1})\in\mathbb{R}^{n\times n},\quad D_{c}=\operatorname{diag}(X^{T}\mathbf{1})\in\mathbb{R}^{m\times m}.$
Then we can express the system as:
$\begin{pmatrix}D_{r}&X\\ X^{T}&D_{c}\end{pmatrix}\begin{pmatrix}\phi\\ \psi\end{pmatrix}=\begin{pmatrix}\Delta a\\ \Delta b\end{pmatrix}$.
Let:
$M=\begin{pmatrix}D_{r}&X\\ X^{T}&D_{c}\end{pmatrix}$.
Thus:
$\begin{pmatrix}\phi\\ \psi\end{pmatrix}=M^{-1}\begin{pmatrix}\Delta_{a}\\ \Delta_{b}\end{pmatrix}$.
This implies:
$\left\|\begin{pmatrix}\phi\\ \psi\end{pmatrix}\right\|\leq\|M^{-1}\|\cdot\left\|\begin{pmatrix}\Delta_{a}\\ \Delta_{b}\end{pmatrix}\right\|$.
Since $\alpha=\exp(\phi)$ and $\beta=\exp(\psi)$, by assumption:
$|\alpha_{i}-1|\approx|\exp(\phi_{i})-1|\approx\phi_{i},$
$|\beta_{j}-1|\approx|\exp(\psi_{j})-1|\approx\psi_{j}.$
Therefore:
$\max_{i}|\alpha_{i}-1|,\max_{j}|\beta_{j}-1|\leq\|M^{-1}\|_{\infty}\cdot\max(\|\Delta a\|_{\infty},\|\Delta b\|_{\infty}).$
Theorem 1.
Let $\mathbf{C}\in\mathbb{R}^{n_{0}\times n_{1}}$ be a general cost matrix and $\mathbf{M}\in\mathbb{R}^{n_{0}\times n_{1}}$ be a prior transition probability matrix. Suppose $\Pi^{*}_{\mathrm{CTF-H}}$ is the solution to the following prior-aware optimal transport problem:
$\Pi^{*}_{\mathrm{CTF-H}}=\operatorname*{arg\,min}_{\Pi\in\mathbb{R}^{n_{0}\times n_{1}}}$ $\sum_{k,l}\Pi_{kl}C_{kl}+\epsilon\sum_{k,l}\Pi_{kl}\log(\Pi_{kl}/{M}_{kl}),$
where $\epsilon>0$ is the regularization parameter. Then, we can show that $\Pi^{*}_{\mathrm{CTF-H}}$ can be computed by the Sinkhorn algorithm and takes the form $\mathrm{diag}(\bm{u})\cdot\mathbf{M}\odot\exp(-\mathbf{C}/\epsilon)\cdot\mathrm{diag}(\bm{v})$, where $\odot$ stands for the elementwise multiplication, and $\bm{u}\in\mathbb{R}^{n_{0}},\bm{v}\in\mathbb{R}^{n_{1}}$ are vectors satisfying the marginalization constraints.
Proof.
We have that:
$\Pi^{*}_{\mathrm{CTF-H}}=\operatorname*{arg\,min}_{\Pi\in\mathbb{R}^{n_{0}\times n_{1}}}$ $\sum_{k,l}\Pi_{kl}C_{kl}+\epsilon\sum_{k,l}\Pi_{kl}\log(\Pi_{kl}/{M}_{kl}),$
Subject to:
$\Pi\mathbf{1}=a,\quad\Pi^{\top}\mathbf{1}=b.$
This formulation is a standard convex optimization setting with constraints. The Lagrangian of this setting is:
$\mathcal{L}(\Pi,f,g)=\sum_{k,l}C_{kl}\Pi_{kl}+\epsilon\sum_{k,l}\Pi_{kl}\left(\log\left(\frac{\Pi_{kl}}{M_{kl}}\right)-1\right)-\sum_{k}f_{k}\left(\sum_{l}\Pi_{kl}-a_{k}\right)-\sum_{l}g_{l}\left(\sum_{k}\Pi_{kl}-b_{l}\right)$
Differentiating with respect to $\Pi_{kl},f_{k},g_{l}$, we get:
$\frac{\partial\mathcal{L}}{\partial\Pi_{kl}}=C_{kl}+\epsilon\log\left(\frac{\Pi_{kl}}{M_{kl}}\right)-f_{k}-g_{l}$
Setting the derivative to zero:
$\epsilon\log\left(\frac{\Pi^{*}_{kl}}{M_{kl}}\right)=f_{k}-C_{kl}-g_{l}$
$\implies\frac{\Pi^{*}_{kl}}{M_{kl}}=e^{\frac{f_{k}}{\epsilon}}e^{-\frac{C_{kl}}{\epsilon}}e^{\frac{g_{l}}{\epsilon}}$
$\implies\Pi^{*}_{kl}=e^{\frac{f_{k}}{\epsilon}}M_{kl}e^{-\frac{C_{kl}}{\epsilon}}e^{\frac{g_{l}}{\epsilon}}$
Let $u\in\mathbb{R}^{n}$ and $v\in\mathbb{R}^{m}$ such that:
$u_{k}=e^{\frac{f_{k}}{\epsilon}},\quad v_{l}=e^{\frac{g_{l}}{\epsilon}}$
Let $K_{kl}$ be the kernel $M_{kl}e^{-C_{kl}/\epsilon}$.
Then, we have:
$\Pi^{*}_{kl}=u_{k}K_{kl}v_{l}$
$\Pi^{*}=\operatorname{diag}(u)\cdot K\cdot\operatorname{diag}(v)$ (13)
Differentiating the Lagrangian with respect to $f_{k}$ and $g_{l}$, we get:
$\frac{\partial\mathcal{L}}{\partial f_{k}}=1\cdot\left(\sum_{l}\Pi^{*}_{kl}-a_{k}\right)=0$
$\implies\Pi^{*}\mathbf{1}=a$ (14)
$\frac{\partial\mathcal{L}}{\partial g_{l}}=1\cdot\left(\sum_{i}\Pi^{*}_{kl}-b_{l}\right)=0$
$\implies\Pi^{*\top}\mathbf{1}=b$ (15)
From equations 16 13, 17 14, and 18 15 above, we get:
$\operatorname{diag}(u)\cdot K\cdot\operatorname{diag}(v)\cdot\mathbf{1}$ $=a$
$(\operatorname{diag}(u)\cdot K\cdot\operatorname{diag}(v))^{\top}\mathbf{1}$ $=b$
Which can be rewritten as:
$u\odot(Kv)$ $=a$
$K^{\top}u\odot v$ $=b$
This is the usual matrix scaling formulation for which the Iterative Proportional Fitting (IPF) updates are:
$u_{k}^{t+1}=\frac{a_{k}}{(Kv^{t})_{k}},\quad v_{l}^{t+1}=\frac{b_{l}}{(K^{\top}u^{t+1})_{l}}$
Sinkhorn Algorithm uses these updates, iteratively, and these updates are shown to converge in Franklin & Lorenz (1989). Thus, Sinkhorn Algorithm can be used for the ContextFlow’s Prior Aware Entropy Regularized (PAER) (CTF-H) formulation.
From equation (9) 13, we get:
$\Pi^{*}_{kl}=e^{f_{k}/\epsilon}M_{kl}e^{-C_{kl}/\epsilon}e^{g_{l}/\epsilon}$
When $\epsilon\to\infty$, we have $C_{kl}/\epsilon\to 0$.
$\quad e^{-C_{kl}/\epsilon}\to 1$
$\implies$ $\Pi^{*}_{kl}\to u_{k}M_{kl}v_{l}$
$\implies$ $\Pi^{*}_{\text{CTF-H}}\to\mathrm{diag}(\bm{u})\cdot\mathbf{M}\cdot\mathrm{diag}(\bm{v})$
Such that marginal constraints, $\Pi^{*}_{\text{CTF-H}}\mathbf{1}=a$ and $\Pi^{*\top}_{\text{CTF-H}}\mathbf{1}=b$ are satisfied. ∎
Appendix C Effects of normalization on Prior aware cost matrix
From Peyré et al. (2019), we know that optimal MOTFM coupling takes the form $\Pi^{*}_{\text{EOT}}=\text{diag}(u)\cdot K\cdot\text{diag}(v)$, where $K$ is the kernel matrix such that $K]_{ij}=\exp(\frac{-c_{ij}}{\epsilon})$, with $u,v$ satisfying marginalization constraints $u\odot Kv=a$ and $K^{T}u\odot v=b$. Sinkhorn updates are given by:
$u^{l+1}=\frac{a}{Kv^{l}};v^{l+1}=\frac{b}{K^{T}u^{l+1}}.$
In cases where the OT cost function consists of information from different modalities the distances are usually normalized to have distances of a similar scale. Normalizing the cost results $\tilde{c}_{ij}=\frac{c_{ij}}{\epsilon}$ such that the new kernel matrix $K_{\text{norm}}]_{ij}=\exp(\frac{-c_{ij}}{C_{\text{max}}\epsilon})$ can cause numerical issues if $C_{\text{max}}\gg 1$. The cost normalization should be performed mindfully, when considering different pairwise distances, as in PACM Section 3. Intuitively, scaling the cost has the same effect as that of increasing $\epsilon$, making solutions more diffused.
Proposition 1.
Let $\mathbf{C}\in\mathbb{R}^{n_{0}\times n_{1}}$ be a cost matrix and $\mathbf{M}\in\mathbb{R}^{n_{0}\times n_{1}}$ a prior transition matrix with positive entries. Consider the entropy-regularized OT formulation:
$\Pi^{*}$ $={\arg\!\min}_{\mathbf{\Pi}\geq 0}\;\sum_{k,l}\Pi_{kl}C_{kl}+\epsilon\sum_{k,l}\Pi_{kl}\log\!\left(\Pi_{kl}\right).$
Let $\tilde{\Pi}^{*}$ be the EOT-coupling where the cost is scaled by a normalization constant $c$ or $\tilde{C}_{ij}=\frac{C_{ij}}{c}$. Let the regularization parameter $\epsilon>0$ be the same in both cases. Then, for any indices $(i,j)$ and $(k,l)$ we have
$\frac{\widetilde{\Pi}^{*}_{ij}}{\widetilde{\Pi}^{*}_{kl}}\leq\gamma\,\left(\frac{\Pi^{*}_{ij}}{\Pi^{*}_{kl}}\right)^{\frac{1}{c}},$
where $\gamma$ depends on $\Pi^{*}_{ij},c$ and OT marginal constraints $a,b$.
From Proposition 1, let $\frac{\Pi^{*}_{ij}}{\Pi^{*}_{kl}}=m$, such that $m>1$ ($\Pi^{*}_{ij}>\Pi^{*}_{kl}$ or entries are faraway) then, for $c>1$, we have $\frac{\tilde{\Pi}^{*}_{ij}}{\tilde{\Pi}^{*}_{kl}}<m^{\frac{1}{c}}<m$, for $\gamma<1$, implying that faraway entries are squeezed together. This results in bringing probabilities that are far apart closer to each other or, in essence, in creating more diffused and less sharp couplings.
Corollary 1.
Let $\mathbf{C}\in\mathbb{R}^{n_{0}\times n_{1}}$ be a cost matrix and $\mathbf{M}\in\mathbb{R}^{n_{0}\times n_{1}}$ a prior transition matrix with positive entries. Consider the entropy-regularized OT formulation:
$\mathbf{\Pi}^{*}$ $={\arg\!\min}_{\mathbf{\Pi}\geq 0}\;\sum_{k,l}\Pi_{kl}C_{kl}+\epsilon\sum_{k,l}\Pi_{kl}\log\!\left(\Pi_{kl}\right)$
and $\tilde{\Pi^{*}}$ be EOT-coupling in the case when cost is scaled by a normalization constant $c$ or $\tilde{C}_{ij}=\frac{C_{ij}}{c}$. Let the regularization parameter $\epsilon>0$ be the same in both cases. Then we have:
$H(\tilde{\Pi}_{ij})\geq mH(\Pi_{ij})-s$
where $m$ and $s$ are constants, that depend on $\Pi^{*}$, marginalization constants $a,b$ and normalization constant $c$.
Corollary 1 can also be interpreted as supporting the results of Proposition 1 and our intuition that normalizing has the same effect on the kernel matrix as increasing $\epsilon$, leading to more diffused couplings or couplings with increased entropy.
Appendix D ContextFlow Algorithm
Algorithm 1 ContextFlow (CTF): Flow Matching with Spatial-Context-Aware OT Couplings
1:Input: gene data $\{\mathbf{X}_{t_{1}},\cdots,\mathbf{X}_{t_{m+1}}\}$, spatial data $\{\mathbf{S}_{t_{1}},\ldots,\mathbf{S}_{t_{m+1}}\}$, parameters $\lambda$, $\alpha$, $\epsilon$, $\sigma$, $\eta$
2:Output: neural velocity vector field $u_{\theta}$
3:Initialize $\theta$
4:while training do
5:  for $i=1,2,\ldots,m$ do
6:   Sample a batch $\mathcal{B}=\{(\bm{x}_{i},\bm{x}_{i+1}):(\bm{x}_{i},\bm{x}_{i+1})\sim(\mathbf{X}_{t_{i}},\mathbf{X}_{t_{i+1}})\}$
7:   Construct TPM: $\mathbf{M}_{i,i+1}(\mathcal{B})\triangleright$ $\mathbf{M}_{i,i+1}$ is defined in Equation 10
8:   if “prior-aware cost matrix” then
9:     $C_{kl}\leftarrow\alpha\cdot\|\bm{x}_{i}(k)-\bm{x}_{i+1}(l)\|_{2}^{2}+(1-\alpha)\cdot[\mathbf{M}_{i,i+1}]_{kl}$ for any pair $(k,l)$
10:     $\mathbf{K}\leftarrow\exp(-\mathbf{C}/\epsilon)$
11:   else if “prior-aware entropy regularization” then
12:     $C_{kl}\leftarrow\|\bm{x}_{i}(k)-\bm{x}_{i+1}(l)\|_{2}^{2}$ for any pair $(k,l)$
13:     $\mathbf{K}\leftarrow\widehat{\mathbf{M}}_{i,i+1}\odot\exp(-\mathbf{C}/\epsilon)$ $\triangleright$ $\widehat{\mathbf{M}}_{i,i+1}$ is defined in Equation 12
14:   end if
15:   Initialize $\bm{a}\leftarrow\frac{1}{n_{i}}\mathbf{1}_{n_{i}}$, $\bm{b}\leftarrow\frac{1}{n_{i+1}}\mathbf{1}_{n_{i+1}}$, $\bm{u}\leftarrow\mathbf{1}_{n_{i}}$, $\bm{v}\leftarrow\mathbf{1}_{n_{i+1}}$
16:   while not converged do
17:     $\bm{u}\leftarrow\bm{a}\oslash(\mathbf{K}\bm{v}),\>\bm{v}\leftarrow\bm{b}\oslash(\mathbf{K}^{\top}\bm{u})$ $\triangleright$ Run Sinkhorn algorithm
18:   end while
19:   Obtain spatial-prior-aware OT couplings $\mathbf{\Pi}^{\mathrm{CTF}}_{i,i+1}\leftarrow\operatorname{diag}(\bm{u})\mathbf{K}\operatorname{diag}(\bm{v})$
20:   Sample $t\sim\mathcal{U}(t_{i},t_{i+1})$ and $\{(\bm{x}_{i},\bm{x}_{i+1}):(\bm{x}_{i},\bm{x}_{i+1})\sim\mathbf{\Pi}^{\mathrm{CTF}}_{i,i+1}\}$
21:   Sample $\bm{x}_{t}\sim\mathcal{N}\left(\frac{t_{i+1}-t}{t_{i+1}-t_{i}}\bm{x}_{i}+\frac{t-t_{i}}{t_{i+1}-t_{i}}\bm{x}_{i+1},\sigma^{2}\mathbf{I}\right)$
22:   $L_{\mathrm{CFM}}\leftarrow\frac{1}{|\mathcal{B}|}\sum_{t,(\bm{x}_{i},\bm{x}_{i+1})}\left\|u_{\theta}(\bm{x}_{t},t)-\frac{\bm{x}_{i+1}-\bm{x}_{i}}{t_{i+1}-t_{i}}\right\|_{2}^{2}$
23:  end for
24:  $\theta\leftarrow\theta-\eta\cdot\nabla_{\theta}L_{\mathrm{CFM}}$
25:end while
Appendix E Kantarovich-OT Formulation
Kantorovich’s formulation (Peyré et al., 2019) is a classical definition of the optimal transport (OT) problem that seeks a joint coupling to move a probability measure to another that minimizes the Euclidean distance cost, corresponding to the following minimization problem with respect to the 2-Wasserstein distance:
$\pi^{*}_{\mathrm{ot}}:={\arg\!\min}_{\pi\in\Pi(q_{0},q_{1})}\int_{\mathbb{R}^{d}\times\mathbb{R}^{d}}\|\bm{x}_{0}-\bm{x}_{1}\|_{2}^{2}\>d\pi(\bm{x}_{0},\bm{x}_{1}),$ (16)
where $\Pi(q_{0},q_{1})$ denotes the set of joint probability measures such that the left and right marginals are $q_{0}$ and $q_{1}$. Equation 16 can be solved in a mini-batch fashion using standard solvers such as POT (Flamary et al., 2021); however, the computational complexity is cubic in batch size.
Appendix F Evaluation Metrics
F.1 2-Wasserstein
The 2-Wasserstein distance ($\mathcal{W}_{2}$ between empirical distributions $\mu,\nu$ is defined as:
$\mathcal{W}_{2}(\mu,\nu)=\inf_{\gamma\in\Pi(\mu,\nu)}\bigg(\sum_{(\bm{x},\bm{y})}\gamma(\bm{x},\bm{y})\cdot\|\bm{x}-\bm{y}\|_{2}^{2}\bigg)^{1/2},$
where $\Pi(\mu,\nu)$ denotes the set of couplings between $\mu$ and $\nu$.
F.2 Weighted 2-Wasserstein
Implausible velocity fields can steer a cell’s transcriptional trajectory in unrealistic directions, potentially leading to entirely different terminal cell types. We thus employ the weighted 2-Wasserstein metric, which ensures the evaluation accounts for both transcriptional similarity and the distributional balance of cell types. We define the weighted 2-Wasserstein distance (Weighted $\mathcal{W}_{2}$) between true and predicted distributions as:
$\text{Weighted-$\mathcal{W}_{2}$}(\mu,\nu)=\sum_{i=1}^{C}\frac{n_{i}^{\text{true}}}{N}\cdot\mathcal{W}_{2}\!\bigg(\frac{1}{n_{i}^{\text{true}}}\!\!\sum_{j:y_{j}=i}\delta_{\bm{x}_{j}},\frac{1}{n_{i}^{\text{pred}}}\!\!\sum_{j:\hat{y}_{j}=i}\delta_{\bm{x}_{j}}\bigg),$
where $n_{i}^{\text{true}},n_{i}^{\text{pred}}$ are the number of true and predicted cells of type $i$, and $N$ is the total number of samples. To determine the cell type of generated trajectories, we employ a multi-class classifier $M_{\phi}$, implemented as an XGBoost model (Chen & Guestrin, 2016) trained for each dataset.
F.3 Energy Distance
Let $\mu$ and $\nu$ be probability distributions with samples $X=\{\bm{x}_{i}\}_{i=1}^{m}\sim\mu$ and $Y=\{\bm{y}_{j}\}_{j=1}^{n}\sim\nu$. The squared empirical energy distance (Energy) is defined as:
$\mathrm{ED}(\mu,\nu)=\frac{2}{mn}\sum_{i=1}^{m}\sum_{j=1}^{n}\|\bm{x}_{i}-\bm{y}_{j}\|-\frac{1}{m^{2}}\sum_{i=1}^{m}\sum_{i^{\prime}=1}^{m}\|\bm{x}_{i}-\bm{x}_{i^{\prime}}\|-\frac{1}{n^{2}}\sum_{j=1}^{n}\sum_{j^{\prime}=1}^{n}\|\bm{y}_{j}-\bm{y}_{j^{\prime}}\|,$
where $\|\cdot\|$ is the Euclidean norm. The distance is non-negative and equals zero if and only if $\mu=\nu$.
F.4 Maximum Mean Discrepancy
For the same samples, the unbiased empirical estimate of the squared maximum mean discrepancy (MMD) with kernel $\kappa$ is defined as:
$\mathrm{MMD}(\mu,\nu;\kappa)=\frac{1}{m(m-1)}\sum_{i\neq i^{\prime}}\kappa(\bm{x}_{i},\bm{x}_{i^{\prime}})+\frac{1}{n(n-1)}\sum_{j\neq j^{\prime}}\kappa(\bm{y}_{j},\bm{y}_{j^{\prime}})-\frac{2}{mn}\sum_{i=1}^{m}\sum_{j=1}^{n}\kappa(\bm{x}_{i},\bm{y}_{j}).$
In our evaluations, we use a multi-kernel variant with radial basis function (RBF) kernels $\kappa_{\gamma}(\bm{x},y)=\exp(-\gamma\|\bm{x}-\bm{y}\|^{2})$, and average over $\gamma\in[2,1,0.5,0.1,0.01,0.005]$.
Appendix G Biological Priors and Dataset Visualizations
G.1 (Im-)Plausibility of OT-Couplings
To demonstrate the need for integrating biological priors within a generative framework, we computed the Entropic-OT plan (Section 2.3) for the MOTFM framework and the PAER-OT plan (Section 3.3) for the ContextFlow framework. From these transport plans, we sampled couplings corresponding to the first two stages of the Brain Regeneration dataset (Wei et al., 2022) together with their associated cell types. Figures 3(a) and 3(b) illustrate the Excitatory–Inhibitory lineage switches present in these sampled couplings. Since excitatory and inhibitory neurons have mutually exclusive neurotransmitter functions and originate from distinct progenitor populations with different transcription factor profiles, a transition from excitatory to inhibitory identity is considered biologically implausible.
In our transport plan couplings, we observed the following cell type lineage switches:
  • Immature MSN $\rightarrow$ Immature nptxEX
  • Immature MSN $\rightarrow$ Immature dpEX
  • Immature MSN $\rightarrow$ Immature CMPN
  • Immature nptxEX $\rightarrow$ Immature cckIN
  • Immature nptxEX $\rightarrow$ Immature MSN
(a) MOTFM Implausible Transitions
(b) ContextFlow Implausible Transitions
Figure 3: Comparison of biologically implausible cell type couplings between Stage 0 and Stage 1 of the Brain Regeneration Dataset (Wei et al., 2022), under the Entropic-OT and ContextFlow Regularized-OT formulations. Biological implausibility is defined here as transitions involving excitatory–inhibitory lineage switches. Our formulation produces substantially fewer biologically implausible couplings (24) compared to MOTFM (54).
Of these, $54$ implausible transitions arose from the Entropic-OT plan compared to the $24$ under the PAER-OT plan, with the specific transitions detailed in the figure legends. We also observed that the Entropic-OT formulation produced implausible transitions across brain hemispheres, for example, coupling cells from the left hemisphere with those from the right. In contrast, the PAER-OT formulation typically restricted transitions to within the same hemisphere, reflecting its integration of spatially aware contextual information. These observations provide strong motivation for incorporating biological priors through ContextFlow as a principled approach to learning biologically consistent developmental trajectories.
G.2 Cell type distributions over time
Figures 46 present the spatial maps of the transcriptomics datasets across different time points, illustrating how tissue organization and cell type distributions evolve during development and regeneration. These maps highlight not only changes in cellular composition but also the preservation of spatial neighborhoods and geometrical arrangements of specific cell types over time. Such contextual information, specific to spatial transcriptomics, remains inaccessible to standard flow-matching frameworks. By contrast, ContextFlow is designed to exploit these spatial features, enabling the inference of trajectories that are both temporally smooth and spatially coherent.
G.2.1 Brain Regeneration
Figure 4: Temporal progression of spatial distribution of different cell types for Brain Regeneration.
G.2.2 Mouse Embryo Organogenesis
Figure 5: Temporal progression of spatial distribution of different cell types for Mouse Organogenesis.
G.2.3 Liver Regeneration
Figure 6: Temporal progression of spatial distribution of fibrogenic states for Liver Regeneration. Here, $0/1$ refers to the absence or presence of fibrogenic spots.
G.3 Ligand Receptor Interactions
Figure 7 shows the ligand-receptor score of the NPTX2-NPTXR pair in two consecutive slides from the Brain regeneration dataset (Wei et al., 2022). Similar activities are visible bilaterally in the cerebral cortex, suggesting that ligand–receptor interactions are preserved across time and spatially aligned with underlying tissue structure. This observation provides strong evidence that including LR interactions as contextual priors is biologically meaningful, as they capture functional communication signals between cells that remain stable across short time intervals.
(a) NPTX2-NPTXR LR pair activation on Stage 3
(b) NPTX2-NPTXR LR pair activation on Stage 4
Figure 7: Spatial distributions of LR activation for NPTX2-NPTXR in two consecutive slides from the Brain regeneration dataset. Similar activations are visible at structurally equal positions.
Based on the activation of NPTX2–NPTXR in Figure 7, we observe that the corresponding communication pattern naturally biases the optimal couplings towards transitions such as Immature dpEX $\rightarrow$ dpEX and Immature nptxEX $\rightarrow$ nptxEX (Figure 8). These transitions are biologically plausible, as they preserve cell type identity within excitatory neuronal lineages while reflecting maturation within the same functional context. This example highlights the richness of the contextual information captured by our proposed biological prior, and demonstrates how incorporating such ligand–receptor–driven cues into the coupling process leads to more interpretable and biologically consistent trajectories.
Figure 8: Visual translation of the bias that NPTX2–NPTXR LR pattern provides in terms of cell type coupling for the two consecutive slides.
Appendix H Additional Experiments & Ablations
H.1 Next Step Sampling for Axolotl Brain Regeneration
Table 5: Interpolation via Next Step Sampling at holdout time 3 for the Brain Regeneration dataset.
Sampling Method $\bm{\lambda}$ $\bm{\alpha}$ Weighted $\bm{\mathcal{W}_{2}}$ $\bm{\mathcal{W}_{2}}$ MMD Energy
Next Step CFM $2.618\pm 0.142$ $2.579\pm 0.197$ $0.043\pm 0.003$ $12.505\pm 1.271$
MOTFM $2.567\pm 0.088$ $2.476\pm 0.161$ $0.040\pm 0.003$ $11.269\pm 1.388$
CTF-C 11 $0.2$ $2.503\pm 0.071$ $2.425\pm 0.239$ $0.037\pm 0.003$ $9.868\pm 1.293$
11 $0.5$ $2.467\pm 0.107$ $2.301\pm 0.163$ $0.037\pm 0.002$ $9.532\pm 1.093$
11 $0.8$ $423\pm 0.164 2.2$$293\pm 0.103$ $0.037\pm 0.001$ $9.874\pm 0.659$
0 $0.2$ $2.396\pm 0.028$ $2.100\pm 0.102$ $0.033\pm 0.003$ $8.577\pm 0.976$
0 $0.5$ $2.447\pm 0.142$ $2.337\pm 0.216$ $0.036\pm 0.005$ $9.696\pm 1.882$
0 $0.8$ $413\pm 0.099 2.2$$293\pm 0.161$ $0.036\pm 0.002$ $9.114\pm 1.092$
$0.5$ $0.2$ $2.460\pm 0.118$ $2.342\pm 0.144$ $0.036\pm 0.003$ $9.500\pm 1.067$
$0.5$ $0.5$ $2.504\pm 0.094$ $2.309\pm 0.139$ $0.036\pm 0.003$ $9.394\pm 1.431$
$0.5$ $0.8$ $2.442\pm 0.173$ $2.353\pm 0.241$ $0.035\pm 0.004$ $9.008\pm 2.094$
CTF-H 0 $2.528\pm 0.143$ $2.534\pm 0.180$ $0.040\pm 0.004$ $11.192\pm 1.304$
11 $\bm{2.316\pm 0.141}$ $\bm{1.969\pm 0.221}$ $\bm{0.030\pm 0.004}$ $\bm{6.359\pm 1.336}$
$0.5$ $2.519\pm 0.167$ $2.412\pm 0.158$ $0.039\pm 0.004$ $10.304\pm 1.808$
Table 6: Extrapolation via Next Step Sampling at holdout time 5 for the Brain Regeneration dataset.
Sampling Method $\bm{\lambda}$ $\bm{\alpha}$ Weighted $\bm{\mathcal{W}_{2}}$ $\bm{\mathcal{W}_{2}}$ MMD Energy
Next Step CFM $7.124\pm 0.443$ $7.133\pm 0.533$ $0.276\pm 0.011$ $76.947\pm 5.661$
MOTFM $7.487\pm 0.698$ $7.449\pm 0.931$ $0.266\pm 0.010$ $81.965\pm 9.812$
CTF-C 11 $0.2$ $7.257\pm 0.597$ $7.077\pm 0.473$ $0.257\pm 0.004$ $79.562\pm 7.787$
11 $0.5$ $68\pm 0.608 6.969$$69\pm 0.628$ $0.265\pm 0.009$ $77.025\pm 6.056$
11 $0.8$ $695\pm 0.443 7.7$$792\pm 0.463$ $0.266\pm 0.007$ $87.179\pm 6.690$
0 $0.2$ $8.170\pm 0.663$ $8.079\pm 0.723$ $0.269\pm 0.008$ $91.572\pm 8.802$
0 $0.5$ $7.244\pm 0.804$ $7.146\pm 0.775$ $0.265\pm 0.003$ $80.424\pm 10.376$
0 $0.8$ $7.382\pm 1.068$ $7.234\pm 0.852$ $0.267\pm 0.009$ $81.635\pm 14.135$
$0.5$ $0.2$ $94\pm 0.239 7.171$$71\pm 0.422$ $0.266\pm 0.001$ $78.924\pm 3.715$
$0.5$ $0.5$ $7.188\pm 0.391$ $\bm{6.931\pm 0.260}$ $0.267\pm 0.005$ $78.992\pm 6.195$
$0.5$ $0.8$ $7.242\pm 0.804$ $7.166\pm 0.980$ $0.267\pm 0.006$ $80.509\pm 10.304$
CTF-H 0 $\bm{6.914\pm 0.471}$ $7.198\pm 0.726$ $0.266\pm 0.009$ $\bm{76.149\pm 8.436}$
11 $7.505\pm 0.667$ $7.338\pm 0.601$ $\bm{0.263\pm 0.006}$ $83.425\pm 8.793$
$0.5$ $7.243\pm 0.479$ $7.157\pm 0.641$ $0.270\pm 0.007$ $79.826\pm 8.067$
H.2 IVP Sampling on Axolotl Brain Regeneration
Table 7: Interpolation via IVP Sampling at time point 3 for the Brain Regeneration dataset.
Sampling Method $\bm{\lambda}$ $\bm{\alpha}$ Weighted $\bm{\mathcal{W}_{2}}$ $\bm{\mathcal{W}_{2}}$ MMD Energy
IVP CFM $4.216\pm 0.463$ $4.266\pm 0.308$ $0.170\pm 0.029$ $32.413\pm 5.122$
MOTFM $198\pm 0.319 4.4$$452\pm 0.243$ $0.173\pm 0.017$ $33.149\pm 3.321$
CTF-C 11 $0.2$ $4.011\pm 0.276$ $4.048\pm 0.321$ $0.147\pm 0.021$ $30.337\pm 4.713$
11 $0.5$ $3.932\pm 0.377$ $4.356\pm 0.398$ $0.156\pm 0.025$ $31.524\pm 4.875$
11 $0.8$ $3.603\pm 0.300$ $3.816\pm 0.310$ $0.127\pm 0.018$ $24.271\pm 3.992$
0 $0.2$ $\bm{3.465\pm 0.232}$ $\bm{3.641\pm 0.320}$ $0.119\pm 0.025$ $23.055\pm 5.939$
0 $0.5$ $3.943\pm 0.413$ $4.241\pm 0.435$ $0.150\pm 0.039$ $29.221\pm 5.713$
0 $0.8$ $3.881\pm 0.368$ $4.094\pm 0.551$ $0.139\pm 0.026$ $27.941\pm 6.676$
$0.5$ $0.2$ $4.152\pm 0.341$ $4.322\pm 0.291$ $0.166\pm 0.014$ $33.299\pm 3.629$
$0.5$ $0.5$ $4.013\pm 0.187$ $4.138\pm 0.297$ $0.153\pm 0.020$ $30.941\pm 3.685$
$0.5$ $0.8$ $4.015\pm 0.351$ $3.974\pm 0.442$ $0.140\pm 0.038$ $27.592\pm 6.669$
CTF-H 0 $3.925\pm 0.267$ $4.375\pm 0.297$ $0.164\pm 0.013$ $32.034\pm 3.270$
11 $3.905\pm 0.395$ $4.188\pm 0.685$ $\bm{0.074\pm 0.014}$ $\bm{18.728\pm 2.689}$
$0.5$ $3.917\pm 0.343$ $4.159\pm 0.455$ $0.147\pm 0.022$ $29.613\pm 4.822$
Table 8: Extrapolation via IVP Sampling at holdout time 5 for the Brain Regeneration dataset.
Sampling Method $\bm{\lambda}$ $\bm{\alpha}$ Weighted $\bm{\mathcal{W}_{2}}$ $\bm{\mathcal{W}_{2}}$ MMD Energy
IVP CFM $6.633\pm 1.312$ $7.116\pm 1.084$ $0.143\pm 0.037$ $60.573\pm 21.756$
MOTFM $6.503\pm 0.720$ $6.352\pm 0.592$ $0.162\pm 0.038$ $56.452\pm 15.932$
CTF-C 11 $0.2$ $6.403\pm 0.959$ $6.558\pm 1.297$ $0.160\pm 0.024$ $61.051\pm 16.594$
11 $0.5$ $6.260\pm 0.616$ $7.681\pm 4.003$ $0.157\pm 0.039$ $52.478\pm 12.010$
11 $0.8$ $6.875\pm 0.643$ $6.920\pm 0.796$ $0.159\pm 0.045$ $62.838\pm 16.897$
0 $0.2$ $6.722\pm 0.905$ $6.782\pm 1.003$ $0.154\pm 0.034$ $53.996\pm 15.617$
0 $0.5$ $6.614\pm 0.710$ $6.854\pm 0.740$ $0.201\pm 0.023$ $70.370\pm 9.099$
0 $0.8$ $6.504\pm 0.925$ $6.744\pm 1.336$ $0.174\pm 0.037$ $56.687\pm 18.118$
$0.5$ $0.2$ $6.514\pm 0.504$ $5.998\pm 0.803$ $0.155\pm 0.032$ $51.329\pm 15.080$
$0.5$ $0.5$ $6.696\pm 0.427$ $6.481\pm 0.387$ $0.195\pm 0.024$ $66.212\pm 3.542$
$0.5$ $0.8$ $6.550\pm 0.975$ $6.563\pm 1.029$ $0.188\pm 0.037$ $63.014\pm 14.173$
CTF-H 0 $6.243\pm 0.760$ $6.220\pm 0.751$ $0.195\pm 0.020$ $61.316\pm 10.288$
11 $\bm{5.277\pm 0.936}$ $6.021\pm 1.192$ $\bm{0.099\pm 0.007}$ $\bm{27.777\pm 8.621}$
$0.5$ $6.254\pm 0.819$ $\bm{5.973\pm 0.757}$ $0.156\pm 0.025$ $54.330\pm 12.089$
H.3 Next Step Sampling for Mouse Embryo Organogenesis
Table 9: Interpolation via Next Step Sampling at holdout time 5 for the Mouse Organogenesis dataset.
Sampling Method $\bm{\lambda}$ $\bm{\alpha}$ Weighted $\bm{\mathcal{W}_{2}}$ $\bm{\mathcal{W}_{2}}$ MMD Energy
Next Step MOTFM $1.892\pm 0.028$ $1.873\pm 0.086$ $0.164\pm 0.002$ $11.615\pm 0.092$
CTF-C 11 $0.2$ $1.881\pm 0.020$ $1.922\pm 0.078$ $0.158\pm 0.003$ $11.529\pm 0.197$
11 $0.5$ $\bm{1.865\pm 0.030}$ $1.852\pm 0.093$ $0.159\pm 0.001$ $11.482\pm 0.108$
11 $0.8$ $1.889\pm 0.024$ $1.888\pm 0.082$ $0.161\pm 0.002$ $11.552\pm 0.166$
0 $0.2$ $1.893\pm 0.035$ $1.912\pm 0.057$ $0.159\pm 0.001$ $11.462\pm 0.154$
0 $0.5$ $1.877\pm 0.039$ $1.933\pm 0.088$ $0.162\pm 0.002$ $11.528\pm 0.110$
0 $0.8$ $1.882\pm 0.022$ $1.869\pm 0.049$ $0.161\pm 0.001$ $\bm{11.399\pm 0.119}$
$0.5$ $0.2$ $1.886\pm 0.022$ $1.927\pm 0.111$ $\bm{0.157\pm 0.002}$ $11.430\pm 0.131$
$0.5$ $0.5$ $1.899\pm 0.027$ $1.899\pm 0.072$ $0.160\pm 0.002$ $11.517\pm 0.097$
$0.5$ $0.8$ $1.888\pm 0.033$ $\bm{1.839\pm 0.134}$ $0.161\pm 0.002$ $11.475\pm 0.159$
CTF-H 0 $1.884\pm 0.027$ $1.862\pm 0.123$ $0.164\pm 0.001$ $11.499\pm 0.123$
11 $1.898\pm 0.029$ $1.866\pm 0.097$ $0.167\pm 0.002$ $11.795\pm 0.170$
$0.5$ $1.871\pm 0.030$ $1.919\pm 0.067$ $0.164\pm 0.002$ $11.639\pm 0.182$
Table 10: Extrapolation via Next Step Sampling at holdout time 8 for Mouse Organogenesis.
Sampling Method $\bm{\lambda}$ $\bm{\alpha}$ Weighted $\bm{\mathcal{W}_{2}}$ $\bm{\mathcal{W}_{2}}$ MMD Energy
Next Step MOTFM $1.626\pm 0.066$ $1.682\pm 0.096$ $0.084\pm 0.007$ $7.418\pm 0.749$
CTF-C 11 $0.2$ $1.683\pm 0.058$ $1.803\pm 0.117$ $0.087\pm 0.006$ $7.830\pm 0.551$
11 $0.5$ $1.685\pm 0.096$ $1.714\pm 0.159$ $0.089\pm 0.006$ $8.056\pm 1.033$
11 $0.8$ $1.703\pm 0.063$ $1.830\pm 0.131$ $0.095\pm 0.005$ $8.928\pm 0.723$
0 $0.2$ $1.715\pm 0.123$ $1.860\pm 0.267$ $0.094\pm 0.009$ $9.021\pm 1.740$
0 $0.5$ $1.725\pm 0.082$ $1.856\pm 0.191$ $0.093\pm 0.006$ $8.806\pm 0.749$
0 $0.8$ $1.774\pm 0.053$ $1.897\pm 0.175$ $0.094\pm 0.007$ $9.466\pm 0.957$
$0.5$ $0.2$ $1.818\pm 0.096$ $2.089\pm 0.222$ $0.084\pm 0.008$ $8.875\pm 0.976$
$0.5$ $0.5$ $1.774\pm 0.104$ $1.899\pm 0.280$ $0.093\pm 0.007$ $9.139\pm 1.437$
$0.5$ $0.8$ $1.768\pm 0.058$ $1.858\pm 0.120$ $0.101\pm 0.006$ $9.303\pm 0.634$
CTF-H 0 $\bm{1.505\pm 0.057}$ $\bm{1.397\pm 0.088}$ $0.087\pm 0.005$ $\bm{5.954\pm 0.492}$
11 $1.890\pm 0.046$ $1.877\pm 0.103$ $0.147\pm 0.006$ $10.752\pm 0.405$
$0.5$ $1.636\pm 0.060$ $1.684\pm 0.099$ $\bm{0.081\pm 0.005}$ $7.088\pm 0.692$
H.4 IVP Sampling for Mouse Embryo Organogenesis
Extrapolating to the last holdout time point of the mouse organogenesis dataset (Chen et al., 2022), particularly under IVP-Sampling, represents the most challenging setting among all our experiments. This difficulty arises because the target time point lies entirely outside the training horizon, requiring integration from the initial samples through to the end. As a result, the velocity field has more opportunity to drift in incorrect directions, often leading to generations that deviate substantially from the true dynamics. In our experiments, this instability was evident: across 10 runs, several produced highly unstable trajectories, reflecting the sensitivity of the system to initial conditions and numerical solvers. This variability is also captured in the performance metrics reported in Table 12.
Table 11: Interpolation via IVP Sampling at holdout time 5 for the Mouse Organogenesis dataset.
Sampling Method $\bm{\lambda}$ $\bm{\alpha}$ Weighted $\bm{\mathcal{W}_{2}}$ $\bm{\mathcal{W}_{2}}$ MMD Energy
IVP MOTFM $3.251\pm 0.676$ $3.418\pm 0.727$ $0.090\pm 0.003$ $9.226\pm 0.648$
CTF-C 11 $0.2$ $3.261\pm 0.880$ $5.264\pm 3.060$ $0.089\pm 0.003$ $10.724\pm 1.288$
11 $0.5$ $3.137\pm 0.407$ $4.093\pm 1.187$ $0.086\pm 0.004$ $11.948\pm 1.393$
11 $0.8$ $3.392\pm 0.757$ $4.716\pm 2.079$ $0.089\pm 0.005$ $9.547\pm 0.752$
0 $0.2$ $2.953\pm 0.425$ $3.816\pm 0.973$ $0.083\pm 0.002$ $9.816\pm 0.715$
0 $0.5$ $2.938\pm 0.476$ $3.904\pm 1.120$ $0.088\pm 0.005$ $9.864\pm 0.764$
0 $0.8$ $3.101\pm 0.539$ $3.855\pm 0.946$ $0.087\pm 0.004$ $9.280\pm 0.551$
$0.5$ $0.2$ $3.771\pm 0.862$ $5.457\pm 1.704$ $\bm{0.079\pm 0.004}$ $9.262\pm 1.134$
$0.5$ $0.5$ $3.090\pm 0.635$ $4.596\pm 2.357$ $0.084\pm 0.005$ $9.786\pm 1.067$
$0.5$ $0.8$ $3.200\pm 0.403$ $3.555\pm 0.637$ $0.084\pm 0.004$ $9.269\pm 0.541$
CTF-H 0 $3.244\pm 0.713$ $3.946\pm 1.671$ $0.089\pm 0.005$ $\bm{8.797\pm 0.612}$
11 $\bm{2.814\pm 0.414}$ $\bm{3.233\pm 0.567}$ $0.093\pm 0.005$ $10.319\pm 0.817$
$0.5$ $5.200\pm 0.799$ $6.306\pm 1.037$ $0.123\pm 0.008$ $45.862\pm 13.765$
Table 12: Extrapolation via IVP Sampling at holdout time 8 for the Mouse Organogenesis dataset.
Sampling Method $\bm{\lambda}$ $\bm{\alpha}$ Weighted $\bm{\mathcal{W}_{2}}$ $\bm{\mathcal{W}_{2}}$ MMD Energy
IVP MOTFM $110835\pm 211671$ $1021005\pm 2063905$ $0.086\pm 0.002$ $14178\pm 29475$
CTF-C 11 $0.2$ $785586\pm 1318212$ $7598321\pm 13497483$ $0.088\pm 0.002$ $98199\pm 150412$
11 $0.5$ $2691\pm 3931$ $28480\pm 36483$ $0.087\pm 0.002$ $1632\pm 2090$
11 $0.8$ $2473\pm 3349$ $19537\pm 26306$ $0.087\pm 0.003$ $517\pm 616$
0 $0.2$ $1493\pm 2497$ $14563\pm 24858$ $0.087\pm 0.001$ $800\pm 1158$
0 $0.5$ $218018\pm 471298$ $1820788\pm 3994886$ $0.086\pm 0.001$ $2170\pm 4697$
0 $0.8$ $2736\pm 34766 11$$18089\pm 310135$ $0.084\pm 0.002$ $27013\pm 60065$
$0.5$ $0.2$ $8114720\pm 16270274$ $69458305\pm 140579849$ $0.088\pm 0.002$ $901074\pm 1775139$
$0.5$ $0.5$ $2414338\pm 6009993$ $23103811\pm 56863018$ $0.086\pm 0.001$ $261335\pm 663279$
$0.5$ $0.8$ $158\pm 3023 11$$1138\pm 30025$ $0.084\pm 0.002$ $445\pm 1085$
CTF-H 0 $353428\pm 952168$ $3011396\pm 8057131$ $0.095\pm 0.004$ $22990\pm 58936$
11 $\bm{15\pm 10}$ $\bm{53\pm 53}$ $0.098\pm 0.006$ $\bm{48\pm 32}$
$0.5$ $107889\pm 275882$ $994606\pm 2772756$ $0.087\pm 0.002$ $8875\pm 24264$
H.5 Liver Regeneration
Table 13: Wasserstein distances for different model configurations
Variant $\bm{\lambda}$ $\bm{\alpha}$ $\bm{\mathcal{W}_{2}}$
EOT $34.30348\pm 1.44797$
CTF-C 11 $0.2$ $34.44455\pm 1.19306$
CTF-C 11 $0.5$ $33.95671\pm 1.64415$
CTF-C 11 $0.8$ $34.62812\pm 0.98181$
CTF-C 0 $0.2$ $34.24147\pm 1.16930$
CTF-C 0 $0.5$ $32.74147\pm 1.86351$
CTF-C 0 $0.8$ $33.71729\pm 1.23057$
CTF-C $0.5$ $0.2$ $33.56646\pm 1.04376$
CTF-C $0.5$ $0.5$ $33.84199\pm 1.71408$
CTF-C $0.5$ $0.8$ $33.04534\pm 1.64399$
CTF-H 0 $\bm{32.68215\pm 1.47185}$
CTF-H 11 $33.48050\pm 1.00149$
CTF-H $0.5$ $33.41444\pm 0.99501$
H.6 IVP Cell Type Progression over time
Figure 9: Temporal cell type predictions from ContextFlow for the major cell types in the Organogenesis dataset (Chen et al., 2022). Early progenitor populations (neural crest and mesenchyme) progressively diminish as development advances, while terminal fates (muscle, cartilage primordium, and liver) emerge at later stages. Major lineages such as brain, heart, and connective tissue remain continuous throughout. Overall, ContextFlow captures biologically coherent and temporally consistent developmental dynamics.

### v2

@@
 By embedding these contextual constraints, ContextFlow generates trajectories that are not only statistically consistent but also biologically meaningful, making it a generalizable framework for modeling spatiotemporal dynamics from longitudinal, spatially resolved omics data.
-Evaluated on three datasets, ContextFlow consistently outperforms state-of-the-art flow matching methods across multiple quantitative and qualitative metrics of inference accuracy and biological coherence.
-Our code is available at: https://github.com/santanurathod/ContextFlow
+Comprehensive experiments show that ContextFlow consistently outperforms state-of-the-art flow matching methods across multiple metrics of inference accuracy and biological coherence.
+Our code is available at: https://github.com/santanurathod/ContextFlow.
+Machine Learning, ICML
 1 Introduction
-Flow matching (Lipman et al., 2023) is an emerging paradigm that provides an efficient approach for learning the complex latent dynamics, or normalizing flows (Papamakarios et al., 2021), of a system of variables, while enabling parametric flexibility to model data distributions.
-Inferring the underlying dynamics from sparse and noisy observations is a central challenge in many domains (Gontis et al., 2010; Brunton et al., 2016; Pandarinath et al., 2018; Li et al., 2025), where continuous trajectories are rarely captured; instead, cross-sectional snapshots, collected at discrete time points, are typically available.
+Flow matching (Lipman et al., 2023) is an emerging generative paradigm that provides an efficient approach for learning the complex latent dynamics, or normalizing flows (Papamakarios et al., 2021), of a system of variables, while enabling parametric flexibility to model data distributions.
+Inferring the underlying dynamics from sparse and noisy observations is a central challenge in many application domains (Gontis et al., 2010; Brunton et al., 2016; Pandarinath et al., 2018; Li et al., 2025), where continuous trajectories are rarely captured; instead, cross-sectional snapshots, collected at discrete time points, are typically available.
 In single-cell RNA sequencing (scRNA-seq), this challenge becomes especially critical as the destructive nature of profiling technologies yields only unpaired population-level snapshots over time.
-Uncovering temporal dynamics from such snapshot data is essential for understanding developmental processes, disease progression, treatment and perturbation responses (Wagner & Klein, 2020).
+Uncovering temporal dynamics from such snapshot data is essential for understanding developmental processes, disease progression, treatment, and perturbation responses (Wagner and Klein, 2020).
 Traditional approaches often rely on heuristics or computationally intensive likelihood-based generative models, which struggle with scalability and flexibility in high-dimensional single-cell data.
-Flow matching overcomes these challenges by directly learning continuous latent dynamics that are constrained to match observed population-level distributions at sampled time points.
+In contrast, flow matching overcomes these challenges by directly learning continuous latent dynamics that are constrained to match observed population-level distributions at sampled time points.
+Table 1: Comparisons of state-of-the-art spatiotemporal OT and flow matching methods.
+In particular, ContextFlow excels at efficiently inferring biologically plausible trajectories by incorporating spatial priors into minibatch-OT-based flow-matching frameworks.
+Method Generative Prior Knowledge Unified Velocity Field OT Runtime
+DeST-OT (Halmos et al., 2025) $O(N^{3})$
+TOAST (Ceccarelli et al., 2025) $O(N^{3})$
+PASTE (Zeira et al., 2022) $O(N^{3})$
+CFM (Lipman et al., 2022)
+MOTFM (Tong et al., 2024) $O(N^{2})$
+ContextFlow (ours) $O(N^{2})$
 The state and function of cells within a tissue are affected by interactions with neighboring cells, extracellular matrix components, and local signaling gradients (Rao et al., 2021).
 Recent advances in spatial omics technologies, particularly spatial transcriptomics (ST), allow gene expression profiling without tissue dissociation, thereby preserving spatial context and providing a complementary view of cellular organization.
-The dynamics of complex cellular processes is affected by the tissue microenvironment, where cells engage in reciprocal communication with their neighbors (Dimitrov et al., 2022; Tanevski et al., 2025).
+The dynamics of complex cellular processes are affected by the tissue microenvironment, where cells engage in reciprocal communication with their neighbors (Dimitrov et al., 2022; Tanevski et al., 2025).
 A growing body of work highlights the critical role of spatial cell–cell communication patterns in shaping cellular phenotypes (Armingol et al., 2021).
@@
 These insights, made possible by the spatiotemporal resolution of transcriptomics data, pave the way for understanding the mechanisms by which cellular interactions drive tissue organization and function in organogenesis (Chen et al., 2022), regeneration (Ben-Moshe et al., 2021; Wei et al., 2022), disease progression (Kukanja et al., 2024), and treatment response (Liu et al., 2024).
-Optimal transport (OT) has become a foundational framework to align spatially resolved samples and infer putative developmental or temporal couplings  (Zeira et al., 2022; Liu et al., 2023).
-As a result, state-of-the-art flow matching frameworks such as minibatch-OT flow matching (MOTFM) (Tong et al., 2024) use OT-derived couplings to define conditional paths to train velocity fields, thus overcoming the lack of generative capabilities in optimal transport.
-The OT formulation adopted in MOTFM, however, does not account for the contextual richness present in spatial transcriptomics and can result in trajectories that are statistically optimal yet biologically implausible (see Figure 3(a) in Appendix G.1 for an illustration).
-While recent studies have extended widely-used OT objectives (Halmos et al., 2025; Ceccarelli et al., 2025) for spatial transcriptomics, they primarily focus on pairwise alignment of populations across conditions or modalities and do not explicitly incorporate the cell–cell communication patterns that drive cellular state transitions.
-To address the above limitations, we introduce a novel flow matching-based framework, ContextFlow, that incorporates spatial priors for modeling temporal tissue dynamics (Figure 1).
+Optimal transport (OT) has become a foundational framework for aligning spatially resolved samples and inferring putative developmental or temporal couplings  (Zeira et al., 2022; Liu et al., 2023).
+As a result, state-of-the-art flow matching frameworks such as minibatch-OT flow matching (MOTFM) (Tong et al., 2024) use OT-derived couplings to define conditional paths for training velocity fields, thereby overcoming the lack of generative capabilities in optimal transport.
+The OT formulation adopted in MOTFM, however, does not account for the contextual richness of spatial transcriptomics and can yield trajectories that are statistically optimal yet biologically implausible (see Figure 4(a) in Appendix G.1 for an illustration).
+While recent studies (Halmos et al., 2025; Ceccarelli et al., 2025) have extended OT objectives to spatial transcriptomics, they primarily focus on pairwise alignment of populations across conditions or modalities and do not explicitly incorporate the cell–cell communication patterns that drive cellular state transitions.
+Figure 1: ContextFlow integrates local tissue organization and ligand-receptor communications to learn biologically meaningful trajectories from spatial omics data.
+Prior knowledge acts as a soft filter, discouraging implausible transitions while preserving flexibility in trajectory inference.
+To address the above limitations, we introduce a novel flow matching-based framework, ContextFlow, that incorporates spatial priors to model temporal tissue dynamics (Figure 1).
 By encoding local tissue organization and ligand-receptor-derived spatial communication patterns into prior-regularized optimal transport formulations, ContextFlow fully exploits the contextual richness of spatial omics data and embeds both structural and functional aspects of tissue organization into its objective, thereby generating more biologically informed trajectories.
+Table 1 highlights the key advantages of ContextFlow, compared with the aforementioned existing methods, where detailed discussions on related works are postponed to Appendix A.
 In summary, our contributions are as follows:
-  • We leverage local tissue organization and local ligand–receptor communication patterns to extract biologically meaningful features from spatial omics data, and encode them into a biologically-informed transition plausibility matrix to constrain temporal dynamics (Section 3.2).
-  • We design two novel integration schemes—cost-based and entropy-based—that incorporate the prior knowledge into an OT-coupled flow matching framework, both amenable to efficient Sinkhorn optimization and scalable on modern hardware (Section 3.3).
-  • Comprehensive experiments on regeneration and developmental datasets demonstrate that ContextFlow consistently outperforms baseline methods under both interpolation and extrapolation settings across metrics that capture biological plausibility and statistical fidelity (Section 4).
-Figure 1: ContextFlow integrates local tissue organization and ligand-receptor communications to learn biologically meaningful trajectories from spatial omics data.
-Prior knowledge acts as a soft filter that discourages implausible transitions while preserving flexibility in trajectory inference.
+  • We leverage the structure of local tissue organization and ligand–receptor communication to extract biologically meaningful features from spatial omics data and encode them into an informed transition plausibility matrix to constrain temporal dynamics (Section 3.2).
+  • We design two novel integration schemes—cost-based and entropy-based—that incorporate prior knowledge into an OT-coupled flow matching framework, both amenable to efficient Sinkhorn optimization and scalable on modern hardware (Section 3.3).
+  • Comprehensive experiments on three regeneration and developmental datasets demonstrate that ContextFlow consistently outperforms baseline flow matching methods under both interpolation and extrapolation settings across multiple evaluation metrics that capture biological plausibility and statistical fidelity (Section 4).
 2 Preliminaries
@@
 $\min_{\theta}\>\mathbb{E}_{t\sim\mathcal{U}(0,1),\bm{x}\sim p_{t}(\bm{x})}\big\|u_{\theta}(t,\bm{x})-u_{t}(\bm{x})\big\|^{2}.$ (1)
-Here, $\mathcal{U}(0,1)$ is the uniform distribution over $[0,1]$, and $p_{t}:[0,1]\times\mathbb{R}^{d}\rightarrow\mathbb{R}_{+}$ denotes a time-varying probability path induced by $u_{t}$ such that (i) $p_{t}$ is a probability density function for any $t\in[0,1]$, (ii) $p_{t}$ satisfies the two boundary conditions: $p_{t=0}=q_{0}$ and $p_{t=1}=q_{1}$, and (iii) the connection between $p_{t}$ and $u_{t}$ can be characterized by the transport equation (Villani et al., 2008): $\frac{\partial p_{t}(\bm{x})}{\partial t}=-\nabla\cdot(u_{t}(\bm{x})p_{t}(\bm{x}))$, where $\nabla$ is the divergence operator.
+Here, $\mathcal{U}(0,1)$ is the uniform distribution over $[0,1]$, and $p_{t}:[0,1]\times\mathbb{R}^{d}\rightarrow\mathbb{R}_{+}$ denotes a time-varying probability path induced by $u_{t}$ such that (i) $p_{t}$ is a probability density function for any $t\in[0,1]$, (ii) $p_{t}$ satisfies the two boundary conditions: $p_{t=0}=q_{0}$ and $p_{t=1}=q_{1}$, and (iii) the connection between $p_{t}$ and $u_{t}$ can be characterized by the transport equation (Villani and others, 2008): $\frac{\partial p_{t}(\bm{x})}{\partial t}=-\nabla\cdot(u_{t}(\bm{x})p_{t}(\bm{x}))$, where $\nabla$ is the divergence operator.
 From a dynamical system’s view, $u_{t}$ defines an ODE system $d\bm{x}=u_{t}(\bm{x})dt$.
@@
 2.2 Conditional Flow Matching
-The central idea of conditional flow matching is to express the target probability path $p_{t}$ via a mixture of more manageable conditional probability paths (Lipman et al., 2023).
-By marginalizing over some conditioning variable $z$, both $p_{t}$ and $u_{t}$ can be constructed using their conditional counterparts:
-$p_{t}(\bm{x})=\int p_{t}(\bm{x}|\bm{z})q(\bm{z})d\bm{z},\quad u_{t}(\bm{x})=\int u_{t}(\bm{x}|\bm{z})\frac{p_{t}(\bm{x}|\bm{z})q(\bm{z})}{p_{t}(\bm{x})}d\bm{z},$ (2)
-where $q(\bm{z})$ denotes the distribution of the conditioning variable $\bm{z}$, and $p_{t}(\bm{x}|\bm{z})$ is selected such that the boundary conditions are satisfied: $\int p_{t=0}(\bm{x}|\bm{z})q(\bm{z})=q_{0}$ and $\int p_{t=1}(\bm{x}|\bm{z})q(\bm{z})=q_{1}$.
+The central idea of conditional flow matching is to express the target probability path via a mixture of more manageable conditional probability paths (Lipman et al., 2023).
+By marginalizing over a conditioning variable $z$, both $p_{t}$ and $u_{t}$ can be constructed using their conditional counterparts:
+$p_{t}(\bm{x})$ $=\int p_{t}(\bm{x}|\bm{z})q(\bm{z})d\bm{z},$ (2)
+$u_{t}(\bm{x})$ $=\int u_{t}(\bm{x}|\bm{z})\frac{p_{t}(\bm{x}|\bm{z})q(\bm{z})}{p_{t}(\bm{x})}d\bm{z},$
+where $q(\bm{z})$ stands for the distribution of the conditioning variable $\bm{z}$, and the conditional probability path $p_{t}(\bm{x}|\bm{z})$ is constructed such that the boundary conditions are satisfied: $\int p_{t=0}(\bm{x}|\bm{z})q(\bm{z})=q_{0}$ and $\int p_{t=1}(\bm{x}|\bm{z})q(\bm{z})=q_{1}$.
 Theorem 1 of Lipman et al.
 (2023) proves that $p_{t}$ and $u_{t}$ defined by Equation 2 satisfy the transport equation, suggesting that $p_{t}$ is a valid probability path generated by $u_{t}$.
-To avoid the intractable integrals, Lipman et al.
-(2023) proposed the following optimization of conditional flow matching (CFM), and proved its equivalence to the original flow matching objective in terms of gradient computation:
+To avoid the intractable integrals, (Lipman et al., 2023) proposed the following optimization of conditional flow matching (CFM), and proved its equivalence to the original FM objective in terms of gradient computation:
 $\min_{\theta}\>\mathbb{E}_{t\sim\mathcal{U}(0,1),z\sim q(\bm{z}),\bm{x}\sim p_{t}(\bm{x}|\bm{z})}\big\|u_{\theta}(t,\bm{x})-u_{t}(\bm{x}|\bm{z})\big\|^{2}.$ (3)
-By choosing an appropriate conditional velocity vector field $u_{t}(\bm{x}|\bm{z})$, we can train the neural network using Equation 3 without requiring a closed-form solution of the conditional probability path $p_{t}(\bm{x}|\bm{z})$, thus avoiding the intractable integration operation.
+By choosing an appropriate conditional velocity vector field $u_{t}(\bm{x}|\bm{z})$, we can train the neural network using Equation 3 without requiring a closed-form solution of the conditional probability path $p_{t}(\bm{x}|\bm{z})$, thereby avoiding the intractable integration operation.
 Therefore, the remaining task is to define the conditional probability path and velocity vector field properly such that we can sample from $p_{t}(\bm{x}|\bm{z})$ and compute $u_{t}(\bm{x}|\bm{z})$ efficiently for solving the optimization problem in Equation 3.
-Gaussian Conditional Probability Paths. A specific choice proposed in Lipman et al. (2023) is Gaussian conditional probability paths and their corresponding conditional velocity vector fields:
-$p_{t}(\bm{x}|\bm{z})=\mathcal{N}(\bm{x}\>|\>\mu_{t}(\bm{z}),\sigma_{t}(\bm{z})^{2}\mathbf{I}),\quad u_{t}(\bm{x}|\bm{z})=\frac{\sigma_{t}^{\prime}(\bm{z})}{\sigma_{t}(\bm{z})}(\bm{z}-\mu_{t}(\bm{z}))+\mu_{t}^{\prime}(\bm{z}),$ (4)
-where $\mu_{t}:[0,1]\times\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$ denotes the time-varying mean of the Gaussian distribution, $\mu_{t}^{\prime}$ is its derivative with respect to time, $\sigma_{t}:[0,1]\times\mathbb{R}^{d}\rightarrow\mathbb{R}_{+}$ stands for the time-varying scalar standard deviation, and $\sigma_{t}^{\prime}$ stand for the corresponding derivative.
+A specific choice proposed in Lipman et al. (2023) is Gaussian conditional probability paths and their corresponding conditional velocity vector fields, defined as follows:
+$p_{t}(\bm{x}|\bm{z})$ $=\mathcal{N}(\bm{x}\>|\>\mu_{t}(\bm{z}),\sigma_{t}(\bm{z})^{2}\mathbf{I}),$ (4)
+$u_{t}(\bm{x}|\bm{z})$ $=\frac{\sigma_{t}^{\prime}(\bm{z})}{\sigma_{t}(\bm{z})}(\bm{x}-\mu_{t}(\bm{z}))+\mu_{t}^{\prime}(\bm{z}),$ (5)
+where $\mu_{t}:[0,1]\times\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$ stands for the time-varying mean of the Gaussian distribution, $\mu_{t}^{\prime}$ is its derivative with respect to time, $\sigma_{t}:[0,1]\times\mathbb{R}^{d}\rightarrow\mathbb{R}_{+}$ is the time-varying scalar standard deviation, and $\sigma_{t}^{\prime}$ denotes its derivative.
 In particular, Lipman et al.
-(2023) set $q(\bm{z})=q(\bm{x}_{1})$, $\mu_{t}(\bm{z})=t\bm{x}_{1}$, and $\sigma_{t}(\bm{z})=1-(1-\sigma)t$.
-Then, we can see that $u_{t}(\bm{x}|\bm{z})$ transports the standard Gaussian distribution $p_{t=0}(\bm{x}|\bm{z})=\mathcal{N}(\bm{x};0,I)$ to a Gaussian distribution with mean $x_{1}$ and standard deviation $\sigma$, namely $p_{t=1}(\bm{x}|\bm{z})=\mathcal{N}(\bm{x};\bm{x}_{1},\sigma^{2})$ for any target point $\bm{x}_{1}$.
-By letting $\sigma\rightarrow 0$, the marginal boundary conditions can easily be verified.
+(2023) selected $q(\bm{z})=q(\bm{x}_{1})$, $\mu_{t}(\bm{z})=t\bm{x}_{1}$, and $\sigma_{t}(\bm{z})=1-(1-\sigma)t$.
+Then, we can see that $u_{t}(\bm{x}|\bm{z})$ transports the standard Gaussian distribution $p_{t=0}(\bm{x}|\bm{z})=\mathcal{N}(\bm{x};\bm{0},\mathbf{I})$ to a Gaussian distribution with mean $\bm{x}_{1}$ and covariance $\sigma^{2}\mathbf{I}$, namely $p_{t=1}(\bm{x}|\bm{z})=\mathcal{N}(\bm{x};\bm{x}_{1},\sigma^{2}\mathbf{I})$ for any target point $\bm{x}_{1}$.
+By letting $\sigma\rightarrow 0$, the marginal boundary conditions can easily be validated.
 Tong et al.
-(2024) further generalized the application scope to arbitrary source distributions, by setting
-$q(\bm{z})=q(\bm{x}_{0})q(\bm{x}_{1}),\quad\mu_{t}(\bm{z})=(1-t)\bm{x}_{0}+t\bm{x}_{1},\quad\sigma_{t}(\bm{z})=\sigma.$ (5)
-This choice satisfies the boundary conditions $p_{t=0}(\bm{x})=q_{0}$ and $p_{t=1}(\bm{x})=q_{1}$ when $\sigma\rightarrow 0$.
-Based on Equation 4, the conditional velocity vector field has a simple analytical form $u_{t}(\bm{x}|\bm{z})=\bm{x}_{1}-\bm{x}_{0}$.
+(2024) further generalized the application scope to arbitrary source distributions by setting:
+$q(\bm{z})=q(\bm{x}_{0})q(\bm{x}_{1}),\mu_{t}(\bm{z})=(1-t)\bm{x}_{0}+t\bm{x}_{1},\sigma_{t}(\bm{z})=\sigma.$ (6)
+Note that the choice in Equation 6 satisfies the boundary conditions $p_{t=0}(\bm{x})=q_{0}$ and $p_{t=1}(\bm{x})=q_{1}$ when $\sigma\rightarrow 0$.
+According to Equation 4, the conditional velocity vector field has a simple analytical form $u_{t}(\bm{x}|\bm{z})=\bm{x}_{1}-\bm{x}_{0}$.
 2.3 Flow Matching with Optimal Transport Couplings
 The conditionals construction specified by Equation 6 corresponds to the simplest choice of independent coupling, where $\bm{z}=(\bm{x}_{0},\bm{x}_{1})$ with source $\bm{x}_{0}$ and target $\bm{x}_{1}$ are independently sampled from $q(\bm{z})=q(\bm{x}_{0})q(\bm{x}_{1})$.
-The use of couplings for constructing the sampling paths in the CFM framework naturally connects to the optimal transport theory (Villani et al., 2008).
+The use of couplings for constructing the sampling paths in the CFM framework naturally connects to the optimal transport theory (Villani and others, 2008).
 Choosing OT-based couplings has several advantages over independent coupling, including smaller training variance and more efficient sampling (Pooladian et al., 2023; Tong et al., 2024).
-Since the classical Kantorovich’s formulation (refer Appendix E) has computational complexity that is cubic with respect to the sample size, a popular alternative is to add an extra regularization term, resulting in entropic optimal transport (EOT), to approximately solve the optimal transport problem while reducing the computational costs from cubic to quadratic:
-$\pi^{*}_{\mathrm{eot}}(\epsilon):={\arg\!\min}_{\pi\in\Pi(q_{0},q_{1})}\int_{\mathbb{R}^{d}\times\mathbb{R}^{d}}\|\bm{x}_{0}-\bm{x}_{1}\|_{2}^{2}\>d\pi(\bm{x}_{0},\bm{x}_{1})+\epsilon H(\pi\>|\>q_{0}\otimes q_{1}),$ (6)
+Since the classical Kantorovich’s formulation has computational complexity that is cubic in terms of the sample size (see Appendix B.1 for its formal definition), a popular alternative is to add an extra regularization term, resulting in entropic optimal transport (EOT), to approximately solve the optimal transport problem while reducing the computational costs from cubic to quadratic:
+${\min}_{\pi\in\Pi(q_{0},q_{1})}\int_{\mathbb{R}^{d}\times\mathbb{R}^{d}}$ $\|\bm{x}_{0}-\bm{x}_{1}\|_{2}^{2}\>d\pi(\bm{x}_{0},\bm{x}_{1})$ (7)
+$\qquad\qquad+\epsilon H(\pi\>|\>q_{0}\otimes q_{1}),$
 where $\epsilon>0$ is the regularization parameter, and $H(\pi\>|\>q_{0}\otimes q_{1})$ denotes the relative entropy (or Kullback-Leibler divergence) with respect to $\pi$ and the product measure $q_{0}\otimes q_{1}$.
-The optimization problem in Equation 6 can be viewed as a special case of the static Schrödinger bridge problem (Bernton et al., 2022), which can be efficiently solved in a mini-batch fashion via the Sinkhorn algorithm (Cuturi, 2013).
-Theoretically, one can prove that $\pi^{*}_{\mathrm{eot}}(\epsilon)$ recovers the Kantorovich’s OT coupling $\pi^{*}_{\mathrm{ot}}$ when $\epsilon\rightarrow 0$ (see Equation 16 in Appendix E for its formal definition) and $\pi^{*}_{\mathrm{eot}}(\epsilon)$ corresponds to the independent coupling $q_{0}\otimes q_{1}$ when $\epsilon\rightarrow\infty$.
+For simplicity, we use $\pi^{*}_{\mathrm{eot}}(\epsilon)$ to denote the optimal solution to the minimization problem in Equation 7.
+Note that this optimization problem can also be viewed as a special case of the static Schrödinger bridge problem (Bernton et al., 2022), which can be efficiently solved in a mini-batch fashion via the Sinkhorn algorithm (Cuturi, 2013).
+Theoretically, one can prove that $\pi^{*}_{\mathrm{eot}}(\epsilon)$ recovers the Kantorovich’s OT coupling $\pi^{*}_{\mathrm{ot}}$ when $\epsilon\rightarrow 0$ (see Equation 14 in Appendix B.1 for its formal definition) and $\pi^{*}_{\mathrm{eot}}(\epsilon)$ corresponds to the independent coupling $q_{0}\otimes q_{1}$ when $\epsilon\rightarrow\infty$.
 3 Regularizing the Flow with Spatial Priors
@@
 We focus on the task of inferring spatiotemporal trajectories, i.e., inferring the dynamic evolution of the cell states across time from spatially resolved gene expression data.
-Let $0=t_{1}<t_{2}<\ldots<t_{m+1}=1$ be a sequence of normalized time points.
-For simplicity, we use $[m]$ to denote the set $\{1,2,\ldots,m\}$.
-For any $i\in[m+1]$, let $q_{i}$ be the data distribution at time $t_{i}$ over $\mathbb{R}^{d}$.
-Given $\{\mathbf{X}_{t_{i}}\}_{i\in[m+1]}$, where $\mathbf{X}_{t_{i}}=[\bm{x}_{i}(k)]_{k\in[n_{i}]}$ is the expression matrix at time $t_{i}$ consisting of $n_{i}$ snapshot data sampled from $q_{i}$, the objective is to learn a neural velocity vector field $u_{\theta}:[0,1]\times\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$ to faithfully characterize the temporal evolution of spatially resolved tissues over time, such that the induced probability path $p_{t}$ can describe the state of each cell at time $t\in[0,1]$.
+Let $0=t_{1}<t_{2}<\ldots<t_{m+1}=1$ be a sequence of normalized time points.For any $i\in\{1,2,\ldots,m+1\}$, let $q_{i}$ be the data distribution over $\mathbb{R}^{d}$ at time point $t_{i}$.
+Given $\{\mathbf{X}_{t_{i}}\}_{i\in\{1,2,\ldots,m+1\}}$, where $\mathbf{X}_{t_{i}}=\{\bm{x}_{i}(k)\}_{k=1}^{n_{i}}$ is the gene expressions at time $t_{i}$ consisting of $n_{i}$ snapshot data sampled from $q_{i}$, the objective is to learn a neural velocity vector field $u_{\theta}:[0,1]\times\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$ to faithfully characterize the temporal evolution of spatially resolved tissues over time, such that the induced probability path $p_{t}$ can describe the state of each cell at time $t\in[0,1]$.
 This task can be viewed as a continuous temporal generalization of the pairwise generative modeling task described in Section 2.1.
 A promising candidate solution is conditional flow matching with entropic OT couplings (EOT-CFM), by targeting linear conditional velocity vector fields for each pair of consecutive time points.
-Specifically, for any $t\in[0,1]$ satisfying $t\in[t_{i},t_{i+1}]$, define
-$p(\bm{x}|\bm{z})=\mathcal{N}\left(\frac{t_{i+1}-t}{t_{i+1}-t_{i}}\bm{x}_{i}+\frac{t-t_{i}}{t_{i+1}-t_{i}}\bm{x}_{i+1},\sigma^{2}\mathbf{I}\right),\quad u_{t}(\bm{x}|\bm{z})=\frac{\bm{x}_{i+1}-\bm{x}_{i}}{t_{i+1}-t_{i}},$ (7)
-where the conditioning variable is selected as $\bm{z}=(\bm{x}_{i},\bm{x}_{i+1})$, and $p(\bm{z})$ is the joint probability measure with marginals $q_{i}$ and $q_{i+1}$ corresponding to the EOT coupling $\pi^{*}_{\mathrm{eot}}(\epsilon)$ defined in Equation 6.
+Specifically, for any $t\in[0,1]$ satisfying $t\in[t_{i},t_{i+1}]$, the conditional velocity field $u_{t}(\bm{x}|\bm{z})$ and its associated probability path $p(\bm{x}|\bm{z})$ are defined as:
+$p(\bm{x}|\bm{z})=\mathcal{N}\left(\frac{t_{i+1}-t}{t_{i+1}-t_{i}}\bm{x}_{i}+\frac{t-t_{i}}{t_{i+1}-t_{i}}\bm{x}_{i+1},\sigma^{2}\mathbf{I}\right),$
+$u_{t}(\bm{x}|\bm{z})=\frac{\bm{x}_{i+1}-\bm{x}_{i}}{t_{i+1}-t_{i}},$ (8)
+where the conditioning variable $\bm{z}=(\bm{x}_{i},\bm{x}_{i+1})$, and $p(\bm{z})$ is the joint probability measure with marginals $q_{i}$ and $q_{i+1}$ corresponding to the EOT coupling $\pi^{*}_{\mathrm{eot}}(\epsilon)$ defined in Equation 7.
 It can be easily verified that the above construction satisfies the boundary condition at each time point $t_{i}$.
 To train $u_{\theta}$, we can randomly sample a mini-batch of data at each time, run the Sinkhorn algorithm (Cuturi, 2013) to obtain the entropic OT couplings for each consecutive pair, and iteratively update the model weights $\theta$ using stochastic gradient descent with CFM regression loss (Equation 3).
-Despite their enhanced ability to model system dynamics, state-of-the-art OT-CFM frameworks lack provisions to fully exploit the contextual richness and integrate the biological prior knowledge that can be inferred from other associated data modalities.
+Despite their enhanced ability to model system dynamics, state-of-the-art OT-CFM frameworks lack provisions to fully exploit the contextual richness and to integrate biological prior knowledge inferred from other associated data modalities.
 Existing approaches can generate statistically optimal trajectories by targeting probability paths induced by (entropic) OT couplings along the temporal dimension.
 However, they may overlook important functional or structural prior information, leading to biologically implausible trajectories (see Figure 4(a) in Appendix G.1 for an illustration).
-3.2 Introducing Spatial Priors & Transitional Plausibility
-To faithfully model the spatial context and cellular organization of spatial omics data, we introduce two types of spatial priors and explain how they relate to the transitional plausibility between locations and cell states at different time points.
+3.2 Spatial Priors and Transitional Plausibility
+To faithfully model the context and cellular organization of spatial omics data, we introduce two types of spatial priors and explain how they relate to the plausibility of transitions between locations and cell states across time points.
 Spatial Smoothness.
 Tissues are well-organized systems.
-Within a microenvironment, neighboring cells respond to the same set of external mechanical stimuli and intercellular communication, which affects their states in a similar manner and results in local smoothness of cell-type-specific expression.
+Within a microenvironment, neighboring cells respond to the same set of external mechanical stimuli and intercellular communication, which affect their states similarly and result in local smoothness in cell-type-specific expression.
 Due to tissue heterogeneity, we cannot assume a common reference coordinate frame across tissue samples or even slices at $t_{i}$ and $t_{j}$ at a larger scale.
@@
 Spatial smoothness suggests that $\mathrm{TP}(c_{i},c_{j})$ is inversely related to the difference between the average expression profiles of their local neighborhoods:
-$\mathrm{SS}(c_{i},c_{j})=\bigg\|\>\frac{1}{|\mathcal{N}_{r}(c_{i})|}\sum_{c\in\mathcal{N}_{r}(c_{i})}\bm{x}(c)-\frac{1}{|\mathcal{N}_{r}(c_{j})|}\sum_{c\in\mathcal{N}_{r}(c_{j})}\bm{x}(c)\>\bigg\|_{2}^{2},$ (8)
-where $\mathcal{N}_{r}(c_{i})=\{c:\|\bm{s}(c)-\bm{s}(c_{i})\|_{2}\leq r\}$ denotes the set of neighboring cells of $c_{i}$ in the same tissue slice, $|\mathcal{N}_{r}(c_{i})|$ is the cardinality of $\mathcal{N}_{r}(c_{i})$, and $\bm{x}(c)$ is the gene expression profile of cell $c$.
+$\mathrm{SS}(c_{i},c_{j})$ $=\bigg\|\frac{1}{|\mathcal{N}_{r}(c_{i})|}\sum_{c\in\mathcal{N}_{r}(c_{i})}\bm{x}(c)$ (9)
+$\qquad\qquad-\frac{1}{|\mathcal{N}_{r}(c_{j})|}\sum_{c\in\mathcal{N}_{r}(c_{j})}\bm{x}(c)\bigg\|_{2}^{2},$
+where $\mathcal{N}_{r}(c_{i})=\{c:\|\bm{s}(c)-\bm{s}(c_{i})\|_{2}\leq r\}$ denotes the set of neighboring cells of $c_{i}$ in the same tissue slice, $|\mathcal{N}_{r}(c_{i})|$ stands for the cardinality of $\mathcal{N}_{r}(c_{i})$, and $\bm{x}(c)$ is the gene expression profile with respect to cell $c$.
 Cell-Cell Communication Patterns.
@@
 There are numerous databases of prior knowledge of ligand-receptor binding and computational methods that use these databases to systematically link gene expression with the activity of ligand-receptor-mediated communication.
-Specifically, we can represent each cell $c_{i}$ by a vector $f_{\mathrm{LR}}\in\mathbb{R}^{p}$, where each entry corresponds to one of $p$ possible ligand–receptor pairs and encodes the extent of $c_{i}$’s participation in communication through that pair.
-The $\mathrm{TP}(c_{i},c_{j})$ between cells in different tissue slices is higher when they exhibit similar ligand-receptor communication patterns $f_{\mathrm{LR}}$ (see Figure 7 for an illustration).
-We define $\mathrm{LR}(c_{i},c_{j})$, the dissimilarity between the ligand–receptor communication patterns in the microenvironments of cells $c_{i}$ and $c_{j}$, as:
-$\mathrm{LR}(c_{i},c_{j})=\left\|\>f_{\mathrm{LR}}(\mathcal{N}_{r}(c_{i}))-f_{\mathrm{LR}}(\mathcal{N}_{r}(c_{j}))\right\|_{2}^{2},$ (9)
-3.3 ContextFlow: CFM with Context-Aware OT Couplings
+Specifically, we represent each cell $c_{i}$ by a feature vector $f_{\mathrm{LR}}\in\mathbb{R}^{p}$, where each entry corresponds to one of $p$ possible ligand–receptor pairs and encodes the extent of $c_{i}$’s participation in communication through that pair.
+The transitional plausibility $\mathrm{TP}(c_{i},c_{j})$ between cells in different tissue slices is higher when they exhibit similar ligand-receptor communication patterns $f_{\mathrm{LR}}$ (see Figure 8 for an illustration).
+Formally, we define $\mathrm{LR}(c_{i},c_{j})$, the dissimilarity between the ligand–receptor communication patterns in the microenvironments of cells $c_{i}$ and $c_{j}$, as:
+$\mathrm{LR}(c_{i},c_{j})=\left\|\>f_{\mathrm{LR}}(\mathcal{N}_{r}(c_{i}))-f_{\mathrm{LR}}(\mathcal{N}_{r}(c_{j}))\right\|_{2}^{2}.$ (10)
+3.3 Flow Matching with Context-Aware OT Couplings
 Our proposed framework, graphically depicted in Figure 1, consists of the following three main steps:
 Transitional Plausibility Matrix.
-First, we create a sequence of transitional plausibility matrices (TPMs) to encode the biological priors for each pair of consecutive time points.
-Specifically, let $\mathbf{M}_{i,i+1}\in\mathbb{R}^{n_{i}\times n_{i+1}}$ be the TPM with respect to the set of cells measured at time $t_{i}$ and at time $t_{i+1}$, with size $n_{i}$ and $n_{i+1}$ respectively, where the $(k,l)$-th entry of $\mathbf{M}_{i,i+1}$ indicates how plausibly the $k$-th cell measured at $t_{i}$ will evolve to the $l$-th cell measured at $t_{i+1}$, defined as follows:
-$\big[\mathbf{M}_{i,i+1}\big]_{kl}=\lambda\cdot\mathrm{SS}\left(c_{i}(k),c_{i+1}(l)\right)+(1-\lambda)\cdot\mathrm{LR}\left(c_{i}(k),c_{i+1}(l)\right),$ (10)
-where $\lambda\in[0,1]$ is a trade-off hyperparameter that balances the contribution of the spatial smoothness prior ($\mathrm{SS}$) and the ligand–receptor communication prior ($\mathrm{LR}$).
+To begin with, we create a sequence of transitional plausibility matrices (TPMs) to encode the biological priors for each pair of consecutive time points.
+Specifically, denote by $\mathbf{M}_{i,i+1}\in\mathbb{R}^{n_{i}\times n_{i+1}}$ the TPM with respect to the set of cells measured at time $t_{i}$ and at time $t_{i+1}$, with size $n_{i}$ and $n_{i+1}$ respectively, where the $(k,l)$-th entry of $\mathbf{M}_{i,i+1}$ indicates how plausibly the $k$-th cell measured at $t_{i}$ will evolve to the $l$-th cell measured at $t_{i+1}$, defined as follows:
+$\big[\mathbf{M}_{i,i+1}\big]_{kl}=\lambda\cdot\mathrm{SS}$ $\left(c_{i}(k),c_{i+1}(l)\right)$ (11)
+$+$ $(1-\lambda)\cdot\mathrm{LR}\left(c_{i}(k),c_{i+1}(l)\right).$
+Here, $\lambda\in[0,1]$ denotes a trade-off hyperparameter that balances the contribution of the spatial smoothness prior ($\mathrm{SS}$) and the ligand–receptor communication prior ($\mathrm{LR}$).
 Prior-Regularized OT Couplings.
-The transitional plausibility matrices capture our spatially informed prior on cell-cell transitions between consecutive time points, which can naturally be incorporated in the EOT formulation (Equation 6) to promote couplings that maintain the structural and functional properties of the tissue organization.
+The transitional plausibility matrices capture our spatially informed prior over cell-cell transitions between consecutive time points, which can naturally be incorporated into the EOT formulation (Equation 7) to promote couplings that maintain the structural and functional properties of tissue organization.
 We propose two techniques for prior integration:
@@
 Consider the empirical counterpart of Equation 7 with respect to time $t_{i}$ and time $t_{i+1}$.
-Our first approach incorporates the transitional plausibility matrix directly into the transport cost:
-$\min_{\mathbf{\Pi}\in\mathbb{R}^{n_{i}\times n_{i+1}}}\sum_{k,l}\Pi_{kl}\underbrace{\Big[\alpha\cdot\|\bm{x}_{i}(k)-\bm{x}_{i+1}(l)\|_{2}^{2}+(1-\alpha)\cdot\big[\mathbf{M}_{i,i+1}\big]_{kl}\Big]}_{\text{Prior-Aware Cost Function}}-\epsilon\sum_{k,l}\Pi_{kl}\log\Pi_{kl},$ (11)
-where the transport plan $\Pi$ satisfies the boundary conditions: $\sum_{l}\Pi_{kl}=1/n_{i}$ for any $k\in[n_{i}]$, and $\sum_{k}\Pi_{kl}=1/n_{i+1}$ for any $l\in[n_{i+1}]$, and $\alpha\in[0,1]$ controls the trade-off between the original Euclidean cost and the prior-aware cost derived from the transitional plausibility.
+Our first PACM approach incorporates the transitional plausibility matrix directly into the transport cost:
+$\min_{\mathbf{\Pi}\in\mathbb{R}^{n_{i}\times n_{i+1}}}\;\langle\mathbf{\Pi},\mathbf{C}_{i,i+1}\rangle-\epsilon\sum_{k,l}\Pi_{kl}(\log\Pi_{kl}-1),$ (12)
+where the $(k,l)$-th entry of the new prior-aware cost matrix is $\mathbf{C}_{i,i+1}]_{kl}=\alpha\|\bm{x}_{i}(k)-\bm{x}_{i+1}(l)\|_{2}^{2}+(1-\alpha)[\mathbf{M}_{i,i+1}]_{kl}$, the transport plan $\Pi$ needs to satisfy the boundary conditions: $\sum_{l}\Pi_{kl}=1/n_{i}$ for any $k\in[n_{i}]$, and $\sum_{k}\Pi_{kl}=1/n_{i+1}$ for any $l\in[n_{i+1}]$, and $\alpha\in[0,1]$ controls the trade-off between the original Euclidean cost and the prior-aware cost derived from the transitional plausibility.
 If $\mathbf{M}_{i,i+1}]_{kl}$ is high, Equation 12 will impose a higher transport cost between the $k$-cell at time $i$ to the $j$-cell at time $i+1$.
-This aligns with our assumption that such transitions are implausible.
+This observation aligns with our assumption that such transitions should be more biologically implausible.
 Prior-Aware Entropy Regularization (PAER).
-While the prior-aware cost matrix approach penalizes couplings in accordance with our spatial priors, it defines a different OT problem characterized by a modified cost function.
+While the previous PACM approach penalizes couplings according to our spatial priors, it defines a different OT problem with a modified cost function.
 Consequently, the standard interpretation of OT as minimizing the transport energy between two transcriptomic distributions no longer holds.
 Since the scales of the pairwise distances often differ, normalization of the cost terms is required to enable meaningful comparison.
-This normalization, however, may result in couplings that deviate from their original counterparts (Proposition 1 and Corollary 1 in the Appendix C).
+This normalization, however, may result in couplings that deviate from their original counterparts (see Proposition 1 and Corollary 1 in Appendix D).
 Besides, selecting an appropriate $\alpha$ in Equation 12 introduces an additional layer of tuning, increasing computational overhead.
-Therefore, we propose a second approach to integrate the biological priors without introducing additional hyperparameters:
-$\min_{\mathbf{\Pi}\in\mathbb{R}^{n_{i}\times n_{i+1}}}$ $\sum_{k,l}\Pi_{kl}\|\bm{x}_{i}(k)-\bm{x}_{i+1}(l)\|_{2}^{2}-\epsilon\sum_{k,l}\underbrace{\Pi_{kl}\log(\Pi_{kl}/[\widehat{\mathbf{M}}_{i,i+1}]_{kl})}_{\text{Prior-Aware Entropy Regularization}},$ (12)
-where $\widehat{\mathbf{M}}_{i,i+1}]_{kl}=\exp(-[\mathbf{M}_{i,i+1}]_{kl})/\sum_{l}\exp(-[\mathbf{M}_{i,i+1}]_{kl})$ denotes the prior joint probability matrix induced by $\mathbf{M}_{i,i+1}$.
-Intuitively, the lower the cost $\mathbf{M}{i,i+1}]_{kl}$, the larger the entry $\widehat{\mathbf{M}}{i,i+1}]_{kl}$, reflecting a higher plausibility of the transition from cell $k$ at $t_{i}$ to cell $l$ at $t_{i+1}$.
+Therefore, we propose the second PAER approach to integrate the biological priors without introducing additional hyperparameters:
+$\min_{\mathbf{\Pi}\in\mathbb{R}^{n_{i}\times n_{i+1}}}\sum_{k,l}\Pi_{kl}\|\bm{x}_{i}(k)-\bm{x}_{i+1}(l)\|_{2}^{2}$ (13)
+$\qquad\qquad-\epsilon\sum_{k,l}\underbrace{\Pi_{kl}\big(\log\big(\Pi_{kl}/[\widehat{\mathbf{M}}_{i,i+1}]_{kl}\big)-1\big)}_{\text{Prior-Aware Entropy Regularization}},$
+where $\widehat{\mathbf{M}}_{i,i+1}]_{kl}=\dfrac{\exp(-[\mathbf{M}_{i,i+1}]_{kl})}{\sum_{l}\exp(-[\mathbf{M}_{i,i+1}]_{kl})}$ denotes the $(k,l)$-th entry of the prior joint probability matrix induced by the originally constructed $\mathbf{M}_{i,i+1}$.
+Note that the lower the cost $\mathbf{M}_{i,i+1}]_{kl}$, the larger the entry $\widehat{\mathbf{M}}_{i,i+1}]_{kl}$, suggesting a higher plausibility of the transition from cell $k$ at $t_{i}$ to cell $l$ at $t_{i+1}$.
 The entropy regularization term in Equation 13 thus biases the learned transport plan toward the prior $\widehat{\mathbf{M}}_{i,i+1}$ rather than a uniform baseline, providing a soft mechanism for incorporating biological prior knowledge.
-ContextFlow.
-Finally, we apply the Sinkhorn algorithm (Cuturi, 2013) to solve the optimization problem in Equation 11 or Equation 12 to obtain the spatial context-aware EOT couplings, and train the neural velocity vector field $u_{\theta}$ based on stochastic gradient descent by minimizing the multi-time generalization of Equation 3 with respect to conditionals $p_{t}(\bm{x}|\bm{z})$ and $u_{t}(\bm{x}|\bm{z})$ defined according to Equation 7.
-The pseudocode for the proposed method, named Conditional Flow Matching with Context-Aware OT Couplings (ContextFlow), is detailed in Algorithm 1 in Appendix D.
-In particular, to apply the Sinkhorn algorithm to solve our prior-aware entropy regularization problem in Equation 12, we make use of the following theorem, a generalized result of Peyré et al.
-(2019).
-Theorem 1.
-Let $\mathbf{C}\in\mathbb{R}^{n_{0}\times n_{1}}$ be a cost matrix and $\mathbf{M}\in\mathbb{R}^{n_{0}\times n_{1}}$ be a prior transition probability matrix.
-Suppose $\mathbf{\Pi}^{*}_{\mathrm{CTF-H}}$ is the solution to the following prior-aware optimal transport problem:
-$\mathbf{\Pi}^{*}_{\mathrm{CTF-H}}={\arg\!\min}_{\mathbf{\Pi}\in\mathbb{R}^{n_{0}\times n_{1}}}$ $\sum_{k,l}\Pi_{kl}C_{kl}+\epsilon\sum_{k,l}\Pi_{kl}\log(\Pi_{kl}/{M}_{kl}),$
-where $\epsilon>0$ is the regularization parameter.
-Then, we can show that $\mathbf{\Pi}^{*}_{\mathrm{CTF-H}}$ can be computed by Sinkhorn and takes the form $\mathrm{diag}(\bm{u})\cdot\mathbf{M}\odot\exp(-\mathbf{C}/\epsilon)\cdot\mathrm{diag}(\bm{v})$, where $\odot$ denotes element-wise multiplication, and $\bm{u}\in\mathbb{R}^{n_{0}},\bm{v}\in\mathbb{R}^{n_{1}}$ are vectors satisfying the marginalization constraints.
-Theorem 1, proven in Appendix B, suggests a new Gibbs kernel $\mathbf{K}=\mathbf{M}\odot\exp(-\mathbf{C}/\epsilon)$, which combines both the transport cost and the prior joint probability matrices.
-When $\epsilon\rightarrow 0$, $\mathbf{\Pi}^{*}_{\mathrm{CTF}}\rightarrow\mathbf{\Pi}_{\mathrm{ot}}^{*}$, thereby recovering the standard OT couplings in Equation 16.
-When $\epsilon\rightarrow\infty$, the optimal coupling $\mathbf{\Pi}^{*}_{\mathrm{CTF}}\rightarrow\mathrm{diag}(\bm{u})\cdot\mathbf{M}\cdot\mathrm{diag}(\bm{v})$, which corresponds to a plan that aligns with the prior defined by $\mathbf{M}$ rather than the independent couplings obtained with Entropic-OT (Section 2.3).
-This has the same effect as constraining our transport plan through the proposed prior and, by extension, the flow.
-By varying the parameter $\epsilon$, we can thus efficiently optimize for a desirable coupling via the Sinkhorn algorithm.
-4 Experiments
-Datasets.
-We evaluate ContextFlow on three longitudinal spatial transcriptomics datasets: Axolotl Brain Regeneration (Wei et al., 2022), Mouse Embryo Organogenesis (Chen et al., 2022), and Liver Regeneration (Ben-Moshe et al., 2021).
-For all the datasets, the gene expression values are log-normalized, and we extract the top $50$ principal components (PCs) as feature vectors.
-The strength of ligand–receptor interactions in the microenvironment was inferred using spatially informed bivariate statistics implemented in LIANA+ (Dimitrov et al., 2024), where we applied the cosine similarity metric to gene expression profiles.
-Interaction evidence was aggregated using the consensus of multiple curated ligand–receptor resources, ensuring robustness of the inferred signals.
-Baselines & Metrics.
-We benchmark ContextFlow using its two prior integration strategies—cost-regularized (CTF-C) and entropy-regularized (CTF-H)—against several baselines.
-As a non-spatial baseline, we include conditional flow matching (CFM), which uses only transcriptomic data with random couplings.
-We further compare against minibatch-OT flow matching (MOTFM), which leverages OT-derived couplings but does not incorporate spatial priors.
-For evaluation, we employed 2-Wasserstein distance ($\mathcal{W}_{2}$), a commonly used OT-based metric, and metrics such as MMD and Energy Distance for statistical fidelity.
-Furthermore, to assess the biological plausibility of our predicted dynamics, we evaluate them using a cell-type-weighted Wasserstein distance (Weighted $\mathcal{W}_{2}$), where the weights correspond to the relative frequency of each cell type in the dataset.
-Exact metric definitions are present in the Appendix F.
-All reported metrics are averaged across $10$ runs.
-Sampling.
-A trained velocity field can be evaluated through the samples it generates.
-We consider two variants.
-Initial value problem sampling (IVP) integrates the learned gradient starting from the first observed batch of cells and evolves them toward a later time point.
-IVP provides the most comprehensive evaluation of flow quality, as errors can accumulate across steps.
-In contrast, next-step sampling (Next Step) integrates the gradient only from the most recently observed batch of cells, thus limiting error propagation but providing a less stringent test of long-term trajectory fidelity.
 Table 2: Interpolation at the middle holdout time point for the Brain Regeneration dataset.
@@
 $0.5$ $3.917\pm 0.343$ $4.159\pm 0.455$ $0.147\pm 0.022$ $29.613\pm 4.822$
-4.1 Axolotl Brain Regeneration
-We first evaluate ContextFlow on longitudinal Stereo-seq spatial transcriptomic data coming from a post-traumatic brain regeneration study of the Salamander (axolotl telencephalon) species (Wei et al., 2022).
-The dataset contains samples from five developmental stages, with replicates collected from different individual organisms at each stage.
-For our CTF-C method, we present the best ablated $\alpha$ in the main text, with full ablation results across different $\alpha$ values provided in Appendix H.
-For interpolation, we hold out the middle time point during training and evaluate it using samples generated by the trained velocity field $u_{\theta}$ via both IVP and next-step sampling.
-Table 1 presents the results.
-Across multiple evaluation metrics, ContextFlow with entropy regularization (CTF-H) produces trajectories that most closely match the ground truth.
-CTF-H consistently achieves the best or comparable performance relative to CTF-C, despite the latter being explicitly tuned across multiple $\alpha$ values.
-This highlights the computational efficiency and superior generalization ability of CTF-H, as it avoids the need for additional hyperparameter tuning while maintaining strong performance.
-For extrapolation, we evaluate generation on the last holdout time point, representing the most challenging test of generalizability for the velocity fields $u_{\theta}$, as it lies outside the training time horizon.
-As shown in Table 2, CTF-H again consistently achieves the best overall performance, particularly under IVP-Sampling, where errors are most likely to accumulate.
-This result further reinforces the robustness and reliability of CTF-H across the entire sampling horizon.
-Finally, Figure 3 (Appendix G.1) demonstrates that incorporating spatial priors enables ContextFlow to produce substantially fewer biologically implausible couplings compared to its context-free counterpart.
+ContextFlow.
+We apply the Sinkhorn algorithm (Cuturi, 2013) to solve the optimization problems defined in Equations 12 and 13 to obtain the spatial context-aware EOT couplings, and train the neural velocity vector field $u_{\theta}$ based on stochastic gradient descent by minimizing the multi-time generalization of Equation 3 with respect to conditionals $p_{t}(\bm{x}|\bm{z})$ and $u_{t}(\bm{x}|\bm{z})$ defined according to Equation 3.1.
+The pseudocode for the proposed method, named Conditional Flow Matching with Context-Aware OT Couplings, is presented in Algorithm 1 in Appendix E.1.
+We also provide time complexity analysis in Appendix E.2.
+In particular, to apply the Sinkhorn algorithm to solve our PAER problem in Equation 13, we make use of the following theorem, a generalized result of Peyré et al.
+(2019).
+Theorem 3.1.
+Let $\mathbf{C}\in\mathbb{R}^{n_{0}\times n_{1}}$ be a cost matrix and $\mathbf{M}\in\mathbb{R}^{n_{0}\times n_{1}}$ be a prior transition probability matrix.
+Suppose $\mathbf{\Pi}^{*}_{\mathrm{CTF-H}}$ is the solution to the following prior-aware optimal transport problem:
+$\min_{\mathbf{\Pi}\in\mathbb{R}^{n_{0}\times n_{1}}}\sum_{k,l}\Pi_{kl}C_{kl}-\epsilon\sum_{k,l}\Pi_{kl}\big(\log(\Pi_{kl}/{M}_{kl})-1\big),$
+where $\epsilon>0$ is the regularization parameter.
+Then, we can show that $\mathbf{\Pi}^{*}_{\mathrm{CTF-H}}$ can be computed by Sinkhorn and takes the form $\mathrm{diag}(\bm{u})\cdot\mathbf{M}\odot\exp(-\mathbf{C}/\epsilon)\cdot\mathrm{diag}(\bm{v})$, where $\odot$ denotes element-wise multiplication, and $\bm{u}\in\mathbb{R}^{n_{0}},\bm{v}\in\mathbb{R}^{n_{1}}$ are vectors satisfying the marginalization constraints.
+Theorem 3.1, proven in Appendix C, suggests a new Gibbs kernel $\mathbf{K}=\mathbf{M}\odot\exp(-\mathbf{C}/\epsilon)$, which combines both the transport cost and the prior joint probability matrices.
+When $\epsilon\rightarrow 0$, $\mathbf{\Pi}^{*}_{\mathrm{CTF}}\rightarrow\mathbf{\Pi}_{\mathrm{ot}}^{*}$, thereby recovering the standard OT couplings in Equation 14.
+When $\epsilon\rightarrow\infty$, the optimal coupling $\mathbf{\Pi}^{*}_{\mathrm{CTF}}\rightarrow\mathrm{diag}(\bm{u})\cdot\mathbf{M}\cdot\mathrm{diag}(\bm{v})$, which corresponds to a plan that aligns with the prior defined by $\mathbf{M}$ rather than the independent couplings obtained with EOT defined by Equation 7.
+This has the same effect as constraining our transport plan through the proposed prior and, by extension, the flow.
+By varying the parameter $\epsilon$, we can thus efficiently optimize for a desirable coupling via the Sinkhorn algorithm.
+4 Experiments
+Datasets.
+We evaluate ContextFlow on three longitudinal spatial transcriptomics datasets: Axolotl Brain Regeneration (Wei et al., 2022), Mouse Embryo Organogenesis (Chen et al., 2022), and Liver Regeneration (Ben-Moshe et al., 2021).
+For all the evaluated datasets, the gene expression values are log-normalized, and we extract the top $50$ principal components (PCs) as feature vectors.
+The strength of ligand–receptor interactions in the microenvironment was inferred using spatially informed bivariate statistics implemented in LIANA+ (Dimitrov et al., 2024), where we applied the cosine similarity metric to gene expression profiles.
+Interaction evidence was aggregated using the consensus of multiple curated ligand–receptor resources, ensuring robustness of the inferred signals.
+Baselines & Metrics.
+We benchmark ContextFlow using its two prior integration strategies—cost-regularized (CTF-C) and entropy-regularized (CTF-H)—against several existing flow matching baselines.
+As a non-spatial baseline, we include conditional flow matching (CFM), which uses only transcriptomic data with random couplings.
+We further compare against minibatch-OT flow matching (MOTFM), which leverages OT-derived couplings but does not incorporate spatial priors.
+For evaluation, we employed the 2-Wasserstein distance ($\mathcal{W}_{2}$), a commonly used OT-based metric, as well as MMD and Energy Distance for statistical fidelity.
+Furthermore, to assess the biological plausibility of our predicted dynamics, we evaluate them using a cell-type-weighted Wasserstein distance (Weighted $\mathcal{W}_{2}$), where the weights correspond to the relative frequency of each cell type in the dataset.
+Exact metric definitions are present in Appendix B.2.
+To account for randomness, all reported metrics are averaged across $10$ runs.
 Table 3: Extrapolation on the last holdout time point for the Brain Regeneration dataset.
@@
 $0.5$ $6.254\pm 0.819$ $\bm{5.973\pm 0.757}$ $0.156\pm 0.025$ $54.330\pm 12.089$
+Sampling.
+A trained velocity field can be evaluated through the samples it generates.
+We consider two variants of sampling configurations: initial value problem sampling (IVP) and next-step sampling (Next Step).
+In particular, IVP integrates the learned gradients from the first observed batch of cells and evolves them to a later time point.
+IVP provides the most comprehensive evaluation of flow quality, as errors can accumulate across steps.
+In contrast, Next Step integrates the gradient only from the most recently observed batch of cells, thus limiting error propagation but providing a less stringent test of long-term trajectory fidelity.
+4.1 Axolotl Brain Regeneration
+We first evaluate ContextFlow on longitudinal Stereo-seq spatial transcriptomic data coming from a post-traumatic brain regeneration study of the Salamander (axolotl telencephalon) species (Wei et al., 2022).
+The dataset contains samples from five developmental stages, with replicates collected from different individual organisms at each stage.
+For our CTF-C method, we present the best ablated $\alpha$, with full ablation results across different $\alpha$ values in Appendix H.
+For interpolation, we hold out the middle time point during training and evaluate it using samples generated by the trained velocity field $u_{\theta}$ via both IVP and next-step sampling.
+Table 2 presents the results.
+Across multiple evaluation metrics, ContextFlow with entropy regularization (CTF-H) produces trajectories that most closely match the ground truth.
+CTF-H consistently achieves the best or comparable performance relative to CTF-C, despite the latter being explicitly tuned across multiple $\alpha$ values.
+This highlights the computational efficiency and superior generalization ability of CTF-H, as it avoids the need for additional hyperparameter tuning while maintaining strong performance.
+For extrapolation, we evaluate generation at the last holdout time point, which represents the most challenging test of generalizability for the velocity fields $u_{\theta}$, as it lies outside the training time horizon.
+As shown in Table 3, CTF-H again consistently achieves the best overall performance, particularly under IVP-Sampling, where errors are most likely to accumulate.
+This result further reinforces the robustness and reliability of CTF-H across the entire sampling horizon.
+Figure 4 in Appendix G.1 further illustrates that incorporating spatial priors enables ContextFlow to produce substantially fewer biologically implausible couplings compared to its context-free counterpart.
 Table 4: Interpolation (time 5) and extrapolation (time 8) results on the Organogenesis dataset.
@@
 $0.5$ $1.871\pm 0.030$ $1.919\pm 0.067$ $\bm{2.814\pm 0.414}$ $\bm{3.233\pm 0.567}$ $1.636\pm 0.060$ $1.684\pm 0.099$
-(a) Brain Regeneration
-(b) Organogenesis
-Figure 2: KL-Divergence between predicted and ground-truth cell type distributions.
 Table 5: Interpolation results on the middle holdout time point for the Liver Regeneration dataset.
@@
 On the extrapolation task, integrating to the final time point, CTF-H again achieves the strongest performance, underscoring that the entropy-regularized formulation not only removes the need for additional parameter tuning but also offers more robust generalization to unseen temporal horizons.
+(a) Brain Regeneration
+(b) Organogenesis
+Figure 2: Comparisons of predicted and true cell type distributions.
 Figure 2 reports the KL-Divergence between normalized histograms of predicted and ground-truth cell types from ContextFlow and MOTFM.
@@
 The cell type progression is further visualized in Figure 10 (Appendix H.6).
-We show that ContextFlow predicts temporal cell type trajectories that evolve smoothly and consistently across consecutive developmental stages.
-Early progenitor populations, such as neural crest and mesenchyme, progressively diminish as development advances, while terminal fates, including muscle, cartilage primordium, and liver, emerge at later stages.
-Major lineages such as brain, heart, and connective tissue remain continuous throughout, demonstrating that ContextFlow captures biologically coherent and temporally consistent developmental dynamics.
+We show that ContextFlow predicts temporal trajectories of cell types that evolve smoothly and consistently across consecutive developmental stages.
+Early progenitor populations, such as neural crest and mesenchyme, progressively diminish as development progresses, while terminal fates, including muscle, the cartilage primordium and the liver, emerge at later stages.
+Major lineages such as brain, heart, and connective tissue remain continuous throughout, confirming that ContextFlow captures biologically coherent and temporally consistent developmental dynamics.
 4.3 Liver Regeneration
 Finally, we evaluate ContextFlow on a Visium spatial transcriptomics dataset profiling the temporal dynamics of mouse liver regeneration following acetaminophen-induced injury (Ben-Moshe et al., 2021), collected across three distinct regeneration stages.
-Unlike the earlier datasets resolved at single-cell resolution, Visium data is captured at the level of $55$ micron diameter spots, capturing the joint expression of multiple cells.
+Unlike earlier datasets resolved at single-cell resolution, Visium data is captured at the level of $55$ micron diameter spots, capturing the joint expression of multiple cells.
 Since direct cell-type information is unavailable, we restrict evaluation to the 2-Wasserstein distance.
 Moreover, since evaluation is performed on the middle of the three time points, IVP and next-step predictions coincide.
-Table 4 presents the results.
+Table 5 summarizes the comparisons between MOTFM and our methods.
 Consistent with the previous findings, CTF-H achieved the lowest reconstruction error, indicating that incorporating contextual information improves trajectory estimation even in aggregated spot-level measurements.
 5 Conclusion
-We introduced ContextFlow, a contextually aware flow matching framework that leverages spatial priors and biologically motivated constraints to learn more plausible trajectories from snapshot spatial transcriptomic data, addressing a central challenge of existing methods.
-The entropic variant of ContextFlow is theoretically grounded, which always yields OT couplings constrained by prior knowledge, promoting stability and consistency with the imposed contextual constraints.
-Across three diverse datasets, we showed that ContextFlow consistently improves over state-of-the-art baselines even in challenging Initial Value Problem sampling settings, underscoring the importance of our contextually informed priors.
-In addition, we demonstrated that our framework reduces the number of biologically implausible couplings and results in coherent and temporally consistent developmental trajectories while maintaining strong quantitative performance across Wasserstein, MMD, and Energy metrics.
+We introduced ContextFlow, a contextually aware flow matching framework that leverages spatial priors and biologically motivated constraints to learn more plausible trajectories from snapshot spatial transcriptomic data.
+The entropic variant of ContextFlow is theoretically grounded, which produces OT couplings that are constrained by prior knowledge, thereby promoting stability and consistency with the imposed contextual constraints.
+Across three diverse datasets, we showed that ContextFlow consistently outperforms state-of-the-art baselines, even in challenging Initial Value Problem sampling settings, underscoring the importance of our contextually informed priors.
+In addition, we demonstrated that our framework reduces the number of biologically implausible couplings and yields coherent, temporally consistent developmental trajectories while maintaining strong quantitative performance across Wasserstein, MMD, and Energy metrics.
 These results highlight the value of embedding biological context into generative flow models.
-Future works can adapt our methods to reconstruct tissues and learn spatial latent dynamics by formulating the flow in space (rather than time), or leverage multi-marginal OT formulations for optimizing temporal flows.
-Looking forward, ContextFlow offers a principled foundation for modeling perturbations and disease progression, bridging generative power with biological interpretability.
+Future work can adapt our methods to reconstruct tissues and learn spatial latent dynamics by formulating the flow in space (rather than time), or by leveraging multi-marginal OT formulations to optimize temporal flows.
+Looking forward, ContextFlow offers a principled foundation for modeling perturbations and disease progression, potentially bridging generative power with biological interpretability.
+Impact Statement
+Our work contributes a new method for modeling dynamical systems from unpaired observations with structural constraints.
+As such, our algorithm can contribute to biomedical domains, for example, by modeling drug perturbation responses in spatially resolved tumor tissues or by analyzing complex biological dynamics with unpaired datasets.
+These applications could positively impact biomedical research by enabling more data-efficient modeling of longitudinal biological processes and expediting hypothesis generation about treatment effects or underlying mechanisms.
+Future work should emphasize broad validation of our method using experimental or clinical ground truth, and the interpretation of results with domain experts to build appropriate trust in predictions.
 Appendix A Related Work
@@
 Normalizing flows provide a parametric framework for characterizing transformations of a random variable into desired distributions (Papamakarios et al., 2021).
-These transformations can be realized through either finite (Rezende & Mohamed, 2015) or continuous compositions (Chen et al., 2018).
+These transformations can be realized through either finite (Rezende and Mohamed, 2015) or continuous compositions (Chen et al., 2018).
 The loss functions used in such formulations typically require computing Jacobians or integrating the flows at each forward pass, making them computationally expensive.
-Flow matching (FM) (Lipman et al., 2023; Albergo & Vanden-Eijnden, 2022; Liu et al., 2022) addresses this limitation by reducing the training of the velocity field to a regression problem, thereby making normalizing flows substantially more scalable.
+Flow matching (FM) (Lipman et al., 2023; Albergo and Vanden-Eijnden, 2022; Liu et al., 2022) addresses this limitation by reducing the training of the velocity field to a regression problem, thereby making normalizing flows substantially more scalable.
 To ensure valid conditional paths at intermediate time points, samples are coupled either randomly or via optimal transport (Pooladian et al., 2023; Tong et al., 2024).
@@
 While these methods are primarily designed to align biological data across space, time, or modality, they do not address the problem of trajectory inference toward biologically plausible solutions, leveraging biological priors to constrain or bias the transport plan.
-Appendix B Proofs of Main Theoretical Results
-Proposition 1.
+Appendix B Additional Definitions
+B.1 Kantarovich’s OT Formulation
+Kantorovich’s formulation (Peyré et al., 2019) is a classical definition of the optimal transport (OT) problem that seeks a joint coupling to move a probability measure to another that minimizes the Euclidean distance cost, corresponding to the following minimization problem with respect to the 2-Wasserstein distance:
+$\pi^{*}_{\mathrm{ot}}:={\arg\!\min}_{\pi\in\Pi(q_{0},q_{1})}\int_{\mathbb{R}^{d}\times\mathbb{R}^{d}}\|\bm{x}_{0}-\bm{x}_{1}\|_{2}^{2}\>d\pi(\bm{x}_{0},\bm{x}_{1}),$ (14)
+where $\Pi(q_{0},q_{1})$ denotes the set of joint probability measures such that the left and right marginals are $q_{0}$ and $q_{1}$.
+Equation 14 can be solved in a mini-batch fashion using standard solvers such as POT (Flamary et al., 2021); however, the computational complexity is cubic in batch size.
+B.2 Evaluation Metrics
+2-Wasserstein. The 2-Wasserstein distance ($\mathcal{W}_{2}$ between empirical distributions $\mu,\nu$ is defined as:
+$\mathcal{W}_{2}(\mu,\nu)=\inf_{\gamma\in\Pi(\mu,\nu)}\bigg(\sum_{(\bm{x},\bm{y})}\gamma(\bm{x},\bm{y})\cdot\|\bm{x}-\bm{y}\|_{2}^{2}\bigg)^{1/2},$
+where $\Pi(\mu,\nu)$ denotes the set of couplings between $\mu$ and $\nu$.
+Weighted 2-Wasserstein.
+Implausible velocity fields can steer a cell’s transcriptional trajectory in unrealistic directions, potentially leading to entirely different terminal cell types.
+We thus employ the weighted 2-Wasserstein metric, which ensures the evaluation accounts for both transcriptional similarity and the distributional balance of cell types.
+We define the weighted 2-Wasserstein distance (Weighted $\mathcal{W}_{2}$) between true and predicted distributions as:
+$\text{Weighted-$\mathcal{W}_{2}$}(\mu,\nu)=\sum_{i=1}^{C}\frac{n_{i}^{\text{true}}}{N}\cdot\mathcal{W}_{2}\!\bigg(\frac{1}{n_{i}^{\text{true}}}\!\!\sum_{j:y_{j}=i}\delta_{\bm{x}_{j}},\frac{1}{n_{i}^{\text{pred}}}\!\!\sum_{j:\hat{y}_{j}=i}\delta_{\bm{x}_{j}}\bigg),$
+where $n_{i}^{\text{true}},n_{i}^{\text{pred}}$ are the number of true and predicted cells of type $i$, and $N$ is the total number of samples.
+To determine the cell type of generated trajectories, we employ a multi-class classifier $M_{\phi}$, implemented as an XGBoost model (Chen and Guestrin, 2016) trained for each dataset.
+Energy Distance.
+Let $\mu$ and $\nu$ be probability distributions with samples $X=\{\bm{x}_{i}\}_{i=1}^{m}\sim\mu$ and $Y=\{\bm{y}_{j}\}_{j=1}^{n}\sim\nu$.
+The squared empirical energy distance (Energy) is defined as:
+$\mathrm{ED}(\mu,\nu)=\frac{2}{mn}\sum_{i=1}^{m}\sum_{j=1}^{n}\|\bm{x}_{i}-\bm{y}_{j}\|-\frac{1}{m^{2}}\sum_{i=1}^{m}\sum_{i^{\prime}=1}^{m}\|\bm{x}_{i}-\bm{x}_{i^{\prime}}\|-\frac{1}{n^{2}}\sum_{j=1}^{n}\sum_{j^{\prime}=1}^{n}\|\bm{y}_{j}-\bm{y}_{j^{\prime}}\|,$
+where $\|\cdot\|$ is the Euclidean norm. The distance is non-negative and equals zero if and only if $\mu=\nu$.
+Maximum Mean Discrepancy. For the same samples, the unbiased empirical estimate of the squared maximum mean discrepancy (MMD) with kernel $\kappa$ is defined as:
+$\mathrm{MMD}(\mu,\nu;\kappa)=\frac{1}{m(m-1)}\sum_{i\neq i^{\prime}}\kappa(\bm{x}_{i},\bm{x}_{i^{\prime}})+\frac{1}{n(n-1)}\sum_{j\neq j^{\prime}}\kappa(\bm{y}_{j},\bm{y}_{j^{\prime}})-\frac{2}{mn}\sum_{i=1}^{m}\sum_{j=1}^{n}\kappa(\bm{x}_{i},\bm{y}_{j}).$
+In our evaluations, we use a multi-kernel variant with radial basis function (RBF) kernels $\kappa_{\gamma}(\bm{x},y)=\exp(-\gamma\|\bm{x}-\bm{y}\|^{2})$, and average over $\gamma\in[2,1,0.5,0.1,0.01,0.005]$.
+Appendix C Proofs of Main Theoretical Results
+Proposition C.1.
 Let $\mathbf{C}\in\mathbb{R}^{n_{0}\times n_{1}}$ be a cost matrix and $\mathbf{M}\in\mathbb{R}^{n_{0}\times n_{1}}$ a prior transition matrix with positive entries.
 Consider the entropy-regularized OT formulation:
-$\Pi^{*}$ $={\arg\!\min}_{\Pi\geq 0}\;\sum_{k,l}\Pi_{kl}C_{kl}+\epsilon\sum_{k,l}\Pi_{kl}\log\!\left(\Pi_{kl}\right).$
+$\Pi^{*}$ $={\arg\!\min}_{\Pi\geq 0}\;\sum_{k,l}\Pi_{kl}C_{kl}+\epsilon\sum_{k,l}\Pi_{kl}(\log\!\left(\Pi_{kl}\right)-1).$
 Let $\tilde{\Pi}^{*}$ be the EOT-coupling where the cost is scaled by a normalization constant $c$ or $\tilde{C}_{ij}=\frac{C_{ij}}{c}$.
@@
 $\left|\log\left(\frac{\tilde{\Pi}^{*}_{ij}}{\tilde{\Pi}^{*}_{kl}}\right)-\frac{1}{c}\log\left(\frac{\Pi^{*}_{ij}}{\Pi^{*}_{kl}}\right)\right|\leq|\phi_{i}|+|\phi_{k}|+|\psi_{j}|+|\psi_{l}|.$
-From Proposition 3 Proposition 2, we have:
+From Proposition 3 C.3, we have:
 $\max_{i}\phi_{i}\leq E,\quad\max_{i}\psi_{i}\leq E.$
@@
 $\frac{\tilde{\Pi}^{*}_{ij}}{\tilde{\Pi}^{*}_{kl}}\leq\gamma\left(\frac{\Pi^{*}_{ij}}{\Pi^{*}_{kl}}\right)^{1/c}.$
-Corollary 1.
+Corollary C.2.
 Let $\mathbf{C}\in\mathbb{R}^{n_{0}\times n_{1}}$ be a cost matrix and $\mathbf{M}\in\mathbb{R}^{n_{0}\times n_{1}}$ a prior transition matrix with positive entries.
 Consider the entropy-regularized OT formulation:
-$\Pi^{*}$ $={\arg\!\min}_{\Pi\geq 0}\;\sum_{k,l}\Pi_{kl}C_{kl}+\epsilon\sum_{k,l}\Pi_{kl}\log\!\left(\Pi_{kl}\right).$
+$\Pi^{*}$ $={\arg\!\min}_{\Pi\geq 0}\;\sum_{k,l}\Pi_{kl}C_{kl}+\epsilon\sum_{k,l}\Pi_{kl}(\log\!\left(\Pi_{kl}\right)-1).$
 Let $\tilde{\Pi}^{*}$ be the EOT-coupling in the case when cost is scaled by a normalization constant $c$ or $\tilde{C}_{ij}=\frac{C_{ij}}{c}$.
@@
 where $m=\frac{e^{2E}}{c\Pi^{*}_{\min}}$ and $s=2E\cdot e^{2E}$. ∎
-Proposition 2.
+Proposition C.3.
 Let $\mathbf{C}\in\mathbb{R}^{n_{0}\times n_{1}}$ be a cost matrix and $\mathbf{M}\in\mathbb{R}^{n_{0}\times n_{1}}$ a prior transition matrix with positive entries.
 Consider the entropy-regularized OT formulation:
-$\Pi^{*}$ $={\arg\!\min}_{\Pi\geq 0}\;\sum_{k,l}\Pi_{kl}C_{kl}+\epsilon\sum_{k,l}\Pi_{kl}\log\!\left(\Pi_{kl}\right).$
+$\Pi^{*}$ $={\arg\!\min}_{\Pi\geq 0}\;\sum_{k,l}\Pi_{kl}C_{kl}+\epsilon\sum_{k,l}\Pi_{kl}(\log\!\left(\Pi_{kl}\right)-1).$
 Let $\tilde{\Pi}^{*}$ be the EOT-coupling in the case when cost is scaled by a normalization constant $c$ or $\tilde{C}_{ij}=\frac{C_{ij}}{c}$.
@@
 $\max_{i}|\alpha_{i}-1|,\max_{j}|\beta_{j}-1|\leq\|M^{-1}\|_{\infty}\cdot\max(\|\Delta a\|_{\infty},\|\Delta b\|_{\infty}).$
-Theorem 1.
+Theorem 3.1.
 Let $\mathbf{C}\in\mathbb{R}^{n_{0}\times n_{1}}$ be a general cost matrix and $\mathbf{M}\in\mathbb{R}^{n_{0}\times n_{1}}$ be a prior transition probability matrix.
 Suppose $\Pi^{*}_{\mathrm{CTF-H}}$ is the solution to the following prior-aware optimal transport problem:
-$\Pi^{*}_{\mathrm{CTF-H}}=\operatorname*{arg\,min}_{\Pi\in\mathbb{R}^{n_{0}\times n_{1}}}$ $\sum_{k,l}\Pi_{kl}C_{kl}+\epsilon\sum_{k,l}\Pi_{kl}\log(\Pi_{kl}/{M}_{kl}),$
+$\Pi^{*}_{\mathrm{CTF-H}}=\operatorname*{arg\,min}_{\Pi\in\mathbb{R}^{n_{0}\times n_{1}}}$ $\sum_{k,l}\Pi_{kl}C_{kl}+\epsilon\sum_{k,l}\Pi_{kl}\log((\Pi_{kl}/{M}_{kl})-1),$
 where $\epsilon>0$ is the regularization parameter.
@@
 Setting the derivative to zero:
-$\epsilon\log\left(\frac{\Pi^{*}_{kl}}{M_{kl}}\right)=f_{k}-C_{kl}-g_{l}$
+$\epsilon\log\left(\frac{\Pi^{*}_{kl}}{M_{kl}}\right)=f_{k}-C_{kl}+g_{l}$
 $\implies\frac{\Pi^{*}_{kl}}{M_{kl}}=e^{\frac{f_{k}}{\epsilon}}e^{-\frac{C_{kl}}{\epsilon}}e^{\frac{g_{l}}{\epsilon}}$
@@
 $\implies\Pi^{*\top}\mathbf{1}=b$ (17)
-From equations 16 13, 17 14, and 18 15 above, we get:
+From equations 16 15, 17 16, and 18 17 above, we get:
 $\operatorname{diag}(u)\cdot K\cdot\operatorname{diag}(v)\cdot\mathbf{1}$ $=a$
@@
 $u_{k}^{t+1}=\frac{a_{k}}{(Kv^{t})_{k}},\quad v_{l}^{t+1}=\frac{b_{l}}{(K^{\top}u^{t+1})_{l}}$
-Sinkhorn Algorithm uses these updates, iteratively, and these updates are shown to converge in Franklin & Lorenz (1989).
+Sinkhorn Algorithm uses these updates, iteratively, and these updates are shown to converge in (Franklin and Lorenz, 1989).
 Thus, Sinkhorn Algorithm can be used for the ContextFlow’s Prior Aware Entropy Regularized (PAER) (CTF-H) formulation.
-From equation (9) 13, we get:
+From equation (9) 15, we get:
 $\Pi^{*}_{kl}=e^{f_{k}/\epsilon}M_{kl}e^{-C_{kl}/\epsilon}e^{g_{l}/\epsilon}$
@@
 Such that marginal constraints, $\Pi^{*}_{\text{CTF-H}}\mathbf{1}=a$ and $\Pi^{*\top}_{\text{CTF-H}}\mathbf{1}=b$ are satisfied. ∎
-Appendix C Effects of normalization on Prior aware cost matrix
-From Peyré et al.
-(2019), we know that optimal MOTFM coupling takes the form $\Pi^{*}_{\text{EOT}}=\text{diag}(u)\cdot K\cdot\text{diag}(v)$, where $K$ is the kernel matrix such that $K]_{ij}=\exp(\frac{-c_{ij}}{\epsilon})$, with $u,v$ satisfying marginalization constraints $u\odot Kv=a$ and $K^{T}u\odot v=b$.
+Appendix D Effects of normalization on Prior aware cost matrix
+From (Peyré et al., 2019), we know that optimal MOTFM coupling takes the form $\Pi^{*}_{\text{EOT}}=\text{diag}(u)\cdot K\cdot\text{diag}(v)$, where $K$ is the kernel matrix such that $K]_{ij}=\exp(\frac{-c_{ij}}{\epsilon})$, with $u,v$ satisfying marginalization constraints $u\odot Kv=a$ and $K^{T}u\odot v=b$.
 Sinkhorn updates are given by:
@@
 Intuitively, scaling the cost has the same effect as that of increasing $\epsilon$, making solutions more diffused.
-Proposition 1.
+Proposition C.1.
 Let $\mathbf{C}\in\mathbb{R}^{n_{0}\times n_{1}}$ be a cost matrix and $\mathbf{M}\in\mathbb{R}^{n_{0}\times n_{1}}$ a prior transition matrix with positive entries.
 Consider the entropy-regularized OT formulation:
-$\Pi^{*}$ $={\arg\!\min}_{\mathbf{\Pi}\geq 0}\;\sum_{k,l}\Pi_{kl}C_{kl}+\epsilon\sum_{k,l}\Pi_{kl}\log\!\left(\Pi_{kl}\right).$
+$\Pi^{*}$ $={\arg\!\min}_{\mathbf{\Pi}\geq 0}\;\sum_{k,l}\Pi_{kl}C_{kl}+\epsilon\sum_{k,l}\Pi_{kl}(\log\!\left(\Pi_{kl}\right)-1).$
 Let $\tilde{\Pi}^{*}$ be the EOT-coupling where the cost is scaled by a normalization constant $c$ or $\tilde{C}_{ij}=\frac{C_{ij}}{c}$.
@@
 This results in bringing probabilities that are far apart closer to each other or, in essence, in creating more diffused and less sharp couplings.
-Corollary 1.
+Proposition C.2.
 Let $\mathbf{C}\in\mathbb{R}^{n_{0}\times n_{1}}$ be a cost matrix and $\mathbf{M}\in\mathbb{R}^{n_{0}\times n_{1}}$ a prior transition matrix with positive entries.
 Consider the entropy-regularized OT formulation:
-$\mathbf{\Pi}^{*}$ $={\arg\!\min}_{\mathbf{\Pi}\geq 0}\;\sum_{k,l}\Pi_{kl}C_{kl}+\epsilon\sum_{k,l}\Pi_{kl}\log\!\left(\Pi_{kl}\right)$
+$\mathbf{\Pi}^{*}$ $={\arg\!\min}_{\mathbf{\Pi}\geq 0}\;\sum_{k,l}\Pi_{kl}C_{kl}+\epsilon\sum_{k,l}\Pi_{kl}(\log\!\left(\Pi_{kl}\right)-1)$
 and $\tilde{\Pi^{*}}$ be EOT-coupling in the case when cost is scaled by a normalization constant $c$ or $\tilde{C}_{ij}=\frac{C_{ij}}{c}$.
@@
 where $m$ and $s$ are constants, that depend on $\Pi^{*}$, marginalization constants $a,b$ and normalization constant $c$.
-Corollary 1 can also be interpreted as supporting the results of Proposition 1 and our intuition that normalizing has the same effect on the kernel matrix as increasing $\epsilon$, leading to more diffused couplings or couplings with increased entropy.
-Appendix D ContextFlow Algorithm
-Algorithm 1 ContextFlow (CTF): Flow Matching with Spatial-Context-Aware OT Couplings
-1:Input: gene data $\{\mathbf{X}_{t_{1}},\cdots,\mathbf{X}_{t_{m+1}}\}$, spatial data $\{\mathbf{S}_{t_{1}},\ldots,\mathbf{S}_{t_{m+1}}\}$, parameters $\lambda$, $\alpha$, $\epsilon$, $\sigma$, $\eta$
-2:Output: neural velocity vector field $u_{\theta}$
-3:Initialize $\theta$
-4:while training do
-5:  for $i=1,2,\ldots,m$ do
-6:   Sample a batch $\mathcal{B}=\{(\bm{x}_{i},\bm{x}_{i+1}):(\bm{x}_{i},\bm{x}_{i+1})\sim(\mathbf{X}_{t_{i}},\mathbf{X}_{t_{i+1}})\}$
-7:   Construct TPM: $\mathbf{M}_{i,i+1}(\mathcal{B})\triangleright$ $\mathbf{M}_{i,i+1}$ is defined in Equation 10
-8:   if “prior-aware cost matrix” then
-9:     $C_{kl}\leftarrow\alpha\cdot\|\bm{x}_{i}(k)-\bm{x}_{i+1}(l)\|_{2}^{2}+(1-\alpha)\cdot[\mathbf{M}_{i,i+1}]_{kl}$ for any pair $(k,l)$
-10:     $\mathbf{K}\leftarrow\exp(-\mathbf{C}/\epsilon)$
-11:   else if “prior-aware entropy regularization” then
-12:     $C_{kl}\leftarrow\|\bm{x}_{i}(k)-\bm{x}_{i+1}(l)\|_{2}^{2}$ for any pair $(k,l)$
-13:     $\mathbf{K}\leftarrow\widehat{\mathbf{M}}_{i,i+1}\odot\exp(-\mathbf{C}/\epsilon)$ $\triangleright$ $\widehat{\mathbf{M}}_{i,i+1}$ is defined in Equation 12
-14:   end if
-15:   Initialize $\bm{a}\leftarrow\frac{1}{n_{i}}\mathbf{1}_{n_{i}}$, $\bm{b}\leftarrow\frac{1}{n_{i+1}}\mathbf{1}_{n_{i+1}}$, $\bm{u}\leftarrow\mathbf{1}_{n_{i}}$, $\bm{v}\leftarrow\mathbf{1}_{n_{i+1}}$
-16:   while not converged do
-17:     $\bm{u}\leftarrow\bm{a}\oslash(\mathbf{K}\bm{v}),\>\bm{v}\leftarrow\bm{b}\oslash(\mathbf{K}^{\top}\bm{u})$ $\triangleright$ Run Sinkhorn algorithm
-18:   end while
-19:   Obtain spatial-prior-aware OT couplings $\mathbf{\Pi}^{\mathrm{CTF}}_{i,i+1}\leftarrow\operatorname{diag}(\bm{u})\mathbf{K}\operatorname{diag}(\bm{v})$
-20:   Sample $t\sim\mathcal{U}(t_{i},t_{i+1})$ and $\{(\bm{x}_{i},\bm{x}_{i+1}):(\bm{x}_{i},\bm{x}_{i+1})\sim\mathbf{\Pi}^{\mathrm{CTF}}_{i,i+1}\}$
-21:   Sample $\bm{x}_{t}\sim\mathcal{N}\left(\frac{t_{i+1}-t}{t_{i+1}-t_{i}}\bm{x}_{i}+\frac{t-t_{i}}{t_{i+1}-t_{i}}\bm{x}_{i+1},\sigma^{2}\mathbf{I}\right)$
-22:   $L_{\mathrm{CFM}}\leftarrow\frac{1}{|\mathcal{B}|}\sum_{t,(\bm{x}_{i},\bm{x}_{i+1})}\left\|u_{\theta}(\bm{x}_{t},t)-\frac{\bm{x}_{i+1}-\bm{x}_{i}}{t_{i+1}-t_{i}}\right\|_{2}^{2}$
-23:  end for
-24:  $\theta\leftarrow\theta-\eta\cdot\nabla_{\theta}L_{\mathrm{CFM}}$
-25:end while
-Appendix E Kantarovich-OT Formulation
-Kantorovich’s formulation (Peyré et al., 2019) is a classical definition of the optimal transport (OT) problem that seeks a joint coupling to move a probability measure to another that minimizes the Euclidean distance cost, corresponding to the following minimization problem with respect to the 2-Wasserstein distance:
-$\pi^{*}_{\mathrm{ot}}:={\arg\!\min}_{\pi\in\Pi(q_{0},q_{1})}\int_{\mathbb{R}^{d}\times\mathbb{R}^{d}}\|\bm{x}_{0}-\bm{x}_{1}\|_{2}^{2}\>d\pi(\bm{x}_{0},\bm{x}_{1}),$ (16)
-where $\Pi(q_{0},q_{1})$ denotes the set of joint probability measures such that the left and right marginals are $q_{0}$ and $q_{1}$.
-Equation 16 can be solved in a mini-batch fashion using standard solvers such as POT (Flamary et al., 2021); however, the computational complexity is cubic in batch size.
-Appendix F Evaluation Metrics
-F.1 2-Wasserstein
-The 2-Wasserstein distance ($\mathcal{W}_{2}$ between empirical distributions $\mu,\nu$ is defined as:
-$\mathcal{W}_{2}(\mu,\nu)=\inf_{\gamma\in\Pi(\mu,\nu)}\bigg(\sum_{(\bm{x},\bm{y})}\gamma(\bm{x},\bm{y})\cdot\|\bm{x}-\bm{y}\|_{2}^{2}\bigg)^{1/2},$
-where $\Pi(\mu,\nu)$ denotes the set of couplings between $\mu$ and $\nu$.
-F.2 Weighted 2-Wasserstein
-Implausible velocity fields can steer a cell’s transcriptional trajectory in unrealistic directions, potentially leading to entirely different terminal cell types.
-We thus employ the weighted 2-Wasserstein metric, which ensures the evaluation accounts for both transcriptional similarity and the distributional balance of cell types.
-We define the weighted 2-Wasserstein distance (Weighted $\mathcal{W}_{2}$) between true and predicted distributions as:
-$\text{Weighted-$\mathcal{W}_{2}$}(\mu,\nu)=\sum_{i=1}^{C}\frac{n_{i}^{\text{true}}}{N}\cdot\mathcal{W}_{2}\!\bigg(\frac{1}{n_{i}^{\text{true}}}\!\!\sum_{j:y_{j}=i}\delta_{\bm{x}_{j}},\frac{1}{n_{i}^{\text{pred}}}\!\!\sum_{j:\hat{y}_{j}=i}\delta_{\bm{x}_{j}}\bigg),$
-where $n_{i}^{\text{true}},n_{i}^{\text{pred}}$ are the number of true and predicted cells of type $i$, and $N$ is the total number of samples.
-To determine the cell type of generated trajectories, we employ a multi-class classifier $M_{\phi}$, implemented as an XGBoost model (Chen & Guestrin, 2016) trained for each dataset.
-F.3 Energy Distance
-Let $\mu$ and $\nu$ be probability distributions with samples $X=\{\bm{x}_{i}\}_{i=1}^{m}\sim\mu$ and $Y=\{\bm{y}_{j}\}_{j=1}^{n}\sim\nu$. The squared empirical energy distance (Energy) is defined as:
-$\mathrm{ED}(\mu,\nu)=\frac{2}{mn}\sum_{i=1}^{m}\sum_{j=1}^{n}\|\bm{x}_{i}-\bm{y}_{j}\|-\frac{1}{m^{2}}\sum_{i=1}^{m}\sum_{i^{\prime}=1}^{m}\|\bm{x}_{i}-\bm{x}_{i^{\prime}}\|-\frac{1}{n^{2}}\sum_{j=1}^{n}\sum_{j^{\prime}=1}^{n}\|\bm{y}_{j}-\bm{y}_{j^{\prime}}\|,$
-where $\|\cdot\|$ is the Euclidean norm. The distance is non-negative and equals zero if and only if $\mu=\nu$.
-F.4 Maximum Mean Discrepancy
-For the same samples, the unbiased empirical estimate of the squared maximum mean discrepancy (MMD) with kernel $\kappa$ is defined as:
-$\mathrm{MMD}(\mu,\nu;\kappa)=\frac{1}{m(m-1)}\sum_{i\neq i^{\prime}}\kappa(\bm{x}_{i},\bm{x}_{i^{\prime}})+\frac{1}{n(n-1)}\sum_{j\neq j^{\prime}}\kappa(\bm{y}_{j},\bm{y}_{j^{\prime}})-\frac{2}{mn}\sum_{i=1}^{m}\sum_{j=1}^{n}\kappa(\bm{x}_{i},\bm{y}_{j}).$
-In our evaluations, we use a multi-kernel variant with radial basis function (RBF) kernels $\kappa_{\gamma}(\bm{x},y)=\exp(-\gamma\|\bm{x}-\bm{y}\|^{2})$, and average over $\gamma\in[2,1,0.5,0.1,0.01,0.005]$.
-Appendix G Biological Priors and Dataset Visualizations
-G.1 (Im-)Plausibility of OT-Couplings
-To demonstrate the need for integrating biological priors within a generative framework, we computed the Entropic-OT plan (Section 2.3) for the MOTFM framework and the PAER-OT plan (Section 3.3) for the ContextFlow framework.
-From these transport plans, we sampled couplings corresponding to the first two stages of the Brain Regeneration dataset (Wei et al., 2022) together with their associated cell types.
-Figures 3(a) and 3(b) illustrate the Excitatory–Inhibitory lineage switches present in these sampled couplings.
-Since excitatory and inhibitory neurons have mutually exclusive neurotransmitter functions and originate from distinct progenitor populations with different transcription factor profiles, a transition from excitatory to inhibitory identity is considered biologically implausible.
-In our transport plan couplings, we observed the following cell type lineage switches:
-  • Immature MSN $\rightarrow$ Immature nptxEX
-  • Immature MSN $\rightarrow$ Immature dpEX
-  • Immature MSN $\rightarrow$ Immature CMPN
-  • Immature nptxEX $\rightarrow$ Immature cckIN
-  • Immature nptxEX $\rightarrow$ Immature MSN
+Corollary C.2 can also be interpreted as supporting the results of Proposition C.1 and our intuition that normalizing has the same effect on the kernel matrix as increasing $\epsilon$, leading to more diffused couplings or couplings with increased entropy.
+Appendix E Detailed Analysis of ContextFlow
+E.1 Algorithm Pseudocode
+Algorithm 1 ContextFlow: Flow Matching with Spatial-Context-Aware OT Couplings
+1:Input: gene data $\{\mathbf{X}_{t_{1}},\cdots,\mathbf{X}_{t_{m+1}}\}$, spatial data $\{\mathbf{S}_{t_{1}},\ldots,\mathbf{S}_{t_{m+1}}\}$, parameters $\lambda$, $\alpha$, $\epsilon$, $\sigma$, $r$, $\eta$
+2:Data-Preprocessing: Compute local neighborhood means using Nearest Neighbor Algorithm and Ligand-Receptor features $f_{\text{LR}}$ using LIANA+ $\triangleright$ As defined in Equation 9 and 10
+3:Output: neural velocity vector field $u_{\theta}$
+4:Initialize $\theta$
+5:while training do
+6:  for $i=1,2,\ldots,m$ do
+7:   Sample a batch $\mathcal{B}=\{(\bm{x}_{i},\bm{x}_{i+1}):(\bm{x}_{i},\bm{x}_{i+1})\sim(\mathbf{X}_{t_{i}},\mathbf{X}_{t_{i+1}})\}$
+8:   Construct TPM: $\mathbf{M}_{i,i+1}(\mathcal{B})\triangleright$ $\mathbf{M}_{i,i+1}$ is defined in Equation 11
+9:   if “prior-aware cost matrix” then
+10:     $C_{kl}\leftarrow\alpha\cdot\|\bm{x}_{i}(k)-\bm{x}_{i+1}(l)\|_{2}^{2}+(1-\alpha)\cdot[\mathbf{M}_{i,i+1}]_{kl}$ for any pair $(k,l)$
+11:     $\mathbf{K}\leftarrow\exp(-\mathbf{C}/\epsilon)$
+12:   else if “prior-aware entropy regularization” then
+13:     $C_{kl}\leftarrow\|\bm{x}_{i}(k)-\bm{x}_{i+1}(l)\|_{2}^{2}$ for any pair $(k,l)$
+14:     $\mathbf{K}\leftarrow\widehat{\mathbf{M}}_{i,i+1}\odot\exp(-\mathbf{C}/\epsilon)$ $\triangleright$ $\widehat{\mathbf{M}}_{i,i+1}$ is defined in Equation 13
+15:   end if
+16:   Initialize $\bm{a}\leftarrow\frac{1}{n_{i}}\mathbf{1}_{n_{i}}$, $\bm{b}\leftarrow\frac{1}{n_{i+1}}\mathbf{1}_{n_{i+1}}$, $\bm{u}\leftarrow\mathbf{1}_{n_{i}}$, $\bm{v}\leftarrow\mathbf{1}_{n_{i+1}}$
+17:   while not converged do
+18:     $\bm{u}\leftarrow\bm{a}\oslash(\mathbf{K}\bm{v}),\>\bm{v}\leftarrow\bm{b}\oslash(\mathbf{K}^{\top}\bm{u})$ $\triangleright$ Run Sinkhorn algorithm
+19:   end while
+20:   Obtain spatial-prior-aware OT couplings $\mathbf{\Pi}^{\mathrm{CTF}}_{i,i+1}\leftarrow\operatorname{diag}(\bm{u})\mathbf{K}\operatorname{diag}(\bm{v})$
+21:   Sample $t\sim\mathcal{U}(t_{i},t_{i+1})$ and $\{(\bm{x}_{i},\bm{x}_{i+1}):(\bm{x}_{i},\bm{x}_{i+1})\sim\mathbf{\Pi}^{\mathrm{CTF}}_{i,i+1}\}$
+22:   Sample $\bm{x}_{t}\sim\mathcal{N}\left(\frac{t_{i+1}-t}{t_{i+1}-t_{i}}\bm{x}_{i}+\frac{t-t_{i}}{t_{i+1}-t_{i}}\bm{x}_{i+1},\sigma^{2}\mathbf{I}\right)$
+23:   $L_{\mathrm{CFM}}\leftarrow\frac{1}{|\mathcal{B}|}\sum_{t,(\bm{x}_{i},\bm{x}_{i+1})}\left\|u_{\theta}(\bm{x}_{t},t)-\frac{\bm{x}_{i+1}-\bm{x}_{i}}{t_{i+1}-t_{i}}\right\|_{2}^{2}$
+24:  end for
+25:  $\theta\leftarrow\theta-\eta\cdot\nabla_{\theta}L_{\mathrm{CFM}}$
+26:end while
+E.2 Time Complexity Analysis
+The training time of ContextFlow is comparable to that of Minibatch-OT FM (Tong et al., 2024) (Figure 3), as both methods solve an entropic variant of optimal transport using the GPU-optimized Sinkhorn algorithm, alongside forward and backward propagation steps that are also GPU-accelerated.
+Although ContextFlow incorporates prior knowledge, such as spatial smoothness (Equation 9) and cell-cell communication patterns (Equation 10), their corresponding features are computed once during preprocessing, resulting in a one-time cost.
+The precomputed features can be reused across multiple hyperparameter settings and model variants, making ContextFlow highly scalable and efficient.
+Data Preprocessing.
+The following preprocessing steps generate additional biologically informed features that complement the original transcriptomic profiles.
+These features incur a one-time computational cost and can be reused across different experiments and model configurations.
+Spatial Smoothness (SS).
+We employ a nearest neighbor (NN) algorithm for calculating the mean of local transcriptomic features for each cell.
+The computational complexity of the NN search is known to be $O(N^{2}d)$, where $N$ denotes the total number of points considered and $d$ denotes the data dimension.
+Cell-Cell Communication Patterns (LR).
+We employ spatially informed bivariate statistics implemented in LIANA+ (Dimitrov et al., 2024), for computing LR features, where we applied the cosine similarity metric to gene expression profiles and used the recommended hyperparameters.
+The exact runtime complexity for LIANA+ is unknown.
+Table 6 summarizes the total time taken for the Brain Regeneration and Mouse Organogenesis datasets in our case.
+Table 6: Runtime for computing cell-cell communication patterns.
+Dataset Total Number of Cells Runtime (seconds)
+Brain Regeneration (Wei et al., 2022) 28,780 23.35
+Mouse Organogenesis (Chen et al., 2022) 399,248 200.40
+Training of ContextFlow.
+The training time complexity largely depends on the total number of training epochs ($E$), mini-batch size ($B$), time per forward and backward pass ($P$), transcriptomic feature dimension ($d$), and total LR pairs ($l$).
+Below, we compute the time complexity for each step in ContextFlow:
+TPM Construction. The construction of the transition plausibility matrix involves a calculation of pairwise distances for each mini-batch, resulting in the runtime of $O(B^{2}(l+d))$.
+Sinkhorn Algorithm.
+According to Theorem 3.1, we know that Sinkhorn iterations can be adapted to solve the prior-aware entropy regularization problem (Equation 13).
+Since the Sinkhorn algorithm has a well-known quadratic time complexity (Cuturi, 2013), the runtime for computing minibatch OT couplings in ContextFlow is $O(B^{2})$.
+Figure 3: Training time comparisons for different algorithms with a mini-batch size of $256$ on the Brain Regeneration dataset.
+Total Runtime.
+Putting pieces together, across all the training epochs, the total runtime complexity of ContextFlow turns out to be $O(E\times(B^{2}(d+l)+P))$.
+As shown in Figure 3, the runtime is linearly dependent on the total epochs $E$, with different linear rates for various configurations.
+CFM is the fastest because it bypasses the optimal transport coupling step required by the other methods.
+Appendix F Spatiotemporal Optimal Transport
+In this section, we compare the recent state-of-the-art spatiotemporal alignment methods, including DeST-OT (Halmos et al., 2025) and TOAST (Ceccarelli et al., 2025), with our prior-aware entropy regularized (PAER) OT objective used in ContextFlow (CTF-H).
+It is important to note that these OT methods are not generative models and are only used for pairwise alignment tasks.
+ContextFlow, on the other hand, is a generative model that learns a dynamic flow across the time horizon and utilizes OT couplings to design better conditional paths for regression.
+In particular, we compute metrics described in DeST-OT on the Axolotl Brain Regeneration dataset following the same setup as used in flow matching.
+Specifically, for each time step, we randomly sample a batch of $1000$ cells and compute the corresponding coupling matrix $\mathbf{\Pi}$, which is then used to derive the metrics.
+We use the CTF-H ($\lambda=0.8$) version of ContextFlow for comparison.
+F.1 Metric Comparison
+DeST-OT introduces an OT objective for aligning spatial transcriptomic tissue slices from different developmental timesteps, with an emphasis on modeling cell growth and tissue expansion/contraction.
+The growth distortion metric is designed to assess whether the inferred growth pattern aligns with the changes in cell-type abundance across timesteps.
+As shown in Table 7, for the growth distortion metric, we find that our CTF-H OT is competitive with DeST-OT and TOAST, despite DeST-OT being specifically developed with consideration for cell growth.
+Table 7: Comparison on growth distortion.
+Static Pair DeST-OT TOAST CTF-H OT
+$1/2$ $0.0007$ $0.0000$ $0.0000$
+$2/3$ $0.0042$ $0.0000$ $0.0000$
+$3/4$ $0.0027$ $0.0000$ $0.0000$
+$4/5$ $0.0009$ $0.0000$ $0.0000$
+The migration metric is another important metric introduced in DeST-OT, which measures whether the coupling results in realistic cell movements between timesteps.
+As seen in Table 8, DeST-OT achieves the best performance, highlighting the advantage of its growth-aware objective compared to TOAST and CTF-H OT, which do not explicitly model tissue expansion or contraction.
+Table 8: Comparison on migration.
+Static Pair DeST-OT TOAST CTF-H OT
+$1/2$ $308.97$ $793.71$ $780.16$
+$2/3$ $551.53$ $1103.29$ $1037.29$
+$3/4$ $1015.73$ $2052.65$ $1991.43$
+$4/5$ $777.13$ $2257.02$ $2222.07$
+Lastly, we compute how similar the transcriptomic values of coupled cells are using a Coupled Transcriptomic Distance metric, which is defined as $\sum_{k=1}^{N}\sum_{l=1}^{M}\big\|X_{t_{i}}[k,:]-X_{t_{i+1}}[l,:]\big\|^{2}\times\Pi_{i,j}$, where $X_{t_{i}}[k,:]$ represents the transcriptomic feature of cell $k$ from timestep $t_{i}$ and $X_{t_{i+1}}[l,:]$ represents the transcriptomic feature of cell $l$ from timestep $t_{i+1}$, and $\Pi$ is the OT coupling matrix.
+From Table 9, we can observe that CTF-H OT is competitive with both DeST and TOAST.
+Table 9: On coupled transcriptomic distance.
+Static Pair DeST-OT TOAST CTF-H OT
+$1/2$ $32.64$ $34.13$ $33.58$
+$2/3$ $14.22$ $21.87$ $18.01$
+$3/4$ $42.26$ $43.89$ $43.26$
+$4/5$ $18.47$ $20.54$ $20.14$
+F.2 Runtime Analysis with Varying Sample Size
+We also compare the runtime complexity of the above-mentioned OT methods, as shown in Table 10.
+CTF-H OT is the fastest of the three, followed by DeST-OT and TOAST, and is competitive in the metrics above.
+We also observe that DeST-OT is the slowest, as expected, since its OT objective includes a Gromov-Wasserstein term, which has $O(n^{3})$ runtime, along with other growth- and tissue-distortion-specific terms.
+Table 10: Runtime (s) with varying sample size.
+Sample Size DeST-OT TOAST CTF-OT
+$10$ $0.0111$ $0.0227$ $0.0009$
+$50$ $0.0342$ $0.0114$ $0.0018$
+$100$ $0.1892$ $0.0167$ $0.0024$
+$150$ $0.4035$ $0.0247$ $0.0053$
+$200$ $0.4913$ $0.0337$ $0.0059$
+$250$ $0.4571$ $0.0426$ $0.0061$
+$300$ $0.6252$ $0.0543$ $0.0074$
+$350$ $0.6974$ $0.0656$ $0.0078$
+$400$ $0.8612$ $0.0817$ $0.0117$
+$450$ $1.1028$ $0.0983$ $0.0110$
+$500$ $1.5478$ $0.1197$ $0.0107$
+$550$ $1.7448$ $0.1468$ $0.0165$
+$600$ $1.9295$ $0.1905$ $0.0209$
+$650$ $2.1282$ $0.2077$ $0.0201$
+$700$ $2.7013$ $0.2309$ $0.0235$
+$750$ $3.2951$ $0.2574$ $0.0327$
+$800$ $4.0964$ $0.3001$ $0.0382$
+$850$ $4.2001$ $0.3229$ $0.0339$
+$900$ $4.4582$ $0.3798$ $0.0483$
+$950$ $5.0965$ $0.4206$ $0.0509$
+$1000$ $6.1452$ $0.4931$ $0.0375$
+Tables 7-10 demonstrate that the design choices of ContextFlow enable it to be highly scalable compared to state-of-the-art spatiotemporal alignment methods, while remaining competitive across several spatiotemporal OT alignment metrics.
 (a) MOTFM Implausible Transitions
@@
 Our formulation produces substantially fewer biologically implausible couplings (24) compared to MOTFM (54).
-Of these, $54$ implausible transitions arose from the Entropic-OT plan compared to the $24$ under the PAER-OT plan, with the specific transitions detailed in the figure legends.
+Appendix G Biological Priors and Dataset Visualizations
+G.1 (Im-)Plausibility of OT-Couplings
+To demonstrate the need of integrating biological priors within a generative framework, we computed the Entropic-OT plan (Section 2.3) for the MOTFM framework and the PAER-OT plan (Section 3.3) for the ContextFlow framework.
+From these transport plans, we sampled couplings corresponding to the first two stages of the Brain Regeneration dataset (Wei et al., 2022) together with their associated cell types.
+Figures 4(a) and 4(b) illustrate the Excitatory–Inhibitory lineage switches present in these sampled couplings.
+Since excitatory and inhibitory neurons have mutually exclusive neurotransmitter functions and originate from distinct progenitor populations with different transcription factor profiles, a transition from excitatory to inhibitory identity is considered biologically implausible.
+In our transport plan couplings, we observed the following cell type lineage switches: (i) Immature MSN $\rightarrow$ Immature nptxEX, (ii) Immature MSN $\rightarrow$ Immature dpEX, (iii) Immature MSN $\rightarrow$ Immature CMPN, (iv) Immature nptxEX $\rightarrow$ Immature cckIN, and (v) Immature nptxEX $\rightarrow$ Immature MSN
+Of these, $54$ implausible transitions arose from the Entropic-OT plan, compared to $24$ from the PAER-OT plan, with the specific transitions detailed in the figure legends.
 We also observed that the Entropic-OT formulation produced implausible transitions across brain hemispheres, for example, coupling cells from the left hemisphere with those from the right.
@@
 G.2 Cell type distributions over time
-Figures 46 present the spatial maps of the transcriptomics datasets across different time points, illustrating how tissue organization and cell type distributions evolve during development and regeneration.
+Figures 57 present spatial maps of transcriptomic datasets across different time points, illustrating how tissue organization and cell-type distributions evolve during development and regeneration.
 These maps highlight not only changes in cellular composition but also the preservation of spatial neighborhoods and geometrical arrangements of specific cell types over time.
@@
 These transitions are biologically plausible, as they preserve cell type identity within excitatory neuronal lineages while reflecting maturation within the same functional context.
-This example highlights the richness of the contextual information captured by our proposed biological prior, and demonstrates how incorporating such ligand–receptor–driven cues into the coupling process leads to more interpretable and biologically consistent trajectories.
-Figure 8: Visual translation of the bias that NPTX2–NPTXR LR pattern provides in terms of cell type coupling for the two consecutive slides.
+This example highlights the richness of the contextual information captured by our proposed biological prior and demonstrates how incorporating such ligand–receptor–driven cues into the coupling process yields more interpretable, biologically consistent trajectories.
+Figure 9: Visual translation of the bias that NPTX2–NPTXR LR pattern provides in terms of cell type coupling for two consecutive slides.
 Appendix H Additional Experiments & Ablations
@@
 CTF-H 0 $3.244\pm 0.713$ $3.946\pm 1.671$ $0.089\pm 0.005$ $\bm{8.797\pm 0.612}$
-11 $\bm{2.814\pm 0.414}$ $\bm{3.233\pm 0.567}$ $0.093\pm 0.005$ $10.319\pm 0.817$
-$0.5$ $5.200\pm 0.799$ $6.306\pm 1.037$ $0.123\pm 0.008$ $45.862\pm 13.765$
+11 $5.200\pm 0.799$ $6.306\pm 1.037$ $0.123\pm 0.008$ $45.862\pm 13.765$
+$0.5$ $\bm{2.814\pm 0.414}$ $\bm{3.233\pm 0.567}$ $0.093\pm 0.005$ $10.319\pm 0.817$
 Table 18: Extrapolation via IVP Sampling at holdout time 8 for the Mouse Organogenesis dataset.
@@
 Overall, ContextFlow captures biologically coherent and temporally consistent developmental dynamics.
+Appendix I Hyperparameter Sensitivity Analysis
+I.1 Ablation on $\lambda$
+First, we conduct a sensitivity analysis on all three datasets for the trade-off parameter $\lambda$, which controls the relative importance of spatial smoothness (SS) and cell-cell communication (LR) priors.
+All the other hyperparameters are kept constant.
+From Tables 20-24 and their corresponding Figures 11-15, we observe that the best performance is usually achieved towards the extremities, at $\lambda=0$ or $\lambda=1$, with values near the latter dominating more often.
+We hypothesize that the SS prior (at $\lambda=1$ only the SS prior is considered), acting as a proxy for spatial distance between cells from different slices, always carries relevant information, encoding the structural information present in the data.
+On the other hand, the informativeness of cell–cell communication patterns depends on how distinct the ligand–receptor features are at a given time step compared with those of neighboring ones.
+When LR features remain highly similar across consecutive timesteps, they contribute little to the discriminative signal that the OT objective can leverage.
+Consequently, the influence of communication priors is strongly dataset- and timestep-dependent.
+This effect is also reflected in the observation that settings with $\lambda=0$ (where only the LR prior is used) tend to perform worse when using Next-Step sampling—where local, immediate effects dominate—than under IVP sampling, which integrates information over the entire preceding trajectory.
+We therefore recommend experimenting with different values of $\lambda$ (e.g., 0, $0.8$, or $1$) depending on the specific use case and context.
+Due to the scalability of ContextFlow, hyperparameter exploration can be performed efficiently, allowing for a rapid assessment of the effect of $\lambda$ on model performance.
+I.2 Ablation on $r$
+We conduct an ablation study on the Brain Regeneration dataset, examining the effect of varying the neighborhood radius $r$ used to define the boundary for computing the Spatial Smoothness Prior.
+We evaluate two settings: $\lambda=1$ and $\lambda=0.8$, corresponding to the use of only the spatial prior and to a setting with a modest contribution from the cell–cell communication prior, respectively.
+From Tables 25-28 and their corresponding Figures 16-19, we observe that the optimal neighborhood radius tends to lie toward the lower end of the tested range.
+Radii smaller than this optimum degrade performance by failing to capture sufficient local context, resulting in neighborhood means that are overly similar to individual cellular profiles.
+Conversely, increasing the radius beyond the optimal range also reduces performance, as the neighborhood begins to include cells from distinct types or spatial regions, thereby diluting the local signal.
+While certain deviations from this trend occur, likely reflecting underlying biological complexity, this behavior is consistent with the trade-off between spatial specificity and contextual coverage inherent to neighborhood-based priors.
+For our case, we set the radius by considering the timestep with the least number of cells, dividing it by half (to account for different hemispheres), and dividing by the order of cell types present in that timestep.
+For the dataset considered in this study, Stage 44 had the fewest number of $1400$ cells, with approximately $10$ cell types present.
+We thus set the radius at $50$ in our studies.
+I.3 Ablation on $\epsilon$
+Additionally, we conduct an ablation study on the Brain Regeneration dataset by varying the parameter $\epsilon$, which is used to weigh the entropic term in the OT objective.
+The ContextFlow configuration we consider here is CTF-H ($\lambda=1$), which includes only the spatial smoothness prior.
+As observed from Tables 29-30 and their corresponding Figures 20-21, drastically decreasing $\epsilon$ results in the OT formulation to ignore the relative entropic term containing the prior information and only to consider the transport cost resulting in higher Wasserstein values.
+Furthermore, in accordance with results from Theorem 3.1, increasing $\epsilon$ too much still does not drastically degrade the performance, as the prior matrix $M$ acts as a soft filter and prohibits uniform couplings.
+While setting $\epsilon$ values, one must look at the Gibbs kernel used in the Sinkhorn Algorithm $\exp(-\mathbf{C}/\epsilon)$, since lower $\epsilon$ values can cause potential numerical issues.
+We thus set $\epsilon$ by examining the order of the cost matrix $\mathbf{C}$, and for the studies above, we set it to $100$ after considering the order of the median of all the elements in the cost matrix.
+Table 20: Extrapolation on the last holdout timestep on the Brain Regeneration dataset.
+$\bm{\lambda}$ Next Step Sampling IVP Sampling
+Weighted $\bm{\mathcal{W}_{2}}$ $\bm{\mathcal{W}_{2}}$ Weighted $\bm{\mathcal{W}_{2}}$ $\bm{\mathcal{W}_{2}}$
+0 $6.968\pm 0.608$ $7.198\pm 0.726$ $6.243\pm 0.760$ $6.220\pm 0.751$
+0.2 $7.313\pm 0.384$ $7.331\pm 0.467$ $6.502\pm 0.634$ $6.039\pm 0.733$
+0.5 $7.243\pm 0.479$ $7.157\pm 0.641$ $6.254\pm 0.819$ $5.973\pm 0.757$
+0.8 $7.333\pm 0.605$ $7.334\pm 0.622$ $6.598\pm 0.892$ $6.402\pm 1.039$
+1 $7.505\pm 0.667$ $7.338\pm 0.601$ $5.277\pm 0.936$ $6.021\pm 1.192$
+(a) Next Step Sampling
+(b) IVP Sampling
+Figure 11: Performance variation with $\lambda$ for extrapolation on the Brain Regeneration dataset.
+Table 21: Interpolation on the middle holdout timestep 3 on the Brain Regeneration dataset.
+$\bm{\lambda}$ Next Step Sampling IVP Sampling
+Weighted $\bm{\mathcal{W}_{2}}$ $\bm{\mathcal{W}_{2}}$ Weighted $\bm{\mathcal{W}_{2}}$ $\bm{\mathcal{W}_{2}}$
+0 $2.528\pm 0.143$ $2.534\pm 0.180$ $3.925\pm 0.267$ $4.375\pm 0.297$
+0.2 $2.544\pm 0.093$ $2.389\pm 0.183$ $4.153\pm 0.432$ $4.393\pm 0.369$
+0.5 $2.519\pm 0.167$ $2.412\pm 0.158$ $3.917\pm 0.343$ $4.159\pm 0.455$
+0.8 $2.533\pm 0.137$ $2.352\pm 0.142$ $151\pm 0.193 4.4$$408\pm 0.285$
+1 $2.316\pm 0.141$ $1.969\pm 0.221$ $3.905\pm 0.395$ $4.188\pm 0.685$
+(a) Next Step Sampling
+(b) IVP Sampling
+Figure 12: Performance variation with $\lambda$ for interpolation on the Brain Regeneration dataset.
+Table 22: Interpolation for the holdout timestep 5 on the Mouse Organogenesis dataset.
+$\bm{\lambda}$ Next Step Sampling IVP Sampling
+Weighted $\bm{\mathcal{W}_{2}}$ $\bm{\mathcal{W}_{2}}$ Weighted $\bm{\mathcal{W}_{2}}$ $\bm{\mathcal{W}_{2}}$
+0 $1.884\pm 0.027$ $1.862\pm 0.123$ $3.244\pm 0.713$ $3.946\pm 1.671$
+0.2 $1.896\pm 0.028$ $1.899\pm 0.078$ $2.990\pm 0.205$ $3.273\pm 0.518$
+0.5 $1.871\pm 0.030$ $1.919\pm 0.067$ $2.814\pm 0.414$ $3.233\pm 0.567$
+0.8 $1.878\pm 0.031$ $1.890\pm 0.064$ $2.966\pm 0.411$ $3.345\pm 0.508$
+1 $1.898\pm 0.029$ $1.866\pm 0.097$ $5.200\pm 0.799$ $6.306\pm 1.037$
+(a) Next Step Sampling
+(b) IVP Sampling
+Figure 13: Performance variation with $\lambda$ for interpolation on the Mouse Organogenesis dataset.
+Table 23: Extrapolation for holdout timestep 5 on the Mouse Organogenesis dataset.
+$\bm{\lambda}$ Next Step Sampling IVP Sampling
+Weighted $\bm{\mathcal{W}_{2}}$ $\bm{\mathcal{W}_{2}}$ Weighted $\bm{\mathcal{W}_{2}}$ $\bm{\mathcal{W}_{2}}$
+0 $1.508\pm 0.047$ $1.386\pm 0.088$ $6.33\times 10^{5}\pm 1.72\times 10^{6}$ $3.95\times 10^{6}\pm 1.05\times 10^{7}$
+0.2 $1.614\pm 0.081$ $1.642\pm 0.136$ $9.18\times 10^{2}\pm 1.69\times 10^{3}$ $4.29\times 10^{4}\pm 1.05\times 10^{5}$
+0.5 $1.638\pm 0.069$ $1.676\pm 0.114$ $1.08\times 10^{5}\pm 2.76\times 10^{5}$ $9.95\times 10^{5}\pm 2.77\times 10^{6}$
+0.8 $1.617\pm 0.042$ $80\pm 0.094 1.66×1$$6\times 10^{5}\pm 3.92\times 10^{5}$ $1.43\times 10^{6}\pm 3.43\times 10^{6}$
+1 $1.906\pm 0.071$ $1.892\pm 0.092$ $1.89\times 10^{1}\pm 1.77\times 10^{1}$ $6.89\times 10^{1}\pm 7.36\times 10^{1}$
+(a) Next Step Sampling
+(b) IVP Sampling
+Figure 14: Performance variation with $\lambda$ for extrapolation on the Mouse Organogenesis dataset.
+Table 24: Interpolation for holdout timestep 3 with IVP Sampling on the Liver Regeneration dataset.
+$\bm{\lambda}$ $\bm{\mathcal{W}_{2}}$
+0 $32.682\pm 1.472$
+$0.2$ $34.647\pm 1.461$
+$0.5$ $33.414\pm 0.995$
+$0.8$ $33.512\pm 0.786$
+11 $33.481\pm 1.001$
+Figure 15: Performance variation with $\lambda$ for interpolation on the Liver Regeneration dataset.
+Table 25: Extrapolation with CTF-H at $\lambda=1$ (only using the spatial smoothness prior) for the last holdout timestep on the Brain Regeneration dataset.
+Radius Next Step Sampling IVP Sampling
+Weighted $\bm{\mathcal{W}_{2}}$ $\bm{\mathcal{W}_{2}}$ Weighted $\bm{\mathcal{W}_{2}}$ $\bm{\mathcal{W}_{2}}$
+12 $6.228\pm 1.276$ $6.163\pm 1.490$ $4.415\pm 0.580$ $6.843\pm 4.812$
+25 $6.244\pm 1.066$ $6.231\pm 1.043$ $6.500\pm 1.751$ $5.613\pm 1.561$
+50 $7.505\pm 0.667$ $7.338\pm 0.601$ $5.277\pm 0.936$ $6.021\pm 1.192$
+100 $6.892\pm 0.930$ $6.702\pm 0.631$ $7.061\pm 1.677$ $6.860\pm 1.880$
+150 $747\pm 0.923 7.7$$793\pm 0.934$ $9.796\pm 3.847$ $10.656\pm 6.591$
+200 $6.039\pm 0.282$ $5.764\pm 0.272$ $5.630\pm 0.793$ $5.000\pm 0.735$
+250 $6.804\pm 1.011$ $6.834\pm 1.124$ $6.578\pm 1.611$ $7.379\pm 2.864$
+(a) Next Step Sampling
+(b) IVP Sampling
+Figure 16: Performance variation with radius for extrapolation on the Brain Regeneration dataset.
+Table 26: Interpolation for the middle holdout timestep 3 for CTF-H at $\lambda=1$ (only using the spatial smoothness prior) on the Brain Regeneration dataset.
+Radius Next Step Sampling IVP Sampling
+Weighted $\bm{\mathcal{W}_{2}}$ $\bm{\mathcal{W}_{2}}$ Weighted $\bm{\mathcal{W}_{2}}$ $\bm{\mathcal{W}_{2}}$
+12 $4.293\pm 0.318$ $3.547\pm 0.343$ $2.650\pm 0.204$ $2.346\pm 0.251$
+25 $5.019\pm 0.270$ $3.968\pm 0.274$ $2.408\pm 0.239$ $1.808\pm 0.257$
+50 $2.316\pm 0.141$ $1.969\pm 0.221$ $3.905\pm 0.395$ $4.188\pm 0.685$
+100 $4.590\pm 0.360$ $3.359\pm 0.166$ $812\pm 0.240 2.2$$220\pm 0.231$
+150 $4.731\pm 0.424$ $3.819\pm 0.239$ $3.533\pm 0.220$ $3.290\pm 0.778$
+200 $4.548\pm 0.780$ $4.249\pm 1.315$ $3.751\pm 0.725$ $3.677\pm 1.016$
+250 $4.768\pm 1.994$ $4.782\pm 4.129$ $4.281\pm 0.985$ $4.103\pm 1.081$
+(a) Next Step Sampling
+(b) IVP Sampling
+Figure 17: Performance variation with radius for interpolation on the Brain Regeneration dataset.
+Table 27: Extrapolation for the last holdout timestep with CTF-H at $\lambda=0.8$ on the Brain Regeneration dataset.
+Radius Next Step Sampling IVP Sampling
+Weighted $\bm{\mathcal{W}_{2}}$ $\bm{\mathcal{W}_{2}}$ Weighted $\bm{\mathcal{W}_{2}}$ $\bm{\mathcal{W}_{2}}$
+12 $6.924\pm 1.178$ $6.831\pm 1.132$ $8.049\pm 2.162$ $13.223\pm 12.164$
+25 $6.633\pm 0.780$ $350\pm 0.654 6.6$$621\pm 1.608$ $9.223\pm 8.396$
+50 $7.130\pm 0.389$ $7.260\pm 0.632$ $5.971\pm 0.461$ $5.836\pm 1.181$
+100 $6.411\pm 0.522$ $6.351\pm 0.456$ $5.932\pm 0.264$ $6.434\pm 0.840$
+150 $6.498\pm 1.056$ $6.501\pm 1.098$ $6.033\pm 0.882$ $7.203\pm 2.443$
+200 $6.052\pm 0.873$ $6.129\pm 1.052$ $5.852\pm 1.085$ $6.247\pm 1.731$
+250 $6.449\pm 0.909$ $6.278\pm 0.726$ $6.151\pm 0.986$ $11.261\pm 7.063$
+(a) Next Step Sampling
+(b) IVP Sampling
+Figure 18: Performance variation with radius for extrapolation on the Brain Regeneration dataset.
+Table 28: Interpolation for the middle holdout timestep 3 with CTF-H at $\lambda=0.8$ on the Brain Regeneration dataset.
+Radius Next Step Sampling IVP Sampling
+Weighted $\bm{\mathcal{W}_{2}}$ $\bm{\mathcal{W}_{2}}$ Weighted $\bm{\mathcal{W}_{2}}$ $\bm{\mathcal{W}_{2}}$
+12 $5.320\pm 1.714$ $4.709\pm 2.260$ $3.722\pm 1.114$ $3.656\pm 1.327$
+25 $943\pm 1.384 4.4$$467\pm 1.821$ $3.350\pm 1.548$ $3.112\pm 1.418$
+50 $2.440\pm 0.090$ $2.302\pm 0.137$ $4.181\pm 0.035$ $4.238\pm 0.068$
+100 $4.028\pm 0.648$ $3.417\pm 0.869$ $2.956\pm 0.580$ $2.678\pm 0.535$
+150 $5.408\pm 0.889$ $4.669\pm 1.364$ $4.535\pm 0.823$ $4.209\pm 0.884$
+200 $7.110\pm 2.581$ $6.490\pm 3.543$ $4.043\pm 1.441$ $3.754\pm 1.350$
+250 $4.502\pm 0.573$ $3.689\pm 1.204$ $3.532\pm 1.148$ $3.457\pm 1.217$
+(a) Next Step Sampling
+(b) IVP Sampling
+Figure 19: Performance variation with radius for interpolation on the Brain Regeneration dataset
+Table 29: Extrapolation for the last holdout timestep on the Brain Regeneration dataset.
+$\bm{\epsilon}$ Next Step Sampling IVP Sampling
+Weighted $\bm{\mathcal{W}_{2}}$ $\bm{\mathcal{W}_{2}}$ Weighted $\bm{\mathcal{W}_{2}}$ $\bm{\mathcal{W}_{2}}$
+0.001 $6.007\pm 0.516$ $5.939\pm 0.286$ $6.260\pm 1.123$ $7.301\pm 2.935$
+0.01 $6.240\pm 0.870$ $6.254\pm 1.111$ $6.396\pm 0.236$ $7.231\pm 0.968$
+0.1 $6.579\pm 0.744$ $6.861\pm 0.845$ $6.758\pm 1.826$ $7.283\pm 2.068$
+1 $5.648\pm 0.471$ $5.721\pm 0.595$ $6.010\pm 0.674$ $5.905\pm 0.737$
+10 $6.841\pm 0.597$ $6.940\pm 0.671$ $5.532\pm 1.775$ $6.646\pm 1.926$
+100 $7.166\pm 0.991$ $7.094\pm 1.148$ $6.455\pm 3.047$ $5.650\pm 1.928$
+1000 $6.291\pm 1.041$ $6.300\pm 1.052$ $7.382\pm 2.553$ $7.626\pm 3.204$
+10000 $587\pm 0.805 6.6$$641\pm 1.083$ $5.754\pm 0.741$ $7.546\pm 3.599$
+(a) Next Step Sampling
+(b) IVP Sampling
+Figure 20: Performance variation with $\epsilon$ for extrapolation on the Brain Regeneration Dataset.
+Table 30: Interpolation for the middle holdout timestep on the Brain Regeneration dataset.
+$\bm{\epsilon}$ Next Step Sampling IVP Sampling
+Weighted $\bm{\mathcal{W}_{2}}$ $\bm{\mathcal{W}_{2}}$ Weighted $\bm{\mathcal{W}_{2}}$ $\bm{\mathcal{W}_{2}}$
+0.001 $2.899\pm 0.582$ $2.715\pm 0.653$ $4.056\pm 0.542$ $3.286\pm 0.289$
+0.01 $4.520\pm 2.066$ $4.589\pm 2.298$ $6.915\pm 3.573$ $7.125\pm 5.289$
+0.1 $2.573\pm 0.476$ $2.472\pm 0.507$ $3.772\pm 0.642$ $3.046\pm 0.537$
+1 $2.865\pm 0.612$ $2.785\pm 0.576$ $4.255\pm 0.679$ $3.355\pm 0.584$
+10 $2.899\pm 0.865$ $2.833\pm 0.984$ $4.908\pm 1.130$ $4.159\pm 1.526$
+100 $2.338\pm 0.101$ $1.835\pm 0.171$ $5.069\pm 0.985$ $4.322\pm 1.461$
+1000 $3.104\pm 0.663$ $2.321\pm 0.521$ $5.109\pm 0.948$ $3.974\pm 1.227$
+10000 $2.838\pm 0.281$ $2.176\pm 0.315$ $4.557\pm 0.710$ $3.373\pm 0.833$
+(a) Next Step Sampling
+(b) IVP Sampling
+Figure 21: Performance variation with $\epsilon$ for interpolation on the Brain Regeneration dataset.

### v3

@@
 Abstract
-Inferring trajectories from longitudinal spatially-resolved omics data is fundamental to understanding the dynamics of structural and functional tissue changes in development, regeneration and repair, disease progression, and response to treatment.
-We propose ContextFlow, a novel context-aware flow matching framework that incorporates prior knowledge to guide the inference of structural tissue dynamics from spatially resolved omics data.
+Inferring trajectories from longitudinal, spatially resolved omics data is fundamental to understanding the dynamics of structural and functional tissue changes in development, regeneration and repair, disease progression, and response to treatment.
+We propose ContextFlow, a context-aware flow matching framework that leverages prior knowledge to guide the inference of structural tissue dynamics from spatially resolved omics data.
+Existing approaches rely on optimal-transport-derived couplings that are determined solely by transcriptomic similarity, which can lead to biologically implausible transitions and, in turn, inconsistent learned dynamics.
 Specifically, ContextFlow integrates local tissue organization and ligand-receptor communication patterns into a transition plausibility matrix that regularizes the optimal transport objective.
-By embedding these contextual constraints, ContextFlow generates trajectories that are not only statistically consistent but also biologically meaningful, making it a generalizable framework for modeling spatiotemporal dynamics from longitudinal, spatially resolved omics data.
-Comprehensive experiments show that ContextFlow consistently outperforms state-of-the-art flow matching methods across multiple metrics of inference accuracy and biological coherence.
+By embedding these contextual constraints, ContextFlow generates trajectories that are not only statistically consistent but also biologically meaningful, making it a generalizable framework for modeling cellular dynamics from longitudinal, spatially resolved omics data.
+Comprehensive experiments show that ContextFlow consistently improves state-of-the-art flow matching methods across multiple metrics of inference accuracy and biological coherence.
 Our code is available at: https://github.com/santanurathod/ContextFlow.
-Machine Learning, ICML
 1 Introduction
-Flow matching (Lipman et al., 2023) is an emerging generative paradigm that provides an efficient approach for learning the complex latent dynamics, or normalizing flows (Papamakarios et al., 2021), of a system of variables, while enabling parametric flexibility to model data distributions.
-Inferring the underlying dynamics from sparse and noisy observations is a central challenge in many application domains (Gontis et al., 2010; Brunton et al., 2016; Pandarinath et al., 2018; Li et al., 2025), where continuous trajectories are rarely captured; instead, cross-sectional snapshots, collected at discrete time points, are typically available.
-In single-cell RNA sequencing (scRNA-seq), this challenge becomes especially critical as the destructive nature of profiling technologies yields only unpaired population-level snapshots over time.
+Inferring the underlying dynamics from sparse and noisy observations is a central challenge in many applications (Gontis et al., 2010; Brunton et al., 2016; Pandarinath et al., 2018; Li et al., 2025), where continuous trajectories are rarely captured; instead, cross-sectional snapshots, collected at discrete time points, are available.
+In single-cell RNA sequencing, this challenge becomes especially critical, as the destructive nature of profiling technologies yields only unpaired population-level snapshots over time.
 Uncovering temporal dynamics from such snapshot data is essential for understanding developmental processes, disease progression, treatment, and perturbation responses (Wagner and Klein, 2020).
-Traditional approaches often rely on heuristics or computationally intensive likelihood-based generative models, which struggle with scalability and flexibility in high-dimensional single-cell data.
-In contrast, flow matching overcomes these challenges by directly learning continuous latent dynamics that are constrained to match observed population-level distributions at sampled time points.
-Table 1: Comparisons of state-of-the-art spatiotemporal OT and flow matching methods.
-In particular, ContextFlow excels at efficiently inferring biologically plausible trajectories by incorporating spatial priors into minibatch-OT-based flow-matching frameworks.
-Method Generative Prior Knowledge Unified Velocity Field OT Runtime
-DeST-OT (Halmos et al., 2025) $O(N^{3})$
-TOAST (Ceccarelli et al., 2025) $O(N^{3})$
-PASTE (Zeira et al., 2022) $O(N^{3})$
-CFM (Lipman et al., 2022)
-MOTFM (Tong et al., 2024) $O(N^{2})$
-ContextFlow (ours) $O(N^{2})$
-The state and function of cells within a tissue are affected by interactions with neighboring cells, extracellular matrix components, and local signaling gradients (Rao et al., 2021).
+Traditional methods (Trapnell et al., 2017; Chen et al., 2018; Rubanova et al., 2019; Bergen et al., 2020) rely on heuristics or computationally intensive likelihood-based generative models, which struggle with flexibility and scalability in handling high-dimensional data.
+In contrast, flow matching (Lipman et al., 2023), an emerging generative paradigm, overcomes these issues by directly learning continuous latent dynamics constrained to match population-level distributions at observational time points, while enabling parametric flexibility to model data distributions.
+Cellular dynamics are affected by the tissue microenvironment, in which cells engage in reciprocal communication with their neighbors (Dimitrov et al., 2022; Tanevski et al., 2025).
+State-of-the-art flow matching frameworks (Tong et al., 2024; Kapuśniak et al., 2024; Atanackovic et al., 2024; Rohbeck et al., 2025) use optimal transport (OT) to derive couplings between consecutive observational time points and design conditional paths for regression over the velocity fields.
+Although OT couplings improve training stability and inference efficiency over vanilla independent couplings (Lipman et al., 2023), they are determined solely by transcriptomic similarity and do not account for the contextual richness of spatial transcriptomics (Armingol et al., 2021; Mayer et al., 2023; Aguadé-Gorgorió et al., 2024; Zhang et al., 2025b), thus yielding couplings that are statistically optimal but biologically implausible.
+Since OT-based flow matching algorithms directly regress on these couplings, these errors propagate into the learned velocity field, leading to globally inconsistent inferred dynamics.
+To address these limitations, we introduce ContextFlow, a generative framework that incorporates explicit spatial priors into OT-coupled flow matching to model temporal tissue dynamics (see Figure 2 for an illustration).
+In contrast to existing approaches, ContextFlow integrates prior knowledge directly into the construction of OT couplings to avoid biologically implausible transitions as the regression target.
+By encoding local tissue organization and ligand–receptor-derived spatial communication patterns into prior-regularized optimal transport formulations, ContextFlow exploits the contextual richness of spatial omics data and embeds both structural and functional aspects of tissue organization into the training objective, thereby generating more biologically informed trajectories.
+In summary, our contributions are as follows:
+  • We define transitional plausibility to capture local tissue organization and ligand–receptor communication patterns, which provides a flexible and interpretable mechanism for encoding spatial and functional priors into couplings used to train the velocity field (Section 4.1).
+Our definition allows priors to be explicitly and flexibly adapted in a domain-specific manner, enabling the incorporation of external biological knowledge and improving interpretability.
+  • We introduce principled methods for integrating explicit biological priors into the state-of-the-art OT-based flow matching framework, instantiated via two novel schemes—cost-based and entropy-based—both amenable to efficient Sinkhorn optimization (Section 4.2).
+  • Comprehensive experiments on regeneration and developmental datasets show that ContextFlow consistently improves existing methods for both interpolation and extrapolation tasks across metrics that capture biological plausibility and statistical fidelity (Section 5).
+2 Related Work
+2.1 Flow Matching and its Applications in Biology
+Flow matching (FM) (Lipman et al., 2023; Albergo and Vanden-Eijnden, 2022; Liu et al., 2022) is a generative modeling framework that trains the velocity field via a tractable regression objective for learning continuous normalizing flows (Papamakarios et al., 2021), which is more efficient and scalable when dealing with high-dimensional data than prior approaches, such as neural ODE (Chen et al., 2018; Rubanova et al., 2019).
+Compared with score-based diffusion models (Song et al., 2020) and Schrödinger Bridges (SB) (Bunne et al., 2023), FM is faster and more stable to train and sample from during inference, as it characterizes the latent continuous dynamics via an ODE.
+Therefore, a growing body of literature has adapted and improved the vanilla FM framework to model dynamic scientific processes, among others, especially in biology and the life sciences (Li et al., 2025).
+Instead of independent couplings, (Pooladian et al., 2023; Tong et al., 2024) used optimal transport (OT) maps to define the conditional coupling between the source and the target, and approximated it using minibatch samples, leading to more stable training and faster inference via straightening the flows.
+Based on this framework,  (Kapuśniak et al., 2024) introduced metric flow matching to learn the flows over a non-Euclidean manifold to make a case against straighter flows in the context of transcriptomics.
+(Atanackovic et al., 2024) proposed meta flow matching, which incorporates implicit cell-cell interaction prior via GNN conditioning.
+Apart from these trajectory inference-centric studies, FM has also been utilized to generate imaging-based cell morphology changes (Zhang et al., 2025a), simulate spatial transcriptomics data from histology images (Huang et al., 2025), and approximate OT maps for drug response modeling and cross-modal translation tasks Klein et al.
+(2024).
+Despite these advances, existing work does not address how to meaningfully incorporate explicit biological prior knowledge to constrain the velocity field, thereby limiting the biological plausibility of inferred trajectories, interpretability, and control over the learned dynamics.
+2.2 Advancement in Spatial Transcriptomics
+The state and function of cells within a tissue are influenced by interactions with neighboring cells, extracellular matrix components, and local signaling gradients (Rao et al., 2021).
 Recent advances in spatial omics technologies, particularly spatial transcriptomics (ST), allow gene expression profiling without tissue dissociation, thereby preserving spatial context and providing a complementary view of cellular organization.
-The dynamics of complex cellular processes are affected by the tissue microenvironment, where cells engage in reciprocal communication with their neighbors (Dimitrov et al., 2022; Tanevski et al., 2025).
-A growing body of work highlights the critical role of spatial cell–cell communication patterns in shaping cellular phenotypes (Armingol et al., 2021).
+The dynamics of cellular processes are regulated by the tissue microenvironment and by cell communications with their neighbours (Dimitrov et al., 2022; Tanevski et al., 2025).
+A growing body of literature highlights the critical role of spatial cell–cell communication patterns in shaping cellular phenotypes (Armingol et al., 2021).
 In particular, location-specific communication circuits between distinct cell types dynamically interact to reprogram cellular states and influence tissue-level behavior (Mayer et al., 2023; Aguadé-Gorgorió et al., 2024; Zheng et al., 2025).
 These insights, made possible by the spatiotemporal resolution of transcriptomics data, pave the way for understanding the mechanisms by which cellular interactions drive tissue organization and function in organogenesis (Chen et al., 2022), regeneration (Ben-Moshe et al., 2021; Wei et al., 2022), disease progression (Kukanja et al., 2024), and treatment response (Liu et al., 2024).
-Optimal transport (OT) has become a foundational framework for aligning spatially resolved samples and inferring putative developmental or temporal couplings  (Zeira et al., 2022; Liu et al., 2023).
-As a result, state-of-the-art flow matching frameworks such as minibatch-OT flow matching (MOTFM) (Tong et al., 2024) use OT-derived couplings to define conditional paths for training velocity fields, thereby overcoming the lack of generative capabilities in optimal transport.
-The OT formulation adopted in MOTFM, however, does not account for the contextual richness of spatial transcriptomics and can yield trajectories that are statistically optimal yet biologically implausible (see Figure 4(a) in Appendix G.1 for an illustration).
-While recent studies (Halmos et al., 2025; Ceccarelli et al., 2025) have extended OT objectives to spatial transcriptomics, they primarily focus on pairwise alignment of populations across conditions or modalities and do not explicitly incorporate the cell–cell communication patterns that drive cellular state transitions.
-Figure 1: ContextFlow integrates local tissue organization and ligand-receptor communications to learn biologically meaningful trajectories from spatial omics data.
-Prior knowledge acts as a soft filter, discouraging implausible transitions while preserving flexibility in trajectory inference.
-To address the above limitations, we introduce a novel flow matching-based framework, ContextFlow, that incorporates spatial priors to model temporal tissue dynamics (Figure 1).
-By encoding local tissue organization and ligand-receptor-derived spatial communication patterns into prior-regularized optimal transport formulations, ContextFlow fully exploits the contextual richness of spatial omics data and embeds both structural and functional aspects of tissue organization into its objective, thereby generating more biologically informed trajectories.
-Table 1 highlights the key advantages of ContextFlow, compared with the aforementioned existing methods, where detailed discussions on related works are postponed to Appendix A.
-In summary, our contributions are as follows:
-  • We leverage the structure of local tissue organization and ligand–receptor communication to extract biologically meaningful features from spatial omics data and encode them into an informed transition plausibility matrix to constrain temporal dynamics (Section 3.2).
-  • We design two novel integration schemes—cost-based and entropy-based—that incorporate prior knowledge into an OT-coupled flow matching framework, both amenable to efficient Sinkhorn optimization and scalable on modern hardware (Section 3.3).
-  • Comprehensive experiments on three regeneration and developmental datasets demonstrate that ContextFlow consistently outperforms baseline flow matching methods under both interpolation and extrapolation settings across multiple evaluation metrics that capture biological plausibility and statistical fidelity (Section 4).
-2 Preliminaries
-2.1 Flow Matching Basics
-Flow matching (Lipman et al., 2023) is a simulation-free and sample-efficient generative framework for training continuous normalizing flows (Chen et al., 2018).
-Given a pair of source and target data distributions over $\mathbb{R}^{d}$ with densities $q_{0}=q(\bm{x}_{0})$ and $q_{1}=q(\bm{x}_{1})$, the problem task is to learn a time-varying velocity vector field $u_{\theta}:[0,1]\times\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$, whose continuous evolution is captured by a function in the form of a neural-net-based model with weights $\theta$, that can transform $q_{0}$ to $q_{1}$ through integration via an ordinary differential equation (ODE).
-To be more specific, flow matching (FM) seeks to optimize $\theta$ by minimizing a simple regression loss between $u_{\theta}$ and a target time-varying velocity vector field $u_{t}:[0,1]\times\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$ as follows:
-$\min_{\theta}\>\mathbb{E}_{t\sim\mathcal{U}(0,1),\bm{x}\sim p_{t}(\bm{x})}\big\|u_{\theta}(t,\bm{x})-u_{t}(\bm{x})\big\|^{2}.$ (1)
-Here, $\mathcal{U}(0,1)$ is the uniform distribution over $[0,1]$, and $p_{t}:[0,1]\times\mathbb{R}^{d}\rightarrow\mathbb{R}_{+}$ denotes a time-varying probability path induced by $u_{t}$ such that (i) $p_{t}$ is a probability density function for any $t\in[0,1]$, (ii) $p_{t}$ satisfies the two boundary conditions: $p_{t=0}=q_{0}$ and $p_{t=1}=q_{1}$, and (iii) the connection between $p_{t}$ and $u_{t}$ can be characterized by the transport equation (Villani and others, 2008): $\frac{\partial p_{t}(\bm{x})}{\partial t}=-\nabla\cdot(u_{t}(\bm{x})p_{t}(\bm{x}))$, where $\nabla$ is the divergence operator.
-From a dynamical system’s view, $u_{t}$ defines an ODE system $d\bm{x}=u_{t}(\bm{x})dt$.
-The corresponding solution to the ODE, usually termed as the probability flow, can then transport any $\bm{x}_{0}\sim q_{0}$ to a point $\bm{x}_{1}\sim q_{1}$ along $u_{t}$ from $t=0$ to $t=1$.
-While the flow matching objective in Equation 1 is simple and intuitive, it is generally intractable in practice: the closed-form velocity vector field $u_{t}$ is unknown for arbitrary source and target distributions ($q_{0}$ and $q_{1}$), and multiple valid probability paths $p_{t}$ may exist between them.
-2.2 Conditional Flow Matching
-The central idea of conditional flow matching is to express the target probability path via a mixture of more manageable conditional probability paths (Lipman et al., 2023).
-By marginalizing over a conditioning variable $z$, both $p_{t}$ and $u_{t}$ can be constructed using their conditional counterparts:
-$p_{t}(\bm{x})$ $=\int p_{t}(\bm{x}|\bm{z})q(\bm{z})d\bm{z},$ (2)
-$u_{t}(\bm{x})$ $=\int u_{t}(\bm{x}|\bm{z})\frac{p_{t}(\bm{x}|\bm{z})q(\bm{z})}{p_{t}(\bm{x})}d\bm{z},$
-where $q(\bm{z})$ stands for the distribution of the conditioning variable $\bm{z}$, and the conditional probability path $p_{t}(\bm{x}|\bm{z})$ is constructed such that the boundary conditions are satisfied: $\int p_{t=0}(\bm{x}|\bm{z})q(\bm{z})=q_{0}$ and $\int p_{t=1}(\bm{x}|\bm{z})q(\bm{z})=q_{1}$.
-Theorem 1 of Lipman et al.
-(2023) proves that $p_{t}$ and $u_{t}$ defined by Equation 2 satisfy the transport equation, suggesting that $p_{t}$ is a valid probability path generated by $u_{t}$.
-To avoid the intractable integrals, (Lipman et al., 2023) proposed the following optimization of conditional flow matching (CFM), and proved its equivalence to the original FM objective in terms of gradient computation:
-$\min_{\theta}\>\mathbb{E}_{t\sim\mathcal{U}(0,1),z\sim q(\bm{z}),\bm{x}\sim p_{t}(\bm{x}|\bm{z})}\big\|u_{\theta}(t,\bm{x})-u_{t}(\bm{x}|\bm{z})\big\|^{2}.$ (3)
-By choosing an appropriate conditional velocity vector field $u_{t}(\bm{x}|\bm{z})$, we can train the neural network using Equation 3 without requiring a closed-form solution of the conditional probability path $p_{t}(\bm{x}|\bm{z})$, thereby avoiding the intractable integration operation.
-Therefore, the remaining task is to define the conditional probability path and velocity vector field properly such that we can sample from $p_{t}(\bm{x}|\bm{z})$ and compute $u_{t}(\bm{x}|\bm{z})$ efficiently for solving the optimization problem in Equation 3.
-A specific choice proposed in Lipman et al. (2023) is Gaussian conditional probability paths and their corresponding conditional velocity vector fields, defined as follows:
-$p_{t}(\bm{x}|\bm{z})$ $=\mathcal{N}(\bm{x}\>|\>\mu_{t}(\bm{z}),\sigma_{t}(\bm{z})^{2}\mathbf{I}),$ (4)
-$u_{t}(\bm{x}|\bm{z})$ $=\frac{\sigma_{t}^{\prime}(\bm{z})}{\sigma_{t}(\bm{z})}(\bm{x}-\mu_{t}(\bm{z}))+\mu_{t}^{\prime}(\bm{z}),$ (5)
-where $\mu_{t}:[0,1]\times\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$ stands for the time-varying mean of the Gaussian distribution, $\mu_{t}^{\prime}$ is its derivative with respect to time, $\sigma_{t}:[0,1]\times\mathbb{R}^{d}\rightarrow\mathbb{R}_{+}$ is the time-varying scalar standard deviation, and $\sigma_{t}^{\prime}$ denotes its derivative.
-In particular, Lipman et al.
-(2023) selected $q(\bm{z})=q(\bm{x}_{1})$, $\mu_{t}(\bm{z})=t\bm{x}_{1}$, and $\sigma_{t}(\bm{z})=1-(1-\sigma)t$.
-Then, we can see that $u_{t}(\bm{x}|\bm{z})$ transports the standard Gaussian distribution $p_{t=0}(\bm{x}|\bm{z})=\mathcal{N}(\bm{x};\bm{0},\mathbf{I})$ to a Gaussian distribution with mean $\bm{x}_{1}$ and covariance $\sigma^{2}\mathbf{I}$, namely $p_{t=1}(\bm{x}|\bm{z})=\mathcal{N}(\bm{x};\bm{x}_{1},\sigma^{2}\mathbf{I})$ for any target point $\bm{x}_{1}$.
-By letting $\sigma\rightarrow 0$, the marginal boundary conditions can easily be validated.
-Tong et al.
-(2024) further generalized the application scope to arbitrary source distributions by setting:
-$q(\bm{z})=q(\bm{x}_{0})q(\bm{x}_{1}),\mu_{t}(\bm{z})=(1-t)\bm{x}_{0}+t\bm{x}_{1},\sigma_{t}(\bm{z})=\sigma.$ (6)
-Note that the choice in Equation 6 satisfies the boundary conditions $p_{t=0}(\bm{x})=q_{0}$ and $p_{t=1}(\bm{x})=q_{1}$ when $\sigma\rightarrow 0$.
-According to Equation 4, the conditional velocity vector field has a simple analytical form $u_{t}(\bm{x}|\bm{z})=\bm{x}_{1}-\bm{x}_{0}$.
-2.3 Flow Matching with Optimal Transport Couplings
-The conditionals construction specified by Equation 6 corresponds to the simplest choice of independent coupling, where $\bm{z}=(\bm{x}_{0},\bm{x}_{1})$ with source $\bm{x}_{0}$ and target $\bm{x}_{1}$ are independently sampled from $q(\bm{z})=q(\bm{x}_{0})q(\bm{x}_{1})$.
-The use of couplings for constructing the sampling paths in the CFM framework naturally connects to the optimal transport theory (Villani and others, 2008).
-Choosing OT-based couplings has several advantages over independent coupling, including smaller training variance and more efficient sampling (Pooladian et al., 2023; Tong et al., 2024).
-Since the classical Kantorovich’s formulation has computational complexity that is cubic in terms of the sample size (see Appendix B.1 for its formal definition), a popular alternative is to add an extra regularization term, resulting in entropic optimal transport (EOT), to approximately solve the optimal transport problem while reducing the computational costs from cubic to quadratic:
-${\min}_{\pi\in\Pi(q_{0},q_{1})}\int_{\mathbb{R}^{d}\times\mathbb{R}^{d}}$ $\|\bm{x}_{0}-\bm{x}_{1}\|_{2}^{2}\>d\pi(\bm{x}_{0},\bm{x}_{1})$ (7)
-$\qquad\qquad+\epsilon H(\pi\>|\>q_{0}\otimes q_{1}),$
-where $\epsilon>0$ is the regularization parameter, and $H(\pi\>|\>q_{0}\otimes q_{1})$ denotes the relative entropy (or Kullback-Leibler divergence) with respect to $\pi$ and the product measure $q_{0}\otimes q_{1}$.
-For simplicity, we use $\pi^{*}_{\mathrm{eot}}(\epsilon)$ to denote the optimal solution to the minimization problem in Equation 7.
-Note that this optimization problem can also be viewed as a special case of the static Schrödinger bridge problem (Bernton et al., 2022), which can be efficiently solved in a mini-batch fashion via the Sinkhorn algorithm (Cuturi, 2013).
-Theoretically, one can prove that $\pi^{*}_{\mathrm{eot}}(\epsilon)$ recovers the Kantorovich’s OT coupling $\pi^{*}_{\mathrm{ot}}$ when $\epsilon\rightarrow 0$ (see Equation 14 in Appendix B.1 for its formal definition) and $\pi^{*}_{\mathrm{eot}}(\epsilon)$ corresponds to the independent coupling $q_{0}\otimes q_{1}$ when $\epsilon\rightarrow\infty$.
+Optimal transport (Bunne et al., 2024; Klein et al., 2025) has been increasingly used to analyze omics data, as they often exhibit uncoupled measurements across modalities, conditions, or time points.
+In ST, several OT formulations have been introduced depending on context.
+For instance, Zeira et al.
+(2022) and Liu et al.
+(2023) proposed PASTE and PASTE2 to align ST data from adjacent tissue slices, while DeST-OT (Halmos et al., 2025) integrates spatio-temporal slices by modeling cell growth and differentiation.
+In addition, Rahimi et al.
+(2024) developed DOT, a multi-objective OT framework for mapping features across scRNA-seq and spatially resolved assays, and Ceccarelli et al.
+(2026) introduced TOAST, a spatially regularized OT framework for slice alignment and annotation transfer.
+While these OT techniques can successfully leverage spatial or structural information to improve alignment quality across space, time, and modalities, they do not learn a velocity field governing temporal evolution, as in generative modeling frameworks like flow matching.
+Consequently, while these techniques can model temporal relationships and align distributions across time, they do not learn a continuous-time generative model and therefore cannot directly simulate intermediate or future cellular states.
+Moreover, although some of the aforementioned approaches incorporate spatial or structural constraints, these are typically introduced in a task-specific or implicit manner, rather than as general mechanisms to encode explicit biological priors directly into the transport plan.
+3 Problem Formulation
+We aim to infer the temporal dynamics of cell states from spatially resolved gene expression data.
+Let $0=t_{1}<t_{2}<\ldots<t_{m+1}=1$ be a sequence of normalized time points and $q_{i}$ be the underlying distribution over $\mathbb{R}^{d}$ at time $t_{i}$.
+Given $\{\mathbf{X}_{t_{i}}\}_{i\in\{1,2,\ldots,m+1\}}$, where $\mathbf{X}_{t_{i}}=\{\bm{x}_{i}(k)\}_{k=1}^{n_{i}}$ corresponds to the gene expressions at time $t_{i}$ consisting of $n_{i}$ snapshot data sampled from $q_{i}$, our goal is to learn a neural velocity vector field $u_{\theta}:[0,1]\times\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$ to faithfully characterize the temporal evolution of spatially resolved tissues over time, such that the induced probability path $p_{t}$ can describe the state of each cell at any time $t\in[0,1]$.
+A standard approach is to apply conditional flow matching with minibatch OT couplings (MOTFM) (Tong et al., 2024) for each pair of consecutive time points.
+Specifically, for any $t\in[0,1]$ satisfying $t\in[t_{i},t_{i+1})$, the following linear conditional velocity field $u_{t}(\bm{x}|\bm{z})$ and its associated probability path $p_{t}(\bm{x}|\bm{z})$ are defined as the regression target in MOTFM:
+$u_{t}(\bm{x}|\bm{z}_{i})=\frac{\bm{x}_{i+1}-\bm{x}_{i}}{t_{i+1}-t_{i}}\quad\text{and}\quad p_{t}(\bm{x}|\bm{z}_{i})=\mathcal{N}\left(\frac{t_{i+1}-t}{t_{i+1}-t_{i}}\bm{x}_{i}+\frac{t-t_{i}}{t_{i+1}-t_{i}}\bm{x}_{i+1},\sigma^{2}\mathbf{I}\right),$ (1)
+where $\bm{z}_{i}=(\bm{x}_{i},\bm{x}_{i+1})$.
+Let $q(\bm{z}_{i})$ be the joint probability measure with marginals $q_{i}$ and $q_{i+1}$, corresponding to the entropic optimal transport (EOT) coupling, defined as:
+$q(\bm{z}_{i})=\operatorname*{arg\,min}_{\pi\in\Pi(q_{i},q_{i+1})}\int_{\mathbb{R}^{d}\times\mathbb{R}^{d}}\|\bm{x}_{0}-\bm{x}_{1}\|_{2}^{2}\>d\pi(\bm{x}_{0},\bm{x}_{1})+\epsilon H(\pi\>|\>q_{i}\otimes q_{i+1}),$ (2)
+where $\epsilon>0$ is the regularization parameter, and $H(\pi\>|\>q_{i}\otimes q_{i+1})$ is the KL divergence with respect to the transport plan $\pi$ and the product measure $q_{i}\otimes q_{i+1}$.
+To train the velocity vector field $u_{\theta}$, MOTFM randomly samples a mini-batch of data at each time, uses the Sinkhorn algorithm (Cuturi, 2013) to obtain the EOT couplings, and optimizes the following regression loss using SGD:
+$\min_{\theta}\>\mathbb{E}_{t\sim\mathcal{U}(0,1),\bm{z}\sim q(\bm{z}_{i}),\bm{x}\sim p_{t}(\bm{x}|\bm{z}_{i})}\big\|u_{\theta}(t,\bm{x})-u_{t}(\bm{x}|\bm{z}_{i})\big\|^{2},$ (3)
+where $\mathcal{U}(0,1)$ is the uniform distribution over $[0,1]$, and $i=i(t)$ denotes the integer that satisfies $t_{i}\leq t<t_{i+1}$ for each sampled $t$.
+Full preliminaries of flow matching are provided in Appendix B.
+Despite their enhanced ability to model system dynamics, state-of-the-art OT-based flow matching frameworks lack provisions to fully exploit the contextual richness and integrate biological prior knowledge inferred from other associated data modalities.
+Existing approaches can generate statistically optimal trajectories by targeting probability paths induced by OT couplings along the temporal dimension.
+However, because they overlook important functional or structural prior information, their velocity fields can be trained on couplings that violate domain principles, leading to biologically implausible trajectories.
+Figure 1 presents an example of statistically optimal yet biologically implausible couplings between inhibitory-excitatory neuron cells coming from the EOT plan used to train MOTFM (Tong et al., 2024) (see Appendix G.1 for detailed discussions).
+This motivates our work to develop a mechanism to integrate domain prior knowledge into the OT-based flow matching framework, thereby encouraging domain-aligned training of the velocity field.
+Figure 1: Visualization of inhibitory-exitatory coupled cells sampled from the EOT plan used in MOTFM on temporally consecutive spatial omics slides from the Brain Regeneration dataset (Wei et al., 2022).
+Here, different colors correspond to different cell types.
+Since inhibitory/excitatory cells serve distinct biological functions, their coupling and thus the implied state transition are infeasible.
 4 Regularizing the Flow with Spatial Priors
-3.1 Problem Formulation
-We focus on the task of inferring spatiotemporal trajectories, i.e., inferring the dynamic evolution of the cell states across time from spatially resolved gene expression data.
-Let $0=t_{1}<t_{2}<\ldots<t_{m+1}=1$ be a sequence of normalized time points.For any $i\in\{1,2,\ldots,m+1\}$, let $q_{i}$ be the data distribution over $\mathbb{R}^{d}$ at time point $t_{i}$.
-Given $\{\mathbf{X}_{t_{i}}\}_{i\in\{1,2,\ldots,m+1\}}$, where $\mathbf{X}_{t_{i}}=\{\bm{x}_{i}(k)\}_{k=1}^{n_{i}}$ is the gene expressions at time $t_{i}$ consisting of $n_{i}$ snapshot data sampled from $q_{i}$, the objective is to learn a neural velocity vector field $u_{\theta}:[0,1]\times\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$ to faithfully characterize the temporal evolution of spatially resolved tissues over time, such that the induced probability path $p_{t}$ can describe the state of each cell at time $t\in[0,1]$.
-This task can be viewed as a continuous temporal generalization of the pairwise generative modeling task described in Section 2.1.
-A promising candidate solution is conditional flow matching with entropic OT couplings (EOT-CFM), by targeting linear conditional velocity vector fields for each pair of consecutive time points.
-Specifically, for any $t\in[0,1]$ satisfying $t\in[t_{i},t_{i+1}]$, the conditional velocity field $u_{t}(\bm{x}|\bm{z})$ and its associated probability path $p(\bm{x}|\bm{z})$ are defined as:
-$p(\bm{x}|\bm{z})=\mathcal{N}\left(\frac{t_{i+1}-t}{t_{i+1}-t_{i}}\bm{x}_{i}+\frac{t-t_{i}}{t_{i+1}-t_{i}}\bm{x}_{i+1},\sigma^{2}\mathbf{I}\right),$
-$u_{t}(\bm{x}|\bm{z})=\frac{\bm{x}_{i+1}-\bm{x}_{i}}{t_{i+1}-t_{i}},$ (8)
-where the conditioning variable $\bm{z}=(\bm{x}_{i},\bm{x}_{i+1})$, and $p(\bm{z})$ is the joint probability measure with marginals $q_{i}$ and $q_{i+1}$ corresponding to the EOT coupling $\pi^{*}_{\mathrm{eot}}(\epsilon)$ defined in Equation 7.
-It can be easily verified that the above construction satisfies the boundary condition at each time point $t_{i}$.
-To train $u_{\theta}$, we can randomly sample a mini-batch of data at each time, run the Sinkhorn algorithm (Cuturi, 2013) to obtain the entropic OT couplings for each consecutive pair, and iteratively update the model weights $\theta$ using stochastic gradient descent with CFM regression loss (Equation 3).
-Despite their enhanced ability to model system dynamics, state-of-the-art OT-CFM frameworks lack provisions to fully exploit the contextual richness and to integrate biological prior knowledge inferred from other associated data modalities.
-Existing approaches can generate statistically optimal trajectories by targeting probability paths induced by (entropic) OT couplings along the temporal dimension.
-However, they may overlook important functional or structural prior information, leading to biologically implausible trajectories (see Figure 4(a) in Appendix G.1 for an illustration).
 4.1 Spatial Priors and Transitional Plausibility
-To faithfully model the context and cellular organization of spatial omics data, we introduce two types of spatial priors and explain how they relate to the plausibility of transitions between locations and cell states across time points.
+To address the limitations, we propose modeling the context and cellular organization of spatial omics data using two types of spatial priors (spatial smoothness and cell-cell communication patterns), strong indicators of transitional plausibility between locations and cell states across time points.
 Spatial Smoothness.
@@
 Due to tissue heterogeneity, we cannot assume a common reference coordinate frame across tissue samples or even slices at $t_{i}$ and $t_{j}$ at a larger scale.
-However, the same heterogeneity allows us to consider the spatial coherence and neighborhood consistency (Greenwald et al., 2024; Ceccarelli et al., 2025) as a proxy for relative cell localization, which cannot change significantly across short time intervals.
-Therefore the aggregate expression within the microenvironment of each cell can be used to quantify the transitional plausibility in consecutive time points.
+However, the same heterogeneity allows us to consider spatial coherence and neighborhood consistency (Greenwald et al., 2024; Ceccarelli et al., 2026) as proxies for relative cell localization, which cannot change significantly over short time intervals.
+Therefore, the aggregate expression within each cell’s microenvironment can be used to quantify the plausibility of transition across consecutive time points.
 Specifically, let $c_{i}=(\bm{x}_{i},\bm{s}_{i})$ and $c_{j}=(\bm{x}_{j},\bm{s}_{j})$ be cells at time points $t_{i}$ and $t_{j}$, respectively, where $\bm{x}_{i},\bm{x}_{j}\in\mathbb{R}^{d}$ denote their gene expression profiles, and $\bm{s}_{i},\bm{s}_{j}\in\mathbb{R}^{2}$ denote their spatial coordinates in the relative tissue reference frame.
@@
 Spatial smoothness suggests that $\mathrm{TP}(c_{i},c_{j})$ is inversely related to the difference between the average expression profiles of their local neighborhoods:
-$\mathrm{SS}(c_{i},c_{j})$ $=\bigg\|\frac{1}{|\mathcal{N}_{r}(c_{i})|}\sum_{c\in\mathcal{N}_{r}(c_{i})}\bm{x}(c)$ (9)
-$\qquad\qquad-\frac{1}{|\mathcal{N}_{r}(c_{j})|}\sum_{c\in\mathcal{N}_{r}(c_{j})}\bm{x}(c)\bigg\|_{2}^{2},$
-where $\mathcal{N}_{r}(c_{i})=\{c:\|\bm{s}(c)-\bm{s}(c_{i})\|_{2}\leq r\}$ denotes the set of neighboring cells of $c_{i}$ in the same tissue slice, $|\mathcal{N}_{r}(c_{i})|$ stands for the cardinality of $\mathcal{N}_{r}(c_{i})$, and $\bm{x}(c)$ is the gene expression profile with respect to cell $c$.
+$\mathrm{SS}(c_{i},c_{j})$ $=\bigg\|\frac{1}{|\mathcal{N}_{r}(c_{i})|}\sum_{c\in\mathcal{N}_{r}(c_{i})}\bm{x}(c)-\frac{1}{|\mathcal{N}_{r}(c_{j})|}\sum_{c\in\mathcal{N}_{r}(c_{j})}\bm{x}(c)\bigg\|_{2}^{2},$ (4)
+where $\mathcal{N}_{r}(c_{i})=\{c:\|\bm{s}(c)-\bm{s}(c_{i})\|_{2}\leq r\}$ denotes the set of neighboring cells of $c_{i}$ in the same tissue slice, $|\mathcal{N}_{r}(c_{i})|$ is the cardinality of $\mathcal{N}_{r}(c_{i})$, and $\bm{x}(c)$ is the gene expression profile of cell $c$.
 Cell-Cell Communication Patterns.
-Cell–cell communication (CCC) has a critical role in the regulation of numerous biological processes, including development, apoptosis, and the maintenance of homeostasis in health and disease (Armingol et al., 2024).
-A major type of CCC is ligand–receptor (LR) signaling, in which ligands expressed by one cell bind to cognate receptors on another, initiating intracellular cascades that ultimately affect the state of the cell (i.e., its expression profile) (Armingol et al., 2021).
-There are numerous databases of prior knowledge of ligand-receptor binding and computational methods that use these databases to systematically link gene expression with the activity of ligand-receptor-mediated communication.
+Cell–cell communication plays a critical role in regulating biological processes, including development, apoptosis, and homeostasis in health and disease (Armingol et al., 2024).
+A major type of cell-cell communication is ligand–receptor (LR) signaling, in which ligands expressed by one cell bind to cognate receptors on another, initiating intracellular cascades that ultimately affect the expression profiles of the cell (Armingol et al., 2021).
+Numerous databases of prior knowledge of LR binding exist, and computational methods use them to systematically link gene expression with the activity of ligand-receptor-mediated communication.
 Specifically, we represent each cell $c_{i}$ by a feature vector $f_{\mathrm{LR}}\in\mathbb{R}^{p}$, where each entry corresponds to one of $p$ possible ligand–receptor pairs and encodes the extent of $c_{i}$’s participation in communication through that pair.
-The transitional plausibility $\mathrm{TP}(c_{i},c_{j})$ between cells in different tissue slices is higher when they exhibit similar ligand-receptor communication patterns $f_{\mathrm{LR}}$ (see Figure 8 for an illustration).
+The transitional plausibility $\mathrm{TP}(c_{i},c_{j})$ between cells in different tissue slices is higher when they exhibit similar ligand-receptor communication patterns $f_{\mathrm{LR}}$ (see Figure 9 in Appendix G.3 for an illustration).
 Formally, we define $\mathrm{LR}(c_{i},c_{j})$, the dissimilarity between the ligand–receptor communication patterns in the microenvironments of cells $c_{i}$ and $c_{j}$, as:
 $\mathrm{LR}(c_{i},c_{j})=\left\|\>f_{\mathrm{LR}}(\mathcal{N}_{r}(c_{i}))-f_{\mathrm{LR}}(\mathcal{N}_{r}(c_{j}))\right\|_{2}^{2}.$ (5)
+Figure 2: ContextFlow integrates local tissue organization and ligand-receptor communications to learn biologically meaningful trajectories from spatial omics data.
+Prior knowledge acts as a soft filter, discouraging implausible transitions while preserving flexibility in trajectory inference.
 4.2 Flow Matching with Context-Aware OT Couplings
@@
 To begin with, we create a sequence of transitional plausibility matrices (TPMs) to encode the biological priors for each pair of consecutive time points.
-Specifically, denote by $\mathbf{M}_{i,i+1}\in\mathbb{R}^{n_{i}\times n_{i+1}}$ the TPM with respect to the set of cells measured at time $t_{i}$ and at time $t_{i+1}$, with size $n_{i}$ and $n_{i+1}$ respectively, where the $(k,l)$-th entry of $\mathbf{M}_{i,i+1}$ indicates how plausibly the $k$-th cell measured at $t_{i}$ will evolve to the $l$-th cell measured at $t_{i+1}$, defined as follows:
-$\big[\mathbf{M}_{i,i+1}\big]_{kl}=\lambda\cdot\mathrm{SS}$ $\left(c_{i}(k),c_{i+1}(l)\right)$ (11)
-$+$ $(1-\lambda)\cdot\mathrm{LR}\left(c_{i}(k),c_{i+1}(l)\right).$
+Specifically, denote by $\mathbf{M}_{i}\in\mathbb{R}^{n_{i}\times n_{i+1}}$ the TPM with respect to the set of cells measured at time $t_{i}$ and at time $t_{i+1}$, with size $n_{i}$ and $n_{i+1}$ respectively, where the $(k,l)$-th entry of $\mathbf{M}_{i}$ indicates how plausibly the $k$-th cell measured at $t_{i}$ will evolve to the $l$-th cell measured at $t_{i+1}$, defined as follows:
+$\big[\mathbf{M}_{i}\big]_{kl}=\lambda\cdot\mathrm{SS}$ $\left(c_{i}(k),c_{i+1}(l)\right)+(1-\lambda)\cdot\mathrm{LR}\left(c_{i}(k),c_{i+1}(l)\right).$ (6)
 Here, $\lambda\in[0,1]$ denotes a trade-off hyperparameter that balances the contribution of the spatial smoothness prior ($\mathrm{SS}$) and the ligand–receptor communication prior ($\mathrm{LR}$).
 Prior-Regularized OT Couplings.
-The transitional plausibility matrices capture our spatially informed prior over cell-cell transitions between consecutive time points, which can naturally be incorporated into the EOT formulation (Equation 7) to promote couplings that maintain the structural and functional properties of tissue organization.
-We propose two techniques for prior integration:
+The transitional plausibility matrix encodes our spatially informed prior over cell-cell transitions between consecutive time points, which can naturally be incorporated into the EOT formulation (Equation 2) to promote couplings that maintain the structural and functional properties of tissue organization.
+In particular, we propose two techniques for prior integration:
 Prior-Aware Cost Matrix (PACM).
-Consider the empirical counterpart of Equation 7 with respect to time $t_{i}$ and time $t_{i+1}$.
-Our first PACM approach incorporates the transitional plausibility matrix directly into the transport cost:
+Consider $(t_{i},t_{i+1})$ and the empirical counterpart of Equation 2.
+Our first approach incorporates the transitional plausibility matrix directly into the transport cost:
 $\min_{\mathbf{\Pi}\in\mathbb{R}^{n_{i}\times n_{i+1}}}\;\langle\mathbf{\Pi},\mathbf{C}_{i,i+1}\rangle-\epsilon\sum_{k,l}\Pi_{kl}(\log\Pi_{kl}-1),$ (7)
-where the $(k,l)$-th entry of the new prior-aware cost matrix is $\mathbf{C}_{i,i+1}]_{kl}=\alpha\|\bm{x}_{i}(k)-\bm{x}_{i+1}(l)\|_{2}^{2}+(1-\alpha)[\mathbf{M}_{i,i+1}]_{kl}$, the transport plan $\Pi$ needs to satisfy the boundary conditions: $\sum_{l}\Pi_{kl}=1/n_{i}$ for any $k\in[n_{i}]$, and $\sum_{k}\Pi_{kl}=1/n_{i+1}$ for any $l\in[n_{i+1}]$, and $\alpha\in[0,1]$ controls the trade-off between the original Euclidean cost and the prior-aware cost derived from the transitional plausibility.
-If $\mathbf{M}_{i,i+1}]_{kl}$ is high, Equation 12 will impose a higher transport cost between the $k$-cell at time $i$ to the $j$-cell at time $i+1$.
-This observation aligns with our assumption that such transitions should be more biologically implausible.
+where the $(k,l)$-th entry $\mathbf{C}_{i,i+1}]_{kl}=\alpha\|\bm{x}_{i}(k)-\bm{x}_{i+1}(l)\|_{2}^{2}+(1-\alpha)[\mathbf{M}_{i,i+1}]_{kl}$, the transport plan $\mathbf{\Pi}$ needs to satisfy the boundary conditions: $\sum_{l}\Pi_{kl}=1/n_{i}$ for any $k\in[n_{i}]$, and $\sum_{k}\Pi_{kl}=1/n_{i+1}$ for any $l\in[n_{i+1}]$, and $\alpha\in[0,1]$ controls the trade-off between the original Euclidean cost and the prior-aware cost derived from the transitional plausibility.
+If $\mathbf{M}_{i}]_{kl}$ is high, Equation 7 will impose a higher transport cost between the $k$-cell at time $i$ to the $j$-cell at time $i+1$.
+This observation aligns with our assumption that such transitions are less biologically plausible.
 Prior-Aware Entropy Regularization (PAER).
-While the previous PACM approach penalizes couplings according to our spatial priors, it defines a different OT problem with a modified cost function.
-Consequently, the standard interpretation of OT as minimizing the transport energy between two transcriptomic distributions no longer holds.
+While PACM penalizes couplings according to our spatial priors, it defines a different OT problem with a modified cost function.
+Thus, the standard interpretation of OT as minimizing the transport energy between two transcriptomic distributions no longer holds.
 Since the scales of the pairwise distances often differ, normalization of the cost terms is required to enable meaningful comparison.
-This normalization, however, may result in couplings that deviate from their original counterparts (see Proposition 1 and Corollary 1 in Appendix D).
+This normalization, however, may result in couplings that deviate from their original counterparts (see Theorem C.1 in Appendix C for a formal statement).
 Besides, selecting an appropriate $\alpha$ in Equation 7 introduces an additional layer of tuning, increasing computational overhead.
-Therefore, we propose the second PAER approach to integrate the biological priors without introducing additional hyperparameters:
-$\min_{\mathbf{\Pi}\in\mathbb{R}^{n_{i}\times n_{i+1}}}\sum_{k,l}\Pi_{kl}\|\bm{x}_{i}(k)-\bm{x}_{i+1}(l)\|_{2}^{2}$ (13)
-$\qquad\qquad-\epsilon\sum_{k,l}\underbrace{\Pi_{kl}\big(\log\big(\Pi_{kl}/[\widehat{\mathbf{M}}_{i,i+1}]_{kl}\big)-1\big)}_{\text{Prior-Aware Entropy Regularization}},$
-where $\widehat{\mathbf{M}}_{i,i+1}]_{kl}=\dfrac{\exp(-[\mathbf{M}_{i,i+1}]_{kl})}{\sum_{l}\exp(-[\mathbf{M}_{i,i+1}]_{kl})}$ denotes the $(k,l)$-th entry of the prior joint probability matrix induced by the originally constructed $\mathbf{M}_{i,i+1}$.
-Note that the lower the cost $\mathbf{M}_{i,i+1}]_{kl}$, the larger the entry $\widehat{\mathbf{M}}_{i,i+1}]_{kl}$, suggesting a higher plausibility of the transition from cell $k$ at $t_{i}$ to cell $l$ at $t_{i+1}$.
-The entropy regularization term in Equation 13 thus biases the learned transport plan toward the prior $\widehat{\mathbf{M}}_{i,i+1}$ rather than a uniform baseline, providing a soft mechanism for incorporating biological prior knowledge.
-Table 2: Interpolation at the middle holdout time point for the Brain Regeneration dataset.
-Sampling Method $\bm{\lambda}$ $\bm{\alpha}$ Weighted $\bm{\mathcal{W}_{2}}$ $\bm{\mathcal{W}_{2}}$ MMD Energy
-Next Step CFM $2.618\pm 0.142$ $2.579\pm 0.197$ $0.043\pm 0.003$ $12.505\pm 1.271$
-MOTFM $2.567\pm 0.088$ $2.476\pm 0.161$ $0.040\pm 0.003$ $11.269\pm 1.388$
-CTF-C 11 $0.8$ $423\pm 0.164 2.2$$293\pm 0.103$ $0.037\pm 0.001$ $9.874\pm 0.659$
-0 $0.2$ $2.396\pm 0.028$ $2.100\pm 0.102$ $0.033\pm 0.003$ $8.577\pm 0.976$
-$0.5$ $0.8$ $2.442\pm 0.173$ $2.353\pm 0.241$ $0.035\pm 0.004$ $9.008\pm 2.094$
-CTF-H 0 $2.528\pm 0.143$ $2.534\pm 0.180$ $0.040\pm 0.004$ $11.192\pm 1.304$
-11 $\bm{2.316\pm 0.141}$ $\bm{1.969\pm 0.221}$ $\bm{0.030\pm 0.004}$ $\bm{6.359\pm 1.336}$
-$0.5$ $2.519\pm 0.167$ $2.412\pm 0.158$ $0.039\pm 0.004$ $10.304\pm 1.808$
-IVP CFM $4.216\pm 0.463$ $4.266\pm 0.308$ $0.170\pm 0.029$ $32.413\pm 5.122$
-MOTFM $198\pm 0.319 4.4$$452\pm 0.243$ $0.173\pm 0.017$ $33.149\pm 3.321$
-CTF-C 11 $0.8$ $3.603\pm 0.300$ $3.816\pm 0.310$ $0.127\pm 0.018$ $24.271\pm 3.992$
-0 $0.2$ $\bm{3.465\pm 0.232}$ $\bm{3.641\pm 0.320}$ $0.119\pm 0.025$ $23.055\pm 5.939$
-$0.5$ $0.8$ $4.015\pm 0.351$ $3.974\pm 0.442$ $0.140\pm 0.038$ $27.592\pm 6.669$
-CTF-H 0 $3.925\pm 0.267$ $4.375\pm 0.297$ $0.164\pm 0.013$ $32.034\pm 3.270$
-11 $3.905\pm 0.395$ $4.188\pm 0.685$ $\bm{0.074\pm 0.014}$ $\bm{18.728\pm 2.689}$
-$0.5$ $3.917\pm 0.343$ $4.159\pm 0.455$ $0.147\pm 0.022$ $29.613\pm 4.822$
+To address the above issues, we propose the following PAER approach:
+$\min_{\mathbf{\Pi}\in\mathbb{R}^{n_{i}\times n_{i+1}}}\sum_{k,l}\Pi_{kl}\|\bm{x}_{i}(k)-\bm{x}_{i+1}(l)\|_{2}^{2}-\epsilon\sum_{k,l}\Pi_{kl}\big(\log\big(\Pi_{kl}/[\widehat{\mathbf{M}}_{i}]_{kl}\big)-1\big),$ (8)
+where $\widehat{\mathbf{M}}_{i}]_{kl}=\exp(-[\mathbf{M}_{i}]_{kl})/{\sum_{l}\exp(-[\mathbf{M}_{i}]_{kl})}$ denotes the $(k,l)$-th entry of the prior joint probability matrix induced by $\mathbf{M}_{i}$.
+Note that the lower the cost $\mathbf{M}_{i}]_{kl}$, the higher plausibility of the transition from cell $k$ at $t_{i}$ to cell $l$ at $t_{i+1}$.
+Consequently, the proposed entropy regularization term in Equation 8 biases the learned transport plan toward the prior $\widehat{\mathbf{M}}_{i}$ rather than a uniform baseline, providing a soft mechanism for incorporating the intended biological prior knowledge.
 ContextFlow.
-We apply the Sinkhorn algorithm (Cuturi, 2013) to solve the optimization problems defined in Equations 12 and 13 to obtain the spatial context-aware EOT couplings, and train the neural velocity vector field $u_{\theta}$ based on stochastic gradient descent by minimizing the multi-time generalization of Equation 3 with respect to conditionals $p_{t}(\bm{x}|\bm{z})$ and $u_{t}(\bm{x}|\bm{z})$ defined according to Equation 3.1.
-The pseudocode for the proposed method, named Conditional Flow Matching with Context-Aware OT Couplings, is presented in Algorithm 1 in Appendix E.1.
-We also provide time complexity analysis in Appendix E.2.
-In particular, to apply the Sinkhorn algorithm to solve our PAER problem in Equation 13, we make use of the following theorem, a generalized result of Peyré et al.
+Finally, we apply the Sinkhorn algorithm (Cuturi, 2013) to obtain the spatial context-aware EOT couplings by solving optimization problems 7 and 8 for minibatch samples, and train the neural velocity vector field $u_{\theta}$ by minimizing the regression loss defined by Equation 3 with respect to conditionals $p_{t}(\bm{x}|\bm{z}_{i})$ and $u_{t}(\bm{x}|\bm{z}_{i})$ defined according to Equation 1.
+The pseudocode of ContextFlow is presented in Algorithm 1 in Appendix E.1, and we analyze its time complexity in Appendix E.3.
+In particular, to apply the Sinkhorn algorithm to solve our PAER problem, we make use of the following theorem, a generalization of Peyré et al.
 (2019).
@@
 Then, we can show that $\mathbf{\Pi}^{*}_{\mathrm{CTF-H}}$ can be computed by Sinkhorn and takes the form $\mathrm{diag}(\bm{u})\cdot\mathbf{M}\odot\exp(-\mathbf{C}/\epsilon)\cdot\mathrm{diag}(\bm{v})$, where $\odot$ denotes element-wise multiplication, and $\bm{u}\in\mathbb{R}^{n_{0}},\bm{v}\in\mathbb{R}^{n_{1}}$ are vectors satisfying the marginalization constraints.
-Theorem 3.1, proven in Appendix C, suggests a new Gibbs kernel $\mathbf{K}=\mathbf{M}\odot\exp(-\mathbf{C}/\epsilon)$, which combines both the transport cost and the prior joint probability matrices.
-When $\epsilon\rightarrow 0$, $\mathbf{\Pi}^{*}_{\mathrm{CTF}}\rightarrow\mathbf{\Pi}_{\mathrm{ot}}^{*}$, thereby recovering the standard OT couplings in Equation 14.
-When $\epsilon\rightarrow\infty$, the optimal coupling $\mathbf{\Pi}^{*}_{\mathrm{CTF}}\rightarrow\mathrm{diag}(\bm{u})\cdot\mathbf{M}\cdot\mathrm{diag}(\bm{v})$, which corresponds to a plan that aligns with the prior defined by $\mathbf{M}$ rather than the independent couplings obtained with EOT defined by Equation 7.
+Theorem 4.1, proven in Appendix D.1, suggests a new Gibbs kernel $\mathbf{K}=\mathbf{M}\odot\exp(-\mathbf{C}/\epsilon)$, which combines both the transport cost matrix $\mathbf{C}$ and the prior joint probability matrix $\mathbf{M}$.
+When $\epsilon\rightarrow 0$, $\mathbf{\Pi}^{*}_{\mathrm{CTF-H}}\rightarrow\mathbf{\Pi}_{\mathrm{OT}}^{*}$, suggesting it recovers the standard OT couplings (Equation 15).
+When $\epsilon\rightarrow\infty$, $\mathbf{\Pi}^{*}_{\mathrm{CTF-H}}\rightarrow\mathrm{diag}(\bm{u})\cdot\mathbf{M}\cdot\mathrm{diag}(\bm{v})$, suggesting a plan that aligns with the prior defined by $\mathbf{M}$ rather than the independent couplings obtained with EOT defined by Equation 2.
 This has the same effect as constraining our transport plan through the proposed prior and, by extension, the flow.
@@
 Baselines & Metrics.
-We benchmark ContextFlow using its two prior integration strategies—cost-regularized (CTF-C) and entropy-regularized (CTF-H)—against several existing flow matching baselines.
-As a non-spatial baseline, we include conditional flow matching (CFM), which uses only transcriptomic data with random couplings.
-We further compare against minibatch-OT flow matching (MOTFM), which leverages OT-derived couplings but does not incorporate spatial priors.
-For evaluation, we employed the 2-Wasserstein distance ($\mathcal{W}_{2}$), a commonly used OT-based metric, as well as MMD and Energy Distance for statistical fidelity.
-Furthermore, to assess the biological plausibility of our predicted dynamics, we evaluate them using a cell-type-weighted Wasserstein distance (Weighted $\mathcal{W}_{2}$), where the weights correspond to the relative frequency of each cell type in the dataset.
-Exact metric definitions are present in Appendix B.2.
+We evaluate ContextFlow with PACM and PAER (CTF-C and CTF-H) against several state-of-the-art FM baselines, including conditional flow matching (CFM) (Lipman et al., 2023), minibatch-OT flow matching (MOTFM) (Tong et al., 2024), metric flow matching (Metric-FM) (Kapuśniak et al., 2024), and meta flow matching (MFM) (Atanackovic et al., 2024).
+To evaluate statistical fidelity, we measure $2$-Wasserstein distance ($\mathcal{W}_{2}$), maximum mean discrepancy (MMD), and energy distance (Energy).
+To assess the biological plausibility of inferred dynamics, we employ a cell-type-weighted $2$-Wasserstein distance (Weighted $\mathcal{W}_{2}$), where the weights correspond to the relative frequency of each cell type in the dataset.
+Formal definitions of these evaluation metrics are detailed in Appendix A.
 To account for randomness, all reported metrics are averaged across $10$ runs.
-Table 3: Extrapolation on the last holdout time point for the Brain Regeneration dataset.
-Sampling Method $\bm{\lambda}$ $\bm{\alpha}$ Weighted $\bm{\mathcal{W}_{2}}$ $\bm{\mathcal{W}_{2}}$ MMD Energy
-Next Step CFM $7.124\pm 0.443$ $7.133\pm 0.533$ $0.275\pm 0.011$ $76.947\pm 5.661$
-MOTFM $619\pm 0.611 7.7$$769\pm 0.763$ $0.272\pm 0.007$ $85.352\pm 8.140$
-CTF-C 11 $0.5$ $68\pm 0.608 6.969$$69\pm 0.628$ $0.265\pm 0.009$ $77.025\pm 6.056$
-0 $0.5$ $7.244\pm 0.804$ $7.146\pm 0.775$ $0.265\pm 0.003$ $80.424\pm 10.376$
-$0.5$ $0.5$ $7.188\pm 0.391$ $\bm{6.931\pm 0.260}$ $0.267\pm 0.005$ $78.992\pm 6.195$
-CTF-H 0 $\bm{6.914\pm 0.471}$ $7.198\pm 0.726$ $0.266\pm 0.009$ $\bm{76.149\pm 8.436}$
-11 $7.505\pm 0.667$ $7.338\pm 0.601$ $\bm{0.263\pm 0.006}$ $83.425\pm 8.793$
-$0.5$ $7.243\pm 0.479$ $7.157\pm 0.641$ $0.270\pm 0.007$ $79.826\pm 8.067$
-IVP CFM $6.633\pm 1.312$ $7.116\pm 1.084$ $0.143\pm 0.037$ $60.573\pm 21.756$
-MOTFM $6.503\pm 0.720$ $6.352\pm 0.592$ $0.162\pm 0.038$ $56.452\pm 15.932$
-CTF-C 11 $0.5$ $6.260\pm 0.616$ $7.681\pm 4.003$ $0.157\pm 0.039$ $52.478\pm 12.010$
-0 $0.5$ $6.614\pm 0.710$ $6.854\pm 0.740$ $0.201\pm 0.023$ $70.370\pm 9.099$
-$0.5$ $0.5$ $6.696\pm 0.427$ $6.481\pm 0.387$ $0.195\pm 0.024$ $66.212\pm 3.542$
-CTF-H 0 $6.243\pm 0.760$ $6.220\pm 0.751$ $0.195\pm 0.020$ $61.316\pm 10.288$
-11 $\bm{5.277\pm 0.936}$ $6.021\pm 1.192$ $\bm{0.099\pm 0.007}$ $\bm{27.777\pm 8.621}$
-$0.5$ $6.254\pm 0.819$ $\bm{5.973\pm 0.757}$ $0.156\pm 0.025$ $54.330\pm 12.089$
+Table 1: Comparisons under interpolation at the middle holdout time point on Brain Regeneration.
+Sampling Method Weighted $\bm{\mathcal{W}_{2}}$ $\bm{\mathcal{W}_{2}}$ MMD Energy
+Next Step CFM $2.618\pm 0.142$ $2.579\pm 0.197$ $0.043\pm 0.003$ $12.505\pm 1.271$
+MOTFM $2.567\pm 0.088$ $2.476\pm 0.161$ $0.040\pm 0.003$ $11.269\pm 1.388$
+Metric-FM $2.697\pm 0.044$ $2.688\pm 0.024$ $\bm{0.029\pm 0.001}$ $9.262\pm 0.311$
+MFM $5.809\pm 0.140$ $5.851\pm 0.156$ $0.124\pm 0.002$ $49.770\pm 0.479$
+CTF-C ($\lambda=0,\alpha=0.2$) $2.396\pm 0.028$ $2.100\pm 0.102$ $0.033\pm 0.003$ $8.577\pm 0.976$
+CTF-H ($\lambda=1$) $\bm{2.316\pm 0.141}$ $\bm{1.969\pm 0.221}$ $0.030\pm 0.004$ $\bm{6.359\pm 1.336}$
+IVP CFM $4.216\pm 0.463$ $4.266\pm 0.308$ $0.170\pm 0.029$ $32.413\pm 5.122$
+MOTFM $198\pm 0.319 4.4$$452\pm 0.243$ $0.173\pm 0.017$ $33.149\pm 3.321$
+Metric-FM $4.319\pm 0.299$ $4.501\pm 0.409$ $0.226\pm 0.007$ $34.941\pm 5.063$
+MFM $5.976\pm 0.133$ $5.699\pm 0.311$ $0.177\pm 0.010$ $51.024\pm 0.859$
+CTF-C ($\lambda=0,\alpha=0.2$) $\bm{3.465\pm 0.232}$ $\bm{3.641\pm 0.320}$ $0.119\pm 0.025$ $23.055\pm 5.939$
+CTF-H ($\lambda=1$) $3.905\pm 0.395$ $4.188\pm 0.685$ $\bm{0.074\pm 0.014}$ $\bm{18.728\pm 2.689}$
 Sampling.
-A trained velocity field can be evaluated through the samples it generates.
-We consider two variants of sampling configurations: initial value problem sampling (IVP) and next-step sampling (Next Step).
-In particular, IVP integrates the learned gradients from the first observed batch of cells and evolves them to a later time point.
-IVP provides the most comprehensive evaluation of flow quality, as errors can accumulate across steps.
-In contrast, Next Step integrates the gradient only from the most recently observed batch of cells, thus limiting error propagation but providing a less stringent test of long-term trajectory fidelity.
-4.1 Axolotl Brain Regeneration
-We first evaluate ContextFlow on longitudinal Stereo-seq spatial transcriptomic data coming from a post-traumatic brain regeneration study of the Salamander (axolotl telencephalon) species (Wei et al., 2022).
-The dataset contains samples from five developmental stages, with replicates collected from different individual organisms at each stage.
-For our CTF-C method, we present the best ablated $\alpha$, with full ablation results across different $\alpha$ values in Appendix H.
+A learned velocity field can be evaluated through the samples it generates.
+We consider two configurations: initial-value-problem (IVP) and next-step (Next Step) sampling.
+In particular, IVP integrates the gradients learned from the first observed batch of cells and evolves them to a later time point, providing the most comprehensive evaluation of flow quality, as errors can accumulate across steps.
+Next Step integrates the gradient only from the most recent batch of cells, thus limiting error propagation but providing a less stringent test of long-term trajectory fidelity.
+5.1 Main Results on Axolotl Brain Regeneration
+First, we evaluate our ContextFlow method on longitudinal Stereo-seq spatial transcriptomic data from a post-traumatic brain regeneration study of the Salamander (axolotl telencephalon) species (Wei et al., 2022).
+The dataset contains samples from $5$ developmental stages, with replicates collected from different individual organisms at each stage.
+For our CTF-C method, we present the best ablated $\alpha$, with full ablation results across different $\alpha$ values in Appendix J.
+Interpolation.
 For interpolation, we hold out the middle time point during training and evaluate it using samples generated by the trained velocity field $u_{\theta}$ via both IVP and Next-Step sampling.
@@
 Across multiple evaluation metrics, ContextFlow with entropy regularization (CTF-H) produces trajectories that most closely match the ground truth.
-CTF-H consistently achieves the best or comparable performance relative to CTF-C, despite the latter being explicitly tuned across multiple $\alpha$ values.
-This highlights the computational efficiency and superior generalization ability of CTF-H, as it avoids the need for additional hyperparameter tuning while maintaining strong performance.
-For extrapolation, we evaluate generation at the last holdout time point, which represents the most challenging test of generalizability for the velocity fields $u_{\theta}$, as it lies outside the training time horizon.
-As shown in Table 3, CTF-H again consistently achieves the best overall performance, particularly under IVP-Sampling, where errors are most likely to accumulate.
+CTF-H consistently achieves the best or comparable performance relative to baselines and CTF-C, despite the latter being explicitly tuned across multiple $\alpha$ values.
+This highlights the computational efficiency and superior generalization of CTF-H, which avoids the need for additional hyperparameter tuning while maintaining strong performance.
+Table 2: Comparisons under extrapolation on the last holdout time point on Brain Regeneration.
+Sampling Method Weighted $\bm{\mathcal{W}_{2}}$ $\bm{\mathcal{W}_{2}}$ MMD Energy
+Next Step CFM $7.124\pm 0.443$ $7.133\pm 0.533$ $0.275\pm 0.011$ $76.947\pm 5.661$
+MOTFM $619\pm 0.611 7.7$$769\pm 0.763$ $0.272\pm 0.007$ $85.352\pm 8.140$
+Metric-FM $10.125\pm 1.863$ $10.391\pm 1.968$ $0.181\pm 0.022$ $112.807\pm 24.325$
+MFM $\bm{6.868\pm 0.348}$ $6.968\pm 0.051$ $\bm{0.175\pm 0.060}$ $81.158\pm 3.677$
+CTF-C ($\lambda=0.5,\alpha=0.5$) $7.188\pm 0.391$ $\bm{6.931\pm 0.260}$ $0.267\pm 0.005$ $78.992\pm 6.195$
+CTF-H ($\lambda=0$) $6.914\pm 0.471$ $7.198\pm 0.726$ $0.266\pm 0.009$ $\bm{76.149\pm 8.436}$
+IVP CFM $6.633\pm 1.312$ $7.116\pm 1.084$ $0.143\pm 0.037$ $60.573\pm 21.756$
+MOTFM $6.503\pm 0.720$ $6.352\pm 0.592$ $0.162\pm 0.038$ $56.452\pm 15.932$
+Metric-FM $8.213\pm 4.959$ $6.972\pm 2.557$ $0.124\pm 0.034$ $29.155\pm 13.973$
+MFM $13.267\pm 1.802$ $13.513\pm 1.908$ $0.080\pm 0.001$ $109.714\pm 26.407$
+CTF-C ($\lambda=0.5,\alpha=0.5$) $6.696\pm 0.427$ $6.481\pm 0.387$ $0.195\pm 0.024$ $66.212\pm 3.542$
+CTF-H ($\lambda=1$) $\bm{5.277\pm 0.936}$ $\bm{6.021\pm 1.192}$ $\bm{0.099\pm 0.007}$ $\bm{27.777\pm 8.621}$
+Extrapolation.
+For extrapolation, we evaluate the data generation at the last holdout time point, which represents the most challenging test of generalizability for the velocity fields $u_{\theta}$, as it lies outside the training time horizon.
+As shown in Table 2, CTF-H again achieves the best overall performance, particularly under IVP sampling, where errors are most likely to accumulate.
 This result further reinforces the robustness and reliability of CTF-H across the entire sampling horizon.
 Figure 5 in Appendix G.1 further illustrates that incorporating spatial priors enables ContextFlow to produce substantially fewer biologically implausible couplings compared to its context-free counterpart.
-Table 4: Interpolation (time 5) and extrapolation (time 8) results on the Organogenesis dataset.
-Method $\bm{\lambda}$ $\bm{\alpha}$ Next Step (Interpolation) IVP (Interpolation) Next Step (Extrapolation)
+Table 3: Comparisons under interpolation (time 55) and extrapolation (time 88) on Organogenesis.
+Method Next Step (Interpolation) IVP (Interpolation) Next Step (Extrapolation)
 Weighted $\bm{\mathcal{W}_{2}}$ $\bm{\mathcal{W}_{2}}$ Weighted $\bm{\mathcal{W}_{2}}$ $\bm{\mathcal{W}_{2}}$ Weighted $\bm{\mathcal{W}_{2}}$ $\bm{\mathcal{W}_{2}}$
 MOTFM $1.892\pm 0.028$ $1.873\pm 0.086$ $3.251\pm 0.676$ $3.418\pm 0.727$ $1.626\pm 0.066$ $1.682\pm 0.096$
-CTF-C 11 $0.5$ $\bm{1.865\pm 0.030}$ $1.852\pm 0.093$ $3.137\pm 0.407$ $4.093\pm 1.187$ $1.685\pm 0.096$ $1.714\pm 0.160$
-0 $0.8$ $1.882\pm 0.022$ $1.869\pm 0.049$ $2.938\pm 0.476$ $3.904\pm 1.120$ $1.773\pm 0.053$ $1.880\pm 0.180$
-$0.5$ $0.8$ $1.888\pm 0.033$ $\bm{1.839\pm 0.134}$ $3.200\pm 0.403$ $3.555\pm 0.637$ $1.768\pm 0.058$ $1.858\pm 0.120$
-11 $0.2$ $1.880\pm 0.020$ $1.922\pm 0.078$ $3.260\pm 0.880$ $5.264\pm 3.060$ $1.683\pm 0.058$ $1.803\pm 0.117$
-0 $0.2$ $1.900\pm 0.035$ $1.912\pm 0.057$ $2.953\pm 0.425$ $3.816\pm 0.970$ $1.715\pm 0.123$ $1.860\pm 0.267$
-CTF-H 0 $1.884\pm 0.027$ $1.862\pm 0.123$ $3.244\pm 0.713$ $3.946\pm 1.671$ $\bm{1.505\pm 0.057}$ $\bm{1.397\pm 0.088}$
-11 $1.898\pm 0.029$ $1.866\pm 0.097$ $5.200\pm 0.799$ $6.306\pm 1.037$ $1.890\pm 0.046$ $1.877\pm 0.103$
-$0.5$ $1.871\pm 0.030$ $1.919\pm 0.067$ $\bm{2.814\pm 0.414}$ $\bm{3.233\pm 0.567}$ $1.636\pm 0.060$ $1.684\pm 0.099$
-Table 5: Interpolation results on the middle holdout time point for the Liver Regeneration dataset.
+Metric-FM $1.897\pm 0.036$ $1.864\pm 0.084$ $4.030\pm 1.552$ $3.826\pm 0.767$ $1.595\pm 0.114$ $1.692\pm 0.272$
+MFM $\bm{1.839\pm 0.065}$ $1.901\pm 0.099$ $476\pm 0.371 3.3$$319\pm 0.281$ $1.607\pm 0.049$ $1.407\pm 0.072$
+CTF-C ($\lambda=1,\alpha=0.5$) $1.865\pm 0.030$ $\bm{1.852\pm 0.093}$ $3.137\pm 0.407$ $4.093\pm 1.187$ $1.685\pm 0.096$ $1.714\pm 0.160$
+CTF-H ($\lambda=0.5$) $1.871\pm 0.030$ $1.919\pm 0.067$ $\bm{2.814\pm 0.414}$ $\bm{3.233\pm 0.567}$ $1.636\pm 0.060$ $1.684\pm 0.099$
+CTF-H ($\lambda=0$) $1.884\pm 0.027$ $1.862\pm 0.123$ $3.244\pm 0.713$ $3.946\pm 1.671$ $\bm{1.505\pm 0.057}$ $\bm{1.397\pm 0.088}$
+Table 4: Comparisons under interpolation on the middle holdout time point on Liver Regeneration.
 MOTFM CTF-C CTF-H
@@
 $\bm{\mathcal{W}_{2}}$ $4.303\pm 1.448 33$$3.506\pm 1.148$ $2.741\pm 1.864 33$$3.045\pm 1.644$ $\bm{32.682\pm 1.472}$ $3.481\pm 1.001 33$$3.414\pm 0.995$
-4.2 Mouse Embryo Organogenesis
-We further evaluated ContextFlow on the larger Mouse Organogenesis Spatiotemporal Atlas (MOSTA) Stereo-seq dataset (Chen et al., 2022) spanning measurements from $8$ developmental time points.
-For the interpolation study of this dataset, we held out time point 5 during training and evaluated its generation during testing.
-Table 4 shows the evaluation results.
-We observe that ContextFlow, with both integration strategies, outperforms MOTFM across all metrics, showcasing the effectiveness of the contextual information.
-While CTF-C shows stronger performance under next-step sampling—albeit only after fine-tuning the trade-off parameter $\alpha$—CTF-H consistently outperforms it in the more challenging IVP-Sampling setting.
+5.2 Additional Evaluations on Organogenesis and Liver Regeneration
+Organogenesis.
+We further evaluate ContextFlow on the larger Mouse Organogenesis Spatiotemporal Atlas (MOSTA) Stereo-seq dataset (Chen et al., 2022) spanning measurements from $8$ developmental time points.
+For the interpolation study of this dataset, we hold out the time point $5$ during training and evaluate its generation during testing.
+Table 3 reports the evaluation results.
+We observe that ContextFlow, with both integration strategies, performs best overall relative to baselines across all metrics, showcasing the effectiveness of the contextual information.
+While CTF-C shows stronger performance under next-step sampling—albeit only after fine-tuning the trade-off parameter $\alpha$—CTF-H consistently outperforms it in the more challenging IVP sampling setting.
 On the extrapolation task, integrating to the final time point, CTF-H again achieves the strongest performance, underscoring that the entropy-regularized formulation not only removes the need for additional parameter tuning but also offers more robust generalization to unseen temporal horizons.
+Liver Regeneration.
+Lastly, we evaluate ContextFlow on a Visium spatial transcriptomics dataset profiling the temporal dynamics of mouse liver regeneration following acetaminophen-induced injury (Ben-Moshe et al., 2021), which is collected across $3$ distinct regeneration stages.
+Unlike previous datasets resolved at single-cell resolution, Visium data is captured at the level of $55$-micron-diameter spots, revealing joint expression across multiple cells.
+Since direct cell-type information is unavailable, we restrict evaluation to the $2$-Wasserstein distance.
+Moreover, since evaluation is performed midway through the $3$ time points, the IVP and next-step predictions coincide.
+Table 4 summarizes the comparisons between MOTFM and our methods.
+Consistent with our previous findings, CTF-H achieves the lowest reconstruction error, indicating that incorporating contextual information improves trajectory estimation even in aggregated spot-level measurements.
 (a) Brain Regeneration
 (b) Organogenesis
-Figure 2: Comparisons of predicted and true cell type distributions.
-Figure 2 reports the KL-Divergence between normalized histograms of predicted and ground-truth cell types from ContextFlow and MOTFM.
-In both cases, CTF exhibits lower divergence on average across time points, indicating that the trajectories generated by our model better preserve the biological composition of cell types over time.
-The cell type progression is further visualized in Figure 10 (Appendix H.6).
-We show that ContextFlow predicts temporal trajectories of cell types that evolve smoothly and consistently across consecutive developmental stages.
-Early progenitor populations, such as neural crest and mesenchyme, progressively diminish as development progresses, while terminal fates, including muscle, the cartilage primordium and the liver, emerge at later stages.
+(c) Ablations on $\lambda$
+(d) OT Runtime
+Figure 3: (a) and (b) compare ContextFlow with MOTFM in terms of the KL divergence between predicted and true cell type distributions under IVP sampling on two datasets, (c) summarizes the ablations on $\lambda$ in terms of the optimal choice counts, and (d) visualizes the runtime curves with respect to varying sample sizes across different OT-based coupling schemes.
+5.3 Further Analysis
+Cell Type Distribution.
+Figures 3(a) and 3(b) demonstrate the KL divergence between normalized histograms of predicted and ground-truth cell types.
+Compared with MOTFM, ContextFlow exhibits lower average divergence across time points, indicating that the inferred trajectories better preserve the biological composition of cell types.
+The cell type progression is further visualized in Figure 11 in Appendix G.4, showing the temporal trajectories of cell types evolve smoothly and consistently across consecutive developmental stages.
+Early progenitor populations, such as neural crest and mesenchyme, progressively diminish as development progresses, while terminal fates, including muscle, the cartilage primordium, and the liver, emerge at later stages.
 Major lineages such as brain, heart, and connective tissue remain continuous throughout, confirming that ContextFlow captures biologically coherent and temporally consistent developmental dynamics.
-4.3 Liver Regeneration
-Finally, we evaluate ContextFlow on a Visium spatial transcriptomics dataset profiling the temporal dynamics of mouse liver regeneration following acetaminophen-induced injury (Ben-Moshe et al., 2021), collected across three distinct regeneration stages.
-Unlike earlier datasets resolved at single-cell resolution, Visium data is captured at the level of $55$ micron diameter spots, capturing the joint expression of multiple cells.
-Since direct cell-type information is unavailable, we restrict evaluation to the 2-Wasserstein distance.
-Moreover, since evaluation is performed on the middle of the three time points, IVP and next-step predictions coincide.
-Table 5 summarizes the comparisons between MOTFM and our methods.
-Consistent with the previous findings, CTF-H achieved the lowest reconstruction error, indicating that incorporating contextual information improves trajectory estimation even in aggregated spot-level measurements.
+Relative Importance of Priors.
+We conduct extensive ablation studies on CTF-H and report them in Appendix I.
+To summarize trends across datasets and settings, we report the counts of the optimal $\lambda$ values in Figure 3(c).
+For IVP sampling, $\lambda=0.5/1$ (both useful, with a bias towards SS) are more frequently optimal settings, while for next-step sampling, $\lambda=0$ (bias towards LR) is the most frequent.
+This behavior can be explained by the different roles of the priors.
+The SS prior ($\lambda=1$) serves as a proxy for relative spatial consistency across tissue microenvironments, providing a global structural constraint that remains informative over long integration horizons and is therefore beneficial for IVP sampling.
+In contrast, the LR prior is informative depending on how distinct the cell-cell communication patterns are over a given time interval, making it more localized and time-dependent, particularly informative for short-range transitions, and thus more beneficial for next-step sampling.
+Computational Advantage.
+We compare the performance of CTF-H OT, the coupling scheme used in ContextFlow, against state-of-the-art spatiotemporal OT methods, including DeST-OT (Halmos et al., 2025), MOSCOT (Klein et al., 2025), and TOAST (Ceccarelli et al., 2026).
+Figure 3(d) reports the runtime of different OT methods on the Brain Regeneration dataset with respect to varying sample sizes, where CTF-H OT is significantly faster than prior methods (270270x for DeST-OT, 257257x for MOSCOT, and 2424x for TOAST), while achieving comparable OT alignment performance (see Tables 6-8 in Appendix F for comparisons in growth distortion, migration, and coupled transcription distance metrics).
+These results suggest that the proposed prior integration strategy is more suited for flow matching frameworks, where thousands of OT problems are involved during training.
 6 Conclusion
 We introduced ContextFlow, a contextually aware flow matching framework that leverages spatial priors and biologically motivated constraints to learn more plausible trajectories from snapshot spatial transcriptomic data.
-The entropic variant of ContextFlow is theoretically grounded, which produces OT couplings that are constrained by prior knowledge, thereby promoting stability and consistency with the imposed contextual constraints.
-Across three diverse datasets, we showed that ContextFlow consistently outperforms state-of-the-art baselines, even in challenging Initial Value Problem sampling settings, underscoring the importance of our contextually informed priors.
-In addition, we demonstrated that our framework reduces the number of biologically implausible couplings and yields coherent, temporally consistent developmental trajectories while maintaining strong quantitative performance across Wasserstein, MMD, and Energy metrics.
-These results highlight the value of embedding biological context into generative flow models.
+Our experiments showed that the entropic variant of ContextFlow (CTF-H) outperforms existing baselines, even in challenging IVP sampling settings.
+In particular, our method effectively reduces the number of biologically implausible couplings and generates more temporally consistent developmental trajectories, while maintaining strong quantitative performance across $2$-Wasserstein, MMD, and Energy metrics.
+These results highlight the value of embedding proper biological context into flow-based generative models.
 Future work can adapt our methods to reconstruct tissues and learn spatial latent dynamics by formulating the flow in space (rather than time), or by leveraging multi-marginal OT formulations to optimize temporal flows.
-Looking forward, ContextFlow offers a principled foundation for modeling perturbations and disease progression, potentially bridging generative power with biological interpretability.
-Impact Statement
-Our work contributes a new method for modeling dynamical systems from unpaired observations with structural constraints.
-As such, our algorithm can contribute to biomedical domains, for example, by modeling drug perturbation responses in spatially resolved tumor tissues or by analyzing complex biological dynamics with unpaired datasets.
-These applications could positively impact biomedical research by enabling more data-efficient modeling of longitudinal biological processes and expediting hypothesis generation about treatment effects or underlying mechanisms.
-Future work should emphasize broad validation of our method using experimental or clinical ground truth, and the interpretation of results with domain experts to build appropriate trust in predictions.
-Appendix A Related Work
-A.1 Flow Matching
-Normalizing flows provide a parametric framework for characterizing transformations of a random variable into desired distributions (Papamakarios et al., 2021).
-These transformations can be realized through either finite (Rezende and Mohamed, 2015) or continuous compositions (Chen et al., 2018).
-The loss functions used in such formulations typically require computing Jacobians or integrating the flows at each forward pass, making them computationally expensive.
-Flow matching (FM) (Lipman et al., 2023; Albergo and Vanden-Eijnden, 2022; Liu et al., 2022) addresses this limitation by reducing the training of the velocity field to a regression problem, thereby making normalizing flows substantially more scalable.
-To ensure valid conditional paths at intermediate time points, samples are coupled either randomly or via optimal transport (Pooladian et al., 2023; Tong et al., 2024).
-Owing to this scalability, FM has been rapidly adopted across scientific domains, including biology and the life sciences (Li et al., 2025).
-In transcriptomics, for example, Klein et al.
-(2024) employed an FM backbone to approximate OT maps for drug response modeling and cross-modal translation tasks.
-Entropic OT formulations have also been applied to infer cellular trajectories (Tong et al., 2024; Rohbeck et al., 2025), generate imaging-based cell morphology changes (Zhang et al., 2025), and simulate spatial transcriptomics data from histology images (Huang et al., 2025).
-Despite these advances, existing work does not address how to meaningfully incorporate biological prior knowledge to constrain the velocity field, limiting the biological plausibility of inferred trajectories.
-A.2 Optimal Transport
-Omics studies frequently generate uncoupled measurements across conditions, modalities, or time points, which must be integrated into a unified representation to provide a more comprehensive view of the underlying biology.
-Optimal transport (OT) has recently gained popularity for this task, as it provides a geometry-based approach to couple probability distributions (Bunne et al., 2024; Klein et al., 2025).
-In spatial transcriptomics (ST), several OT formulations have been introduced depending on context.
-For instance, Zeira et al.
-(2022) and Liu et al.
-(2023) proposed PASTE and PASTE2 to align ST data from adjacent tissue slices, while DeST-OT (Halmos et al., 2025) integrates spatio-temporal slices by modeling cell growth and differentiation.
-Rahimi et al.
-(2024) developed DOT, a multi-objective OT framework for mapping features across scRNA-seq and spatially resolved assays, and Ceccarelli et al.
-(2025) introduced TOAST, a spatially regularized OT framework for slice alignment and annotation transfer.
-While these methods are primarily designed to align biological data across space, time, or modality, they do not address the problem of trajectory inference toward biologically plausible solutions, leveraging biological priors to constrain or bias the transport plan.
-Appendix B Additional Definitions
-B.1 Kantarovich’s OT Formulation
-Kantorovich’s formulation (Peyré et al., 2019) is a classical definition of the optimal transport (OT) problem that seeks a joint coupling to move a probability measure to another that minimizes the Euclidean distance cost, corresponding to the following minimization problem with respect to the 2-Wasserstein distance:
-$\pi^{*}_{\mathrm{ot}}:={\arg\!\min}_{\pi\in\Pi(q_{0},q_{1})}\int_{\mathbb{R}^{d}\times\mathbb{R}^{d}}\|\bm{x}_{0}-\bm{x}_{1}\|_{2}^{2}\>d\pi(\bm{x}_{0},\bm{x}_{1}),$ (14)
-where $\Pi(q_{0},q_{1})$ denotes the set of joint probability measures such that the left and right marginals are $q_{0}$ and $q_{1}$.
-Equation 14 can be solved in a mini-batch fashion using standard solvers such as POT (Flamary et al., 2021); however, the computational complexity is cubic in batch size.
-B.2 Evaluation Metrics
+Looking forward, our ContextFlow framework offers a principled foundation for modeling perturbations and disease progression, potentially bridging generative power with biological interpretability.
+NeurIPS Paper Checklist
+  1. 1.
+    Claims
+  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
+  3. Answer: [Yes]
+  4.
+Justification: Our main claims focus on demonstrating the limitations of prior flow matching techniques and on our proposed solution to mitigate them.
+Section 3 discusses the limitations of prior techniques, Section 4 presents the proposed methodology, and Section 5 reports comprehensive experimental results against baselines to validate our claims.
+Additionally, we defer more analysis to Appendix C (effect of normalization in OT cost), Appendix G.1 (biological plausibility analysis), and Appendices I and J (full ablation studies).
+  5. Guidelines:
+    • The answer [N/A] means that the abstract and introduction do not include the claims made in the paper.
+    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations.
+A [No] or [N/A] answer to this question will not be perceived well by the reviewers.
+    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
+    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
+  6. 2.
+    Limitations
+  7. Question: Does the paper discuss the limitations of the work performed by the authors?
+  8. Answer: [Yes]
+  9.
+Justification: Main limitations of our work include poorly specified or noisy priors, which may bias the learned dynamics.
+Furthermore, while our evaluation demonstrates strong statistical and biological coherence, validation using experimental or longitudinal ground truth data would strengthen confidence in the inferred trajectories.
+  10. Guidelines:
+    • The answer [N/A] means that the paper has no limitation while the answer [No] means that the paper has limitations, but those are not discussed in the paper.
+    • The authors are encouraged to create a separate “Limitations” section in their paper.
+    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally).
+The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
+    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs.
+In general, empirical results often depend on implicit assumptions, which should be articulated.
+    • The authors should reflect on the factors that influence the performance of the approach.
+For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting.
+Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
+    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
+    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
+    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper.
+The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community.
+Reviewers will be specifically instructed to not penalize honesty concerning limitations.
+  11. 3.
+    Theory assumptions and proofs
+  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
+  13. Answer: [Yes]
+  14. Justification: Yes, the formal statements of our theorems list all the necessary assumptions. Appendix D provides their detailed proofs.
+  15. Guidelines:
+    • The answer [N/A] means that the paper does not include theoretical results.
+    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
+    • All assumptions should be clearly stated or referenced in the statement of any theorems.
+    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
+    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
+    • Theorems and Lemmas that the proof relies upon should be properly referenced.
+  16. 4.
+    Experimental result reproducibility
+  17.
+Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
+  18. Answer: [Yes]
+  19.
+Justification: We’ve included the information to fully describe the algorithm in Appendix E, as well as the data processing, splits, and the settings of the packages used in our studies in Section 5 and Appendix E.2.
+  20. Guidelines:
+    • The answer [N/A] means that the paper does not include experiments.
+    • If the paper includes experiments, a [No] answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
+    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
+    • Depending on the contribution, reproducibility can be accomplished in various ways.
+For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model.
+In general.
+releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
+    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution.
+For example
+      1. (a)
+        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
+      2. (b)
+        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
+      3. (c)
+        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
+      4. (d)
+        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility.
+In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
+  21. 5.
+    Open access to data and code
+  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
+  23. Answer: [No]
+  24. Justification: The refactored source code will be released after the review process.
+  25. Guidelines:
+    • The answer [N/A] means that paper does not include experiments requiring code.
+    • Please see the NeurIPS code and data submission guidelines (https://neurips.cc/public/guides/CodeSubmissionPolicy) for more details.
+    • While we encourage the release of code and data, we understand that this might not be possible, so [No] is an acceptable answer.
+Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
+    • The instructions should contain the exact command and environment needed to run to reproduce the results.
+See the NeurIPS code and data submission guidelines (https://neurips.cc/public/guides/CodeSubmissionPolicy) for more details.
+    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
+    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines.
+If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
+    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
+    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
+  26. 6.
+    Experimental setting/details
+  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer) necessary to understand the results?
+  28. Answer: [Yes]
+  29.
+Justification: We’ve included the information to fully describe the algorithm in Appendix E, as well as the data processing, splits, and the settings of the packages used in our studies in Section 5 and Appendix E.2.
+  30. Guidelines:
+    • The answer [N/A] means that the paper does not include experiments.
+    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
+    • The full details can be provided either with the code, in appendix, or as supplemental material.
+  31. 7.
+    Experiment statistical significance
+  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
+  33. Answer: [Yes]
+  34. Justification: We’ve provided both the mean and standard deviation statistics across random seeds for all the reported metrics, and extensive ablations in Appendices I and J.
+  35. Guidelines:
+    • The answer [N/A] means that the paper does not include experiments.
+    • The authors should answer [Yes] if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
+    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
+    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
+    • The assumptions made should be given (e.g., Normally distributed errors).
+    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
+    • It is OK to report 1-sigma error bars, but one should state it.
+The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
+    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g., negative error rates).
+    • If error bars are reported in tables or plots, the authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
+  36. 8.
+    Experiments compute resources
+  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
+  38. Answer: [Yes]
+  39.
+Justification: We’ve included information in terms of data processing, splits, settings of packages used in our studies, runtime, etc.
+Importantly, all our experiments were run on a single Nvidia DGX A100.
+  40. Guidelines:
+    • The answer [N/A] means that the paper does not include experiments.
+    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
+    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
+    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
+  41. 9.
+    Code of ethics
+  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
+  43. Answer: [Yes]
+  44. Justification: We’ve read the guidelines and confirm that we’re not in violation of any code of ethics.
+  45. Guidelines:
+    • The answer [N/A] means that the authors have not reviewed the NeurIPS Code of Ethics.
+    • If the authors answer [No] , they should explain the special circumstances that require a deviation from the Code of Ethics.
+    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
+  46. 10.
+    Broader impacts
+  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
+  48. Answer: [No]
+  49.
+Justification: The paper presents a general principled methodology to propose a flow matching model for modeling spatial omics in general, rather than specific applications with direct societal implications.
+  50. Guidelines:
+    • The answer [N/A] means that there is no societal impact of the work performed.
+    • If the authors answer [N/A] or [No] , they should explain why their work has no societal impact or why the paper does not address societal impact.
+    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
+    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments.
+However, if there is a direct path to any negative applications, the authors should point it out.
+For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate Deepfakes for disinformation.
+On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
+    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
+    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
+  51. 11.
+    Safeguards
+  52.
+Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pre-trained language models, image generators, or scraped datasets)?
+  53. Answer: [N/A]
+  54. Justification: [N/A]
+  55. Guidelines:
+    • The answer [N/A] means that the paper poses no such risks.
+    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
+    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
+    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
+  56. 12.
+    Licenses for existing assets
+  57.
+Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
+  58. Answer: [N/A]
+  59. Justification: [N/A]
+  60. Guidelines:
+    • The answer [N/A] means that the paper does not use existing assets.
+    • The authors should cite the original paper that produced the code package or dataset.
+    • The authors should state which version of the asset is used and, if possible, include a URL.
+    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
+    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
+    • If assets are released, the license, copyright information, and terms of use in the package should be provided.
+For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets.
+Their licensing guide can help determine the license of a dataset.
+    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
+    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
+  61. 13.
+    New assets
+  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
+  63. Answer: [N/A]
+  64. Justification: [N/A]
+  65. Guidelines:
+    • The answer [N/A] means that the paper does not release new assets.
+    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
+    • The paper should discuss whether and how consent was obtained from people whose asset is used.
+    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
+  66. 14.
+    Crowdsourcing and research with human subjects
+  67.
+Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
+  68. Answer: [N/A]
+  69. Justification: [N/A]
+  70. Guidelines:
+    • The answer [N/A] means that the paper does not involve crowdsourcing nor research with human subjects.
+    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
+    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
+  71. 15.
+    Institutional review board (IRB) approvals or equivalent for research with human subjects
+  72.
+Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
+  73. Answer: [N/A]
+  74. Justification: [N/A]
+  75. Guidelines:
+    • The answer [N/A] means that the paper does not involve crowdsourcing nor research with human subjects.
+    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research.
+If you obtained IRB approval, you should clearly state this in the paper.
+    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
+    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
+  76. 16.
+    Declaration of LLM usage
+  77.
+Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research?
+Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigor, or originality of the research, declaration is not required.
+  78. Answer: [N/A]
+  79. Justification: [N/A]
+  80. Guidelines:
+    • The answer [N/A] means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
+    • Please refer to our LLM policy in the NeurIPS handbook for what should or should not be described.
+Appendix A Formal Definitions of Evaluation Metrics
 2-Wasserstein. The 2-Wasserstein distance ($\mathcal{W}_{2}$ between empirical distributions $\mu,\nu$ is defined as:
@@
 where $n_{i}^{\text{true}},n_{i}^{\text{pred}}$ are the number of true and predicted cells of type $i$, and $N$ is the total number of samples.
-To determine the cell type of generated trajectories, we employ a multi-class classifier $M_{\phi}$, implemented as an XGBoost model (Chen and Guestrin, 2016) trained for each dataset.
+To determine the cell type of generated trajectories, we employ a multi-class classifier $M_{\phi}$, implemented as an XGBoost model [Chen and Guestrin, 2016] trained for each dataset.
 Energy Distance.
@@
 In our evaluations, we use a multi-kernel variant with radial basis function (RBF) kernels $\kappa_{\gamma}(\bm{x},y)=\exp(-\gamma\|\bm{x}-\bm{y}\|^{2})$, and average over $\gamma\in[2,1,0.5,0.1,0.01,0.005]$.
-Appendix C Proofs of Main Theoretical Results
-Proposition C.1.
+Appendix B Detailed Preliminaries on Flow Matching
+Flow Matching Basics.
+Flow matching [Lipman et al., 2023] is a simulation-free and sample-efficient generative framework for training continuous normalizing flows [Chen et al., 2018].
+Given a pair of source and target distributions over $\mathbb{R}^{d}$ with probability densities $q_{0}=q(\bm{x}_{0})$ and $q_{1}=q(\bm{x}_{1})$, the problem task is to learn a time-varying velocity vector field $u_{\theta}:[0,1]\times\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$, whose continuous evolution is captured by a function in the form of a neural-net-based model with weights $\theta$, that can transform $q_{0}$ to $q_{1}$ through integration via an ordinary differential equation (ODE).
+To be more specific, flow matching (FM) seeks to optimize $\theta$ by minimizing a simple regression loss between $u_{\theta}$ and a target time-varying velocity vector field $u_{t}:[0,1]\times\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$ as follows:
+$\min_{\theta}\>\mathbb{E}_{t\sim\mathcal{U}(0,1),\bm{x}\sim p_{t}(\bm{x})}\big\|u_{\theta}(t,\bm{x})-u_{t}(\bm{x})\big\|^{2}.$ (9)
+Here, $\mathcal{U}(0,1)$ is the uniform distribution over $[0,1]$, and $p_{t}:[0,1]\times\mathbb{R}^{d}\rightarrow\mathbb{R}_{+}$ denotes a time-varying probability path induced by $u_{t}$ such that (i) $p_{t}$ is a probability density function for any $t\in[0,1]$, (ii) $p_{t}$ satisfies the two boundary conditions: $p_{t=0}=q_{0}$ and $p_{t=1}=q_{1}$, and (iii) the connection between $p_{t}$ and $u_{t}$ can be characterized by the transport equation [Villani and others, 2008]:
+$\frac{\partial p_{t}(\bm{x})}{\partial t}=-\nabla\cdot(u_{t}(\bm{x})p_{t}(\bm{x})),$ (10)
+where $\nabla$ is the divergence operator.
+From a dynamical system’s view, $u_{t}$ defines an ODE system $d\bm{x}=u_{t}(\bm{x})dt$.
+The corresponding solution to the ODE, usually termed as the probability flow, can then transport any $\bm{x}_{0}\sim q_{0}$ to a point $\bm{x}_{1}\sim q_{1}$ along $u_{t}$ from $t=0$ to $t=1$.
+While the FM objective in Equation 9 is simple and intuitive, it is generally intractable in practice: the closed-form velocity vector field $u_{t}$ is unknown for arbitrary source and target distributions ($q_{0}$ and $q_{1}$), and multiple valid probability paths $p_{t}$ may exist between them.
+Conditional Flow Matching.
+The central idea is to express the target probability path via a mixture of more manageable conditional probability paths [Lipman et al., 2023].
+By marginalizing over a conditioning variable $z$, both $p_{t}$ and $u_{t}$ can be constructed using their conditional counterparts:
+$p_{t}(\bm{x})$ $=\int p_{t}(\bm{x}|\bm{z})q(\bm{z})d\bm{z}\quad\text{and}\quad u_{t}(\bm{x})$ $=\int u_{t}(\bm{x}|\bm{z})\frac{p_{t}(\bm{x}|\bm{z})q(\bm{z})}{p_{t}(\bm{x})}d\bm{z},$ (11)
+where $q(\bm{z})$ is the distribution of the conditioning variable $\bm{z}$, and the conditional probability path $p_{t}(\bm{x}|\bm{z})$ is constructed such that the boundary conditions are satisfied: $\int p_{t=0}(\bm{x}|\bm{z})q(\bm{z})=q_{0}$ and $\int p_{t=1}(\bm{x}|\bm{z})q(\bm{z})=q_{1}$.
+Theorem 1 of Lipman et al.
+[2023] proves that $p_{t}$ and $u_{t}$ defined by Equation 11 satisfy the transport equation, suggesting that $p_{t}$ is a valid probability path generated by $u_{t}$.
+To avoid the intractable integrals, Lipman et al.
+[2023] proposed the following training objective of conditional flow matching, and proved its equivalence to Equation 9 in terms of gradient computation:
+$\min_{\theta}\>\mathbb{E}_{t\sim\mathcal{U}(0,1),z\sim q(\bm{z}),\bm{x}\sim p_{t}(\bm{x}|\bm{z})}\big\|u_{\theta}(t,\bm{x})-u_{t}(\bm{x}|\bm{z})\big\|^{2}.$ (12)
+By choosing an appropriate conditional velocity vector field $u_{t}(\bm{x}|\bm{z})$, we can train the neural network using Equation 12 without requiring a closed-form solution of the conditional probability path $p_{t}(\bm{x}|\bm{z})$, thereby avoiding the intractable integration operation.
+Therefore, the remaining task is to define the conditional probability path and velocity vector field properly such that we can sample from $p_{t}(\bm{x}|\bm{z})$ and compute $u_{t}(\bm{x}|\bm{z})$ efficiently for solving the optimization problem in Equation 12.
+Gaussian Conditional Probability Paths. A specific choice proposed in Lipman et al. [2023] is Gaussian conditional paths and their corresponding velocity vector fields, which are defined as:
+$p_{t}(\bm{x}|\bm{z})=\mathcal{N}(\bm{x}\>|\>\mu_{t}(\bm{z}),\sigma_{t}(\bm{z})^{2}\mathbf{I})\quad\text{and}\quad u_{t}(\bm{x}|\bm{z})=\frac{\sigma_{t}^{\prime}(\bm{z})}{\sigma_{t}(\bm{z})}(\bm{x}-\mu_{t}(\bm{z}))+\mu_{t}^{\prime}(\bm{z}),$ (13)
+where $\mu_{t}:[0,1]\times\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$ stands for the time-varying mean of the Gaussian distribution, $\mu_{t}^{\prime}$ is its derivative with respect to time, $\sigma_{t}:[0,1]\times\mathbb{R}^{d}\rightarrow\mathbb{R}_{+}$ is the time-varying scalar standard deviation, and $\sigma_{t}^{\prime}$ denotes its derivative.
+We refer to Theorem 3 of Lipman et al.
+[2023] for a formal argument that $u_{t}(\bm{x}|\bm{z})$ uniquely generates the Gaussian conditional probability path $p_{t}(\bm{x}|\bm{z})$ for any differentiable $\mu_{t}(\bm{z})$ and $\sigma_{t}(\bm{z})$ in Equation 13 satisfying the boundary conditions.
+In particular, Lipman et al.
+[2023] selected $q(\bm{z})=q(\bm{x}_{1})$, $\mu_{t}(\bm{z})=t\bm{x}_{1}$, and $\sigma_{t}(\bm{z})=1-(1-\sigma)t$.
+Then, we can see that $u_{t}(\bm{x}|\bm{z})$ transports the standard Gaussian distribution $p_{t=0}(\bm{x}|\bm{z})=\mathcal{N}(\bm{x};\bm{0},\mathbf{I})$ to a Gaussian distribution with mean $\bm{x}_{1}$ and covariance $\sigma^{2}\mathbf{I}$, namely $p_{t=1}(\bm{x}|\bm{z})=\mathcal{N}(\bm{x};\bm{x}_{1},\sigma^{2}\mathbf{I})$ for any target point $\bm{x}_{1}$.
+By letting $\sigma\rightarrow 0$, the marginal boundary conditions can easily be validated.
+Tong et al.
+[2024] further generalized the application scope to arbitrary source distributions by setting:
+$q(\bm{z})=q(\bm{x}_{0})q(\bm{x}_{1}),\quad\mu_{t}(\bm{z})=(1-t)\bm{x}_{0}+t\bm{x}_{1},\quad\sigma_{t}(\bm{z})=\sigma.$ (14)
+The above choice satisfies the boundary conditions $p_{t=0}(\bm{x})=q_{0}$ and $p_{t=1}(\bm{x})=q_{1}$ when $\sigma\rightarrow 0$.
+Based on Equation 13, the conditional velocity vector field has an analytical form $u_{t}(\bm{x}|\bm{z})=\bm{x}_{1}-\bm{x}_{0}$.
+Conditional Flow Matching with OT-Couplings.
+The above construction in Equation 14 corresponds to the simplest choice of independent coupling, where $\bm{z}=(\bm{x}_{0},\bm{x}_{1})$ with source $\bm{x}_{0}$ and target $\bm{x}_{1}$ are independently sampled from $q(\bm{z})=q(\bm{x}_{0})q(\bm{x}_{1})$.
+The use of couplings to construct sampling paths in the CFM framework naturally connects to optimal transport theory [Villani and others, 2008].
+Choosing OT-based couplings has several advantages over independent coupling, including smaller training variance and more efficient sampling [Pooladian et al., 2023, Tong et al., 2024].
+The most classical definition of the optimal transport (OT) problem seeks a joint coupling to move a probability measure to another that minimizes the Euclidean distance cost, corresponding to the following minimization problem with respect to the Wasserstein-2 distance:
+$\Pi^{*}_{\mathrm{OT}}={\arg\!\min}_{\pi\in\Pi(q_{0},q_{1})}\int_{\mathbb{R}^{d}\times\mathbb{R}^{d}}\|\bm{x}_{0}-\bm{x}_{1}\|_{2}^{2}\>d\pi(\bm{x}_{0},\bm{x}_{1}),$ (15)
+where $\Pi(q_{0},q_{1})$ denotes the set of joint probability measures such that the left and right marginals are $q_{0}$ and $q_{1}$.
+The above formulation is also known as Kantorovich’s formulation [Peyré et al., 2019].
+Equation 15 can be solved in a mini-batch fashion using standard solvers such as POT [Flamary et al., 2021]; however, the computational complexity is cubic in batch size.
+A more efficient alternative is entropic optimal transport (EOT), which approximately solves the OT problem using entropic regularization, reducing the computational costs from cubic to quadratic:
+$\Pi^{*}_{\mathrm{EOT}}(\epsilon)=\operatorname*{arg\,min}_{\pi\in\Pi(q_{0},q_{1})}\int_{\mathbb{R}^{d}\times\mathbb{R}^{d}}\|\bm{x}_{0}-\bm{x}_{1}\|_{2}^{2}\>d\pi(\bm{x}_{0},\bm{x}_{1})+\epsilon H(\pi\>|\>q_{0}\otimes q_{1}),$ (16)
+where $\epsilon>0$ is the regularization parameter, and $H(\pi\>|\>q_{0}\otimes q_{1})$ denotes the relative entropy (or KL divergence) with respect to $\pi$ and the product measure $q_{0}\otimes q_{1}$.
+Note that this optimization problem can also be viewed as a special case of the static Schrödinger bridge problem [Bernton et al., 2022], which can be efficiently solved in a mini-batch fashion via the Sinkhorn algorithm [Cuturi, 2013].
+Theoretically, one can prove that $\Pi^{*}_{\mathrm{EOT}}(\epsilon)$ recovers the Kantorovich’s OT coupling $\Pi^{*}_{\mathrm{OT}}$ when $\epsilon\rightarrow 0$ and $\Pi^{*}_{\mathrm{EOT}}(\epsilon)$ corresponds to the independent coupling $q_{0}\otimes q_{1}$ when $\epsilon\rightarrow\infty$.
+Appendix C Effect of Normalization in PACM Optimization
+We elaborate on why normalizing the cost matrix in the EOT optimization objective results in suboptimal couplings, as discussed in Section 4.2 when motivating our PAER design.
+According to Peyré et al.
+[2019], we know that optimal EOT coupling takes the form $\Pi^{*}_{\text{EOT}}=\text{diag}(u)\cdot K\cdot\text{diag}(v)$, where $K$ is the kernel matrix such that $K]_{ij}=\exp(-c_{ij}/\epsilon)$, and $u,v$ satisfies marginalization constraints $u\odot Kv=a$ and $K^{T}u\odot v=b$.
+Thus, the Sinkhorn updates are given by:
+$u^{l+1}=\frac{a}{Kv^{l}}\quad\text{and}\quad v^{l+1}=\frac{b}{K^{T}u^{l+1}}.$
+In cases where the OT cost function involves information from different modalities, the distances are usually normalized to a similar scale.
+Normalizing the cost results $\tilde{c}_{ij}=c_{ij}/{\epsilon}$ such that the new kernel matrix $K_{\text{norm}}]_{ij}=\exp(\frac{-c_{ij}}{C_{\text{max}}\epsilon})$ can cause numerical issues if $C_{\text{max}}\gg 1$.
+Therefore, cost normalization should be performed carefully when considering different pairwise distances.
+Scaling the cost has the same effect as that of increasing $\epsilon$, making solutions more diffuse.
+The following theorem, proven in Appendix D.2, formalizes how normalization affects the optimal transport plan.
+Theorem C.1.
+Let $\mathbf{C}\in\mathbb{R}^{n_{0}\times n_{1}}$ be a cost matrix and $\mathbf{M}\in\mathbb{R}^{n_{0}\times n_{1}}$ a prior transition matrix with positive entries.
+Consider the following optimization problem with respect to entropy-regularized OT:
+$\Pi^{*}$ $={\arg\!\min}_{\Pi}\sum_{k,l}\Pi_{kl}C_{kl}+\epsilon\sum_{k,l}\Pi_{kl}(\log\!\left(\Pi_{kl}\right)-1).$
+Let $\tilde{\Pi}^{*}$ be the EOT coupling where the cost is scaled by a normalization constant $c$ (i.e., $\tilde{C}_{ij}=\frac{C_{ij}}{c}$) with the same regularization parameter $\epsilon>0$.
+Then, for any $(i,j)$ and $(k,l)$, we have
+$\frac{\tilde{\Pi}^{*}_{ij}}{\tilde{\Pi}^{*}_{kl}}\leq\gamma\,\left(\frac{\Pi^{*}_{ij}}{\Pi^{*}_{kl}}\right)^{\frac{1}{c}}\quad\text{and}\quad H\big(\tilde{\Pi}_{ij}^{*}\big)\geq m\cdot H(\Pi_{ij}^{*})-s,$
+where $\gamma$, $m$ and $s$ are constants depending on $\Pi^{*}_{ij}$, the normalization scaler $c$, and the parameters $a,b$ in the marginalization constraints, $H(\cdot)$ stands for the entropy function.
+Suppose $\Pi^{*}_{ij}=m\cdot\Pi^{*}_{kl}$ with $m>1$.
+Then, from Theorem C.1 we know $\tilde{\Pi}^{*}_{ij}<m\cdot\tilde{\Pi}^{*}_{kl}$ for any $c>1$.
+This suggests that normalization brings probabilities that are far apart under the optimal transport plan closer together, leading to more diffuse couplings and increased entropy.
+Appendix D Technical Proofs of Our Theoretical Results
+D.1 Proof of Theorem 4.1
+Proof.
+We have that:
+$\Pi^{*}_{\mathrm{CTF-H}}=\operatorname*{arg\,min}_{\Pi\in\mathbb{R}^{n_{0}\times n_{1}}}$ $\sum_{k,l}\Pi_{kl}C_{kl}+\epsilon\sum_{k,l}\Pi_{kl}(\log(\Pi_{kl}/{M}_{kl})-1),$
+subject to:
+$\Pi\mathbf{1}=a,\quad\Pi^{\top}\mathbf{1}=b.$
+This formulation is a standard convex optimization setting with constraints. The Lagrangian of this setting is:
+$\mathcal{L}(\Pi,f,g)=\sum_{k,l}C_{kl}\Pi_{kl}$ $+\epsilon\sum_{k,l}\Pi_{kl}\left(\log\left(\frac{\Pi_{kl}}{M_{kl}}\right)-1\right)$
+$\qquad-\sum_{k}f_{k}\left(\sum_{l}\Pi_{kl}-a_{k}\right)-\sum_{l}g_{l}\left(\sum_{k}\Pi_{kl}-b_{l}\right).$
+Differentiating with respect to $\Pi_{kl},f_{k},g_{l}$, we get:
+$\frac{\partial\mathcal{L}}{\partial\Pi_{kl}}=C_{kl}+\epsilon\log\left(\frac{\Pi_{kl}}{M_{kl}}\right)-f_{k}-g_{l}.$
+Setting the derivative to zero:
+$\epsilon\log\left(\frac{\Pi^{*}_{kl}}{M_{kl}}\right)=f_{k}-C_{kl}+g_{l}$
+$\implies\frac{\Pi^{*}_{kl}}{M_{kl}}=e^{\frac{f_{k}}{\epsilon}}e^{-\frac{C_{kl}}{\epsilon}}e^{\frac{g_{l}}{\epsilon}}$
+$\implies\Pi^{*}_{kl}=e^{\frac{f_{k}}{\epsilon}}M_{kl}e^{-\frac{C_{kl}}{\epsilon}}e^{\frac{g_{l}}{\epsilon}}$
+Let $u\in\mathbb{R}^{n}$ and $v\in\mathbb{R}^{m}$ such that
+$u_{k}=e^{\frac{f_{k}}{\epsilon}},\quad v_{l}=e^{\frac{g_{l}}{\epsilon}}.$
+Let $K_{kl}$ be the kernel $M_{kl}e^{-C_{kl}/\epsilon}$. Then, we have:
+$\Pi^{*}_{kl}=u_{k}K_{kl}v_{l},$
+suggesting that
+$\Pi^{*}=\operatorname{diag}(u)\cdot K\cdot\operatorname{diag}(v).$ (17)
+Differentiating the Lagrangian with respect to $f_{k}$ and $g_{l}$, we get:
+$\frac{\partial\mathcal{L}}{\partial f_{k}}=1\cdot\left(\sum_{l}\Pi^{*}_{kl}-a_{k}\right)=0$
+$\implies\Pi^{*}\mathbf{1}=a$ (18)
+$\frac{\partial\mathcal{L}}{\partial g_{l}}=1\cdot\left(\sum_{i}\Pi^{*}_{kl}-b_{l}\right)=0$
+$\implies\Pi^{*\top}\mathbf{1}=b$ (19)
+From Equations 17, 18 and 19 above, we get:
+$\operatorname{diag}(u)\cdot K\cdot\operatorname{diag}(v)\cdot\mathbf{1}$ $=a,$
+$(\operatorname{diag}(u)\cdot K\cdot\operatorname{diag}(v))^{\top}\mathbf{1}$ $=b,$
+which can be rewritten as:
+$u\odot(Kv)$ $=a$
+$K^{\top}u\odot v$ $=b.$
+This is the usual matrix scaling formulation for which the Iterative Proportional Fitting (IPF) updates are:
+$u_{k}^{t+1}=\frac{a_{k}}{(Kv^{t})_{k}},\quad v_{l}^{t+1}=\frac{b_{l}}{(K^{\top}u^{t+1})_{l}}.$
+The Sinkhorn Algorithm uses these updates iteratively, and these updates are shown to converge in Franklin and Lorenz [1989].
+Thus, the Sinkhorn Algorithm can be used for the ContextFlow’s Prior Aware Entropy Regularized (PAER) (CTF-H) formulation.
+From Equation 17, we get:
+$\Pi^{*}_{kl}=e^{f_{k}/\epsilon}\cdot M_{kl}\cdot e^{-C_{kl}/\epsilon}e^{g_{l}/\epsilon}.$
+When $\epsilon\to\infty$, we have $C_{kl}/\epsilon\to 0$.
+$\quad e^{-C_{kl}/\epsilon}\to 1$
+$\implies$ $\Pi^{*}_{kl}\to u_{k}M_{kl}v_{l}$
+$\implies$ $\Pi^{*}_{\text{CTF-H}}\to\mathrm{diag}(\bm{u})\cdot\mathbf{M}\cdot\mathrm{diag}(\bm{v})$
+Therefore, both marginal constraints, $\Pi^{*}_{\text{CTF-H}}\mathbf{1}=a$ and $\Pi^{*\top}_{\text{CTF-H}}\mathbf{1}=b$, are satisfied. ∎
+D.2 Proof of Theorem C.1
+To derive Theorem C.1, we make use of the following lemma, which is proven in Appendix D.3.
+Lemma D.1.
 Let $\mathbf{C}\in\mathbb{R}^{n_{0}\times n_{1}}$ be a cost matrix and $\mathbf{M}\in\mathbb{R}^{n_{0}\times n_{1}}$ a prior transition matrix with positive entries.
@@
 $\Pi^{*}$ $={\arg\!\min}_{\Pi\geq 0}\;\sum_{k,l}\Pi_{kl}C_{kl}+\epsilon\sum_{k,l}\Pi_{kl}(\log\!\left(\Pi_{kl}\right)-1).$
-Let $\tilde{\Pi}^{*}$ be the EOT-coupling where the cost is scaled by a normalization constant $c$ or $\tilde{C}_{ij}=\frac{C_{ij}}{c}$.
+Let $\tilde{\Pi}^{*}$ be the EOT-coupling in the case when cost is scaled by a normalization constant $c$ or $\tilde{C}_{ij}=\frac{C_{ij}}{c}$.
 Let the regularization parameter $\epsilon>0$ be the same in both cases.
-Then, for indices $(i,j)$ and $(k,l)$,
-$\frac{\tilde{\Pi}^{*}_{ij}}{\tilde{\Pi}^{*}_{kl}}\leq\gamma\,\left(\frac{\Pi^{*}_{ij}}{\Pi^{*}_{kl}}\right)^{\frac{1}{c}},$
-where $\gamma$ depends on $\Pi^{*}_{ij},c$ and the OT marginal constraints $a,b$.
-Proof.
+Consider the scaling factors $\alpha,\beta$ such that: $\tilde{u}_{i}=\alpha_{i}u_{i}^{1/c}$, $\tilde{v}_{j}=\beta_{j}v_{j}^{1/c}$ where $u,v$ are the Sinkhorn algorithm converged vectors for the original setting and $\tilde{u},\tilde{v}$ are for the cost-scaled version.
+Then, we have
+$\max\{\left\|\phi\right\|_{\infty},\left\|\psi\right\|_{\infty}\}\leq\|M^{-1}\|_{\infty}\cdot\left\|\begin{pmatrix}\Delta_{a}\\ \Delta_{b}\end{pmatrix}\right\|_{\infty}$,
+where $\phi=\log(\alpha)$ and $\psi=\log(\beta)$. We also have that,
+$\max_{i}|\alpha_{i}-1|,\>\max_{i}|\beta_{i}-1|\leq\|M^{-1}\|_{\infty}\cdot\max(\|\Delta_{a}\|_{\infty},\|\Delta_{b}\|_{\infty}),$
+where $M,\Delta_{a},\Delta_{b}$ depend on $\Pi^{*}$, marginalization constants $a,b$ and normalization constant $c$.
+Now, we are ready to prove Theorem C.1.
+Proof of Theorem C.1.
 For the original optimal transport (OT) formulation, we note:
@@
 Thus, there exist scaling factors $\alpha_{i},\beta_{j}\in\mathbb{R}$ such that:
-$\tilde{u}_{i}$ $=\alpha_{i}u^{\frac{1}{c}}_{i},$
-$\tilde{v}_{j}$ $=\beta_{j}v^{\frac{1}{c}}_{j}.$
+$\tilde{u}_{i}=\alpha_{i}u^{\frac{1}{c}}_{i},\quad\tilde{v}_{j}=\beta_{j}v^{\frac{1}{c}}_{j}.$
 This implies:
 $\tilde{\Pi}^{*}_{ij}=(\alpha_{i}u_{i}^{1/c})\tilde{K}_{ij}(\beta_{j}v_{j}^{1/c}),$
-$\implies\tilde{\Pi}^{*}=\operatorname{diag}(\alpha u^{1/c})\tilde{K}\operatorname{diag}(\beta v^{1/c}),$ (g1)
-$\implies\tilde{\Pi}^{*}=\operatorname{diag}(\alpha)\Pi^{1/c}\operatorname{diag}(\beta).$
+$\implies\tilde{\Pi}^{*}=\operatorname{diag}(\alpha u^{1/c})\tilde{K}\operatorname{diag}(\beta v^{1/c}),$ (20)
+$\implies\tilde{\Pi}^{*}=\operatorname{diag}(\alpha)\Pi^{1/c}\operatorname{diag}(\beta),$
 subject to the constraints:
@@
 $\left|\log\left(\frac{\tilde{\Pi}^{*}_{ij}}{\tilde{\Pi}^{*}_{kl}}\right)-\frac{1}{c}\log\left(\frac{\Pi^{*}_{ij}}{\Pi^{*}_{kl}}\right)\right|\leq|\phi_{i}|+|\phi_{k}|+|\psi_{j}|+|\psi_{l}|.$
-From Proposition 3 C.3, we have:
+According to Lemma D.1, we have:
 $\max_{i}\phi_{i}\leq E,\quad\max_{i}\psi_{i}\leq E.$
@@
 $\frac{\tilde{\Pi}^{*}_{ij}}{\tilde{\Pi}^{*}_{kl}}\leq\gamma\left(\frac{\Pi^{*}_{ij}}{\Pi^{*}_{kl}}\right)^{1/c}.$
-Corollary C.2.
-Let $\mathbf{C}\in\mathbb{R}^{n_{0}\times n_{1}}$ be a cost matrix and $\mathbf{M}\in\mathbb{R}^{n_{0}\times n_{1}}$ a prior transition matrix with positive entries.
-Consider the entropy-regularized OT formulation:
-$\Pi^{*}$ $={\arg\!\min}_{\Pi\geq 0}\;\sum_{k,l}\Pi_{kl}C_{kl}+\epsilon\sum_{k,l}\Pi_{kl}(\log\!\left(\Pi_{kl}\right)-1).$
-Let $\tilde{\Pi}^{*}$ be the EOT-coupling in the case when cost is scaled by a normalization constant $c$ or $\tilde{C}_{ij}=\frac{C_{ij}}{c}$.
-Let the regularization parameter $\epsilon>0$ be the same in both cases.
-Then:
-$H(\tilde{\Pi}_{ij})\geq mH(\Pi_{ij})-s,$
-where $m$ and $s$ are constants that depend on $\Pi^{*}$, the marginalization constants $a,b$ and the normalization constant $c$.
-Proof.
-From equation (g1) in Proposition 1 above, we know that:
-$\tilde{\Pi}^{*}_{ij}=(\Pi^{*}_{ij})^{1/c}\cdot\exp(\phi_{i},\psi_{j})$
-and from Proposition 2, we have that,
+Now the remaining task is to prove the second inequality. From Equation 20, we know that:
+$\tilde{\Pi}^{*}_{ij}=(\Pi^{*}_{ij})^{1/c}\cdot\exp(\phi_{i},\psi_{j}),$
+and from Lemma D.1, we have
 $\tilde{\Pi}^{*}_{ij}\leq(\Pi^{*}_{ij})^{1/c}\cdot e^{2E}$
@@
 $H(\tilde{\Pi^{*}})\geq mH(\Pi^{*})-s,$
-where $m=\frac{e^{2E}}{c\Pi^{*}_{\min}}$ and $s=2E\cdot e^{2E}$. ∎
-Proposition C.3.
-Let $\mathbf{C}\in\mathbb{R}^{n_{0}\times n_{1}}$ be a cost matrix and $\mathbf{M}\in\mathbb{R}^{n_{0}\times n_{1}}$ a prior transition matrix with positive entries.
-Consider the entropy-regularized OT formulation:
-$\Pi^{*}$ $={\arg\!\min}_{\Pi\geq 0}\;\sum_{k,l}\Pi_{kl}C_{kl}+\epsilon\sum_{k,l}\Pi_{kl}(\log\!\left(\Pi_{kl}\right)-1).$
-Let $\tilde{\Pi}^{*}$ be the EOT-coupling in the case when cost is scaled by a normalization constant $c$ or $\tilde{C}_{ij}=\frac{C_{ij}}{c}$.
-Let the regularization parameter $\epsilon>0$ be the same in both cases.
-Consider the scaling factors $\alpha,\beta$ such that: $\tilde{u}_{i}=\alpha_{i}u_{i}^{1/c}$, $\tilde{v}_{j}=\beta_{j}v_{j}^{1/c}$ where $u,v$ are the Sinkhorn algorithm converged vectors for the original setting and $\tilde{u},\tilde{v}$ are for the cost-scaled version.
-Then, we have
-$\max\{\left\|\phi\right\|_{\infty},\left\|\psi\right\|_{\infty}\}\leq\|M^{-1}\|_{\infty}\cdot\left\|\begin{pmatrix}\Delta_{a}\\ \Delta_{b}\end{pmatrix}\right\|_{\infty}$,
-where $\phi=\log(\alpha)$ and $\psi=\log(\beta)$. We also have that,
-$\max_{i}|\alpha_{i}-1|,\max_{i}|\beta_{i}-1|\leq\|M^{-1}\|_{\infty}\max(\|\Delta_{a}\|_{\infty},\|\Delta_{b}\|_{\infty}),$
-where $M,\Delta_{a},\Delta_{b}$ depend on $\Pi^{*}$, marginalization constants $a,b$ and normalization constant $c$.
+where $m=\frac{e^{2E}}{c\Pi^{*}_{\min}}$ and $s=2E\cdot e^{2E}$. Thus, we complete the proof. ∎
+D.3 Proof of Lemma D.1
 Proof.
@@
 $|\beta_{j}-1|\approx|\exp(\psi_{j})-1|\approx\psi_{j}.$
-Therefore:
-$\max_{i}|\alpha_{i}-1|,\max_{j}|\beta_{j}-1|\leq\|M^{-1}\|_{\infty}\cdot\max(\|\Delta a\|_{\infty},\|\Delta b\|_{\infty}).$
-Theorem 3.1.
-Let $\mathbf{C}\in\mathbb{R}^{n_{0}\times n_{1}}$ be a general cost matrix and $\mathbf{M}\in\mathbb{R}^{n_{0}\times n_{1}}$ be a prior transition probability matrix.
-Suppose $\Pi^{*}_{\mathrm{CTF-H}}$ is the solution to the following prior-aware optimal transport problem:
-$\Pi^{*}_{\mathrm{CTF-H}}=\operatorname*{arg\,min}_{\Pi\in\mathbb{R}^{n_{0}\times n_{1}}}$ $\sum_{k,l}\Pi_{kl}C_{kl}+\epsilon\sum_{k,l}\Pi_{kl}\log((\Pi_{kl}/{M}_{kl})-1),$
-where $\epsilon>0$ is the regularization parameter.
-Then, we can show that $\Pi^{*}_{\mathrm{CTF-H}}$ can be computed by the Sinkhorn algorithm and takes the form $\mathrm{diag}(\bm{u})\cdot\mathbf{M}\odot\exp(-\mathbf{C}/\epsilon)\cdot\mathrm{diag}(\bm{v})$, where $\odot$ stands for the elementwise multiplication, and $\bm{u}\in\mathbb{R}^{n_{0}},\bm{v}\in\mathbb{R}^{n_{1}}$ are vectors satisfying the marginalization constraints.
-Proof.
-We have that:
-$\Pi^{*}_{\mathrm{CTF-H}}=\operatorname*{arg\,min}_{\Pi\in\mathbb{R}^{n_{0}\times n_{1}}}$ $\sum_{k,l}\Pi_{kl}C_{kl}+\epsilon\sum_{k,l}\Pi_{kl}\log(\Pi_{kl}/{M}_{kl}),$
-Subject to:
-$\Pi\mathbf{1}=a,\quad\Pi^{\top}\mathbf{1}=b.$
-This formulation is a standard convex optimization setting with constraints. The Lagrangian of this setting is:
-$\mathcal{L}(\Pi,f,g)=\sum_{k,l}C_{kl}\Pi_{kl}+\epsilon\sum_{k,l}\Pi_{kl}\left(\log\left(\frac{\Pi_{kl}}{M_{kl}}\right)-1\right)-\sum_{k}f_{k}\left(\sum_{l}\Pi_{kl}-a_{k}\right)-\sum_{l}g_{l}\left(\sum_{k}\Pi_{kl}-b_{l}\right)$
-Differentiating with respect to $\Pi_{kl},f_{k},g_{l}$, we get:
-$\frac{\partial\mathcal{L}}{\partial\Pi_{kl}}=C_{kl}+\epsilon\log\left(\frac{\Pi_{kl}}{M_{kl}}\right)-f_{k}-g_{l}$
-Setting the derivative to zero:
-$\epsilon\log\left(\frac{\Pi^{*}_{kl}}{M_{kl}}\right)=f_{k}-C_{kl}+g_{l}$
-$\implies\frac{\Pi^{*}_{kl}}{M_{kl}}=e^{\frac{f_{k}}{\epsilon}}e^{-\frac{C_{kl}}{\epsilon}}e^{\frac{g_{l}}{\epsilon}}$
-$\implies\Pi^{*}_{kl}=e^{\frac{f_{k}}{\epsilon}}M_{kl}e^{-\frac{C_{kl}}{\epsilon}}e^{\frac{g_{l}}{\epsilon}}$
-Let $u\in\mathbb{R}^{n}$ and $v\in\mathbb{R}^{m}$ such that:
-$u_{k}=e^{\frac{f_{k}}{\epsilon}},\quad v_{l}=e^{\frac{g_{l}}{\epsilon}}$
-Let $K_{kl}$ be the kernel $M_{kl}e^{-C_{kl}/\epsilon}$.
-Then, we have:
-$\Pi^{*}_{kl}=u_{k}K_{kl}v_{l}$
-$\Pi^{*}=\operatorname{diag}(u)\cdot K\cdot\operatorname{diag}(v)$ (15)
-Differentiating the Lagrangian with respect to $f_{k}$ and $g_{l}$, we get:
-$\frac{\partial\mathcal{L}}{\partial f_{k}}=1\cdot\left(\sum_{l}\Pi^{*}_{kl}-a_{k}\right)=0$
-$\implies\Pi^{*}\mathbf{1}=a$ (16)
-$\frac{\partial\mathcal{L}}{\partial g_{l}}=1\cdot\left(\sum_{i}\Pi^{*}_{kl}-b_{l}\right)=0$
-$\implies\Pi^{*\top}\mathbf{1}=b$ (17)
-From equations 16 15, 17 16, and 18 17 above, we get:
-$\operatorname{diag}(u)\cdot K\cdot\operatorname{diag}(v)\cdot\mathbf{1}$ $=a$
-$(\operatorname{diag}(u)\cdot K\cdot\operatorname{diag}(v))^{\top}\mathbf{1}$ $=b$
-Which can be rewritten as:
-$u\odot(Kv)$ $=a$
-$K^{\top}u\odot v$ $=b$
-This is the usual matrix scaling formulation for which the Iterative Proportional Fitting (IPF) updates are:
-$u_{k}^{t+1}=\frac{a_{k}}{(Kv^{t})_{k}},\quad v_{l}^{t+1}=\frac{b_{l}}{(K^{\top}u^{t+1})_{l}}$
-Sinkhorn Algorithm uses these updates, iteratively, and these updates are shown to converge in (Franklin and Lorenz, 1989).
-Thus, Sinkhorn Algorithm can be used for the ContextFlow’s Prior Aware Entropy Regularized (PAER) (CTF-H) formulation.
-From equation (9) 15, we get:
-$\Pi^{*}_{kl}=e^{f_{k}/\epsilon}M_{kl}e^{-C_{kl}/\epsilon}e^{g_{l}/\epsilon}$
-When $\epsilon\to\infty$, we have $C_{kl}/\epsilon\to 0$.
-$\quad e^{-C_{kl}/\epsilon}\to 1$
-$\implies$ $\Pi^{*}_{kl}\to u_{k}M_{kl}v_{l}$
-$\implies$ $\Pi^{*}_{\text{CTF-H}}\to\mathrm{diag}(\bm{u})\cdot\mathbf{M}\cdot\mathrm{diag}(\bm{v})$
-Such that marginal constraints, $\Pi^{*}_{\text{CTF-H}}\mathbf{1}=a$ and $\Pi^{*\top}_{\text{CTF-H}}\mathbf{1}=b$ are satisfied. ∎
-Appendix D Effects of normalization on Prior aware cost matrix
-From (Peyré et al., 2019), we know that optimal MOTFM coupling takes the form $\Pi^{*}_{\text{EOT}}=\text{diag}(u)\cdot K\cdot\text{diag}(v)$, where $K$ is the kernel matrix such that $K]_{ij}=\exp(\frac{-c_{ij}}{\epsilon})$, with $u,v$ satisfying marginalization constraints $u\odot Kv=a$ and $K^{T}u\odot v=b$.
-Sinkhorn updates are given by:
-$u^{l+1}=\frac{a}{Kv^{l}};v^{l+1}=\frac{b}{K^{T}u^{l+1}}.$
-In cases where the OT cost function consists of information from different modalities the distances are usually normalized to have distances of a similar scale.
-Normalizing the cost results $\tilde{c}_{ij}=\frac{c_{ij}}{\epsilon}$ such that the new kernel matrix $K_{\text{norm}}]_{ij}=\exp(\frac{-c_{ij}}{C_{\text{max}}\epsilon})$ can cause numerical issues if $C_{\text{max}}\gg 1$.
-The cost normalization should be performed mindfully, when considering different pairwise distances, as in PACM Section 3.
-Intuitively, scaling the cost has the same effect as that of increasing $\epsilon$, making solutions more diffused.
-Proposition C.1.
-Let $\mathbf{C}\in\mathbb{R}^{n_{0}\times n_{1}}$ be a cost matrix and $\mathbf{M}\in\mathbb{R}^{n_{0}\times n_{1}}$ a prior transition matrix with positive entries.
-Consider the entropy-regularized OT formulation:
-$\Pi^{*}$ $={\arg\!\min}_{\mathbf{\Pi}\geq 0}\;\sum_{k,l}\Pi_{kl}C_{kl}+\epsilon\sum_{k,l}\Pi_{kl}(\log\!\left(\Pi_{kl}\right)-1).$
-Let $\tilde{\Pi}^{*}$ be the EOT-coupling where the cost is scaled by a normalization constant $c$ or $\tilde{C}_{ij}=\frac{C_{ij}}{c}$.
-Let the regularization parameter $\epsilon>0$ be the same in both cases.
-Then, for any indices $(i,j)$ and $(k,l)$ we have
-$\frac{\widetilde{\Pi}^{*}_{ij}}{\widetilde{\Pi}^{*}_{kl}}\leq\gamma\,\left(\frac{\Pi^{*}_{ij}}{\Pi^{*}_{kl}}\right)^{\frac{1}{c}},$
-where $\gamma$ depends on $\Pi^{*}_{ij},c$ and OT marginal constraints $a,b$.
-From Proposition 1, let $\frac{\Pi^{*}_{ij}}{\Pi^{*}_{kl}}=m$, such that $m>1$ ($\Pi^{*}_{ij}>\Pi^{*}_{kl}$ or entries are faraway) then, for $c>1$, we have $\frac{\tilde{\Pi}^{*}_{ij}}{\tilde{\Pi}^{*}_{kl}}<m^{\frac{1}{c}}<m$, for $\gamma<1$, implying that faraway entries are squeezed together.
-This results in bringing probabilities that are far apart closer to each other or, in essence, in creating more diffused and less sharp couplings.
-Proposition C.2.
-Let $\mathbf{C}\in\mathbb{R}^{n_{0}\times n_{1}}$ be a cost matrix and $\mathbf{M}\in\mathbb{R}^{n_{0}\times n_{1}}$ a prior transition matrix with positive entries.
-Consider the entropy-regularized OT formulation:
-$\mathbf{\Pi}^{*}$ $={\arg\!\min}_{\mathbf{\Pi}\geq 0}\;\sum_{k,l}\Pi_{kl}C_{kl}+\epsilon\sum_{k,l}\Pi_{kl}(\log\!\left(\Pi_{kl}\right)-1)$
-and $\tilde{\Pi^{*}}$ be EOT-coupling in the case when cost is scaled by a normalization constant $c$ or $\tilde{C}_{ij}=\frac{C_{ij}}{c}$.
-Let the regularization parameter $\epsilon>0$ be the same in both cases.
-Then we have:
-$H(\tilde{\Pi}_{ij})\geq mH(\Pi_{ij})-s$
-where $m$ and $s$ are constants, that depend on $\Pi^{*}$, marginalization constants $a,b$ and normalization constant $c$.
-Corollary C.2 can also be interpreted as supporting the results of Proposition C.1 and our intuition that normalizing has the same effect on the kernel matrix as increasing $\epsilon$, leading to more diffused couplings or couplings with increased entropy.
+Therefore, we obtain
+$\max_{i}|\alpha_{i}-1|,\max_{j}|\beta_{j}-1|\leq\|M^{-1}\|_{\infty}\cdot\max(\|\Delta a\|_{\infty},\|\Delta b\|_{\infty}),$
+which completes the proof. ∎
 Appendix E Detailed Analysis of ContextFlow
@@
 1:Input: gene data $\{\mathbf{X}_{t_{1}},\cdots,\mathbf{X}_{t_{m+1}}\}$, spatial data $\{\mathbf{S}_{t_{1}},\ldots,\mathbf{S}_{t_{m+1}}\}$, parameters $\lambda$, $\alpha$, $\epsilon$, $\sigma$, $r$, $\eta$
-2:Data-Preprocessing: Compute local neighborhood means using Nearest Neighbor Algorithm and Ligand-Receptor features $f_{\text{LR}}$ using LIANA+ $\triangleright$ As defined in Equation 9 and 10
+2:Data-Preprocessing: Compute local neighborhood means using Nearest Neighbor Algorithm and Ligand-Receptor features $f_{\text{LR}}$ using LIANA+ $\triangleright$ As defined in Equation 4 and 5
 3:Output: neural velocity vector field $u_{\theta}$
@@
 26:end while
-E.2 Time Complexity Analysis
-The training time of ContextFlow is comparable to that of Minibatch-OT FM (Tong et al., 2024) (Figure 3), as both methods solve an entropic variant of optimal transport using the GPU-optimized Sinkhorn algorithm, alongside forward and backward propagation steps that are also GPU-accelerated.
+E.2 Training Details and Hyperparameters
+Similar neural network architectures and training settings are used for all the evaluated baselines and ContextFlow. Specifically, we list the training hyperparameters used in our experiments:
+  • Learning rate: $0.0005$
+  • Learning rate scheduler: Cosine
+  • Epochs: $10000$
+  • NN: MLP
+  • Depth: 44
+  • Width: $100$
+  • Radius of SS prior: $50$
+  • LIANA+ [Dimitrov et al., 2024] settings for LR prior:
+    • Resource: Consensus
+    • nzprop: $0.01$
+    • Similarity Metric: Cosine
+  • Hyperparameter $\epsilon$ used in EOT formulation (Equation 2):
+    • $0.1$ for Brain Regeneration experiments
+    • $10$ for Organogenesis and Liver regeneration
+E.3 Time Complexity Analysis
+The training time of ContextFlow is comparable to that of Minibatch-OT FM [Tong et al., 2024] (Figure 4), as both methods solve an entropic variant of optimal transport using the GPU-optimized Sinkhorn algorithm, alongside forward and backward propagation steps that are also GPU-accelerated.
 Although ContextFlow incorporates prior knowledge, such as spatial smoothness (Equation 4) and cell-cell communication patterns (Equation 5), their corresponding features are computed once during preprocessing, resulting in a one-time cost.
@@
 Data Preprocessing.
-The following preprocessing steps generate additional biologically informed features that complement the original transcriptomic profiles.
+The preprocessing steps generate additional biologically informed features that complement the original transcriptomic profiles.
 These features incur a one-time computational cost and can be reused across different experiments and model configurations.
-Spatial Smoothness (SS).
-We employ a nearest neighbor (NN) algorithm for calculating the mean of local transcriptomic features for each cell.
+For the SS prior, we use a nearest-neighbor (NN) algorithm to compute the mean of local transcriptomic features for each cell.
 The computational complexity of the NN search is known to be $O(N^{2}d)$, where $N$ denotes the total number of points considered and $d$ denotes the data dimension.
-Cell-Cell Communication Patterns (LR).
-We employ spatially informed bivariate statistics implemented in LIANA+ (Dimitrov et al., 2024), for computing LR features, where we applied the cosine similarity metric to gene expression profiles and used the recommended hyperparameters.
+For the LR prior, we employ spatially informed bivariate statistics implemented in LIANA+ [Dimitrov et al., 2024], using the cosine similarity metric on gene expression profiles and the recommended hyperparameters.
 The exact runtime complexity for LIANA+ is unknown.
-Table 6 summarizes the total time taken for the Brain Regeneration and Mouse Organogenesis datasets in our case.
+Table 5 summarizes the total time taken for the Brain Regeneration and Mouse Organogenesis datasets.
 Table 5: Runtime for computing cell-cell communication patterns.
 Dataset Total Number of Cells Runtime (seconds)
-Brain Regeneration (Wei et al., 2022) 28,780 23.35
-Mouse Organogenesis (Chen et al., 2022) 399,248 200.40
+Brain Regeneration [Wei et al., 2022] 28,78028{,}780 $23.35$
+Mouse Organogenesis [Chen et al., 2022] 399,248399{,}248 $200.40$
+Figure 4: Runtime comparisons across various algorithms on Brain Regeneration with batch size $256$.
 Training of ContextFlow.
-The training time complexity largely depends on the total number of training epochs ($E$), mini-batch size ($B$), time per forward and backward pass ($P$), transcriptomic feature dimension ($d$), and total LR pairs ($l$).
-Below, we compute the time complexity for each step in ContextFlow:
-TPM Construction. The construction of the transition plausibility matrix involves a calculation of pairwise distances for each mini-batch, resulting in the runtime of $O(B^{2}(l+d))$.
-Sinkhorn Algorithm.
+The time complexity depends on the total number of training epochs ($E$), mini-batch size ($B$), time per forward and backward pass ($P$), transcriptomic feature dimension ($d$), and the total number of LR pairs ($l$).
+Below, we compute the time complexity for each step.
+The construction of the transition plausibility matrix involves a calculation of pairwise distances for each mini-batch, resulting in the runtime of $O(B^{2}(l+d))$.
 According to Theorem 4.1, we know that Sinkhorn iterations can be adapted to solve the prior-aware entropy regularization problem (Equation 8).
-Since the Sinkhorn algorithm has a well-known quadratic time complexity (Cuturi, 2013), the runtime for computing minibatch OT couplings in ContextFlow is $O(B^{2})$.
-Figure 3: Training time comparisons for different algorithms with a mini-batch size of $256$ on the Brain Regeneration dataset.
-Total Runtime.
+Since the Sinkhorn algorithm has a well-known quadratic time complexity [Cuturi, 2013], the runtime for computing minibatch OT couplings in ContextFlow is $O(B^{2})$.
 Putting pieces together, across all the training epochs, the total runtime complexity of ContextFlow turns out to be $O(E\times(B^{2}(d+l)+P))$.
@@
 Appendix F Spatiotemporal Optimal Transport
-In this section, we compare the recent state-of-the-art spatiotemporal alignment methods, including DeST-OT (Halmos et al., 2025) and TOAST (Ceccarelli et al., 2025), with our prior-aware entropy regularized (PAER) OT objective used in ContextFlow (CTF-H).
+In this section, we compare state-of-the-art spatiotemporal alignment methods, including DeST-OT [Halmos et al., 2025], MOSCOT [Klein et al., 2025] and TOAST [Ceccarelli et al., 2026], with our PAER-OT objective used in the entropic variant of ContextFlow (CTF-H).
 It is important to note that these OT methods are not generative models and are only used for pairwise alignment tasks.
-ContextFlow, on the other hand, is a generative model that learns a dynamic flow across the time horizon and utilizes OT couplings to design better conditional paths for regression.
-In particular, we compute metrics described in DeST-OT on the Axolotl Brain Regeneration dataset following the same setup as used in flow matching.
+Our method, on the other hand, is a generative model that learns a dynamic flow across the time horizon and utilizes OT couplings to design better conditional paths for regression.
+In particular, we compute the metrics described in DeST-OT on the Axolotl Brain Regeneration dataset, using the same setup as in flow matching.
 Specifically, for each time step, we randomly sample a batch of $1000$ cells and compute the corresponding coupling matrix $\mathbf{\Pi}$, which is then used to derive the metrics.
 We use the CTF-H ($\lambda=0.8$) version of ContextFlow for comparison.
-F.1 Metric Comparison
+Table 6: Comparisons in terms of growth distortion metric.
+Static Pair CTF-H OT DeST-OT TOAST MOSCOT
+$(1,2)$ $0.0007$ $0.0000$ $0.0000$ $0.0000$
+$(2,3)$ $0.0042$ $0.0000$ $0.0000$ $0.0000$
+$(3,4)$ $0.0027$ $0.0000$ $0.0000$ $0.0000$
+$(4,5)$ $0.0009$ $0.0000$ $0.0000$ $0.0000$
+Table 7: Comparisons in terms of migration metric.
+Static Pair CTF-H OT DeST-OT TOAST MOSCOT
+$(1,2)$ $231.39$ $308.96$ $793.70$ $232.83$
+$(2,3)$ $289.54$ $551.52$ $1103.28$ $395.15$
+$(3,4)$ $2094.11$ $1015.72$ $2052.65$ $950.87$
+$(4,5)$ $1763.51$ $777.13$ $2257.01$ $536.65$
+Growth Distortion.
 DeST-OT introduces an OT objective for aligning spatial transcriptomic tissue slices from different developmental timesteps, with an emphasis on modeling cell growth and tissue expansion/contraction.
-The growth distortion metric is designed to assess whether the inferred growth pattern aligns with the changes in cell-type abundance across timesteps.
-As shown in Table 7, for the growth distortion metric, we find that our CTF-H OT is competitive with DeST-OT and TOAST, despite DeST-OT being specifically developed with consideration for cell growth.
-Table 7: Comparison on growth distortion.
-Static Pair DeST-OT TOAST CTF-H OT
-$1/2$ $0.0007$ $0.0000$ $0.0000$
-$2/3$ $0.0042$ $0.0000$ $0.0000$
-$3/4$ $0.0027$ $0.0000$ $0.0000$
-$4/5$ $0.0009$ $0.0000$ $0.0000$
-The migration metric is another important metric introduced in DeST-OT, which measures whether the coupling results in realistic cell movements between timesteps.
-As seen in Table 8, DeST-OT achieves the best performance, highlighting the advantage of its growth-aware objective compared to TOAST and CTF-H OT, which do not explicitly model tissue expansion or contraction.
-Table 8: Comparison on migration.
-Static Pair DeST-OT TOAST CTF-H OT
-$1/2$ $308.97$ $793.71$ $780.16$
-$2/3$ $551.53$ $1103.29$ $1037.29$
-$3/4$ $1015.73$ $2052.65$ $1991.43$
-$4/5$ $777.13$ $2257.02$ $2222.07$
-Lastly, we compute how similar the transcriptomic values of coupled cells are using a Coupled Transcriptomic Distance metric, which is defined as $\sum_{k=1}^{N}\sum_{l=1}^{M}\big\|X_{t_{i}}[k,:]-X_{t_{i+1}}[l,:]\big\|^{2}\times\Pi_{i,j}$, where $X_{t_{i}}[k,:]$ represents the transcriptomic feature of cell $k$ from timestep $t_{i}$ and $X_{t_{i+1}}[l,:]$ represents the transcriptomic feature of cell $l$ from timestep $t_{i+1}$, and $\Pi$ is the OT coupling matrix.
-From Table 9, we can observe that CTF-H OT is competitive with both DeST and TOAST.
-Table 9: On coupled transcriptomic distance.
-Static Pair DeST-OT TOAST CTF-H OT
-$1/2$ $32.64$ $34.13$ $33.58$
-$2/3$ $14.22$ $21.87$ $18.01$
-$3/4$ $42.26$ $43.89$ $43.26$
-$4/5$ $18.47$ $20.54$ $20.14$
-F.2 Runtime Analysis with Varying Sample Size
-We also compare the runtime complexity of the above-mentioned OT methods, as shown in Table 10.
-CTF-H OT is the fastest of the three, followed by DeST-OT and TOAST, and is competitive in the metrics above.
+The growth distortion metric assesses whether the inferred growth pattern aligns with changes in cell-type abundance across timesteps.
+Table 6 shows that our CTF-H OT method is competitive, despite DeST-OT being specifically designed with cell growth in mind.
+Migration.
+The migration metric introduced in DeST-OT assesses whether the coupling produces realistic cell movements between timesteps.
+As seen in Table 7, DeST-OT and MOSCOT achieve the best performance overall, highlighting the advantage of its growth-aware objective compared to TOAST and CTF-H OT, which do not explicitly model tissue expansion or contraction.
+Nevertheless, CTF-H OT achieves superior performance, especially for the first two developmental stage pairs.
+Coupled Transcriptomic Distance. Finally, coupled transcriptomic distance computes how similar the transcriptomic values of coupled cells are, which is defined as:
+$\sum_{k=1}^{N}\sum_{l=1}^{M}\big\|X_{t_{i}}[k,:]-X_{t_{i+1}}[l,:]\big\|^{2}\times\Pi_{i,j},$
+where $X_{t_{i}}[k,:]$ represents the transcriptomic feature of cell $k$ from timestep $t_{i}$ and $X_{t_{i+1}}[l,:]$ represents the transcriptomic feature of cell $l$ from timestep $t_{i+1}$, and $\Pi$ is the OT coupling matrix.
+From Table 8, we can observe that CTF-H OT is competitive with DeST, MOSCOT, and TOAST.
+Table 8: Comparisons in terms of coupled transcriptomic distance
+Static Pair CTF-H OT DeST-OT TOAST MOSCOT
+$(1,2)$ $33.58$ $32.64$ $34.13$ $33.07$
+$(2,3)$ $18.01$ $14.22$ $21.87$ $15.63$
+$(3,4)$ $43.26$ $42.26$ $43.89$ $43.19$
+$(4,5)$ $20.14$ $18.47$ $20.54$ $18.40$
+Table 9: Comparisons in terms of runtime (s) with varying batch size.
+Batch Size CTF-H OT DeST-OT MOSCOT TOAST
+$10$ $0.0124$ $0.0114$ $0.5199$ $0.0227$
+$50$ $0.0017$ $0.0351$ $0.8479$ $0.0114$
+$100$ $0.0011$ $0.1016$ $2.3381$ $0.0167$
+$150$ $0.0018$ $0.1778$ $1.1979$ $0.0247$
+$200$ $0.0034$ $0.3521$ $1.5328$ $0.0337$
+$250$ $0.0025$ $0.3836$ $1.9100$ $0.0426$
+$300$ $0.0035$ $0.5323$ $1.0772$ $0.0543$
+$350$ $0.0044$ $0.5935$ $1.2983$ $0.0656$
+$400$ $0.0047$ $0.6217$ $1.1782$ $0.0817$
+$450$ $0.0051$ $0.7806$ $1.2357$ $0.0983$
+$500$ $0.0048$ $0.9984$ $1.2243$ $0.1197$
+$550$ $0.0054$ $1.3473$ $1.4695$ $0.1468$
+$600$ $0.0073$ $1.6192$ $1.6680$ $0.1905$
+$650$ $0.0070$ $1.7771$ $2.2559$ $0.2077$
+$700$ $0.0090$ $2.2104$ $1.9901$ $0.2309$
+$750$ $0.0089$ $2.9324$ $2.7181$ $0.2574$
+$800$ $0.0106$ $3.2879$ $1.9840$ $0.3001$
+$850$ $0.0117$ $5.0941$ $2.3511$ $0.3229$
+$900$ $0.0128$ $5.6585$ $2.4864$ $0.3798$
+$950$ $0.0150$ $5.3771$ $3.0964$ $0.4206$
+$1000$ $0.0146$ $5.9733$ $3.3683$ $0.4931$
+Runtime.
+To complement Figure 3(d), we report the runtime complexity of the above-mentioned OT methods in Table 9.
+CTF-H OT is the fastest of the three, followed by TOAST, MOSCOT, and DeST-OT, and is competitive in the metrics above.
 We also observe that DeST-OT is the slowest, as expected, since its OT objective includes a Gromov-Wasserstein term, which has $O(n^{3})$ runtime, along with other growth- and tissue-distortion-specific terms.
-Table 10: Runtime (s) with varying sample size.
-Sample Size DeST-OT TOAST CTF-OT
-$10$ $0.0111$ $0.0227$ $0.0009$
-$50$ $0.0342$ $0.0114$ $0.0018$
-$100$ $0.1892$ $0.0167$ $0.0024$
-$150$ $0.4035$ $0.0247$ $0.0053$
-$200$ $0.4913$ $0.0337$ $0.0059$
-$250$ $0.4571$ $0.0426$ $0.0061$
-$300$ $0.6252$ $0.0543$ $0.0074$
-$350$ $0.6974$ $0.0656$ $0.0078$
-$400$ $0.8612$ $0.0817$ $0.0117$
-$450$ $1.1028$ $0.0983$ $0.0110$
-$500$ $1.5478$ $0.1197$ $0.0107$
-$550$ $1.7448$ $0.1468$ $0.0165$
-$600$ $1.9295$ $0.1905$ $0.0209$
-$650$ $2.1282$ $0.2077$ $0.0201$
-$700$ $2.7013$ $0.2309$ $0.0235$
-$750$ $3.2951$ $0.2574$ $0.0327$
-$800$ $4.0964$ $0.3001$ $0.0382$
-$850$ $4.2001$ $0.3229$ $0.0339$
-$900$ $4.4582$ $0.3798$ $0.0483$
-$950$ $5.0965$ $0.4206$ $0.0509$
-$1000$ $6.1452$ $0.4931$ $0.0375$
-Tables 7-10 demonstrate that the design choices of ContextFlow enable it to be highly scalable compared to state-of-the-art spatiotemporal alignment methods, while remaining competitive across several spatiotemporal OT alignment metrics.
-(a) MOTFM Implausible Transitions
-(b) ContextFlow Implausible Transitions
-Figure 4: Comparison of biologically implausible cell type couplings between Stage 0 and Stage 1 of the Brain Regeneration Dataset (Wei et al., 2022), under the Entropic-OT and ContextFlow Regularized-OT formulations.
-Biological implausibility is defined here as transitions involving excitatory–inhibitory lineage switches.
-Our formulation produces substantially fewer biologically implausible couplings (24) compared to MOTFM (54).
-Appendix G Biological Priors and Dataset Visualizations
+To summarize, our evaluations above demonstrate that ContextFlow’s design choices enable it to be highly scalable compared to state-of-the-art methods, while remaining competitive across several spatiotemporal OT alignment metrics.
+Figure 5: Visualization of biologically implausible cell type couplings resulted from the PAER-OT plan used in ContextFlow on the Brain Regeneration dataset, where implausibility refers to transitions involving excitatory-inhibitory lineage switches.
+This result is in sharp contrast to Figure 1 with 55 percent decrease in implausible couplings in this case.
+Appendix G Visualizations
 G.1 (Im-)Plausibility of OT-Couplings
-To demonstrate the need of integrating biological priors within a generative framework, we computed the Entropic-OT plan (Section 2.3) for the MOTFM framework and the PAER-OT plan (Section 3.3) for the ContextFlow framework.
-From these transport plans, we sampled couplings corresponding to the first two stages of the Brain Regeneration dataset (Wei et al., 2022) together with their associated cell types.
-Figures 4(a) and 4(b) illustrate the Excitatory–Inhibitory lineage switches present in these sampled couplings.
+To demonstrate the need for integrating biological priors within a generative framework, we compute the EOT plan (Equation 2) for the MOTFM framework and the PAER-OT plan (Section 4.2) for the ContextFlow framework.
+From these transport plans, we sample couplings corresponding to the first two stages of the Brain Regeneration dataset [Wei et al., 2022] with their associated cell types.
+Figures 1 and 5 illustrate the excitatory–inhibitory lineage switches present in these couplings.
 Since excitatory and inhibitory neurons have mutually exclusive neurotransmitter functions and originate from distinct progenitor populations with different transcription factor profiles, a transition from excitatory to inhibitory identity is considered biologically implausible.
-In our transport plan couplings, we observed the following cell type lineage switches: (i) Immature MSN $\rightarrow$ Immature nptxEX, (ii) Immature MSN $\rightarrow$ Immature dpEX, (iii) Immature MSN $\rightarrow$ Immature CMPN, (iv) Immature nptxEX $\rightarrow$ Immature cckIN, and (v) Immature nptxEX $\rightarrow$ Immature MSN
-Of these, $54$ implausible transitions arose from the Entropic-OT plan, compared to $24$ from the PAER-OT plan, with the specific transitions detailed in the figure legends.
-We also observed that the Entropic-OT formulation produced implausible transitions across brain hemispheres, for example, coupling cells from the left hemisphere with those from the right.
+In our transport plan couplings, we observed the following cell type lineage switches: (i) Immature MSN $\rightarrow$ Immature nptxEX, (ii) Immature MSN $\rightarrow$ Immature dpEX, (iii) Immature MSN $\rightarrow$ Immature CMPN, (iv) Immature nptxEX $\rightarrow$ Immature cckIN, and (v) Immature nptxEX $\rightarrow$ Immature MSN.
+Of these, $54$ implausible transitions arise from the Entropic-OT plan, compared to $24$ from PAER-OT, with the specific transitions detailed in the figure legends.
+We also observed that the EOT formulation produced implausible transitions across brain hemispheres, for example, coupling cells from the left hemisphere with those from the right.
 In contrast, the PAER-OT formulation typically restricted transitions to within the same hemisphere, reflecting its integration of spatially aware contextual information.
 These observations provide strong motivation for incorporating biological priors through ContextFlow as a principled approach to learning biologically consistent developmental trajectories.
-G.2 Cell type distributions over time
+G.2 Cell Type Distributions across Datasets
 Figures 68 present spatial maps of transcriptomic datasets across different time points, illustrating how tissue organization and cell-type distributions evolve during development and regeneration.
@@
 Figure 8: Temporal progression of spatial distribution of fibrogenic states for Liver Regeneration. Here, $0/1$ refers to the absence or presence of fibrogenic spots.
-G.3 Ligand Receptor Interactions
-Figure 8 shows the ligand-receptor score of the NPTX2-NPTXR pair in two consecutive slides from the Brain regeneration dataset (Wei et al., 2022).
-Similar activities are visible bilaterally in the cerebral cortex, suggesting that ligand–receptor interactions are preserved across time and spatially aligned with underlying tissue structure.
+G.3 Ligand-Receptor Interactions
+Figure 9 illustrates the ligand-receptor score of the NPTX2-NPTXR pair in two consecutive slides from the Brain regeneration dataset.
+Similar activities are bilaterally visible in the cerebral cortex, suggesting that ligand–receptor interactions are preserved over time and spatially aligned with underlying tissue structure.
 This observation provides strong evidence that including LR interactions as contextual priors is biologically meaningful, as they capture functional communication signals between cells that remain stable across short time intervals.
+Based on the activation of NPTX2–NPTXR in Figure 9, we observe that the communication pattern naturally biases the optimal couplings towards transitions, such as Immature dpEX $\rightarrow$ dpEX and Immature nptxEX $\rightarrow$ nptxEX (Figure 10).
+These transitions are biologically plausible, as they preserve cell type identity within excitatory neuronal lineages while reflecting maturation within the same functional context.
+This example highlights the richness of the contextual information captured by our proposed biological prior and demonstrates how incorporating such ligand–receptor–driven cues into the coupling process yields more interpretable, biologically consistent trajectories.
 (a) NPTX2-NPTXR LR pair activation on Stage 3
@@
 Figure 9: Spatial distributions of LR activation for NPTX2-NPTXR in two consecutive slides from the Brain regeneration dataset. Similar activations are visible at structurally equal positions.
-Based on the activation of NPTX2–NPTXR in Figure 8, we observe that the corresponding communication pattern naturally biases the optimal couplings towards transitions such as Immature dpEX $\rightarrow$ dpEX and Immature nptxEX $\rightarrow$ nptxEX (Figure 9).
-These transitions are biologically plausible, as they preserve cell type identity within excitatory neuronal lineages while reflecting maturation within the same functional context.
-This example highlights the richness of the contextual information captured by our proposed biological prior and demonstrates how incorporating such ligand–receptor–driven cues into the coupling process yields more interpretable, biologically consistent trajectories.
-Figure 9: Visual translation of the bias that NPTX2–NPTXR LR pattern provides in terms of cell type coupling for two consecutive slides.
-Appendix H Additional Experiments & Ablations
-H.1 Next Step Sampling for Axolotl Brain Regeneration
-Table 11: Interpolation via Next Step Sampling at holdout time 3 for the Brain Regeneration dataset.
-Sampling Method $\bm{\lambda}$ $\bm{\alpha}$ Weighted $\bm{\mathcal{W}_{2}}$ $\bm{\mathcal{W}_{2}}$ MMD Energy
-Next Step CFM $2.618\pm 0.142$ $2.579\pm 0.197$ $0.043\pm 0.003$ $12.505\pm 1.271$
-MOTFM $2.567\pm 0.088$ $2.476\pm 0.161$ $0.040\pm 0.003$ $11.269\pm 1.388$
-CTF-C 11 $0.2$ $2.503\pm 0.071$ $2.425\pm 0.239$ $0.037\pm 0.003$ $9.868\pm 1.293$
-11 $0.5$ $2.467\pm 0.107$ $2.301\pm 0.163$ $0.037\pm 0.002$ $9.532\pm 1.093$
-11 $0.8$ $423\pm 0.164 2.2$$293\pm 0.103$ $0.037\pm 0.001$ $9.874\pm 0.659$
-0 $0.2$ $2.396\pm 0.028$ $2.100\pm 0.102$ $0.033\pm 0.003$ $8.577\pm 0.976$
-0 $0.5$ $2.447\pm 0.142$ $2.337\pm 0.216$ $0.036\pm 0.005$ $9.696\pm 1.882$
-0 $0.8$ $413\pm 0.099 2.2$$293\pm 0.161$ $0.036\pm 0.002$ $9.114\pm 1.092$
-$0.5$ $0.2$ $2.460\pm 0.118$ $2.342\pm 0.144$ $0.036\pm 0.003$ $9.500\pm 1.067$
-$0.5$ $0.5$ $2.504\pm 0.094$ $2.309\pm 0.139$ $0.036\pm 0.003$ $9.394\pm 1.431$
-$0.5$ $0.8$ $2.442\pm 0.173$ $2.353\pm 0.241$ $0.035\pm 0.004$ $9.008\pm 2.094$
-CTF-H 0 $2.528\pm 0.143$ $2.534\pm 0.180$ $0.040\pm 0.004$ $11.192\pm 1.304$
-11 $\bm{2.316\pm 0.141}$ $\bm{1.969\pm 0.221}$ $\bm{0.030\pm 0.004}$ $\bm{6.359\pm 1.336}$
-$0.5$ $2.519\pm 0.167$ $2.412\pm 0.158$ $0.039\pm 0.004$ $10.304\pm 1.808$
-Table 12: Extrapolation via Next Step Sampling at holdout time 5 for the Brain Regeneration dataset.
-Sampling Method $\bm{\lambda}$ $\bm{\alpha}$ Weighted $\bm{\mathcal{W}_{2}}$ $\bm{\mathcal{W}_{2}}$ MMD Energy
-Next Step CFM $7.124\pm 0.443$ $7.133\pm 0.533$ $0.276\pm 0.011$ $76.947\pm 5.661$
-MOTFM $7.487\pm 0.698$ $7.449\pm 0.931$ $0.266\pm 0.010$ $81.965\pm 9.812$
-CTF-C 11 $0.2$ $7.257\pm 0.597$ $7.077\pm 0.473$ $0.257\pm 0.004$ $79.562\pm 7.787$
-11 $0.5$ $68\pm 0.608 6.969$$69\pm 0.628$ $0.265\pm 0.009$ $77.025\pm 6.056$
-11 $0.8$ $695\pm 0.443 7.7$$792\pm 0.463$ $0.266\pm 0.007$ $87.179\pm 6.690$
-0 $0.2$ $8.170\pm 0.663$ $8.079\pm 0.723$ $0.269\pm 0.008$ $91.572\pm 8.802$
-0 $0.5$ $7.244\pm 0.804$ $7.146\pm 0.775$ $0.265\pm 0.003$ $80.424\pm 10.376$
-0 $0.8$ $7.382\pm 1.068$ $7.234\pm 0.852$ $0.267\pm 0.009$ $81.635\pm 14.135$
-$0.5$ $0.2$ $94\pm 0.239 7.171$$71\pm 0.422$ $0.266\pm 0.001$ $78.924\pm 3.715$
-$0.5$ $0.5$ $7.188\pm 0.391$ $\bm{6.931\pm 0.260}$ $0.267\pm 0.005$ $78.992\pm 6.195$
-$0.5$ $0.8$ $7.242\pm 0.804$ $7.166\pm 0.980$ $0.267\pm 0.006$ $80.509\pm 10.304$
-CTF-H 0 $\bm{6.914\pm 0.471}$ $7.198\pm 0.726$ $0.266\pm 0.009$ $\bm{76.149\pm 8.436}$
-11 $7.505\pm 0.667$ $7.338\pm 0.601$ $\bm{0.263\pm 0.006}$ $83.425\pm 8.793$
-$0.5$ $7.243\pm 0.479$ $7.157\pm 0.641$ $0.270\pm 0.007$ $79.826\pm 8.067$
-H.2 IVP Sampling on Axolotl Brain Regeneration
-Table 13: Interpolation via IVP Sampling at time point 3 for the Brain Regeneration dataset.
-Sampling Method $\bm{\lambda}$ $\bm{\alpha}$ Weighted $\bm{\mathcal{W}_{2}}$ $\bm{\mathcal{W}_{2}}$ MMD Energy
-IVP CFM $4.216\pm 0.463$ $4.266\pm 0.308$ $0.170\pm 0.029$ $32.413\pm 5.122$
-MOTFM $198\pm 0.319 4.4$$452\pm 0.243$ $0.173\pm 0.017$ $33.149\pm 3.321$
-CTF-C 11 $0.2$ $4.011\pm 0.276$ $4.048\pm 0.321$ $0.147\pm 0.021$ $30.337\pm 4.713$
-11 $0.5$ $3.932\pm 0.377$ $4.356\pm 0.398$ $0.156\pm 0.025$ $31.524\pm 4.875$
-11 $0.8$ $3.603\pm 0.300$ $3.816\pm 0.310$ $0.127\pm 0.018$ $24.271\pm 3.992$
-0 $0.2$ $\bm{3.465\pm 0.232}$ $\bm{3.641\pm 0.320}$ $0.119\pm 0.025$ $23.055\pm 5.939$
-0 $0.5$ $3.943\pm 0.413$ $4.241\pm 0.435$ $0.150\pm 0.039$ $29.221\pm 5.713$
-0 $0.8$ $3.881\pm 0.368$ $4.094\pm 0.551$ $0.139\pm 0.026$ $27.941\pm 6.676$
-$0.5$ $0.2$ $4.152\pm 0.341$ $4.322\pm 0.291$ $0.166\pm 0.014$ $33.299\pm 3.629$
-$0.5$ $0.5$ $4.013\pm 0.187$ $4.138\pm 0.297$ $0.153\pm 0.020$ $30.941\pm 3.685$
-$0.5$ $0.8$ $4.015\pm 0.351$ $3.974\pm 0.442$ $0.140\pm 0.038$ $27.592\pm 6.669$
-CTF-H 0 $3.925\pm 0.267$ $4.375\pm 0.297$ $0.164\pm 0.013$ $32.034\pm 3.270$
-11 $3.905\pm 0.395$ $4.188\pm 0.685$ $\bm{0.074\pm 0.014}$ $\bm{18.728\pm 2.689}$
-$0.5$ $3.917\pm 0.343$ $4.159\pm 0.455$ $0.147\pm 0.022$ $29.613\pm 4.822$
-Table 14: Extrapolation via IVP Sampling at holdout time 5 for the Brain Regeneration dataset.
-Sampling Method $\bm{\lambda}$ $\bm{\alpha}$ Weighted $\bm{\mathcal{W}_{2}}$ $\bm{\mathcal{W}_{2}}$ MMD Energy
-IVP CFM $6.633\pm 1.312$ $7.116\pm 1.084$ $0.143\pm 0.037$ $60.573\pm 21.756$
-MOTFM $6.503\pm 0.720$ $6.352\pm 0.592$ $0.162\pm 0.038$ $56.452\pm 15.932$
-CTF-C 11 $0.2$ $6.403\pm 0.959$ $6.558\pm 1.297$ $0.160\pm 0.024$ $61.051\pm 16.594$
-11 $0.5$ $6.260\pm 0.616$ $7.681\pm 4.003$ $0.157\pm 0.039$ $52.478\pm 12.010$
-11 $0.8$ $6.875\pm 0.643$ $6.920\pm 0.796$ $0.159\pm 0.045$ $62.838\pm 16.897$
-0 $0.2$ $6.722\pm 0.905$ $6.782\pm 1.003$ $0.154\pm 0.034$ $53.996\pm 15.617$
-0 $0.5$ $6.614\pm 0.710$ $6.854\pm 0.740$ $0.201\pm 0.023$ $70.370\pm 9.099$
-0 $0.8$ $6.504\pm 0.925$ $6.744\pm 1.336$ $0.174\pm 0.037$ $56.687\pm 18.118$
-$0.5$ $0.2$ $6.514\pm 0.504$ $5.998\pm 0.803$ $0.155\pm 0.032$ $51.329\pm 15.080$
-$0.5$ $0.5$ $6.696\pm 0.427$ $6.481\pm 0.387$ $0.195\pm 0.024$ $66.212\pm 3.542$
-$0.5$ $0.8$ $6.550\pm 0.975$ $6.563\pm 1.029$ $0.188\pm 0.037$ $63.014\pm 14.173$
-CTF-H 0 $6.243\pm 0.760$ $6.220\pm 0.751$ $0.195\pm 0.020$ $61.316\pm 10.288$
-11 $\bm{5.277\pm 0.936}$ $6.021\pm 1.192$ $\bm{0.099\pm 0.007}$ $\bm{27.777\pm 8.621}$
-$0.5$ $6.254\pm 0.819$ $\bm{5.973\pm 0.757}$ $0.156\pm 0.025$ $54.330\pm 12.089$
-H.3 Next Step Sampling for Mouse Embryo Organogenesis
-Table 15: Interpolation via Next Step Sampling at holdout time 5 for the Mouse Organogenesis dataset.
-Sampling Method $\bm{\lambda}$ $\bm{\alpha}$ Weighted $\bm{\mathcal{W}_{2}}$ $\bm{\mathcal{W}_{2}}$ MMD Energy
-Next Step MOTFM $1.892\pm 0.028$ $1.873\pm 0.086$ $0.164\pm 0.002$ $11.615\pm 0.092$
-CTF-C 11 $0.2$ $1.881\pm 0.020$ $1.922\pm 0.078$ $0.158\pm 0.003$ $11.529\pm 0.197$
-11 $0.5$ $\bm{1.865\pm 0.030}$ $1.852\pm 0.093$ $0.159\pm 0.001$ $11.482\pm 0.108$
-11 $0.8$ $1.889\pm 0.024$ $1.888\pm 0.082$ $0.161\pm 0.002$ $11.552\pm 0.166$
-0 $0.2$ $1.893\pm 0.035$ $1.912\pm 0.057$ $0.159\pm 0.001$ $11.462\pm 0.154$
-0 $0.5$ $1.877\pm 0.039$ $1.933\pm 0.088$ $0.162\pm 0.002$ $11.528\pm 0.110$
-0 $0.8$ $1.882\pm 0.022$ $1.869\pm 0.049$ $0.161\pm 0.001$ $\bm{11.399\pm 0.119}$
-$0.5$ $0.2$ $1.886\pm 0.022$ $1.927\pm 0.111$ $\bm{0.157\pm 0.002}$ $11.430\pm 0.131$
-$0.5$ $0.5$ $1.899\pm 0.027$ $1.899\pm 0.072$ $0.160\pm 0.002$ $11.517\pm 0.097$
-$0.5$ $0.8$ $1.888\pm 0.033$ $\bm{1.839\pm 0.134}$ $0.161\pm 0.002$ $11.475\pm 0.159$
-CTF-H 0 $1.884\pm 0.027$ $1.862\pm 0.123$ $0.164\pm 0.001$ $11.499\pm 0.123$
-11 $1.898\pm 0.029$ $1.866\pm 0.097$ $0.167\pm 0.002$ $11.795\pm 0.170$
-$0.5$ $1.871\pm 0.030$ $1.919\pm 0.067$ $0.164\pm 0.002$ $11.639\pm 0.182$
-Table 16: Extrapolation via Next Step Sampling at holdout time 8 for Mouse Organogenesis.
-Sampling Method $\bm{\lambda}$ $\bm{\alpha}$ Weighted $\bm{\mathcal{W}_{2}}$ $\bm{\mathcal{W}_{2}}$ MMD Energy
-Next Step MOTFM $1.626\pm 0.066$ $1.682\pm 0.096$ $0.084\pm 0.007$ $7.418\pm 0.749$
-CTF-C 11 $0.2$ $1.683\pm 0.058$ $1.803\pm 0.117$ $0.087\pm 0.006$ $7.830\pm 0.551$
-11 $0.5$ $1.685\pm 0.096$ $1.714\pm 0.159$ $0.089\pm 0.006$ $8.056\pm 1.033$
-11 $0.8$ $1.703\pm 0.063$ $1.830\pm 0.131$ $0.095\pm 0.005$ $8.928\pm 0.723$
-0 $0.2$ $1.715\pm 0.123$ $1.860\pm 0.267$ $0.094\pm 0.009$ $9.021\pm 1.740$
-0 $0.5$ $1.725\pm 0.082$ $1.856\pm 0.191$ $0.093\pm 0.006$ $8.806\pm 0.749$
-0 $0.8$ $1.774\pm 0.053$ $1.897\pm 0.175$ $0.094\pm 0.007$ $9.466\pm 0.957$
-$0.5$ $0.2$ $1.818\pm 0.096$ $2.089\pm 0.222$ $0.084\pm 0.008$ $8.875\pm 0.976$
-$0.5$ $0.5$ $1.774\pm 0.104$ $1.899\pm 0.280$ $0.093\pm 0.007$ $9.139\pm 1.437$
-$0.5$ $0.8$ $1.768\pm 0.058$ $1.858\pm 0.120$ $0.101\pm 0.006$ $9.303\pm 0.634$
-CTF-H 0 $\bm{1.505\pm 0.057}$ $\bm{1.397\pm 0.088}$ $0.087\pm 0.005$ $\bm{5.954\pm 0.492}$
-11 $1.890\pm 0.046$ $1.877\pm 0.103$ $0.147\pm 0.006$ $10.752\pm 0.405$
-$0.5$ $1.636\pm 0.060$ $1.684\pm 0.099$ $\bm{0.081\pm 0.005}$ $7.088\pm 0.692$
-H.4 IVP Sampling for Mouse Embryo Organogenesis
-Extrapolating to the last holdout time point of the mouse organogenesis dataset (Chen et al., 2022), particularly under IVP-Sampling, represents the most challenging setting among all our experiments.
-This difficulty arises because the target time point lies entirely outside the training horizon, requiring integration from the initial samples through to the end.
-As a result, the velocity field has more opportunity to drift in incorrect directions, often leading to generations that deviate substantially from the true dynamics.
-In our experiments, this instability was evident: across 10 runs, several produced highly unstable trajectories, reflecting the sensitivity of the system to initial conditions and numerical solvers.
-This variability is also captured in the performance metrics reported in Table 18.
-Table 17: Interpolation via IVP Sampling at holdout time 5 for the Mouse Organogenesis dataset.
-Sampling Method $\bm{\lambda}$ $\bm{\alpha}$ Weighted $\bm{\mathcal{W}_{2}}$ $\bm{\mathcal{W}_{2}}$ MMD Energy
-IVP MOTFM $3.251\pm 0.676$ $3.418\pm 0.727$ $0.090\pm 0.003$ $9.226\pm 0.648$
-CTF-C 11 $0.2$ $3.261\pm 0.880$ $5.264\pm 3.060$ $0.089\pm 0.003$ $10.724\pm 1.288$
-11 $0.5$ $3.137\pm 0.407$ $4.093\pm 1.187$ $0.086\pm 0.004$ $11.948\pm 1.393$
-11 $0.8$ $3.392\pm 0.757$ $4.716\pm 2.079$ $0.089\pm 0.005$ $9.547\pm 0.752$
-0 $0.2$ $2.953\pm 0.425$ $3.816\pm 0.973$ $0.083\pm 0.002$ $9.816\pm 0.715$
-0 $0.5$ $2.938\pm 0.476$ $3.904\pm 1.120$ $0.088\pm 0.005$ $9.864\pm 0.764$
-0 $0.8$ $3.101\pm 0.539$ $3.855\pm 0.946$ $0.087\pm 0.004$ $9.280\pm 0.551$
-$0.5$ $0.2$ $3.771\pm 0.862$ $5.457\pm 1.704$ $\bm{0.079\pm 0.004}$ $9.262\pm 1.134$
-$0.5$ $0.5$ $3.090\pm 0.635$ $4.596\pm 2.357$ $0.084\pm 0.005$ $9.786\pm 1.067$
-$0.5$ $0.8$ $3.200\pm 0.403$ $3.555\pm 0.637$ $0.084\pm 0.004$ $9.269\pm 0.541$
-CTF-H 0 $3.244\pm 0.713$ $3.946\pm 1.671$ $0.089\pm 0.005$ $\bm{8.797\pm 0.612}$
-11 $5.200\pm 0.799$ $6.306\pm 1.037$ $0.123\pm 0.008$ $45.862\pm 13.765$
-$0.5$ $\bm{2.814\pm 0.414}$ $\bm{3.233\pm 0.567}$ $0.093\pm 0.005$ $10.319\pm 0.817$
-Table 18: Extrapolation via IVP Sampling at holdout time 8 for the Mouse Organogenesis dataset.
-Sampling Method $\bm{\lambda}$ $\bm{\alpha}$ Weighted $\bm{\mathcal{W}_{2}}$ $\bm{\mathcal{W}_{2}}$ MMD Energy
-IVP MOTFM $110835\pm 211671$ $1021005\pm 2063905$ $0.086\pm 0.002$ $14178\pm 29475$
-CTF-C 11 $0.2$ $785586\pm 1318212$ $7598321\pm 13497483$ $0.088\pm 0.002$ $98199\pm 150412$
-11 $0.5$ $2691\pm 3931$ $28480\pm 36483$ $0.087\pm 0.002$ $1632\pm 2090$
-11 $0.8$ $2473\pm 3349$ $19537\pm 26306$ $0.087\pm 0.003$ $517\pm 616$
-0 $0.2$ $1493\pm 2497$ $14563\pm 24858$ $0.087\pm 0.001$ $800\pm 1158$
-0 $0.5$ $218018\pm 471298$ $1820788\pm 3994886$ $0.086\pm 0.001$ $2170\pm 4697$
-0 $0.8$ $2736\pm 34766 11$$18089\pm 310135$ $0.084\pm 0.002$ $27013\pm 60065$
-$0.5$ $0.2$ $8114720\pm 16270274$ $69458305\pm 140579849$ $0.088\pm 0.002$ $901074\pm 1775139$
-$0.5$ $0.5$ $2414338\pm 6009993$ $23103811\pm 56863018$ $0.086\pm 0.001$ $261335\pm 663279$
-$0.5$ $0.8$ $158\pm 3023 11$$1138\pm 30025$ $0.084\pm 0.002$ $445\pm 1085$
-CTF-H 0 $353428\pm 952168$ $3011396\pm 8057131$ $0.095\pm 0.004$ $22990\pm 58936$
-11 $\bm{15\pm 10}$ $\bm{53\pm 53}$ $0.098\pm 0.006$ $\bm{48\pm 32}$
-$0.5$ $107889\pm 275882$ $994606\pm 2772756$ $0.087\pm 0.002$ $8875\pm 24264$
-H.5 Liver Regeneration
-Table 19: Wasserstein distances for different model configurations
-Variant $\bm{\lambda}$ $\bm{\alpha}$ $\bm{\mathcal{W}_{2}}$
-EOT $34.30348\pm 1.44797$
-CTF-C 11 $0.2$ $34.44455\pm 1.19306$
-CTF-C 11 $0.5$ $33.95671\pm 1.64415$
-CTF-C 11 $0.8$ $34.62812\pm 0.98181$
-CTF-C 0 $0.2$ $34.24147\pm 1.16930$
-CTF-C 0 $0.5$ $32.74147\pm 1.86351$
-CTF-C 0 $0.8$ $33.71729\pm 1.23057$
-CTF-C $0.5$ $0.2$ $33.56646\pm 1.04376$
-CTF-C $0.5$ $0.5$ $33.84199\pm 1.71408$
-CTF-C $0.5$ $0.8$ $33.04534\pm 1.64399$
-CTF-H 0 $\bm{32.68215\pm 1.47185}$
-CTF-H 11 $33.48050\pm 1.00149$
-CTF-H $0.5$ $33.41444\pm 0.99501$
-H.6 IVP Cell Type Progression over time
-Figure 10: Temporal cell type predictions from ContextFlow for the major cell types in the Organogenesis dataset (Chen et al., 2022).
+Figure 10: Visual translation of the bias of NPTX2–NPTXR LR pattern in cell type coupling.
+G.4 Cell Type Predictions from ContextFlow
+Figure 11: Temporal predictions from ContextFlow under IVP sampling for major cell types on the Mouse Organogenesis dataset.
 Early progenitor populations (neural crest and mesenchyme) progressively diminish as development advances, while terminal fates (muscle, cartilage primordium, and liver) emerge at later stages.
-Major lineages such as brain, heart, and connective tissue remain continuous throughout.
+Major lineages such as the brain, heart, and connective tissue remain continuous throughout.
 Overall, ContextFlow captures biologically coherent and temporally consistent developmental dynamics.
+Appendix H Statistical Significance Testing
+We evaluate the statistical significance of ContextFlow’s performance using a paired two-sided $t$-test over repeated runs.
+For each metric, we compute paired differences between the best (bolded) setting from the result tables from Section 5 for CTF and the corresponding MOTFM baseline, let $\{d_{i}\}_{i=1}^{n}$ be these differences; the test statistic is given by $t=\bar{d}/(s_{d}/\sqrt{n})$, where $\bar{d}$ and $s_{d}$ are the sample mean and standard deviation of the differences.
+The $p$-value measures the likelihood of observing such a deviation in metric values by chance, assuming there is no true difference between the methods; the lower the $p$-value, the better.
+Table 10 and Table 11 report the p-value on the Brain Regeneration and Mouse Organogenesis datasets.
+For Liver regeneration, the p-value is computed as $2.74\mathrm{e}{-02}$.
+Table 10: P-values for interpolation and extrapolation on Brain Regeneration from Tables 1 and 2.
+Sampling Interpolation Extrapolation
+Weighted $\bm{\mathcal{W}_{2}}$ $\bm{\mathcal{W}_{2}}$ MMD Energy Weighted $\bm{\mathcal{W}_{2}}$ $\bm{\mathcal{W}_{2}}$ MMD Energy
+Next Step $1.53\mathrm{e}{-02}$ $3.62\mathrm{e}{-03}$ $5.64\mathrm{e}{-04}$ $3.57\mathrm{e}{-04}$ $1.07\mathrm{e}{-02}$ $2.58\mathrm{e}{-02}$ $8.60\mathrm{e}{-02}$ $1.76\mathrm{e}{-02}$
+IVP $1.05\mathrm{e}{-02}$ $1.57\mathrm{e}{-02}$ $1.83\mathrm{e}{-05}$ $6.01\mathrm{e}{-05}$ $1.53\mathrm{e}{-02}$ $3.29\mathrm{e}{-02}$ $5.13\mathrm{e}{-03}$ $6.33\mathrm{e}{-03}$
+Table 11: P-values for interpolation and extrapolation on Mouse Organogenesis from Table 3.
+Next Step (Interpolation) IVP (Interpolation) Next Step (Extrapolation)
+Weighted $\bm{\mathcal{W}_{2}}$ $\bm{\mathcal{W}_{2}}$ Weighted $\bm{\mathcal{W}_{2}}$ $\bm{\mathcal{W}_{2}}$ Weighted $\bm{\mathcal{W}_{2}}$ $\bm{\mathcal{W}_{2}}$
+$5.70\mathrm{e}{-03}$ $1.06\mathrm{e}{-03}$ $2.63\mathrm{e}{-02}$ $2.43\mathrm{e}{-01}$ $1.96\mathrm{e}{-03}$ $1.30\mathrm{e}{-03}$
 Appendix I Hyperparameter Sensitivity Analysis
 I.1 Ablation on $\lambda$
-First, we conduct a sensitivity analysis on all three datasets for the trade-off parameter $\lambda$, which controls the relative importance of spatial smoothness (SS) and cell-cell communication (LR) priors.
+First, we conduct a sensitivity analysis on all three datasets for the parameter $\lambda$, which controls the relative importance of the two priors.
 All the other hyperparameters are kept constant.
-From Tables 20-24 and their corresponding Figures 11-15, we observe that the best performance is usually achieved towards the extremities, at $\lambda=0$ or $\lambda=1$, with values near the latter dominating more often.
-We hypothesize that the SS prior (at $\lambda=1$ only the SS prior is considered), acting as a proxy for spatial distance between cells from different slices, always carries relevant information, encoding the structural information present in the data.
+From Tables 12-16 and Figures 12-16, we observe that the best performance is usually achieved towards the extremities, at $\lambda=0$ or $\lambda=1$, with values near the latter dominating more often.
+We hypothesize that the spatial smoothness prior serves as a proxy for the spatial distance between cells from different slices and always carries relevant information, encoding the structural information present in the data.
 On the other hand, the informativeness of cell–cell communication patterns depends on how distinct the ligand–receptor features are at a given time step compared with those of neighboring ones.
-When LR features remain highly similar across consecutive timesteps, they contribute little to the discriminative signal that the OT objective can leverage.
+When these features remain highly similar across consecutive timesteps, they contribute little to the discriminative signal that the OT objective can leverage.
 Consequently, the influence of communication priors is strongly dataset- and timestep-dependent.
-This effect is also reflected in the observation that settings with $\lambda=0$ (where only the LR prior is used) tend to perform worse when using Next-Step sampling—where local, immediate effects dominate—than under IVP sampling, which integrates information over the entire preceding trajectory.
+This effect is also reflected in the observation that settings with $\lambda=0$ tend to perform worse when using Next-Step sampling—where local, immediate effects dominate—than under IVP sampling, which integrates information over the entire preceding trajectory.
 We therefore recommend experimenting with different values of $\lambda$ (e.g., 0, $0.8$, or $1$) depending on the specific use case and context.
-Due to the scalability of ContextFlow, hyperparameter exploration can be performed efficiently, allowing for a rapid assessment of the effect of $\lambda$ on model performance.
+Due to ContextFlow’s scalability, hyperparameter exploration can be performed efficiently, enabling rapid assessment of the effect of $\lambda$.
 I.2 Ablation on $r$
-We conduct an ablation study on the Brain Regeneration dataset, examining the effect of varying the neighborhood radius $r$ used to define the boundary for computing the Spatial Smoothness Prior.
-We evaluate two settings: $\lambda=1$ and $\lambda=0.8$, corresponding to the use of only the spatial prior and to a setting with a modest contribution from the cell–cell communication prior, respectively.
-From Tables 25-28 and their corresponding Figures 16-19, we observe that the optimal neighborhood radius tends to lie toward the lower end of the tested range.
+We examine the effect of varying the neighborhood radius $r$ used to define the boundary for computing the spatial smoothness prior.
+We evaluate two settings for Brain Regeneration: $\lambda=1$ and $\lambda=0.8$, corresponding to using only the spatial prior and a setting with a modest contribution from the cell–cell communication prior.
+From Tables 17-20 and their corresponding Figures 17-20, we observe that the optimal neighborhood radius tends to lie toward the lower end of the tested range.
 Radii smaller than this optimum degrade performance by failing to capture sufficient local context, resulting in neighborhood means that are overly similar to individual cellular profiles.
@@
 While certain deviations from this trend occur, likely reflecting underlying biological complexity, this behavior is consistent with the trade-off between spatial specificity and contextual coverage inherent to neighborhood-based priors.
-For our case, we set the radius by considering the timestep with the least number of cells, dividing it by half (to account for different hemispheres), and dividing by the order of cell types present in that timestep.
-For the dataset considered in this study, Stage 44 had the fewest number of $1400$ cells, with approximately $10$ cell types present.
+For our case, we set the radius by considering the timestep with the fewest cells, dividing it by $2$ (to account for different hemispheres), and dividing by the number of cell types present in that timestep.
+For the dataset considered in this study, Stage $44$ had the fewest number of cells, $1400$, with approximately $10$ cell types.
 We thus set the radius at $50$ in our studies.
 I.3 Ablation on $\epsilon$
-Additionally, we conduct an ablation study on the Brain Regeneration dataset by varying the parameter $\epsilon$, which is used to weigh the entropic term in the OT objective.
+Additionally, we conduct an ablation study on the Brain Regeneration dataset by varying the parameter $\epsilon$, which weights the entropic term in the OT objective.
 The ContextFlow configuration we consider here is CTF-H ($\lambda=1$), which includes only the spatial smoothness prior.
-As observed from Tables 29-30 and their corresponding Figures 20-21, drastically decreasing $\epsilon$ results in the OT formulation to ignore the relative entropic term containing the prior information and only to consider the transport cost resulting in higher Wasserstein values.
+As observed from Tables 21 and 22, as well as their corresponding Figures 21 and 22, drastically decreasing $\epsilon$ results in the OT formulation to ignore the relative entropic term containing the prior information and only to consider the transport cost, resulting in higher Wasserstein values.
 Furthermore, in accordance with results from Theorem 4.1, increasing $\epsilon$ too much still does not drastically degrade the performance, as the prior matrix $M$ acts as a soft filter and prohibits uniform couplings.
-While setting $\epsilon$ values, one must look at the Gibbs kernel used in the Sinkhorn Algorithm $\exp(-\mathbf{C}/\epsilon)$, since lower $\epsilon$ values can cause potential numerical issues.
-We thus set $\epsilon$ by examining the order of the cost matrix $\mathbf{C}$, and for the studies above, we set it to $100$ after considering the order of the median of all the elements in the cost matrix.
+While setting $\epsilon$, one must look at the Gibbs kernel used in the Sinkhorn Algorithm $\exp(-\mathbf{C}/\epsilon)$, since a small $\epsilon$ can cause potential numerical issues.
+We thus set $\epsilon$ by examining the order of the cost matrix $\mathbf{C}$, and, for the studies above, we set it to $100$ based on the median of all elements in the cost matrix.
 Table 12: Extrapolation on the last holdout timestep on the Brain Regeneration dataset.
@@
 Weighted $\bm{\mathcal{W}_{2}}$ $\bm{\mathcal{W}_{2}}$ Weighted $\bm{\mathcal{W}_{2}}$ $\bm{\mathcal{W}_{2}}$
-0 $1.508\pm 0.047$ $1.386\pm 0.088$ $6.33\times 10^{5}\pm 1.72\times 10^{6}$ $3.95\times 10^{6}\pm 1.05\times 10^{7}$
-0.2 $1.614\pm 0.081$ $1.642\pm 0.136$ $9.18\times 10^{2}\pm 1.69\times 10^{3}$ $4.29\times 10^{4}\pm 1.05\times 10^{5}$
-0.5 $1.638\pm 0.069$ $1.676\pm 0.114$ $1.08\times 10^{5}\pm 2.76\times 10^{5}$ $9.95\times 10^{5}\pm 2.77\times 10^{6}$
-0.8 $1.617\pm 0.042$ $80\pm 0.094 1.66×1$$6\times 10^{5}\pm 3.92\times 10^{5}$ $1.43\times 10^{6}\pm 3.43\times 10^{6}$
-1 $1.906\pm 0.071$ $1.892\pm 0.092$ $1.89\times 10^{1}\pm 1.77\times 10^{1}$ $6.89\times 10^{1}\pm 7.36\times 10^{1}$
+0 $1.508\pm 0.047$ $1.386\pm 0.088$ $6.986\pm 1.296$ $23.807\pm 7.516$
+0.2 $1.614\pm 0.081$ $1.642\pm 0.136$ $6.907\pm 1.471$ $24.675\pm 9.126$
+0.5 $1.638\pm 0.069$ $1.676\pm 0.114$ $7.749\pm 1.377$ $32.480\pm 8.322$
+0.8 $1.617\pm 0.042$ $1.680\pm 0.094$ $7.645\pm 2.376$ $29.854\pm 10.015$
+1 $1.906\pm 0.071$ $1.892\pm 0.092$ $6.378\pm 1.813$ $11.845\pm 6.567$
 (a) Next Step Sampling
@@
 Figure 15: Performance variation with $\lambda$ for extrapolation on the Mouse Organogenesis dataset.
-Table 24: Interpolation for holdout timestep 3 with IVP Sampling on the Liver Regeneration dataset.
+Table 16: Interpolation for holdout timestep 2 with IVP Sampling on the Liver Regeneration dataset.
 $\bm{\lambda}$ $\bm{\mathcal{W}_{2}}$
@@
 Figure 18: Performance variation with radius for interpolation on the Brain Regeneration dataset.
-Table 27: Extrapolation for the last holdout timestep with CTF-H at $\lambda=0.8$ on the Brain Regeneration dataset.
+Table 19: Extrapolation for the last holdout timestep with CTF-H at $\lambda=0.8$ on Brain Regeneration.
 Radius Next Step Sampling IVP Sampling
@@
 Figure 19: Performance variation with radius for extrapolation on the Brain Regeneration dataset.
-Table 28: Interpolation for the middle holdout timestep 3 with CTF-H at $\lambda=0.8$ on the Brain Regeneration dataset.
+Table 20: Interpolation for the middle holdout timestep with CTF-H at $\lambda=0.8$ on Brain Regeneration.
 Radius Next Step Sampling IVP Sampling
@@
 Figure 22: Performance variation with $\epsilon$ for interpolation on the Brain Regeneration dataset.
+Appendix J Full Ablation Results
+J.1 Next Step Sampling for Axolotl Brain Regeneration
+Table 23: Interpolation via Next Step Sampling at holdout time 3 for the Brain Regeneration dataset.
+Sampling Method $\bm{\lambda}$ $\bm{\alpha}$ Weighted $\bm{\mathcal{W}_{2}}$ $\bm{\mathcal{W}_{2}}$ MMD Energy
+Next Step CFM $2.618\pm 0.142$ $2.579\pm 0.197$ $0.043\pm 0.003$ $12.505\pm 1.271$
+MOTFM $2.567\pm 0.088$ $2.476\pm 0.161$ $0.040\pm 0.003$ $11.269\pm 1.388$
+CTF-C 11 $0.2$ $2.503\pm 0.071$ $2.425\pm 0.239$ $0.037\pm 0.003$ $9.868\pm 1.293$
+11 $0.5$ $2.467\pm 0.107$ $2.301\pm 0.163$ $0.037\pm 0.002$ $9.532\pm 1.093$
+11 $0.8$ $423\pm 0.164 2.2$$293\pm 0.103$ $0.037\pm 0.001$ $9.874\pm 0.659$
+0 $0.2$ $2.396\pm 0.028$ $2.100\pm 0.102$ $0.033\pm 0.003$ $8.577\pm 0.976$
+0 $0.5$ $2.447\pm 0.142$ $2.337\pm 0.216$ $0.036\pm 0.005$ $9.696\pm 1.882$
+0 $0.8$ $413\pm 0.099 2.2$$293\pm 0.161$ $0.036\pm 0.002$ $9.114\pm 1.092$
+$0.5$ $0.2$ $2.460\pm 0.118$ $2.342\pm 0.144$ $0.036\pm 0.003$ $9.500\pm 1.067$
+$0.5$ $0.5$ $2.504\pm 0.094$ $2.309\pm 0.139$ $0.036\pm 0.003$ $9.394\pm 1.431$
+$0.5$ $0.8$ $2.442\pm 0.173$ $2.353\pm 0.241$ $0.035\pm 0.004$ $9.008\pm 2.094$
+CTF-H 0 $2.528\pm 0.143$ $2.534\pm 0.180$ $0.040\pm 0.004$ $11.192\pm 1.304$
+11 $\bm{2.316\pm 0.141}$ $\bm{1.969\pm 0.221}$ $\bm{0.030\pm 0.004}$ $\bm{6.359\pm 1.336}$
+$0.5$ $2.519\pm 0.167$ $2.412\pm 0.158$ $0.039\pm 0.004$ $10.304\pm 1.808$
+Table 24: Extrapolation via Next Step Sampling at holdout time 5 for the Brain Regeneration dataset.
+Sampling Method $\bm{\lambda}$ $\bm{\alpha}$ Weighted $\bm{\mathcal{W}_{2}}$ $\bm{\mathcal{W}_{2}}$ MMD Energy
+Next Step CFM $7.124\pm 0.443$ $7.133\pm 0.533$ $0.276\pm 0.011$ $76.947\pm 5.661$
+MOTFM $7.487\pm 0.698$ $7.449\pm 0.931$ $0.266\pm 0.010$ $81.965\pm 9.812$
+CTF-C 11 $0.2$ $7.257\pm 0.597$ $7.077\pm 0.473$ $0.257\pm 0.004$ $79.562\pm 7.787$
+11 $0.5$ $68\pm 0.608 6.969$$69\pm 0.628$ $0.265\pm 0.009$ $77.025\pm 6.056$
+11 $0.8$ $695\pm 0.443 7.7$$792\pm 0.463$ $0.266\pm 0.007$ $87.179\pm 6.690$
+0 $0.2$ $8.170\pm 0.663$ $8.079\pm 0.723$ $0.269\pm 0.008$ $91.572\pm 8.802$
+0 $0.5$ $7.244\pm 0.804$ $7.146\pm 0.775$ $0.265\pm 0.003$ $80.424\pm 10.376$
+0 $0.8$ $7.382\pm 1.068$ $7.234\pm 0.852$ $0.267\pm 0.009$ $81.635\pm 14.135$
+$0.5$ $0.2$ $94\pm 0.239 7.171$$71\pm 0.422$ $0.266\pm 0.001$ $78.924\pm 3.715$
+$0.5$ $0.5$ $7.188\pm 0.391$ $\bm{6.931\pm 0.260}$ $0.267\pm 0.005$ $78.992\pm 6.195$
+$0.5$ $0.8$ $7.242\pm 0.804$ $7.166\pm 0.980$ $0.267\pm 0.006$ $80.509\pm 10.304$
+CTF-H 0 $\bm{6.914\pm 0.471}$ $7.198\pm 0.726$ $0.266\pm 0.009$ $\bm{76.149\pm 8.436}$
+11 $7.505\pm 0.667$ $7.338\pm 0.601$ $\bm{0.263\pm 0.006}$ $83.425\pm 8.793$
+$0.5$ $7.243\pm 0.479$ $7.157\pm 0.641$ $0.270\pm 0.007$ $79.826\pm 8.067$
+J.2 IVP Sampling on Axolotl Brain Regeneration
+Table 25: Interpolation via IVP Sampling at time point 3 for the Brain Regeneration dataset.
+Sampling Method $\bm{\lambda}$ $\bm{\alpha}$ Weighted $\bm{\mathcal{W}_{2}}$ $\bm{\mathcal{W}_{2}}$ MMD Energy
+IVP CFM $4.216\pm 0.463$ $4.266\pm 0.308$ $0.170\pm 0.029$ $32.413\pm 5.122$
+MOTFM $198\pm 0.319 4.4$$452\pm 0.243$ $0.173\pm 0.017$ $33.149\pm 3.321$
+CTF-C 11 $0.2$ $4.011\pm 0.276$ $4.048\pm 0.321$ $0.147\pm 0.021$ $30.337\pm 4.713$
+11 $0.5$ $3.932\pm 0.377$ $4.356\pm 0.398$ $0.156\pm 0.025$ $31.524\pm 4.875$
+11 $0.8$ $3.603\pm 0.300$ $3.816\pm 0.310$ $0.127\pm 0.018$ $24.271\pm 3.992$
+0 $0.2$ $\bm{3.465\pm 0.232}$ $\bm{3.641\pm 0.320}$ $0.119\pm 0.025$ $23.055\pm 5.939$
+0 $0.5$ $3.943\pm 0.413$ $4.241\pm 0.435$ $0.150\pm 0.039$ $29.221\pm 5.713$
+0 $0.8$ $3.881\pm 0.368$ $4.094\pm 0.551$ $0.139\pm 0.026$ $27.941\pm 6.676$
+$0.5$ $0.2$ $4.152\pm 0.341$ $4.322\pm 0.291$ $0.166\pm 0.014$ $33.299\pm 3.629$
+$0.5$ $0.5$ $4.013\pm 0.187$ $4.138\pm 0.297$ $0.153\pm 0.020$ $30.941\pm 3.685$
+$0.5$ $0.8$ $4.015\pm 0.351$ $3.974\pm 0.442$ $0.140\pm 0.038$ $27.592\pm 6.669$
+CTF-H 0 $3.925\pm 0.267$ $4.375\pm 0.297$ $0.164\pm 0.013$ $32.034\pm 3.270$
+11 $3.905\pm 0.395$ $4.188\pm 0.685$ $\bm{0.074\pm 0.014}$ $\bm{18.728\pm 2.689}$
+$0.5$ $3.917\pm 0.343$ $4.159\pm 0.455$ $0.147\pm 0.022$ $29.613\pm 4.822$
+Table 26: Extrapolation via IVP Sampling at holdout time 5 for the Brain Regeneration dataset.
+Sampling Method $\bm{\lambda}$ $\bm{\alpha}$ Weighted $\bm{\mathcal{W}_{2}}$ $\bm{\mathcal{W}_{2}}$ MMD Energy
+IVP CFM $6.633\pm 1.312$ $7.116\pm 1.084$ $0.143\pm 0.037$ $60.573\pm 21.756$
+MOTFM $6.503\pm 0.720$ $6.352\pm 0.592$ $0.162\pm 0.038$ $56.452\pm 15.932$
+CTF-C 11 $0.2$ $6.403\pm 0.959$ $6.558\pm 1.297$ $0.160\pm 0.024$ $61.051\pm 16.594$
+11 $0.5$ $6.260\pm 0.616$ $7.681\pm 4.003$ $0.157\pm 0.039$ $52.478\pm 12.010$
+11 $0.8$ $6.875\pm 0.643$ $6.920\pm 0.796$ $0.159\pm 0.045$ $62.838\pm 16.897$
+0 $0.2$ $6.722\pm 0.905$ $6.782\pm 1.003$ $0.154\pm 0.034$ $53.996\pm 15.617$
+0 $0.5$ $6.614\pm 0.710$ $6.854\pm 0.740$ $0.201\pm 0.023$ $70.370\pm 9.099$
+0 $0.8$ $6.504\pm 0.925$ $6.744\pm 1.336$ $0.174\pm 0.037$ $56.687\pm 18.118$
+$0.5$ $0.2$ $6.514\pm 0.504$ $5.998\pm 0.803$ $0.155\pm 0.032$ $51.329\pm 15.080$
+$0.5$ $0.5$ $6.696\pm 0.427$ $6.481\pm 0.387$ $0.195\pm 0.024$ $66.212\pm 3.542$
+$0.5$ $0.8$ $6.550\pm 0.975$ $6.563\pm 1.029$ $0.188\pm 0.037$ $63.014\pm 14.173$
+CTF-H 0 $6.243\pm 0.760$ $6.220\pm 0.751$ $0.195\pm 0.020$ $61.316\pm 10.288$
+11 $\bm{5.277\pm 0.936}$ $6.021\pm 1.192$ $\bm{0.099\pm 0.007}$ $\bm{27.777\pm 8.621}$
+$0.5$ $6.254\pm 0.819$ $\bm{5.973\pm 0.757}$ $0.156\pm 0.025$ $54.330\pm 12.089$
+J.3 Next Step Sampling for Mouse Embryo Organogenesis
+Table 27: Interpolation via Next Step Sampling at holdout time 5 for the Organogenesis dataset.
+Sampling Method $\bm{\lambda}$ $\bm{\alpha}$ Weighted $\bm{\mathcal{W}_{2}}$ $\bm{\mathcal{W}_{2}}$ MMD Energy
+Next Step MOTFM $1.892\pm 0.028$ $1.873\pm 0.086$ $0.164\pm 0.002$ $11.615\pm 0.092$
+CTF-C 11 $0.2$ $1.881\pm 0.020$ $1.922\pm 0.078$ $0.158\pm 0.003$ $11.529\pm 0.197$
+11 $0.5$ $\bm{1.865\pm 0.030}$ $1.852\pm 0.093$ $0.159\pm 0.001$ $11.482\pm 0.108$
+11 $0.8$ $1.889\pm 0.024$ $1.888\pm 0.082$ $0.161\pm 0.002$ $11.552\pm 0.166$
+0 $0.2$ $1.893\pm 0.035$ $1.912\pm 0.057$ $0.159\pm 0.001$ $11.462\pm 0.154$
+0 $0.5$ $1.877\pm 0.039$ $1.933\pm 0.088$ $0.162\pm 0.002$ $11.528\pm 0.110$
+0 $0.8$ $1.882\pm 0.022$ $1.869\pm 0.049$ $0.161\pm 0.001$ $\bm{11.399\pm 0.119}$
+$0.5$ $0.2$ $1.886\pm 0.022$ $1.927\pm 0.111$ $\bm{0.157\pm 0.002}$ $11.430\pm 0.131$
+$0.5$ $0.5$ $1.899\pm 0.027$ $1.899\pm 0.072$ $0.160\pm 0.002$ $11.517\pm 0.097$
+$0.5$ $0.8$ $1.888\pm 0.033$ $\bm{1.839\pm 0.134}$ $0.161\pm 0.002$ $11.475\pm 0.159$
+CTF-H 0 $1.884\pm 0.027$ $1.862\pm 0.123$ $0.164\pm 0.001$ $11.499\pm 0.123$
+11 $1.898\pm 0.029$ $1.866\pm 0.097$ $0.167\pm 0.002$ $11.795\pm 0.170$
+$0.5$ $1.871\pm 0.030$ $1.919\pm 0.067$ $0.164\pm 0.002$ $11.639\pm 0.182$
+Table 28: Extrapolation via Next Step Sampling at holdout time 8 for the Organogenesis dataset.
+Sampling Method $\bm{\lambda}$ $\bm{\alpha}$ Weighted $\bm{\mathcal{W}_{2}}$ $\bm{\mathcal{W}_{2}}$ MMD Energy
+Next Step MOTFM $1.626\pm 0.066$ $1.682\pm 0.096$ $0.084\pm 0.007$ $7.418\pm 0.749$
+CTF-C 11 $0.2$ $1.683\pm 0.058$ $1.803\pm 0.117$ $0.087\pm 0.006$ $7.830\pm 0.551$
+11 $0.5$ $1.685\pm 0.096$ $1.714\pm 0.159$ $0.089\pm 0.006$ $8.056\pm 1.033$
+11 $0.8$ $1.703\pm 0.063$ $1.830\pm 0.131$ $0.095\pm 0.005$ $8.928\pm 0.723$
+0 $0.2$ $1.715\pm 0.123$ $1.860\pm 0.267$ $0.094\pm 0.009$ $9.021\pm 1.740$
+0 $0.5$ $1.725\pm 0.082$ $1.856\pm 0.191$ $0.093\pm 0.006$ $8.806\pm 0.749$
+0 $0.8$ $1.774\pm 0.053$ $1.897\pm 0.175$ $0.094\pm 0.007$ $9.466\pm 0.957$
+$0.5$ $0.2$ $1.818\pm 0.096$ $2.089\pm 0.222$ $0.084\pm 0.008$ $8.875\pm 0.976$
+$0.5$ $0.5$ $1.774\pm 0.104$ $1.899\pm 0.280$ $0.093\pm 0.007$ $9.139\pm 1.437$
+$0.5$ $0.8$ $1.768\pm 0.058$ $1.858\pm 0.120$ $0.101\pm 0.006$ $9.303\pm 0.634$
+CTF-H 0 $\bm{1.505\pm 0.057}$ $\bm{1.397\pm 0.088}$ $0.087\pm 0.005$ $\bm{5.954\pm 0.492}$
+11 $1.890\pm 0.046$ $1.877\pm 0.103$ $0.147\pm 0.006$ $10.752\pm 0.405$
+$0.5$ $1.636\pm 0.060$ $1.684\pm 0.099$ $\bm{0.081\pm 0.005}$ $7.088\pm 0.692$
+J.4 IVP Sampling for Mouse Embryo Organogenesis
+Extrapolating to the last holdout time point of the mouse organogenesis dataset [Chen et al., 2022], particularly under IVP-Sampling, represents the most challenging setting among all our experiments.
+This difficulty arises because the target time point lies entirely outside the training horizon, requiring integration from the initial samples through to the end.
+As a result, the velocity field has more opportunity to drift in incorrect directions, often leading to generations that deviate substantially from the true dynamics.
+In our experiments, this instability was evident: across 10 runs, several produced highly unstable trajectories, reflecting the sensitivity of the system to initial conditions and numerical solvers.
+This variability is also captured in the performance metrics reported in Table 30.
+Table 29: Interpolation via IVP Sampling at holdout time 5 for the Mouse Organogenesis dataset.
+Sampling Method $\bm{\lambda}$ $\bm{\alpha}$ Weighted $\bm{\mathcal{W}_{2}}$ $\bm{\mathcal{W}_{2}}$ MMD Energy
+IVP MOTFM $3.251\pm 0.676$ $3.418\pm 0.727$ $0.090\pm 0.003$ $9.226\pm 0.648$
+CTF-C 11 $0.2$ $3.261\pm 0.880$ $5.264\pm 3.060$ $0.089\pm 0.003$ $10.724\pm 1.288$
+11 $0.5$ $3.137\pm 0.407$ $4.093\pm 1.187$ $0.086\pm 0.004$ $11.948\pm 1.393$
+11 $0.8$ $3.392\pm 0.757$ $4.716\pm 2.079$ $0.089\pm 0.005$ $9.547\pm 0.752$
+0 $0.2$ $2.953\pm 0.425$ $3.816\pm 0.973$ $0.083\pm 0.002$ $9.816\pm 0.715$
+0 $0.5$ $2.938\pm 0.476$ $3.904\pm 1.120$ $0.088\pm 0.005$ $9.864\pm 0.764$
+0 $0.8$ $3.101\pm 0.539$ $3.855\pm 0.946$ $0.087\pm 0.004$ $9.280\pm 0.551$
+$0.5$ $0.2$ $3.771\pm 0.862$ $5.457\pm 1.704$ $\bm{0.079\pm 0.004}$ $9.262\pm 1.134$
+$0.5$ $0.5$ $3.090\pm 0.635$ $4.596\pm 2.357$ $0.084\pm 0.005$ $9.786\pm 1.067$
+$0.5$ $0.8$ $3.200\pm 0.403$ $3.555\pm 0.637$ $0.084\pm 0.004$ $9.269\pm 0.541$
+CTF-H 0 $3.244\pm 0.713$ $3.946\pm 1.671$ $0.089\pm 0.005$ $\bm{8.797\pm 0.612}$
+11 $5.200\pm 0.799$ $6.306\pm 1.037$ $0.123\pm 0.008$ $45.862\pm 13.765$
+$0.5$ $\bm{2.814\pm 0.414}$ $\bm{3.233\pm 0.567}$ $0.093\pm 0.005$ $10.319\pm 0.817$
+Table 30: Extrapolation via IVP Sampling at holdout time 8 for the Mouse Organogenesis dataset.
+Sampling Method $\bm{\lambda}$ $\bm{\alpha}$ Weighted $\bm{\mathcal{W}_{2}}$ $\bm{\mathcal{W}_{2}}$ MMD Energy
+IVP MOTFM $8.058\pm 2.415$ $27.283\pm 4.350$ $0.090\pm 0.003$ $28.890\pm 6.948$
+CTF-C 11 $0.2$ $9.988\pm 2.116$ $45.117\pm 15.750$ $0.084\pm 0.003$ $50.952\pm 12.577$
+11 $0.5$ $7.929\pm 1.464$ $35.572\pm 13.551$ $0.086\pm 0.002$ $42.040\pm 18.811$
+0 $0.8$ $7.729\pm 2.846$ $26.006\pm 5.753$ $0.086\pm 0.004$ $24.431\pm 3.700$
+11 $0.2$ $10.324\pm 1.979$ $51.492\pm 16.681$ $0.084\pm 0.002$ $45.002\pm 24.327$
+0 $0.5$ $7.799\pm 1.136$ $32.535\pm 9.496$ $0.085\pm 0.002$ $51.080\pm 13.594$
+0 $0.8$ $9.317\pm 5.319$ $32.195\pm 10.782$ $0.087\pm 0.004$ $28.873\pm 6.062$
+$0.5$ $0.2$ $10.762\pm 1.998$ $54.796\pm 13.489$ $0.089\pm 0.002$ $60.875\pm 28.610$
+$0.5$ $0.5$ $7.755\pm 1.519$ $31.069\pm 11.415$ $0.086\pm 0.002$ $39.806\pm 13.866$
+$0.5$ $0.8$ $8.326\pm 2.558$ $24.306\pm 6.339$ $0.086\pm 0.002$ $29.195\pm 8.088$
+CTF-H 0 $6.986\pm 1.295$ $23.806\pm 7.516$ $0.093\pm 0.004$ $\bm{22.466\pm 5.099}$
+11 $\bm{6.378\pm 1.813}$ $\bm{11.844\pm 6.567}$ $0.103\pm 0.008$ $30.752\pm 9.439$
+$0.5$ $7.749\pm 1.376$ $32.480\pm 8.322$ $\bm{0.084\pm 0.002}$ $26.129\pm 7.315$
+J.5 Liver Regeneration
+Table 31: 22-Wasserstein distances for different methods and configurations.
+Variant $\bm{\lambda}$ $\bm{\alpha}$ $\bm{\mathcal{W}_{2}}$
+EOT $34.303\pm 1.447$
+CTF-C 11 $0.2$ $34.444\pm 1.193$
+CTF-C 11 $0.5$ $33.956\pm 1.644$
+CTF-C 11 $0.8$ $34.628\pm 0.981$
+CTF-C 0 $0.2$ $34.241\pm 1.169$
+CTF-C 0 $0.5$ $32.741\pm 1.863$
+CTF-C 0 $0.8$ $33.717\pm 1.230$
+CTF-C $0.5$ $0.2$ $33.566\pm 1.043$
+CTF-C $0.5$ $0.5$ $33.841\pm 1.714$
+CTF-C $0.5$ $0.8$ $33.045\pm 1.643$
+CTF-H 0 $\bm{32.682\pm 1.471}$
+CTF-H 11 $33.480\pm 1.001$
+CTF-H $0.5$ $33.414\pm 0.995$